### v1

# The discrete periodic Pitman transform: invariances, braid relations, and Burke properties

Abstract.
We develop the theory of the discrete periodic Pitman transform, first introduced by Corwin, Gu, and the fifth author. We prove that, for polymers in a periodic environment, single-path and multi-path partition functions are preserved under the action of this transform on the weights in the polymer model. As a corollary, we prove that the discrete periodic Pitman transform satisfies the same braid relations that are satisfied for the full-line Pitman transform, shown by Biane, Bougerol, and O’Connell. Combined with a new inhomogeneous Burke property for the periodic Pitman transform, we prove a multi-path invariance result for the periodic inverse-gamma polymer under permutations of the column parameters. In the limit to the full-line case, we obtain a multi-path extension of a recent invariance result of Bates, Emrah, Martin, Seppäläinen, and the fifth author, in both positive and zero-temperature. Additionally, we give a combinatorial description of the distribution of the $2$-component jointly invariant measures for a discrete-time Markov chain.
1. Introduction
The study of the Pitman transform traces back to Pitman’s celebrated 2MX2M-X theorem [Pit75]. Since then, there have been numerous studies around this transform and its many variants. In the semi-discrete case, the positive-temperature lifting of the full-line Pitman transform was introduced by [MY01, O’C03] in the context of Brownian queuing process and later by [O’C12] in the setting of semi-discrete directed polymers. In the fully discrete case, the discrete geometric Pitman transform was connected to the geometric lifting of the classical Robinson-Schensted-Knuth (RSK) correspondence [Kir01] and studied by [NY04, COSZ14, OSZ14, Cor21] in the context of discrete directed polymer models.
Despite arising in different contexts, these Pitman transforms share a rich set of both algebraic and probabilistic properties that have enabled significant progress in the study of stochastic integrable systems, including directed percolation, random polymers, the KPZ equation, and so on. For example, the Burke property–the analogous product measure preservation property of Burke’s theorem for $M/M/1$ queues [Bur56]–has been proved in several contexts (see [HB76, OY01, OY02, KOR02, Mar09, DMO05, AD95] as well as [EJS24, BEM+25a] in an inhomogeneous setting) and has been used to determine the fluctuation exponents in several last-passage and polymer settings [CG06, BCS06, SV10, Sep12, EJS23, LNS23]. Moreover, the joint invariance of last-passage values/polymer partition functions under the zero-temperature/positive-temperature Pitman transforms provides key tools to the construction of central limiting objects in the KPZ universality class [DOV22] and for proving convergence of many different models to these limiting objects [DV21, Wu23, ACH24, Zha25].
The purpose of this paper is to demonstrate the rich algebraic and probabilistic structure of a periodic version of the Pitman transform, recently introduced in [CGS24]. Our first main result, Theorem 1.4, gives an invariance result for directed polymers in a periodic environment under a periodic Pitman transformation of the column weights. Previous results of this flavor in a non-periodic setting have appeared in several contexts and are discussed after the statement of the theorem. The key machinery is the encoding of geometric RSK in terms of matrix relations from [NY04]. Corwin [Cor21] previously used these tools to prove the analogous invariance for full-space polymers. The next main result, Theorem 1.6, uses this polymer invariance result to prove braid relations for the periodic Pitman transform, analogous to the relations enjoyed by the full-line Pitman transform [BBO05]. More specifically, we consider a collection of operators on the space $(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$ of bi-infinite sequences of $N$-periodic vectors and operators $\mathcal{P}_{k}$, which replace the $k$th and k+1k+1st vector with their Pitman transform. The braid relations, along with the fact that each $\mathcal{P}_{k}$ is an involution (proved in [CGS24], see Lemma A.2) shows that this collection of operators generates a copy of the infinite symmetric group. Both of these main results pass to zero-temperature limits, stated in Theorems 1.8 and 1.9. We then utilize the invariance of partition functions and a new inhomogeneous Burke property (Propositions 5.1, 5.2, and 5.3) to prove a distributional invariance of the inverse-gamma polymer partition function under the permutation of the row and column parameters in both the periodic and full-line cases (Theorems 1.10 and 1.11). In the full-line case, this gives a multi-path and positive temperature extension of a recent result of Bates, Emrah, Martin, Seppäläinen, and the fifth author [BEM+25b]. This result bears similarity to an invariance found previously by Dauvergne [Dau22]. See the discussion before and after Theorem 1.11 for more on the connection to other works.
Beginning with the work of Ferrari and Martin in particle systems [FM07, FM06, FM09], and adapted to last-passage percolation, polymers, and their scaling limits in [FS20, SS23, Bus24, BSS24, BFS25, GRASS25, CGS24], Pitman transforms have been used to describe the distribution of jointly invariant or multi-type invariant measures in these models. Indeed, the work [CGS24] introduced the periodic Pitman transform to describe the jointly invariant measures for the KPZ equation with periodic noise. A similar object was previously used in [FM07] to find the stationary measure for multi-species TASEP on a ring (see [CGS24, Appendix F] for more of a discussion on this connection). In Theorem 1.12, we develop a similar description of the jointly invariant measures for a discrete-time Markov chain in zero-temperature. Then, in Theorem 1.13, we give a combinatorial description of this invariant measure. We can obtain these joint stationary measures in the full-line settings by taking appropriate scaling limits of the measure we study in Theorem 1.13, and we hope that our description will lead to further understanding in future work.
1.1. The positive-temperature discrete periodic Pitman transform
For $N\in\mathbb{N}$, let $\mathbb{R}^{\mathbb{Z}_{N}}$ denote the set of $N$-tuples, together with periodic indexing. That is, for $\mathbf{X}\in\mathbb{R}^{\mathbb{Z}_{N}}$, $\mathbf{X}=(X_{1},X_{2},\ldots,X_{N})$, where the indices represent elements of $\mathbb{Z}_{N}$ (see Figure 1). For a vector $\mathbf{X}\in\mathbb{R}^{\mathbb{Z}_{N}}$ and indices $i,j\in\mathbb{Z}_{N}$, we set
$\mathbf{X}_{(i,j]}:=\sum_{\ell=i+1}^{j}X_{\ell},\quad\text{and}\quad\mathbf{X}_{[i,j]}:=\sum_{\ell=i}^{j}X_{\ell},$ (1.1)
where the sums are taken in cyclic order. We follow the convention that $\mathbf{X}_{(i,i]}=0$.
j][i
Figure 1. Representation of indices $[i,j]$, demonstrating periodicity.
We define the positive-temperature discrete periodic Pitman transform as the map $\mathcal{P}:\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}\to\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}$ defined by
$\mathcal{P}(\mathbf{W}_{1},\mathbf{W}_{2})$ $=(T(\mathbf{W}_{2},\mathbf{W}_{1}),D(\mathbf{W}_{2},\mathbf{W}_{1})),\quad\text{where}$ (1.2)
$T(\mathbf{X}_{1},\mathbf{X}_{2})$ $=X_{1,i-1}+\log\Biggl{(}\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{[i-1,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{[i,j]}}}\Biggr{)},\quad\text{and}$
$D(\mathbf{X}_{1},\mathbf{X}_{2})$ $=X_{2,i+1}+\log\Biggl{(}\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{(i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{(i-1,j]}}}\Biggr{)},\quad\text{where}\quad Y_{\ell}=X_{2,\ell+1}-X_{1,\ell}.$
This map is a shifted version of the map studied in [CGS24]; see Appendix A for the precise connection. Note that we reverse the order of the inputs $\mathbf{W}_{1},\mathbf{W}_{2}$ in the application of $T$ and $D$. Also, note that the maps $\mathcal{P},T,D$ depend on the length $N$ of the periodicity; we will subsume dependence on this value in the notation and fix $N$, unless otherwise noted.
1.2. Polymers on $\mathbb{Z}^{2}$ and periodic environments
Definition 1.1.
Consider the 2-dimensional integer lattice $\mathbb{Z}^{2}$. We say that an up-right path of length $L\in\mathbb{N}$ is an ordered collection of points
$\pi:=\left\{(a_{i},b_{i})_{i\in\llbracket 1,L\rrbracket}\mid(a_{i+1},b_{i+1})-(a_{i},b_{i})\in\{(1,0),(0,1)\}\ \forall 1\leq i<L\right\}$ (1.3)
where $(a_{i},b_{i})\in\mathbb{Z}^{2}\ \forall i$ and the notation $\llbracket a,b\rrbracket$ denotes the set of $k\in\mathbb{Z}$ satisfying $a\leq k\leq b$. We also define the set of all up-right paths to be $\Pi_{\mathbb{Z}^{2}}$.
For a fixed path $\pi$, we say that $(a_{1},b_{1})$ is the initial point of the path while $(a_{L},b_{L})$ is the end point of the path. For given start and end points $(i,j),(n,m)\in\mathbb{Z}^{2}$, we define
$\Pi_{(i,j),(n,m)}:=\{\pi\in\Pi_{\mathbb{Z}^{2}}\mid(a_{1},b_{1})=(i,j)\ ,(a_{L},b_{L})=(n,m)\}$ (1.4)
to be the set of paths starting at $(i,j)$ and ending at $(n,m)$, noting $\Pi_{(i,j),(n,m)}\neq\varnothing\iff m\geq j$ and $n\geq i$. In this case, L=m+njiL=m+n-j-i for any $\pi$ in $\Pi_{(i,j),(n,m)}$. We say that two paths $\pi_{1},\pi_{2}\in\Pi_{\mathbb{Z}^{2}}$ are non-intersecting if $\pi_{1}\cap\pi_{2}=\varnothing$.
Also of interest to us are sets of non-intersecting paths on our graph. We define a multi-path $(\pi_{1},\ldots,\pi_{k})$ to be a non-intersecting $k$-tuple of paths. See Figure 2 for a visual example.
As in the case of single paths, we can consider the set of paths between given start and endpoints. For $k\in\mathbb{N}$ and $k$-tuples of ordered pairs $U=\bigl{(}(a_{i},b_{i})\bigr{)}_{i\in\llbracket 1,k\rrbracket}$, $V=\bigl{(}(c_{i},d_{i})\bigr{)}_{i\in\llbracket 1,k\rrbracket}$, we define the set of multi-paths between them to be
$\Pi_{U,V}:=\{\pi=(\pi_{1},\dots,\pi_{k})\mid\pi_{i}\in\Pi_{(a_{i},b_{i}),(c_{i},d_{i})}\ \forall i\in\llbracket 1,k\rrbracket,\ \pi_{i}\cap\pi_{j}=\varnothing\ \forall i\neq j\in\llbracket 1,k\rrbracket\}$ (1.5)
We will often require that there is a unique pairing of initial points in $U$ and end points in $V$ such that a non-intersecting multi-path exists. In other words, we require that if $U$ and $V$ each have $k$ elements, then there does not exist a nontrivial permutation $\sigma$ in the $k$th-order symmetric group $S_{k}$ such that $\Pi_{U,V_{\sigma}}\neq\varnothing$ where $V_{\sigma}:=\bigl{(}(c_{\sigma(i)},d_{\sigma(i)})\bigr{)}_{i\in\llbracket 1,k\rrbracket}$. We denote the pairs of $k$-tuples of ordered pairs satisfying this property as
$\Psi:=\bigl{\{}(U,V)\mid U,V\subseteq\mathbb{Z}^{2},\ |U|=|V|,\ \Pi_{U,V}\neq\varnothing,\ \text{ and }\Pi_{U,V_{\sigma}}=\varnothing\ \forall\sigma\in S_{|U|}\setminus\{\text{Id}\}\bigr{\}}$ (1.6)
[…]
Figure 2. Example multi-path in $\mathbb{Z}^{2}$ from $\{(0,0),(2,0)\}$ to $\{(2,4),(4,4)\}$
Definition 1.2.
For real-valued weights $\mathbf{W}=\bigl{(}\mathbf{W}_{i}\in\mathbb{R}^{\mathbb{Z}}\bigr{)}_{i\in\mathbb{Z}}=\bigl{(}W_{(i,j)}\in\mathbb{R}\mid i,j\in\mathbb{Z}\bigr{)}$ corresponding to the vertices of the directed lattice $\mathbb{Z}^{2}$, we define the polymer partition function with inverse temperature $\beta>0$ between the initial point $(i,j)$ and endpoint $(n,m)$ in $\mathbb{Z}^{2}$ as
$Z^{\beta,\mathbf{W}}(n,m\mid i,j):=\sum_{\pi\in\Pi_{(i,j),(n,m)}}\prod_{(r,\ell)\in\pi}e^{\beta W_{(r,\ell)}}$ (1.7)
The partition function has a natural multi-path extension. For $(U,V)\in\Psi$, we define the multi-path partition function between $U$ and $V$ to be
$Z^{\beta,\mathbf{W}}(V\mid U):=\sum_{\pi\in\Pi_{U,V}}\prod_{(r,\ell)\in\pi}e^{\beta W_{(r,\ell)}}$ (1.8)
We typically take $\beta=1$, in which case we omit the superscript $\beta$ in the notation.
Remark 1.3.
In Definition 1.2, we defined a polymer for arbitrary weights on $\mathbb{Z}^{2}$. In all results below (with the exception of Theorem 1.11) we specialize to the case where the weights are periodic in the vertical direction.
1.3. Invariance of periodic polymer partition functions
Let $(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$ be the set of bi-infinite sequences of elements of $\mathbb{R}^{\mathbb{Z}_{N}}$. For $k\in\mathbb{Z}$, we define the operator $\mathcal{P}_{k}:(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}\to(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$ by its action:
$\mathcal{P}_{k}(\ldots,\mathbf{W}_{k-1},\mathbf{W}_{k},\mathbf{W}_{k+1},\mathbf{W}_{k+2},\ldots)=(\ldots,\mathbf{W}_{k-1},T(\mathbf{W}_{k+1},\mathbf{W}_{k}),D(\mathbf{W}_{k+1},\mathbf{W}_{k}),\mathbf{W}_{k+2},\ldots)$ (1.9)
Our first main result gives the invariance of the polymer partition functions under the iterated Pitman transform defined in (1.2).
Theorem 1.4.
Let $N\in\mathbb{N}$, and let $\mathbf{W}=(\mathbf{W}_{k}\in\mathbb{R}^{\mathbb{Z}_{N}}\mid k\in\mathbb{Z})=(W_{k,\ell}\mid k\in\mathbb{Z},\ell\in\mathbb{Z}_{N})$ be a collection of real-valued weights. Extend these periodically to weights on $\mathbb{Z}\times\mathbb{Z}$ by the condition $W_{k,\ell}=W_{j,\ell}$ whenever $k\equiv j\mod N$. If $k_{1},k_{2},\ldots,k_{m}\in\mathbb{Z}$ and the sets of initial points $U=\bigl{(}\bigl{(}a_{i},b_{i})_{i\in\llbracket 1,m\rrbracket}\bigr{)}$ and end points $V=\bigl{(}(c_{i},d_{i})_{i\in\llbracket 1,m\rrbracket}\bigr{)}$ satisfy $(U,V)\in\Psi$ and, for all $i,j\in\llbracket 1,m\rrbracket$, $a_{i}\neq k_{j}+1$ and $c_{i}\neq k_{j}$, then,
$Z^{\mathcal{P}_{k_{m}}\ldots\mathcal{P}_{k_{1}}\mathbf{W}}(V\mid U)=Z^{\mathbf{W}}(V\mid U).$ (1.10)
Remark 1.5.
See Figure 3 for a visual depiction of the invariance in Theorem 1.4. In Theorem 1.6 below, we show braid relations for the operators $\mathcal{P}_{k}$. Along with the fact that each $\mathcal{P}_{k}$ is an involution (see Lemma A.2), these relations imply that, if we have the equality of products of transpositions $(k_{m},k_{m}+1)(k_{m-1},k_{m-1}+1)\cdots(k_{1},k_{1}+1)=(k_{m}^{\prime},k_{m}^{\prime}+1)(k_{m-1}^{\prime},k_{m-1}^{\prime}+1)\cdots(k_{1}^{\prime},k_{1}^{\prime}+1)$, then $\mathcal{P}_{k_{m}}\ldots\mathcal{P}_{k_{1}}\mathbf{W}=\mathcal{P}_{k_{m}}^{\prime}\ldots\mathcal{P}_{k_{1}}^{\prime}\mathbf{W}$. Hence, we consider the action of $\mathcal{P}_{k_{m}}\ldots\mathcal{P}_{k_{1}}$ on $\mathbf{W}$ as the action of a general permutation, independent of its choice of decomposition. In this case, (1.12) gives an alternate condition for the set of endpoints whose partition function is preserved. An example is given in Figure 4 below. This connection is made precise in Lemma 5.4.
Figure 3. A depiction of the polymer invariance in Theorem 1.4. Each column is associated to a periodic vector of weights, $\mathbf{W}_{k}$. The horizontal lines are used to indicate a periodic environment for the polymer with $N=4$. In this example, we take $m=1$ and consider the action of the operator $\mathcal{P}_{3}$, which replaces the weights in columns $3$ and $4$ with $T(\mathbf{W}_{4},\mathbf{W}_{3})$ and $D(\mathbf{W}_{4},\mathbf{W}_{3})$, respectively. The partition function for each pair of points and the multi-path partition function for this collection of $3$ pairs are both preserved under this transformation of the weights.
Theorem 1.4 is proved in Section 2. It is the first of its kind for a polymer in a periodic environment. The invariance under Pitman transformations in other contexts has been widely studied. The concept can be traced back to queuing theory from the works [Web79, TW87], where it was shown that the law of the departures process from two $\cdot/M/1$ queues in series is invariant under swapping of the service rates of the two queues. In the polymer context, this type of result can be traced back to the work of Noumi and Yamada [NY04], who gave a description of the geometric Robinson-Schendsted-Knuth correspondence in terms of certain matrix product relations. Corwin [Cor21] used this to give an analogous result to our Theorem 1.4 for a full-line version of the discrete Pitman transform. The proof of Theorem 1.4 likewise uses the machinery from [NY04], but we instead work with $\mathbb{Z}\times\mathbb{Z}$ periodic matrices instead of finite-dimensional matrices.
Theorem 1.4 limits to a zero-temperature analogue in Theorem 1.8. This zero-temperature limit is the periodic analogue of invariances shown in [DOV22, DNV22, DV24, BEM+25a]. The result in [DOV22] is particularly significant, as it was key in constructing the directed landscape, the central limiting object of the KPZ universality class.
It would be interesting to see if Theorem 1.4 could be the starting point for constructing and proving convergence to the periodic directed landscape, but we note that the action of this transform is qualitatively different than that in [DOV22]. Specifically, in the full-line case, the action of the iterated Pitman transforms and independent collection of Brownian motions and transforms them to the Brownian melon–a set of nonintersecting paths whose geodesic structure is more amenable to asymptotic analysis. By contrast, in this periodic setting when we take the weights to be i.i.d. with the inverse-gamma distribution, the transformed weights are equal in distribution to the original weights. This is the Burke property shown in [CGS24, Proposition 5.12], and extended to an inhomogeneous setting in Proposition 5.1. As a corollary, we obtain a permutation invariance result for the multi-path inhomogeneous inverse-gamma polymer below in Theorem 1.10. This extends the previous result of [BEM+25a, Theorem 3.1] to a positive-temperature and multi-path setting. In this sense, the periodic Pitman transform is more analogous to a limiting case of the full-line Pitman transform, where the initial point of a random walk is sent to -$\infty$ and the output of the transform is recentered at 0 (see [DNV22, DV24]). This limiting full-line setting is where we see analogous Burke properties, as in [OY01, OY02].
1.4. Braid relations
We now describe our next result, which gives a braid relation for the Pitman operators defined in (1.9). The next theorem gives a braid relation for these operators, which is the same as that satisfied for the full-line Pitman transform [BBO05]. See also [DV24]. Combined with the fact that the map $\mathcal{P}_{k}$ is an involution (see Lemma A.2), the operators $\mathcal{P}_{k}$ generate a representation of the infinite symmetric group (note that the same braid relations are satisfied by the nearest-neighbor transpositions $(k,k+1)$).
Theorem 1.6.
For $k\in\mathbb{Z}$,
$\mathcal{P}_{k}\mathcal{P}_{k+1}\mathcal{P}_{k}=\mathcal{P}_{k+1}\mathcal{P}_{k}\mathcal{P}_{k+1}.$
Theorem 1.6 is proved in Section 3. The key is the matrix relation in Corollary 2.5 below, which is also used to prove Theorem 1.4. When written out, the braid relation consists of three equalities ((3.1)-(3.3) below). The third of these equalities was previously proved in [CGS24, Proposition 5.13]. With that as our starting point, we iterate the matrix relation in Corollary 2.5 in two different ways to obtain and use the interpretation of the matrix product as the polymer partition function to obtain the other two equalities.
1.5. Zero-temperature limits
We also define the zero-temperature discrete periodic Pitman transform as the map $\widetilde{\mathcal{P}}:\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}\to\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}$ defined by
$\widetilde{\mathcal{P}}(\mathbf{W}_{1},\mathbf{W}_{2})$ $=\bigl{(}\widetilde{T}(\mathbf{W}_{2},\mathbf{W}_{1}),\widetilde{D}(\mathbf{W}_{2},\mathbf{W}_{1})\bigr{)},\quad\text{where}$ (1.11)
$\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2})$ $=X_{1,i-1}+\max_{j\in\mathbb{Z}_{N}}Y_{[i-1,j]}-\max_{j\in\mathbb{Z}_{N}}Y_{[i,j]},\quad\text{and}$
$\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})$ $=X_{2,i+1}+\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}-\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]},\quad\text{where}\quad Y_{\ell}=X_{2,\ell+1}-X_{1,\ell}.$
One immediately sees that we can obtain $\widetilde{\mathcal{P}}$ as a limiting case of the positive-temperature discrete periodic Pitman transform:
$\widetilde{\mathcal{P}}(\mathbf{W}_{1},\mathbf{W}_{2})=\lim_{\beta\to\infty}\frac{1}{\beta}\mathcal{P}(\beta\mathbf{W}_{1},\beta\mathbf{W}_{2})$
Furthermore, this convergence is uniform as a function of the vectors $\mathbf{W}_{1},\mathbf{W}_{2}$. See Lemma A.1 for details.
Definition 1.7.
For real-valued weights $\mathbf{W}=\bigl{(}W_{(i,j)}\in\mathbb{R}\mid i,j\in\mathbb{Z}\bigr{)}$ corresponding to the vertices of the directed lattice $\mathbb{Z}^{2}$, we define the last-passage percolation (LPP) time between the points $(i,j)$ and $(n,m)$ in $\mathbb{Z}^{2}$ as the maximal sum over weights along an up-right path between $(i,j)$ and $(n,m)$. That is,
$G^{\mathbf{W}}(n,m\mid i,j)\coloneqq\max\limits_{\pi\in\Pi_{(i,j),(n,m)}}\sum_{(r,\ell)\in\pi}W_{(r,\ell)}$
We similarly define multi-path last-passage times analogously as in (1.8):
$G^{\mathbf{W}}(V\mid U)=\max_{\pi\in\Pi_{U,V}}\sum_{(r,\ell)\in\pi}W_{(r,\ell)}.$
With this notation defined, we get the following zero-temperature analogues of Theorems 1.4 and 1.6:
Theorem 1.8.
The same invariance in Theorem 1.4 holds, $Z^{\mathbf{W}}$ is replaced by $G^{\mathbf{W}}$, and $T,D$ are replaced by $\widetilde{T},\widetilde{D}$.
Theorem 1.9.
For $k\in\mathbb{Z}$, define the map $\widetilde{\mathcal{P}}_{k}:(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}\to(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$ by (1.9), with $T,D$ replaced by $\widetilde{T},\widetilde{D}$. Then, the same braid relations in Theorem 1.6 hold.
The details of the limit transitions to Theorems 1.8 and 1.9 are handled in Section 4.
1.6. Permutation invariance
In this section, we consider an inhomogeneous, periodic inverse-gamma polymer. Let $\lambda>0$ and $\mathbf{a}\in\mathbb{R}^{\mathbb{Z}}$ and $\mathbf{b}\in\mathbb{R}^{\mathbb{Z}_{N}}$ be sequences of real numbers such that $a_{i}+b_{j}>0$ for $i\in\mathbb{Z}$ and $j\in\mathbb{Z}_{N}$. We take independent weights $\mathbf{W}^{\mathbf{a},\mathbf{b}}=\bigl{(}W_{(i,j)}\mid i\in\mathbb{Z},\ j\in\mathbb{Z}_{N}\bigr{)}$ such that $W_{(i,j)}\sim\text{Log-Inv-Gamma}(a_{i}+b_{j},\lambda)$ for $i\in\mathbb{Z}$ and $j\in\mathbb{Z}_{N}$ (see the notation section below). Extend these periodically to weights on $\mathbb{Z}\times\mathbb{Z}$. We call a permutation $\sigma:\mathbb{Z}\to\mathbb{Z}$ a finite permutation if $\sigma$ fixes all but finitely many coordinates. For a finite permutation $\sigma$, we let $\mathcal{U}^{\sigma}$ be the set of tuples $(U,V)\in\Psi$, such that
[…]
Theorem 1.10.
For vectors $\mathbf{a}\in\mathbb{R}^{\mathbb{Z}}$ and $\mathbf{b}\in\mathbb{R}^{\mathbb{Z}_{N}}$ satisfying $a_{i}+b_{j}>0$ for $i\in\mathbb{Z}$ and $j\in\mathbb{Z}_{N}$, define periodic weights $\mathbf{W}^{\mathbf{a},\mathbf{b}}$ as above and extend them to $\mathbb{Z}\times\mathbb{Z}$. Then, for any finite permutation $\sigma:\mathbb{Z}\to\mathbb{Z}$, we have
$\Bigl{(}Z^{\mathbf{W}^{\mathbf{a},\mathbf{b}}}(V\mid U):(U,V)\in\mathcal{U}^{\sigma}\Bigr{)}\overset{d}{=}\Bigl{(}Z^{\mathbf{W}^{\sigma\mathbf{a},\mathbf{b}}}(V\mid U):(U,V)\in\mathcal{U}^{\sigma}\Bigr{)},$
where $\sigma\mathbf{a}=(a_{\sigma(i)})_{i\in\mathbb{Z}}$. The same invariance holds for last-passage percolation if the weights satisfy either $W_{(i,j)}\sim\text{Exp}(a_{i}+b_{j})$ with $a_{i}+b_{j}>0$ for all $i,j$, or $W_{(i,j)}\sim\text{Geom}(q_{i}r_{j})$ for all $i,j$, where $q_{i}r_{j}\in(0,1)$ for all $i,j$.
The condition (1.12) ensures that, between the pairs of points in $(U,V)$, the column rates are either permuted among themselves or are left unchanged. See Figure 4 for an example.
Figure 4. An example of an element $(U,V)\in\mathcal{U}^{\sigma}$, where $U$ and $V$ each contain 33 elements. The gray horizontal lines are used to indicate a periodic environment with $N=4$. In this figure, the column parameters of columns $3,5$, and 66 (blue) are permuted, and the column parameters in columns $8$ and 99 (red) are permuted. The condition is that, for each of the three endpoints $(x,y)$ in $U$ (lower left), $\sigma(\mathbb{Z}_{<x})=\mathbb{Z}_{<x}$, while for each of the three endpoints $(x,y)$ in $V$ (upper right), $\sigma(\mathbb{Z}_{>x})=\mathbb{Z}_{>x}$. The distributional equality in Theorem 1.10 holds jointly over all endpoint pairs $(U,V)\in\Psi$ satisfying this property.
We can allow for permuting the $\mathbf{b}$ parameters as well, with doubly periodic weights $(W_{(i,j)}\mid i\in\mathbb{Z}_{M},j\in\mathbb{Z}_{N})$ for $M,N\in\mathbb{N}$ when we restrict to points that lie inside the box $\llbracket 0,M\rrbracket\times\llbracket 0,N\rrbracket$. In this case, the model isn’t actually periodic at all; it is just a restriction of the full-space polymer to a finite grid. For finite permutations $\sigma,\tau:\mathbb{Z}\to\mathbb{Z}$, define the set $\mathcal{U}^{\sigma,\tau}$ to be the set of tuples $(U,V)$, where $U,V$ are each sets of endpoint pairs in $\mathbb{Z}\times\mathbb{Z}$ satisfying (1.12) as well as
[…]
The following is an extension of the recent result [BEM+25a, Theorem 3.1] to multi-paths and to the positive-temperature case. The technique follows the same outline as developed in [BEM+25a], as we use a coupling of the weights given by applying a Pitman transform to neighboring rows and columns. However, the coupling we use in this proof is different because we treat the weights as periodic outside a large box and apply the periodic Pitman transform.
Theorem 1.11.
For vectors $\mathbf{a}\in\mathbb{R}^{\mathbb{Z}}$ and $\mathbf{b}\in\mathbb{R}^{\mathbb{Z}}$ satisfying $a_{i}+b_{j}>0$ for $i,j\in\mathbb{Z}$ define independent weights $\mathbf{W}^{\mathbf{a},\mathbf{b}}=\bigl{(}W_{(i,j)}:i,j\in\mathbb{Z}\bigr{)}$ such that $W_{(i,j)}\sim\text{Log-Inv-Gamma}(a_{i}+b_{j},\lambda)$ for all $i,j$. Then, for any finite permutations $\sigma,\tau:\mathbb{Z}\to\mathbb{Z}$, we have
$\Bigl{(}Z^{\mathbf{W}^{\mathbf{a},\mathbf{b}}}(V\mid U):(U,V)\in\mathcal{U}^{\sigma,\tau}\Bigr{)}\overset{d}{=}\Bigl{(}Z^{\mathbf{W}^{\sigma\mathbf{a},\tau\mathbf{b}}}(V\mid U):(U,V)\in\mathcal{U}^{\sigma,\tau}\Bigr{)},$
where $\sigma\mathbf{a}=(a_{\sigma(i)})_{i\in\mathbb{Z}}$ and $\tau\mathbf{b}=(b_{\tau(j)})_{j\in\mathbb{Z}}$. The same invariance holds for last-passage percolation if the weights satisfy either $W_{(i,j)}\sim\text{Exp}(a_{i}+b_{j})$ for all $i,j$ or $W_{(i,j)}\sim\text{Geom}(q_{i}r_{j})$ for all $i,j$, where $q_{i}r_{j}\in(0,1)$ for all $i,j$.
Theorems 1.10 and 1.11 are proved in Section 5.1. When the initial point is fixed, invariances of this flavor can be deduced from the symmetry of Schur functions and Whittaker functions (see, for example, [BP08, COSZ14, OSZ14, BC14]). A similar type of invariance for last-passage percolation was discovered in the work of Dauvergne [Dau22, Theorem 1.5], motivated by a shift-invariance result previously proved in Borodin, Gorin, and Wheeler [BGW22]. Connections between these results are discussed further in [BEM+25a, Section 3].
1.7. Distribution of the jointly invariant measures for a discrete-time Markov chain
The original motivation for introducing the map $D$ in [CGS24] was to understand the jointly invariant measures for the KPZ equation with periodic noise. This was done by first finding jointly invariant measures for the periodic O’Connell-Yor polymer. As a by-product of the analysis, it was shown that the same measures are jointly invariant for a discrete-time Markov chain (see [CGS24, Section 4]). In this section, we study the zero-temperature analogue of that Markov chain. The state space is $\mathbb{R}_{\geq 0}^{\mathbb{Z}_{N}}\times\mathbb{R}_{\geq 0}^{\mathbb{Z}_{N}}$, which is preserved by the chain by Lemma A.3. If the state of the Markov chain at time $m$ is $(\mathbf{U}_{1}^{(m)},\mathbf{U}_{2}^{(m)})$, and $\mathbf{W}\in\mathbb{R}^{\mathbb{Z}_{N}}_{\geq 0}$ is independent of the chain up to time $m$, then the time $m+1$ state is defined as
$\bigl{(}\mathbf{U}^{(m+1)}_{1},\mathbf{U}^{(m+1)}_{2}\bigr{)}=\bigl{(}\widetilde{D}(\mathbf{W},\mathbf{U}_{1}^{(m)}),\widetilde{D}(\mathbf{W},\mathbf{U}_{2}^{(m)})\bigr{)}.$ (1.14)
Marginally, $\bigl{(}\mathbf{U}_{1}^{(m)}\bigr{)}_{m\geq 0}$ and $\bigl{(}\mathbf{U}_{2}^{(m)}\bigr{)}_{m\geq 0}$ are each Markov chains with the same dynamics. We call the invariant measures for this coupled Markov chain jointly invariant measures because their joint distribution is preserved under the Markov chain where each component evolves using the same driving weights.
For $\mathbf{X}\in\mathbb{R}^{\mathbb{Z}_{N}}$, define
$\mathfrak{s}(\mathbf{X})=\sum_{i\in\mathbb{Z}_{N}}X_{i}.$ (1.15)
For $\theta\in\mathbb{R}$, define
$\mathbb{R}^{\mathbb{Z}_{N}}_{\theta}$ $=\{\mathbf{X}\in\mathbb{R}^{\mathbb{Z}_{N}}\mid\mathfrak{s}(\mathbf{X})=\theta\},\quad\text{and}$ (1.16)
$\mathbb{R}^{\mathbb{Z}_{N}}_{\theta,\geq 0}$ $=\{\mathbf{X}\in\mathbb{R}^{\mathbb{Z}_{N}}_{\theta}\mid X_{i}\geq 0\text{ for }i\in\mathbb{Z}_{N}\}$
We also define $\mathbb{Z}^{\mathbb{Z}_{N}}_{\theta}=\mathbb{R}_{\theta}^{\mathbb{Z}_{N}}\cap\mathbb{Z}^{\mathbb{Z}_{N}}$ and $\mathbb{Z}^{\mathbb{Z}_{N}}_{\theta,\geq 0}:=\mathbb{R}^{\mathbb{Z}_{N}}_{\theta,\geq 0}\cap\mathbb{Z}^{\mathbb{Z}_{N}}$. We now state the following theorem.
Theorem 1.12.
Let $\theta_{1},\theta_{2}\geq 0$ and $q\in(0,1)$. At each step $m$, let the driving weights $\mathbf{W}$ in the evolution of the Markov chain (1.14) be an i.i.d. sequence of $\text{Geom}(q)$ random variables. Let $\mathbf{X}_{1}$ be uniformly distributed on the set $\mathbb{Z}^{\mathbb{Z}_{N}}_{\theta_{1},\geq 0}$, and let $\mathbf{X}_{2}$ be uniformly distributed on the set $\mathbb{Z}^{\mathbb{Z}_{N}}_{\theta_{2},\geq 0}$, then the law of $\bigl{(}\mathbf{X}_{1},\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})\bigr{)}$ is a jointly invariant measure for the Markov chain. Furthermore, we have the distributional equality
$\bigl{(}\mathbf{X}_{1},\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})\bigr{)}\overset{d}{=}\bigl{(}\widetilde{D}(\mathbf{X}_{2},\mathbf{X}_{1}),\mathbf{X}_{2}\bigr{)}$ (1.17)
The same result holds when $\mathbf{W}$ is an i.i.d. sequence of $\text{Exp}(\alpha)$ random variables for any $\alpha>0$, if we instead take $\mathbf{X}_{1}$ to be uniformly distributed on $\mathbb{R}_{\theta_{1},\geq 0}^{\mathbb{Z}_{N}}$ and $\mathbf{X}_{2}$ to be uniformly distributed on $\mathbb{R}_{\theta_{2},\geq 0}^{\mathbb{Z}_{N}}$.
Theorems 1.12, along with Theorem 1.13 below, is proved in Section 6. In fact, we prove that the same invariance holds when $\mathbf{X}_{1},\mathbf{X}_{2}$ are sequences of i.i.d. geometric random variables. The uniform distribution is obtained by conditioning on $\mathfrak{s}(\mathbf{X}_{1})$ and $\mathfrak{s}(\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2}))$. These vector sums are preserved by the map $\widetilde{D}$ and hence also the Markov chain (Lemma A.4), so the extremal invariant measures come when we consider $\mathbf{X}_{1},\mathbf{X}_{2}$ to be uniform over all nonnegative vectors with a given sum. As in [CGS24], we can also consider the joint Markov chain with more than two components. In this case, the invariant measures will be built by iterating the map $\widetilde{D}$. Specifically, the invariant initial conditions will be of the form $(\mathbf{X}_{1},\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2}),\widetilde{D}(\mathbf{X}_{1},\widetilde{D}(\mathbf{X}_{2},\mathbf{X}_{3})),\ldots$ For brevity, we include only the case of two components to motivate the next theorem.
In this next result, we give an explicit combinatorial description of these jointly invariant measures from Theorem 1.12. We leave it as an open problem to develop a similar combinatorial description for the jointly invariant measures with more than $2$ components. In analogy with [CGS24], when $\mathbf{X}_{1}$ is uniform on $\mathbb{R}_{\theta_{1},\geq 0}^{\mathbb{Z}_{N}}$ and $\mathbf{X}_{2}$ is uniform on $\mathbb{R}_{\theta_{2},\geq 0}^{\mathbb{Z}_{N}}$, the law of $(\mathbf{X}_{1},\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2}))$ should be a jointly invariant measure for Brownian last-passage percolation in a periodic environment of Brownian motions. It can also be shown (see [CGS24, Appendix F]) that as $N\to\infty$, the invariant measure from Theorem 1.12 converges to the jointly invariant measure for i.i.d. exponential last-passage percolation studied in [FS20], which in turn, converges under diffusive scaling to the $2$-component marginal of the stationary horizon [Bus24]–a coupling of Brownian motions with drift that describes the jointly invariant measures for the KPZ fixed point [BSS24].
Below, we take $\theta_{2}>\theta_{1}$, since we have the distributional equality $\bigl{(}\mathbf{X}_{1},\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})\bigr{)}\overset{d}{=}\bigl{(}\widetilde{D}(\mathbf{X}_{2},\mathbf{X}_{1}),\mathbf{X}_{2}\bigr{)}$ in Theorem 1.12. When $\theta_{1}=\theta_{2}$, we have $\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})=\mathbf{X}_{1}$ via Corollary 6.4 below, so the joint distribution is trivial in this case.
Below, for $\theta_{1}\leq\theta_{2}$, we define the space
$\mathcal{Z}_{\theta_{1},\theta_{2}}:=\{(\mathbf{u}_{1},\mathbf{u}_{2})\in\mathbb{Z}^{\mathbb{Z}_{N}}_{\theta_{1},\geq 0}\times\mathbb{Z}^{\mathbb{Z}_{N}}_{\theta_{2},\geq 0}\mid\mathbf{u}_{1}\leq\mathbf{u}_{2}\},$ (1.18)
where $\mathbf{u}_{1}\leq\mathbf{u}_{2}$ means $u_{1,i}\leq u_{2,i}$ for all $i\in\mathbb{Z}_{N}$.
Theorem 1.13.
[…]
$\mathbb{P}(\mathbf{X}_{1}=\mathbf{u}_{1},\mathbf{\widetilde{D}}=\mathbf{u}_{2})=\frac{\alpha(\mathbf{u}_{1},\mathbf{u}_{2})}{\binom{\theta_{1}+N-1}{N-1}\binom{\theta_{2}+N-1}{N-1}}$
where $\alpha(\mathbf{u}_{1},\mathbf{u}_{2})$ is the following integer, split into two cases:
  1. (i)
    If $u_{1,i}<u_{2,i}$ for all $i\in\mathbb{Z}_{N}$, then $\alpha(\mathbf{u}_{1},\mathbf{u}_{2})=1$.
  2. (ii)
[…]
    $\alpha(\mathbf{u}_{1},\mathbf{u}_{2})=\prod_{n=1}^{s}\sum_{a_{1}=-u_{1,\ell_{m}}}^{0}\sum_{a_{2}=-u_{1,\ell_{m}+1}}^{-a_{1}}\sum_{a_{3}=-u_{1,\ell_{m}+2}}^{-(a_{1}+a_{2})}\cdots\sum_{a_{k_{m+1}-\ell_{m}+N\mathbf{1}_{m=s}}=-u_{1,k_{m+1}-1+N1_{m=s}}}^{-(a_{1}+a_{2}+\cdots+a_{k_{m+1}-\ell_{m}-1+N1_{m=s}})}1.$
    Alternatively, $\alpha(\mathbf{u}_{1},\mathbf{u}_{2})$ is the number of $s$-tuples of vectors $(\mathbf{a}_{1},\mathbf{a}_{2},\ldots,\mathbf{a}_{s})$, where, for $m\in\llbracket 1,s\rrbracket$, the vector $\mathbf{a}_{m}$ lies in the set $\mathbb{Z}^{k_{m+1}-\ell_{m}+N1_{m=s}}$ and satisfies the following conditions:
    $a_{m,1}\leq 0$ (1.19)
    $a_{m,1}+a_{m,2}\leq 0$
    $\vdots$
    $a_{m,1}+a_{m,2}+\cdots+a_{m,k_{m+1}-\ell_{m}+N1_{m=s}}\leq 0,\text{ and}$
    $a_{m,p}\geq-u_{1,\ell_{m}+p-1}\text{ for }1\leq p\leq k_{m+1}-\ell_{m}+N1_{m=s}.$
[…]
Figure 5. Comparison of $\mathbf{u}_{1}$ to $\mathbf{u}_{2}$ with $s$ alternating equality/inequality segments. We place an equal sign at index $i$ if $u_{1,i}=u_{2,i}$ and an inequality sign at index $j$ if $u_{1,j}<u_{2,j}$.
Example 1.14.
We demonstrate a sample application of Theorem 1.13 in the case of $N=2$ and $\theta_{1},\theta_{2}\in\mathbb{Z}_{\geq 0}$ where $\theta_{1}<\theta_{2}$. Take some $x_{1},x_{2},y_{1},y_{2}\in\mathbb{Z}_{\geq 0}$ satisfying $x_{1}\leq y_{1},x_{2}\leq y_{2}$ and $x_{1}+x_{2}=\theta_{1},y_{1}+y_{2}=\theta_{2}$.
If $x_{1}=y_{1}$, then $x_{2}<y_{2}$. In the notation of Theorem 1.13, $s=1,k_{1}=1$, and $\ell_{1}=2$, and the only relevant condition is $x_{2}=-u_{1,2}\leq a_{1,1}\leq 0,$ giving $x_{2}+1$ possible values for $a_{1,1}$. We can symmetrically describe the possible values when $x_{1}<y_{1},x_{2}=y_{2}$. Then, we observe that
$\mathbb{P}(\mathbf{X}_{1}=(x_{1},x_{2}),\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})=(y_{1},y_{2}))=\begin{cases}\frac{x_{1}+1}{(\theta_{1}+1)(\theta_{2}+1)}&x_{1}<y_{1}\text{ and }x_{2}=y_{2}\\ \frac{x_{2}+1}{(\theta_{1}+1)(\theta_{2}+1)}&x_{1}=y_{1}\text{ and }x_{2}<y_{2}\\ \frac{1}{(\theta_{1}+1)(\theta_{2}+1)}&x_{1}<y_{1}\text{ and }x_{2}<y_{2}\end{cases}$
[…]
Example 1.15.
We similarly demonstrate a computation in the case of $N=3$ and $\theta_{1}<\theta_{2}\in\mathbb{Z}_{\geq 0}$. Take $x_{1},x_{2},x_{3},y_{1},y_{2},y_{3}\in\mathbb{Z}_{\geq 0}$ satisfying $x_{1}\leq y_{1}$,  $x_{2}\leq y_{2},\ x_{3}\leq y_{3}$ and $x_{1}+x_{2}+x_{3}=\theta_{1},\ y_{1}+y_{2}+y_{3}=\theta_{2}$. Write $\mathbf{u}_{1}=(x_{1},x_{2},x_{3}),\mathbf{u}_{2}=(y_{1},y_{2},y_{3})$. Then, we consider the cases in which the only equality is $x_{1}=y_{1}$ and when the only equalities are $x_{1}=y_{1},\ x_{2}=y_{2}$. We can achieve all other sets of equalities and inequalities through cyclic permutations of $\mathbf{u}_{1},\mathbf{u}_{2}$ (Lemma A.6).
When we have that $x_{1}=y_{1},\ x_{2}<y_{2},\ x_{3}<y_{3}$, then we have that $k_{1}=1,\ell_{1}=3$, so the conditions (1.19) are exactly $x_{3}=-u_{1,3}\leq a_{1,1}\leq 0$, giving $x_{3}+1$ distinct possible values. If $x_{1}=y_{1},\ x_{2}=y_{2},\ x_{3}<y_{3}$, then we have that $k_{1}=2,\ell_{1}=3$, so the conditions (1.19) are $x_{3}=-u_{1,3}\leq a_{1,1}\leq 0,\ a_{1,1}+a_{1,2}\leq 0,\ -x_{1}=-u_{1,1}\leq a_{1,2}$. This gives a count of
$\alpha(\mathbf{u}_{1},\mathbf{u}_{2})$ $=\sum_{a_{1}=-x_{3}}^{0}\sum_{a_{2}=-x_{1}}^{-a_{1}}1=\sum_{a_{1}=-x_{3}}^{0}(x_{1}-a_{1}+1)=(x_{3}+1)(x_{1}+1)+\frac{x_{3}(x_{3}+1)}{2}.$
Noting that there are $\binom{\theta_{1}+2}{2}\binom{\theta_{2}+2}{2}$ elements in $\mathbb{Z}_{\theta_{1},+}^{\mathbb{Z}_{N}}\times\mathbb{Z}_{\theta_{2},+}^{\mathbb{Z}_{N}}$, we have the computation
($\binom{\theta_{1}+2}{2}\binom{\theta_{2}+2}{2}\cdot\mathbb{P}$( $\mathbf{X}_{1}=(x_{1},x_{2},x_{3}),\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})=(y_{1},y_{2},y_{3}))$
$=\begin{cases}(x_{3}+1)(x_{1}+1)+\frac{x_{3}(x_{3}+1)}{2}&x_{1}=y_{1},x_{2}=y_{2},\text{ and }x_{3}<y_{3}\\ \hfil x_{3}+1&x_{1}=y_{1},x_{2}<y_{2},\text{ and }x_{3}<y_{3}\\ \hfil 1&x_{1}<y_{1},x_{2}<y_{2},\text{ and }x_{3}<y_{3}\end{cases}$
1.8. Organization of the paper
Section 2 is focused on proving the partition function invariance in Theorem 1.4. The braid relations in Theorem 1.6 are proved in Section 3. The details of the zero-temperature limit transitions are handled in Section 4. Section 5 proves inhomogeneous Burke properties for the positive and zero-temperature periodic Pitman transforms and then uses those to prove the permutation invariance in Theorems 1.10 and 1.11. Theorems 1.12 and 1.13 are proved in Section 6. The Appendix contains some prior results from the work [CGS24] and some auxiliary technical results.
1.9. Notation and conventions
We use $\overset{d}{=}$ to indicate that two random variables are equal in distribution. Below, we define our notation for geometric, exponential, and log-inverse-gamma random variables.
Definition 1.16 (Geometric Random Variable).
A random variable $X\in\mathbb{Z}_{\geq 1}$ has a geometric distribution with parameter $q\in[0,1]$ if for all $n\in\mathbb{Z}_{\geq 0}$, we have that
$\mathbb{P}(X=n)=q^{n}(1-q)$
In this case, we write $X\sim\text{Geom}(q)$.
Definition 1.17 (Exponential Random Variable).
A random variable $X\in\mathbb{R}$ has an exponential distribution with rate $\alpha>0$ if
$\mathbb{P}(X\in dx)=\begin{cases}\alpha e^{-\alpha x}&x\geq 0\\ 0&x<0\end{cases}$
In this case, we write $X\sim\text{Exp}(\alpha)$.
Definition 1.18 (Log-Inverse-Gamma Random Variable).
A random variable $X\in\mathbb{R}$ has a log-inverse-gamma distribution with shape $\gamma>0$ and scale $\lambda>0$ if
$\mathbb{P}(X\in dx)=\frac{\lambda^{\gamma}}{\Gamma(\gamma)}e^{-\gamma x}e^{-\lambda e^{-x}}$
In this case, we write $X\sim\text{Log-Inv-Gamma}(\gamma,\lambda)$.
1.10. Acknowledgements
This work was completed during the Columbia Math Summer Undergraduate Research Program in Summer 2025, and we acknowledge support for this program from the Columbia University math department. We thank George Dragomir for organizing this program. During the program, E.S. and X.Z. were partially supported by Ivan Corwin’s Simons Investigator Grant #929852. In addition, X.Z. was supported by Ivan Corwin’s NSF grant DMS:2246576. E.R.E. was supported by Princeton University’s Office of Undergraduate Research OURSIP Program through the Frances Lane Summer Research Fund. We also thank Ivan Corwin and Amol Aggarwal for helpful discussions related to this project.
2. Invariance of polymer partition function
In this section, we prove Theorem 1.4. The key to our proof is a matrix encoding of polymer partition functions, analogous to that developed in [NY04]. Section 2.1 describes the necessary algebraic properties of these matrices, and then Section 2.2 makes the connection to the polymer partition function and completes the proof.
2.1. Matrix encoding
In order to prove our polymer partition function invariance, we follow the matrix encoding method introduced at Noumi and Yamada in [NY04]. We will work with infinite matrices.
Definition 2.1.
For $\mathbf{x}\in\mathbb{R}^{\mathbb{Z}}$, we introduce two functions $E:\mathbb{R}^{\mathbb{Z}}\to\mathbb{R}^{\mathbb{Z}\times\mathbb{Z}}$ and $H:\mathbb{R}^{\mathbb{Z}}\to\mathbb{R}^{\mathbb{Z}\times\mathbb{Z}}$ as follows:
$E(\mathbf{x}):=\sum_{i\in\mathbb{Z}}e^{x_{i}}E_{i,i}+\sum_{i\in\mathbb{Z}}E_{i,i+1}\,,\quad H(\mathbf{x}):=\sum_{i\leq j}e^{x_{i}+x_{i+1}+\ldots+x_{j}}E_{i,j}$ (2.1)
where $E_{p,q}$ is the $\mathbb{Z}\times\mathbb{Z}$ matrix with a $1$ at index $(p,q)$ and 0s elsewhere.
These matrix functions admit natural extensions for $\mathbf{x}\in\mathbb{R}^{\mathbb{Z}_{N}}$, where, for $k\notin\{1,\ldots,N\}$, we simply set $x_{k}\coloneqq x_{\ell}$, where $\ell$ is the unique element of $\{1,\ldots,N\}$ satisfying $\ell\equiv k\pmod{N}$. All of the same properties that hold for $\mathbb{Z}$ trivially also hold for $\mathbb{Z}_{N}$, via the inclusion map $i:\mathbb{R}^{\mathbb{Z}_{N}}\to\mathbb{R}^{\mathbb{Z}}$.
The intuition behind this approach is that matrix multiplication $H(\mathbf{x}_{1})\cdots H(\mathbf{x}_{n})$ naturally encodes the partition function across a certain diagram with vertex weights given by the arrays $\mathbf{x}_{1},\ldots,\mathbf{x}_{n}$ as discussed in Lemma 2.6. Meanwhile, the outputs of map $E$ are more algebraically straightforward and closely related to the matrix inverse of $H$, which we prove in the next Lemma 2.2 as an infinite-dimensional analogue of [NY04, Equation 1.5].
Lemma 2.2.
For all $\mathbf{x}\in\mathbb{R}^{\mathbb{Z}}$, we have
$\Delta H(\mathbf{x})\Delta E(-\mathbf{x})=E(-\mathbf{x})\Delta H(\mathbf{x})\Delta=\text{Id}_{\mathbb{Z}},$ (2.2)
where $\Delta=\operatorname{diag}((-1)^{i-1})_{i\in\mathbb{Z}}$ and $\text{Id}_{\mathbb{Z}}=\operatorname{diag}(1)_{i\in\mathbb{Z}}$.
Proof.
We compute $H(\mathbf{x})\Delta E(-\mathbf{x})$ as follows:
$H(\mathbf{x})\Delta E(-\mathbf{x})$ $=H(\mathbf{x})\left(\sum_{j\in\mathbb{Z}}(-1)^{j-1}e^{-x_{j}}E_{j,j}+\sum_{j\in\mathbb{Z}}(-1)^{j-1}E_{j,j+1}\right)$
$=H(\mathbf{x})\sum_{j\in\mathbb{Z}}(-1)^{j-1}e^{-x_{j}}E_{j,j}+H(\mathbf{x})\sum_{j\in\mathbb{Z}}(-1)^{j-1}E_{j,j+1}$
$=\sum_{i\in\mathbb{Z}}(-1)^{i-1}E_{i,i}+\sum_{i\leq j}\bigl{(}(-1)^{j-1}e^{x_{i}+\cdots+x_{j-1}}+(-1)^{j-2}e^{x_{i}+\cdots+x_{j-1}}\bigr{)}E_{i,k}$
$=\Delta\ ,$
where the third equality is obtained by shifting the index of the second summand from $j$ to j1j-1 and last equality above follows by a telescoping of terms. A similar calculation shows that $E(-\mathbf{x})\Delta H(\mathbf{x})=\Delta$. This concludes the proof, since one can easily check that $\Delta^{2}=Id$. ∎
Remark 2.3.
Noticing that $\Delta H(\mathbf{x})\Delta$ is a left and right inverse for $E(-\mathbf{x})$, we see that it is an inverse in the semi-group of finite valued upper-triangular $\mathbb{Z}\times\mathbb{Z}$ matrices under matrix multiplication. Under this notation, we see that Lemma 2.2 translates to
$H(\mathbf{x})=\Delta E(-\mathbf{x})^{-1}\Delta.$ (2.3)
Lemma 2.4.
For $\mathbf{W}_{1},\mathbf{W}_{2}\in\mathbb{R}^{\mathbb{Z}_{N}}$, we have that
$E(-\mathbf{W}_{2})E(-\mathbf{W}_{1})=E(-D(\mathbf{W}_{2},\mathbf{W}_{1}))E(-T(\mathbf{W}_{2},\mathbf{W}_{1})).$ (2.4)
Proof.
By direct computation and applying Lemma A.5 in the third equality below, we see that
$\quad\,E(-D(\mathbf{W}_{2},\mathbf{W}_{1}))E(-T(\mathbf{W}_{2},\mathbf{W}_{1}))$
$=\left(\sum_{i\in\mathbb{Z}}e^{-D_{i}(\mathbf{W}_{2},\mathbf{W}_{1})}E_{i,i}+\sum_{i\in\mathbb{Z}}E_{i,i+1}\right)\left(\sum_{i\in\mathbb{Z}}e^{-T_{i}(\mathbf{W}_{2},\mathbf{W}_{1})}E_{i,i}+\sum_{i\in\mathbb{Z}}E_{i,i+1}\right)$
$=\sum_{i\in\mathbb{Z}}e^{-(D_{i}(\mathbf{W}_{2},\mathbf{W}_{1})+T_{i}(\mathbf{W}_{2},\mathbf{W}_{1}))}E_{i,i}+\sum_{i\in\mathbb{Z}}\left(e^{-D_{i}(\mathbf{W}_{2},\mathbf{W}_{1})}+e^{-T_{i+1}(\mathbf{W}_{2},\mathbf{W}_{1})}\right)E_{i,i+1}+\sum_{i\in\mathbb{Z}}E_{i,i+2}$
$=\sum_{i\in\mathbb{Z}}e^{-(W_{1,i}+W_{2,i})}E_{i,i}+\sum_{i\in\mathbb{Z}}\left(e^{-W_{1,i+1}}+e^{-W_{2,i}}\right)E_{i,i+1}+\sum_{i\in\mathbb{Z}}E_{i,i+2}$
$=\left(\sum_{i\in\mathbb{Z}}e^{-W_{2,i}}E_{i,i}+\sum_{i\in\mathbb{Z}}E_{i,i+1}\right)\left(\sum_{i\in\mathbb{Z}}e^{-W_{1,i}}E_{i,i}+\sum_{i\in\mathbb{Z}}E_{i,i+1}\right)$
$=E(-\mathbf{W}_{2})E(-\mathbf{W}_{1}).\qed$
Corollary 2.5.
For $\mathbf{W}_{1},\mathbf{W}_{2}\in\mathbb{R}^{\mathbb{Z}_{N}}$, we have
$H(\mathbf{W}_{1})H(\mathbf{W}_{2})=H(T(\mathbf{W}_{2},\mathbf{W}_{1}))H(D(\mathbf{W}_{2},\mathbf{W}_{1})).$ (2.5)
Proof.
First, we justify that this matrix product even makes sense, in that it is finite valued. Because these $H$ matrices (as well as the $E$ matrices) are upper triangular, each individual index in the product consists of a finite sum of products of elements in the respective $H$ matrices. While each row and column has infinitely many non-zero terms, there are only a finite number of non-zero terms in the dot product. Further, the product is upper triangular, so by repeating this argument, we can see that an arbitrary product of $H$ (or $E$) matrices is finite valued.
With this in mind, the proof follows from Lemma 2.2, since:
[…]
$=\Delta E(-\mathbf{W}_{1})^{-1}E(-\mathbf{W}_{2})^{-1}\Delta$
$=\Delta\bigl{(}E(-\mathbf{W}_{2})E(-\mathbf{W}_{1})\bigr{)}^{-1}\Delta$
[…]
$=(\Delta E(-T(\mathbf{W}_{2},\mathbf{W}_{1}))^{-1}\Delta)(\Delta E(-D(\mathbf{W}_{2},\mathbf{W}_{1}))^{-1}\Delta)$
$=H(T(\mathbf{W}_{2},\mathbf{W}_{1}))H(D(\mathbf{W}_{2},\mathbf{W}_{1}))\qed$
2.2. Relation to the periodic polymer
We now show the following relation between the $H$ matrices and the polymer partition function:
Lemma 2.6.
Let $\mathbf{W}=(\mathbf{W}_{i})_{i\in\mathbb{Z}}\in(\mathbb{Z}_{N})^{\mathbb{Z}}$ and $k,\ell\in\mathbb{Z}$ such that $\ell\geq k$, then:
$\Bigl{(}H(\mathbf{W}_{k})H(\mathbf{W}_{k+1})\ldots H(\mathbf{W}_{\ell})\Bigr{)}_{(i,j)}=Z^{\mathbf{W}}(\ell,j\mid k,i)\quad\forall i,j\in\mathbb{Z}.$
In other words, the partition function across any two points $(k,i)$ and $(\ell,j)$ on the directed lattice $\mathbb{Z}^{2}$ with periodic weights $\mathbf{W}$ is given by the $(i,j)$th entry in the matrix product of the $H$ matrices.
Proof.
We prove the result by induction on $\ell\geq k$. For the base case, $\ell=k$, by definition we have that
$H(\mathbf{W}_{k})_{(i,j)}=\begin{cases}\exp\Biggl{(}{\sum\limits_{r=i}^{j}}W_{(k,r)}\Biggr{)},\quad&j\geq i\\ 0,\ &\text{otherwise}\end{cases}$
Similarly, on a single column, we either have a unique up/right path between $(k,i)$ and $(k,j)$ picking up all the weights between $i$ and $j$ or have no path between them, depending on whether $j\geq i$. Thus,
$Z^{W}(k,j\mid k,i)=\begin{cases}\exp\Biggl{(}{\sum\limits_{r=i}^{j}}W_{(k,r)}\Biggr{)},\quad&j\geq i\\ 0,\ &\text{otherwise}\end{cases}=H(\mathbf{W}_{k})_{(i,j)}$
This proves our base case. Now, let us assume that it holds for some $\ell$, and prove it holds for $\ell+1$. We split into two cases: $j\geq i$ and $j<i$.
For $j<i$, notice that $H$ is upper triangular, so their product is as well. Thus, the $(i,j)$-th index of our matrix product is 0 for $j<i$. This matches up with our partition function, as there are no up/right paths whose initial point is lower than its end point.
For the case $j\geq i$, we notice that every path from $(k,i)$ to $(\ell+1,j)$ consists of a path from $(k,i)$ to $(\ell,n)$ concatenated with a path from $(\ell+1,n)$ to $(\ell+1,j)$, for some $i\leq n\leq j$. Thus,
$Z^{\mathbf{W}}(\ell+1,j\mid k,i)$ $=\sum\limits_{r=i}^{j}\Biggl{(}Z^{\mathbf{W}}(\ell,r\mid k,i)\exp\bigl{(}\sum_{s=r}^{j}W_{(\ell+1,s)}\bigr{)}\Biggr{)}$
$=\sum\limits_{r=i}^{j}\Biggl{(}Z^{\mathbf{W}}(\ell,r\mid k,i)\Bigl{(}H(\mathbf{W}_{\ell+1})\Bigr{)}_{(r,j)}\Biggr{)}$
By our inductive hypothesis, we know that this equals
$\sum\limits_{r=i}^{j}\Biggl{(}\Bigl{(}H(\mathbf{W}_{k})H(\mathbf{W}_{k+1})\ldots H(\mathbf{W}_{\ell})\Bigr{)}_{(i,r)}\Bigl{(}H(\mathbf{W}_{\ell+1})\Bigr{)}_{(r,j)}\Biggr{)}$
Remembering now that $\bigl{(}H(\mathbf{W}_{\ell+1})\bigr{)}_{(r,j)}=0$ for $r>j$, and $\bigl{(}H(\mathbf{W}_{k})H(\mathbf{W}_{k+1})\ldots H(\mathbf{W}_{\ell})\bigr{)}_{(i,r)}=0$ for $r<i$, we see that
$Z^{\mathbf{W}}(\ell+1,j\mid k,i)$ $=\sum\limits_{r=-\infty}^{\infty}\Biggl{(}\Bigl{(}H(\mathbf{W}_{k})H(\mathbf{W}_{k+1})\ldots H(\mathbf{W}_{\ell})\Bigr{)}_{(i,r)}\Bigl{(}H(\mathbf{W}_{\ell+1})\Bigr{)}_{(r,j)}\Biggr{)}$
$=\Bigl{(}H(\mathbf{W}_{k})H(\mathbf{W}_{k+1})\ldots H(\mathbf{W}_{\ell})H(\mathbf{W}_{\ell+1})\Bigr{)}_{(i,j)}.\qed$
We now show the invariance of polymer partition function under the periodic Pitman transform. Retaining the periodic weight vectors $\mathbf{W}=(\mathbf{W}_{r}\in\mathbb{R}^{\mathbb{Z}_{N}}:r\in\mathbb{Z})$ as before, for $k\in\mathbb{Z}$, recall the operator $\mathcal{P}_{k}:(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}\to(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$ where
$\mathcal{P}_{k}\mathbf{W}=(\ldots,\mathbf{W}_{k-1},T(\mathbf{W}_{k+1},\mathbf{W}_{k}),D(\mathbf{W}_{k+1},\mathbf{W}_{k}),\mathbf{W}_{k+2,\ldots})$
Proposition 2.7.
Fix $a,c\in\mathbb{Z}$ such that $a\leq c$, and let $k_{1},\dots,k_{m}\in\mathbb{Z}$ such that for all $i\in\llbracket 1,m\rrbracket$, $a\neq k_{i}+1$ and $c\neq k_{i}$. Then, for all $b,d\in\mathbb{Z}$,
$Z^{\mathcal{P}_{k_{m}}\cdots\mathcal{P}_{k_{1}}\mathbf{W}}(c,d\mid a,b)=Z^{\mathbf{W}}(c,d\mid a,b)$
Proof.
We will prove this by induction on $m$. By Lemma 2.6,
$Z^{\mathbf{W}}(c,d\mid a,b)=\bigl{(}H(\mathbf{W}_{a})H(\mathbf{W}_{a+1})\cdots H(\mathbf{W}_{c})\bigr{)}_{(b,d)},$
which we claim is preserved by $\mathcal{P}_{k_{m}}\ldots\mathcal{P}_{k_{1}}$ for any $b,d$. In our base case $n=1$, we split into three sub-cases: $k_{1}<a-1$, $k_{1}>c$, and $a\leq k_{1}<c$. In the first two sub-cases, $\mathbf{W}_{a}\ldots\mathbf{W}_{c}$ are unaffected by $\mathcal{P}_{k_{1}}$, so our invariance holds trivially.
If $a\leq k_{1}<c$, then Corollary 2.5 gives that
$Z^{\mathbf{W}}(c,d\mid a,b)$ $=\bigl{(}H(\mathbf{W}_{a})H(\mathbf{W}_{a+1})\ldots H(\mathbf{W}_{k_{1}})H(\mathbf{W}_{k_{1}+1})\ldots H(\mathbf{W}_{c})\bigr{)}_{(b,d)}$
[…]
$=Z^{\mathcal{P}_{k_{1}}\mathbf{W}}(c,d\mid a,b)$
Now, we assume by way of induction that the invariance holds for m1m-1 transforms. Again, if $k_{m}<a-1$ or $k_{m}>c$ our equivalence holds trivially. Otherwise, let $\mathbf{W}^{\prime}=\mathcal{P}_{k_{m-1}}\ldots\mathcal{P}_{k_{1}}\mathbf{W}$. Then
$Z^{\mathbf{W}}(c,d\mid a,b)$ $=Z^{\mathbf{W}^{\prime}}(c,d\mid a,b)$
$=Z^{\mathcal{P}_{k_{m}}\mathbf{W}^{\prime}}(c,d\mid a,b)$
$=Z^{\mathcal{P}_{k_{m}}\mathcal{P}_{k_{m-1}}\ldots\mathcal{P}_{k_{1}}\mathbf{W}}(c,d\mid a,b)$
This concludes our proof, noting that the second equality follows from the case $m=1$. ∎
With Proposition 2.7 proved, we now prove Theorem 1.4 by extending to the multi-path result.
Proof of Theorem 1.4..
We use the Lindström-Gessel-Viennot (LGV) lemma (cf. [AZ18, Chapter 32], originally from [Lin73, GV85]) to prove the theorem. Because LGV applies only to edge-weighted graphs, we need to transfer our vertex weights to edge weights in the following way: for each vertex $(r,s)$, split it into $(r,s)_{\text{in}}$ and $(r,s)_{\text{out}}$, with edge $\bigl{(}(r,s)_{\text{in}},(r,s)_{\text{out}}\bigr{)}$ given the weight $\omega_{(r,s)}:=e^{W_{(r,s)}}$, and edges $\bigl{(}(r,s)_{\text{out}},(r,s+1)_{\text{in}}\bigr{)}$ and $\bigl{(}(r,s)_{\text{out}},(r+1,s)_{\text{in}}\bigr{)}$ given the weight 1. See Figure 6. Any path which started at $(r,s)$ starts at $(r,s)_{\text{in}}$, while any path ending at $(r,s)$ ends at $(r,s)_{\text{out}}$. Under this graph transformation, our partition function on the pre-transformed graph equals the sum over weight products along possible paths on the transformed diagram, so we can apply LGV.
[…]
Figure 6. Local transformation of a vertex-weighted node (left) with weight $W_{(r,s)}$ into an edge-weighted directed graph (right), where the internal edge carries weight $\omega_{(r,s)}=\exp(W_{(r,s)})$ and external edges carry weight 1. Paths in the original graph correspond to paths from “in” nodes to “out” nodes in the transformed graph, preserving path weights.
Write the path matrix $M=\bigl{(}Z^{\mathbf{W}}(c_{j},d_{j}\mid a_{i},b_{i})\bigr{)}_{(i,j)\in\llbracket 1,m\rrbracket}$. By LGV, we have
$\det M=\sum\limits_{P=(P_{1}\ldots P_{m})\in\Pi_{U,V_{\sigma}}\ \forall\sigma\in S_{m}}\operatorname{sgn}(\sigma)\prod\limits_{(i,j)\in P}e^{W_{(i,j)}}$
By definition of $\Psi$, for $\sigma\neq\text{Id}$, $\Pi_{U,V_{\sigma}}=\varnothing$. Because $\operatorname{sgn}(\text{Id})=1$, we have that
$\det M=\sum\limits_{P=(P_{1}\ldots P_{m})\in\Pi_{U,V}}\prod\limits_{(i,j)\in P}e^{W_{(i,j)}}=Z^{\mathbf{W}}(V\mid U)$
By Proposition 2.7, each $M_{i,j}=Z^{\mathbf{W}}(c_{j},d_{j}\mid a_{i},b_{i})$ remains invariant under $\mathcal{P}_{k_{m}}\cdots\mathcal{P}_{k_{1}}$, so $M$ is unchanged, and so $\det M$ is invariant. Thus, $Z^{\mathbf{W}}(V\mid U)=\det M$ is invariant. ∎
3. Braid relations: proof of Theorem 1.6
In this section, we prove the braid relation stated in Theorem 1.6. Notice that this relation translates into the following three equalities:
$T\Bigl{(}T\bigl{(}\mathbf{W}_{k+2},D(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)},T(\mathbf{W}_{k+1},\mathbf{W}_{k})\biggr{)}=T\Bigl{(}T(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),\mathbf{W}_{k}\Bigr{)},$ (3.1)
$D\Bigl{(}T\bigl{(}\mathbf{W}_{k+2},D(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)},T(\mathbf{W}_{k+1},\mathbf{W}_{k})\Bigr{)}=T\Bigl{(}D(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),D\bigl{(}T(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),\mathbf{W}_{k}\bigr{)}\Bigr{)},$ (3.2)
$D\Bigl{(}\mathbf{W}_{k+2},D(\mathbf{W}_{k+1},\mathbf{W}_{k})\Bigr{)}=D\Bigl{(}D(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),D\bigl{(}T(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),\mathbf{W}_{k}\bigr{)}\Bigr{)}.$ (3.3)
Equation (3.3) follows from [CGS24, Proposition 5.13], as explained in Lemma A.8 of the appendix. We now use the partition function invariance, Theorem 1.4, to demonstrate that (3.1) implies both (3.2) and (3.3). The main tool will be the following lemma:
Lemma 3.1.
For $\mathbf{Y}=(\mathbf{Y}_{i})_{i\in\mathbb{Z}},\mathbf{Y}^{\prime}=(\mathbf{Y}^{\prime}_{i})_{i\in\mathbb{Z}}\in(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$ satisfying $\mathfrak{s}(\mathbf{Y}_{1})=\mathfrak{s}(\mathbf{Y}^{\prime}_{1})$, $\mathfrak{s}(\mathbf{Y}_{2})=\mathfrak{s}(\mathbf{Y}_{2}^{\prime})$, we have that $H(\mathbf{Y}_{1})H(\mathbf{Y}_{2})=H(\mathbf{Y}^{\prime}_{1})H(\mathbf{Y}_{2}^{\prime})$ if and only if $\mathbf{Y}_{1}=\mathbf{Y}^{\prime}_{1},\mathbf{Y}_{2}=\mathbf{Y}_{2}^{\prime}$.
Before proving Lemma 3.1, we first use Lemmas A.8 and 3.1 to complete the proof of Theorem 1.6:
Proof of Theorem 1.6..
Using the partition function invariance under the operator $\mathcal{P}_{i+1}\mathcal{P}_{i}\mathcal{P}_{i+1}$ yields:
$H\bigl{(}\mathbf{W}_{k}\bigr{)}H\bigl{(}\mathbf{W}_{k+1}\bigr{)}H\bigl{(}\mathbf{W}_{k+2}\bigr{)}=$
[…]
$\qquad H\Bigl{(}D\bigl{(}D(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),D\bigl{(}T(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),\mathbf{W}_{k}\bigr{)}\bigr{)}\Bigr{)}$ (3.4)
On the other hand, we can also apply the partition function invariance under the operator $\mathcal{P}_{i}\mathcal{P}_{i+1}\mathcal{P}_{i}$:
$H\bigl{(}\mathbf{W}_{k}\bigr{)}H\bigl{(}\mathbf{W}_{k+1}\bigr{)}H\bigl{(}\mathbf{W}_{k+2}\bigr{)}=$
[…]
$\qquad H\Bigl{(}D\bigl{(}T\bigl{(}\mathbf{W}_{k+2},D(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)},T(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)}\Bigr{)}H\Bigl{(}D(\mathbf{W}_{k+2},D(\mathbf{W}_{k+1},\mathbf{W}_{k}))\Bigr{)}$ (3.5)
From Lemma A.8, we know that the last terms in (3.4) and (3.5) are equal. Because $H$ is invertible (see Lemma 2.2), we can equate both equations and cancel out the last term, obtaining that:
$H\Bigl{(}T\bigl{(}T(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),\mathbf{W}_{k}\bigr{)}\Bigr{)}H\Bigl{(}T\bigl{(}D(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),D\bigl{(}T(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),\mathbf{W}_{k}\bigr{)}\Bigr{)}$
$=H\Bigl{(}T\bigl{(}T\bigl{(}\mathbf{W}_{k+2},D(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)},T(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)}\Bigr{)}H\Bigl{(}D\bigl{(}T\bigl{(}\mathbf{W}_{k+2},D(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)},T(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)}\Bigr{)}$
By repeated application of Lemma A.4, we have that
$\mathfrak{s}(T(T(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),\mathbf{W}_{k}))=\mathfrak{s}(\mathbf{W}_{k+2})$
$=\mathfrak{s}(T(T(\mathbf{W}_{k+2},D(\mathbf{W}_{k+1},\mathbf{W}_{k})),T(\mathbf{W}_{k+1},\mathbf{W}_{k}))),\quad\text{and}$
$\mathfrak{s}(T(D(T(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),\mathbf{W}_{k})),D(\mathbf{W}_{k+2},\mathbf{W}_{k+1})))$
$=\mathfrak{s}(\mathbf{W}_{k+1})=\mathfrak{s}(T\bigl{(}D(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),D\bigl{(}T(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),\mathbf{W}_{k}\bigr{)})).$
Applying Lemma 3.1, we conclude (3.2) and (3.3), as we wanted to prove. ∎
Proof of Lemma 3.1..
Immediately, the backwards direction is clear. Then, it remains to show that, under the assumption that $H(\mathbf{Y}_{1})H(\mathbf{Y}_{2})=H(\mathbf{Y}^{\prime}_{1})H(\mathbf{Y}_{2}^{\prime})\implies\mathbf{Y}_{1}=\mathbf{Y}^{\prime}_{1},\mathbf{Y}_{2}=\mathbf{Y}_{2}^{\prime}$. Let us prove by backwards induction on $k=N,\ldots,1$ that $Y_{1,k}=Y^{\prime}_{1,k}$ and $Y_{2,k}=Y^{\prime}_{2,k}$.
As a base case, let’s show this for $k=N$. From Lemma 2.6, note that
$\left(H(\mathbf{Y}_{1})H(\mathbf{Y}_{2})\right)_{1,N}$ $=Z^{\mathbf{Y}}(2,N\mid 1,1)$
$=e^{Y_{2,N}}\left(Z^{\mathbf{Y}}(1,N\mid 1,1)+Z^{\mathbf{Y}}(2,N-1\mid 1,1)\right)$
$=e^{Y_{2,N}}\Biggl{(}\prod_{i=1}^{N}e^{Y_{1,i}}+\left(H(\mathbf{Y}_{1})H(\mathbf{Y}_{2})\right)_{1,N-1}\Biggr{)}$ (3.6)
Similarly, we conclude that:
$\left(H(\mathbf{Y}^{\prime}_{1})H(\mathbf{Y}^{\prime}_{2})\right)_{1,N}$ $=Z^{\mathbf{Y}^{\prime}}(2,N\mid 1,1)$
$=e^{Y^{\prime}_{2,N}}\left(Z^{\mathbf{Y}^{\prime}}(1,N\mid 1,1)+Z^{\mathbf{Y}^{\prime}}(2,N-1\mid 1,1)\right)$
$=e^{Y^{\prime}_{2,N}}\Biggl{(}\prod_{i=1}^{N}e^{Y^{\prime}_{1,i}}+\left(H(\mathbf{Y}^{\prime}_{1})H(\mathbf{Y}^{\prime}_{2})\right)_{1,N-1}\Biggr{)}.$ (3.7)
By assumption, we have $H(\mathbf{Y}_{1})H(\mathbf{Y}_{2})=H(\mathbf{Y}^{\prime}_{1})H(\mathbf{Y}_{2}^{\prime})$ and $\prod_{i=1}^{N}e^{Y_{1,i}}=e^{\mathfrak{s}(Y_{1})}=e^{\mathfrak{s}(Y^{\prime}_{1})}=\prod_{i=1}^{N}e^{Y^{\prime}_{1,i}}$. Equating (3.6) and (3.7) and canceling the nonzero expressions in parenthesis yields $Y_{2,N}=Y^{\prime}_{2,N}$. Next, we see that $Y^{\prime}_{1,N}=Y_{1,N}$, since:
$(H(\mathbf{Y}_{1})H(\mathbf{Y}_{2}))_{(N,N)}=Z^{\mathbf{Y}}(2,N\mid 1,N)=e^{Y_{1,N}+Y_{2,N}},$ (3.8)
and
$(H(\mathbf{Y}^{\prime}_{1})H(\mathbf{Y}_{2}^{\prime}))_{(N,N)}=Z^{\mathbf{Y}^{\prime}}(2,N\mid 1,N)=e^{Y^{\prime}_{1,N}+Y^{\prime}_{2,N}}$ (3.9)
To finish the proof, simply notice that the proof of the above only required $\prod_{i=1}^{N}e^{Y^{\prime}_{i}}=\prod_{i=1}^{N}e^{Y_{i}}$. Alternatively, if we have $\prod_{i=1}^{k}e^{Y_{1,i}}=\prod_{i=1}^{k}e^{Y_{2,i}}$ for some $k\in\llbracket 1,N\rrbracket$, then the same proof shows that $Y_{1,k}=Y_{1,k}^{\prime}$ and $Y_{2,k}=Y_{2,k}^{\prime}$. This follows by backwards induction, since
$\prod_{i=1}^{k}e^{Y_{1,i}}=\mathfrak{s}(\mathbf{Y}_{1})\Bigl{(}\prod_{i=k+1}^{N}e^{Y_{1,i}}\Bigr{)}^{-1}.\qed$
4. Zero-temperature analogues: proofs of Theorems 1.8 and 1.9
Because polymer models have a strong connection to last-passage percolation (LPP), our algebraic results will naturally generalize to zero-temperature equivalents, through limiting arguments. In these limits, the $(+,\times)$ algebra is swapped for the $(\max,+)$ algebra, and our maps reflect this change. As introduced in [CGS24], the ‘correct’ map to look at here is $\widetilde{D}$, the uniform limit as $\beta$ goes to infinity of $\frac{1}{\beta}D(\beta\mathbf{X}_{1},\beta\mathbf{X}_{2})$, as well as $\widetilde{T}$, which is reached through an identical limit. In this section, we use our algebraic results about the behavior of the polymer model under the periodic Pitman transform (as well as the aforementioned limiting arguments) to prove analogous facts about LPP under the zero-temperature periodic Pitman transform. The following intermediate result is needed.
Lemma 4.1.
The limit as $\beta\to\infty$ of $\frac{1}{\beta}\log\bigl{(}Z^{\beta,\mathbf{W}}(V\mid U)\bigr{)}$ is $G^{\mathbf{W}}(V\mid U)$, the (multi-path)-LPP time between $U$ and $V$. Further, this convergence is uniform as a function of $\mathbf{W}$.
Proof.
By Lemma A.1, this converges uniformly, since there are finitely many multi-paths between any two sets of points. ∎
Note that in the special case where $U,V$ are points, this shows that the single-path polymer partition function converges to single-path LPP. Also, because the convergence to LPP is uniform on the weights, we know that the zero-temperature limit is well-behaved up to weights which depend on $\beta$. If the weights $\mathbf{W}_{\beta}$ converge to $\mathbf{\widetilde{\mathcal{W}}}$, then $\frac{1}{\beta}\log(Z^{\beta,\mathbf{W}_{\beta}}(V\mid U))$ will converge to $G^{\mathbf{\widetilde{\mathcal{W}}}}(V\mid U)$. With this in mind, we can now prove Theorem 1.8.
Proof of Theorem 1.8..
As in Theorem 1.4, we take $k_{1},k_{2},\ldots,k_{m}\in\mathbb{Z},U=\bigl{(}a_{i},b_{i})_{i\in\llbracket 1,m\rrbracket}\bigr{)}$, and $V=\bigl{(}(c_{i},d_{i})_{i\in\llbracket 1,m\rrbracket}\bigr{)}$ satisfying $(U,V)\in\Psi$ and, for all $i,j\in\llbracket 1,n\rrbracket$, $a_{i}\neq k_{j}+1$ and $c_{i}\neq k_{j}$. To prove Theorem 1.8, we must show that $G^{\mathbf{W}}(V\mid U)=G^{\widetilde{\mathcal{P}}_{k_{n}}\ldots\widetilde{\mathcal{P}}_{k_{1}}\mathbf{W}}(V\mid U)$, where the operators $\mathcal{P}_{k}$ are defined in 1.9, with $T,D$ replaced by $\widetilde{T},\widetilde{D}$. Now, for $\beta>0$, let $\mathcal{P}_{k}^{\beta}:(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}\to(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$ be operators defined by their action:
$\quad\,\mathcal{P}_{k}^{\beta}(\ldots,\mathbf{W}_{k-1},\mathbf{W}_{k},\mathbf{W}_{k+1},\mathbf{W}_{k+2},\ldots)$ (4.1)
$=(\ldots,\mathbf{W}_{k-1},\frac{1}{\beta}T(\beta\mathbf{W}_{k+1},\beta\mathbf{W}_{k}),\frac{1}{\beta}D(\beta\mathbf{W}_{k+1},\beta\mathbf{W}_{k}),\mathbf{W}_{k+2},\ldots).$
By Theorem 1.4, we know that $Z^{\mathbf{W}}(V\mid U)=Z^{\mathcal{P}_{k_{m}}\ldots\mathcal{P}_{k_{1}}\mathbf{W}}(V\mid U)$. Then, by definition of the polymer model with inverse temperature $\beta$ (1.8), for all $\beta>0$,
$Z^{\beta,\mathbf{W}}(V\mid U)=Z^{\beta\mathbf{W}}(V\mid U)=Z^{\mathcal{P}_{k_{1}}(\beta\mathbf{W})}(V\mid U)=Z^{\beta(\frac{1}{\beta}\mathcal{P}_{k_{1}}(\beta\mathbf{W}))}(V\mid U)=Z^{\beta,\mathcal{P}_{k_{1}}^{\beta}(\mathbf{W})}(V\mid U).$ (4.2)
By induction, we have $Z^{\beta,\mathbf{W}}(V\mid U)=Z^{\beta,\mathcal{P}_{k_{n}}^{\beta}\ldots\mathcal{P}_{k_{1}}^{\beta}\mathbf{W}}(V\mid U)$.
By Lemma 4, we know that $\frac{1}{\beta}\log\bigl{(}Z^{\beta,\mathbf{W}}(V\mid U)\bigr{)}$ converges to $G^{\mathbf{W}}(V\mid U)$ uniformly on the weights, so $G^{\mathbf{W}}(V\mid U)$ is preserved under the limiting operator $\lim\limits_{\beta\to\infty}\mathcal{P}_{k_{n}}^{\beta}\ldots\mathcal{P}_{k_{1}}^{\beta}$. We claim that each $\mathcal{P}_{k}^{\beta}$ converges to $\widetilde{\mathcal{P}}_{k}$ uniformly as a function of the weights, which implies by induction that
$G^{\mathbf{W}}(V\mid U)=G^{\lim\limits_{\beta\to\infty}\mathcal{P}_{k_{m}}^{\beta}\ldots\mathcal{P}_{k_{1}}^{\beta}\mathbf{W}}(V\mid U)=G^{\widetilde{\mathcal{P}}_{k_{m}}\ldots\widetilde{\mathcal{P}}_{k_{1}}\mathbf{W}}(V\mid U),$ (4.3)
proving the theorem. The necessary uniform convergence is shown component-wise:
$\quad\,\frac{1}{\beta}D_{i}(\beta\mathbf{X}_{1},\beta\mathbf{X}_{2})-\widetilde{D}_{i}(\mathbf{X}_{1},\mathbf{X}_{2})$
$=X_{2,i+1}+\frac{1}{\beta}\log\Biggl{(}\frac{\sum_{j\in\mathbb{Z}_{N}}e^{\beta Y_{(i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{\beta Y_{(i-1,j]}}}\Biggr{)}-(X_{2,i+1}+\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}-\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]})$
$=\Bigl{(}\frac{1}{\beta}\log\bigl{(}\sum_{j\in\mathbb{Z}_{N}}e^{\beta Y_{(i,j]}}\bigr{)}-\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}\Bigr{)}-\Bigl{(}\frac{1}{\beta}\log\bigl{(}\sum_{j\in\mathbb{Z}_{N}}e^{\beta Y_{(i-1,j]}}\bigr{)}-\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]}\Bigr{)}$
By Lemma A.1, we know that $\frac{1}{\beta}\log(N)\leq\frac{1}{\beta}D_{i}(\beta\mathbf{X}_{1},\beta\mathbf{X}_{2})-\widetilde{D}_{i}(\mathbf{X}_{1},\mathbf{X}_{2})\leq\frac{1}{\beta}\log(N)$, proving uniform convergence. An identical argument holds for $T_{i}-\widetilde{T}_{i}$. ∎
We next show the braid relations pass to zero-temperature.
Proof of Theorem 1.9..
Defining $\mathcal{P}_{k}^{\beta}$ as in (4.1), we see that $\mathcal{P}_{k}^{\beta}\mathcal{P}_{k+1}^{\beta}\mathcal{P}_{k}^{\beta}=\mathcal{P}_{k+1}^{\beta}\mathcal{P}_{k}^{\beta}\mathcal{P}_{k+1}^{\beta}$, as an immediate consequence of Theorem 1.6. As noted in the proof of Theorem 1.8, $\mathcal{P}_{k}^{\beta}$ converges to $\widetilde{\mathcal{P}}_{k}$ uniformly. This immediately applies the braid relation for $\widetilde{\mathcal{P}}_{k}$, since we have
$\widetilde{\mathcal{P}}_{k}\widetilde{\mathcal{P}}_{k+1}\widetilde{\mathcal{P}}_{k}=\lim\limits_{\beta\to\infty}\mathcal{P}_{k}^{\beta}\mathcal{P}_{k+1}^{\beta}\mathcal{P}_{k}^{\beta}=\lim\limits_{\beta\to\infty}\mathcal{P}_{k+1}^{\beta}\mathcal{P}_{k}^{\beta}\mathcal{P}_{k+1}^{\beta}=\widetilde{\mathcal{P}}_{k+1}\widetilde{\mathcal{P}}_{k}\widetilde{\mathcal{P}}_{k+1}.\qed$
5. Burke properties and permutation invariance
In this section, we generalize the Burke property in [CGS24, Proposition 5.12] to an inhomogeneous environment and provide geometric and exponential analogues of other Burke property for the zero-temperature discrete periodic Pitman transform.
Proposition 5.1 (Log-inverse-gamma Burke property).
Let $\lambda>0$, and let $\mathbf{a},\mathbf{b}\in\mathbb{R}^{\mathbb{Z}_{N}}$ be sequences such that $a_{i}+b_{j}>0$ for $i,j\in\mathbb{Z}_{N}$. Let $\mathbf{X}_{1},\mathbf{X}_{2}$ be independent sequences of independent random variables such that $X_{i,j}\sim(a_{i}+b_{j},\lambda)$ (see Definition 1.18). Then,
$(\mathbf{X}_{1},\mathbf{X}_{2})\overset{d}{=}(T(\mathbf{X}_{1},\mathbf{X}_{2}),D(\mathbf{X}_{1},\mathbf{X}_{2})).$
Proof.
The joint density $\mathbb{P}(\mathbf{X}_{1}\in d\mathbf{x}_{1},\mathbf{X}_{2}\in d\mathbf{x}_{2})$ is proportional to
$\exp\Biggl{(}-\sum_{i\in\mathbb{Z}_{N}}\Bigl{(}(a_{1}+b_{i})x_{1,i}+(a_{2}+b_{i})x_{2,i}\Bigr{)}-\lambda\sum_{i\in\mathbb{Z}_{N}}(e^{-x_{1,i}}+e^{-x_{2,i}})\Biggr{)}$
$=\exp\Biggl{(}-a_{1}\sum_{i\in\mathbb{Z}_{N}}x_{1,i}-a_{2}\sum_{i\in\mathbb{Z}_{N}}x_{2,i}-\sum_{i\in\mathbb{Z}_{N}}b_{i}(x_{1,i}+x_{2,i})-\lambda\sum_{i\in\mathbb{Z}_{N}}(e^{-x_{1,i}}+e^{-x_{2,i+1}})\Biggr{)}$
$=\exp\Biggl{(}-a_{1}\sum_{i\in\mathbb{Z}_{N}}T_{i}(\mathbf{x}_{1},\mathbf{x}_{2})-a_{2}\sum_{i\in\mathbb{Z}_{N}}D_{i}(\mathbf{x}_{1},\mathbf{x}_{2})-\sum_{i\in\mathbb{Z}_{N}}b_{i}(T_{i}(\mathbf{x}_{1},\mathbf{x}_{2})+D_{i}(\mathbf{x}_{1},\mathbf{x}_{2}))$
$\qquad\qquad\qquad\qquad-\lambda\sum_{i\in\mathbb{Z}_{N}}(e^{-T_{i+1}(\mathbf{x}_{1},\mathbf{x}_{2})}+e^{-D_{i}(\mathbf{x}_{1},\mathbf{x}_{2})})\Biggr{)}$
$=\exp\Biggl{(}-\sum_{i\in\mathbb{Z}_{N}}\Bigl{(}(a_{1}+b_{i})T_{i}(\mathbf{x}_{1},\mathbf{x}_{2})+(a_{2}+b_{i})D_{i}(\mathbf{x}_{1},\mathbf{x}_{2})\Bigr{)}-\lambda\sum_{i\in\mathbb{Z}_{N}}(e^{-T_{i}(\mathbf{x}_{1},\mathbf{x}_{2})}+e^{-D_{i}(\mathbf{x}_{1},\mathbf{x}_{2})})\Biggr{)},$
where we used Lemmas A.4-A.5 in the penultimate equality. Thus, the density is preserved under our transformation, so all we must show is that the determinant of the Jacobian matrix of the transformation has absolute value 1. This follows from Lemma A.7. ∎
Proposition 5.2 (Geometric Burke property).
Take $\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{Z}_{\geq 0}^{\mathbb{Z}_{N}}$ to be independent sequences of independent random variables such that $X_{1,i}\sim\text{Geom}(p_{1}q_{i})$ and $X_{2,i}\sim\text{Geom}(p_{2}q_{i})$ for all $i\in\mathbb{Z}_{N}$. Then,
$(\mathbf{X}_{1},\mathbf{X}_{2})\overset{d}{=}(\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2}),\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2}))$
Proof.
Letting $\mathbf{k}_{1},\mathbf{k}_{2}\in\mathbb{Z}_{\geq 0}^{N}$ where $k_{m,i}$ is the $i$th component of $\mathbf{k}_{m}$, we have that
$\mathbb{P}(\mathbf{X}_{1}=\mathbf{k}_{1},\mathbf{X}_{2}=\mathbf{k}_{2})$ $=\mathbb{P}\left(X_{1,1}=k_{1,1}\right)\cdots\mathbb{P}\left(X_{1,N}=k_{1,N}\right)\mathbb{P}\left(X_{2,1}=k_{2,1}\right)\cdots\mathbb{P}\left(X_{2,N}=k_{2,N}\right)$
$=\left(\prod\limits_{i=1}^{N}(1-p_{1}q_{i})(1-p_{2}q_{i})\right)p_{1}^{\mathfrak{s}(\mathbf{k}_{1})}p_{2}^{\mathfrak{s}(\mathbf{k}_{2})}\prod_{i=1}^{N}q_{i}^{k_{1,i}+k_{2,i}}$
$=\left(\prod\limits_{i=1}^{N}(1-p_{1}q_{i})(1-p_{2}q_{i})\right)p_{1}^{\mathfrak{s}(\widetilde{D}(\mathbf{k}_{2},\mathbf{k}_{1}))}p_{2}^{\mathfrak{s}(\widetilde{T}(\mathbf{k}_{2},\mathbf{k}_{1}))}\prod_{i=1}^{N}q_{i}^{\widetilde{D}_{i}(\mathbf{k}_{2},\mathbf{k}_{1})+\widetilde{T}_{i}(\mathbf{k}_{2},\mathbf{k}_{1})}$
$=\prod\limits_{i=1}^{N}(1-p_{1}q_{i})(p_{1}q_{i})^{\widetilde{D}_{i}(\mathbf{k}_{2},\mathbf{k}_{1})}\prod\limits_{i=1}^{N}(1-p_{2}q_{i})(p_{2}q_{i})^{\widetilde{T}_{i}(\mathbf{k}_{2},\mathbf{k}_{1})}$
$=\mathbb{P}(\mathbf{X}_{1}=\widetilde{D}(\mathbf{k}_{2},\mathbf{k}_{1}),\mathbf{X}_{2}=\widetilde{T}(\mathbf{k}_{2},\mathbf{k}_{1}))$
$=\mathbb{P}(\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})=\mathbf{k}_{2},\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2})=\mathbf{k}_{1})$
where the first equality follows from the independence of $\mathbf{X}_{1},\mathbf{X}_{2}$ and $X_{m,i},X_{m,j}$ for $m\in\{1,2\}$ and $i,j\in\mathbb{Z}_{N}$ where $i\neq j$; the second equality follows from our definition of the geometric distribution; the third equality follows from Lemmas A.4(iii)-(iv) and A.5(i); and the final equality follows from the bijection in Lemma A.3. ∎
Proposition 5.3 (Exponential Burke property).
Take $\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{Z}_{\geq 0}^{\mathbb{Z}_{N}}$ to be independent sequences such that $X_{1,i}\sim\text{Exp}(a_{1}+b_{i})$ and $X_{2,i}\sim\text{Exp}(a_{2}+b_{i})$ for all $i\in\mathbb{Z}_{N}$ where $a_{m}+b_{i}>0$ for $m\in\{1,2\}$ and $i\in\mathbb{Z}_{N}$. Then,
$(\mathbf{X}_{1},\mathbf{X}_{2})\overset{d}{=}(\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2}),\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2}))$
Proof.
For $m\in\{1,2\}$ and $i\in\mathbb{Z}_{N}$, define
$X_{m,i}^{\prime}\sim\text{Geom}\Biggl{(}\left(1-\frac{a_{m}}{n}\right)\left(1-\frac{b_{i}}{n}\right)\Biggr{)}$
Then, since $\left(1-\frac{a_{m}}{n}\right)\left(1-\frac{b_{i}}{n}\right)=1-\frac{a_{m}+b_{i}}{n}+O(n^{-2})$, we have that
$\lim\limits_{n\to\infty}\frac{X_{m,i}^{\prime}}{n}\overset{d}{=}X_{m,i}\sim\text{Exp}(a_{m}+b_{i}),\quad\text{for }m\in\{1,2\},\text{ and }i\in\mathbb{Z}_{N}$
Observe from the definitions of $\widetilde{T}$ and $\widetilde{D}$ (1.11) that
$\frac{1}{n}\widetilde{T}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime})=\widetilde{T}\Bigl{(}\frac{\mathbf{X}_{1}^{\prime}}{n},\frac{\mathbf{X}_{2}^{\prime}}{n}\Bigr{)},\quad\text{and}\quad\frac{1}{n}\widetilde{D}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime})=\widetilde{D}\Bigl{(}\frac{\mathbf{X}_{1}^{\prime}}{n},\frac{\mathbf{X}_{2}^{\prime}}{n}\Bigr{)}$
By Proposition 5.2, we know that $(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime})\overset{d}{=}(\widetilde{T}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime}),\widetilde{D}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime}))$. Since $\max$, addition, and subtraction are continuous, $\widetilde{D}$ and $\widetilde{T}$ are continuous maps. We then have that
$\Bigl{(}\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2}),\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})\Bigr{)}\overset{d}{=}\lim\limits_{n\to\infty}\frac{1}{n}\Bigl{(}\widetilde{T}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime}),\widetilde{D}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime})\Bigr{)}\overset{d}{=}\lim\limits_{n\to\infty}\frac{1}{n}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime})\overset{d}{=}(\mathbf{X}_{1},\mathbf{X}_{2}).\qed$
5.1. Proofs of Theorems 1.10 and 1.11
We first prove a lemma that connects the condition stated in Theorem 1.10 to the condition in Theorem 1.4.
Lemma 5.4.
Let $\sigma:\mathbb{Z}\to\mathbb{Z}$ be a finite permutation, decomposed as a product of nearest-neighbor transpositions $(k_{m},k_{m}+1)(k_{m-1},k_{m-1}+1)\cdots(k_{1},k_{1}+1)$, where $k_{1},\ldots,k_{n}\in\mathbb{Z}$. Assume that this decomposition has the minimal number of terms. Then, if $x\in\mathbb{Z}$ satisfies $\sigma(\mathbb{Z}_{<x})=\mathbb{Z}_{<x}$, then $x\neq k_{\ell}+1$ for all $1\leq\ell\leq n$. Similarly, if $\sigma(\mathbb{Z}_{>x})=\mathbb{Z}_{>x}$, then $x\neq k_{\ell}$ for all $1\leq\ell\leq m$.
Proof.
We prove the first statement; the second has a symmetric proof. If $\sigma(\mathbb{Z}_{<x})=\mathbb{Z}_{<x}$, then $\sigma$ can be decomposed as $\sigma=\sigma_{1}\sigma_{2}$, where $\sigma_{1}$ permutes only elements of $\mathbb{Z}_{<x}$, and $\sigma_{2}$ permutes only elements of $\mathbb{Z}_{\geq x}$. By minimality of the decomposition $\sigma=(k_{m},k_{m}+1)(k_{m-1},k_{m-1}+1)\cdots(k_{1},k_{1}+1)$, we must have that
$\sigma_{1}=(k_{\ell_{r}},k_{\ell_{r}+1})\cdots(k_{\ell_{1}},k_{\ell_{1}+1}),\quad\text{and}\quad\sigma_{2}=(k_{p_{n}},k_{p_{n}+1})\cdots(k_{p_{1}},k_{p_{1}+1})$
for disjoint subsequences $k_{\ell_{1}},\ldots,k_{\ell_{r}}$ and $k_{p_{1}},\ldots,k_{p_{n}}$ with $r+n=m$. Since $\sigma_{1}$ only permutes elements of $\mathbb{Z}_{<x}$, we have that $k_{\ell_{t}}+1<x$ for $1\leq t\leq r$. Since $\sigma_{2}$ only permutes elements of $\mathbb{Z}_{\geq x}$, $k_{p_{t}}+1>x$ for $1\leq t\leq n$. Hence, $x\neq k_{\ell}+1$ for $1\leq\ell\leq m$. ∎
Proof of Theorem 1.10..
For an arbitrary finite permutation $\sigma$, decompose it as a product of nearest-neighbor transpositions $\sigma=(k_{m},k_{m}+1)(k_{m-1},k_{m-1}+1)\cdots(k_{1},k_{1}+1)$. Take this decomposition to have the minimal number of terms. Then, we consider the action of the map $\mathcal{P}_{k_{m}}\cdots\mathcal{P}_{k_{1}}$ on the weights $\mathbf{W}$, as in Theorem 1.4. By Lemma 5.4, the condition (1.12) and the minimality of the decomposition ensures that the conditions of Theorem 1.4 are met, so
$Z^{\mathcal{P}_{k_{m}}\cdots\mathcal{P}_{k_{1}}\mathbf{W}}(V\mid U)=Z^{\mathbf{W}}(V\mid U),\quad\text{for all }(U,V)\in\mathcal{U}^{\sigma}.$
Furthermore, by successively applying the Burke property in Proposition 5.1, we see that each operator $\mathcal{P}_{k}$ swaps the column parameters in columns $k$ and $k+1$, so we have that $\mathcal{P}_{k_{m}}\cdots\mathcal{P}_{k_{1}}\mathbf{W}\overset{d}{=}\mathbf{W}^{\sigma\mathbf{a},\mathbf{b}}$. This completes the proof.
In the last-passage percolation setting, the proof is the same, instead using the invariance statement in Theorem 1.8 and the Burke properties in Propositions 5.2 and 5.3. ∎
Proof of Theorem 1.11..
This is proved similarly as Theorem 1.10. Consider the independent weights in the finite grid $\llbracket-M,M\rrbracket\times\llbracket-M,M\rrbracket$, and extend these periodically. Consider first endpoint pairs $(U,V)$ that consist of points lying in the finite grid $\llbracket-M,M\rrbracket\times\llbracket-M,M\rrbracket$, and apply the partition function invariance in Theorem 1.4 and the Burke property in Proposition 5.1 (with Lemma 5.4 to connect the two conditions) to both rows and columns, with $N=2M+1$. The result follows by sending $M\to\infty$. ∎
6. Jointly invariant measures: proofs of Theorems 1.12 and 1.13
We prove two theorems in this section, starting with Theorem 1.12.
6.1. Proof of Theorem 1.12
Proof of Theorem 1.12..
We prove the case when the driving weights are geometric random variables, as the exponential case is proved the same way. The key is the Burke property in Proposition 5.2. The proof of the case when the driving weights are exponential is the same, instead using the Burke property in Proposition 5.3. One can readily see that, for any $q\in(0,1)$, if $\mathbf{X}\in\mathbb{Z}_{\geq 0}^{\mathbb{Z}_{N}}$ is an i.i.d. collection of $\text{Geom}(r)$ random variables, then the conditional law of $\mathbf{X}$, given $\mathfrak{s}(\mathbf{X})=\theta$ is the uniform distribution on $\mathbb{Z}_{\theta,\geq 0}^{\mathbb{Z}_{N}}$. Since $\mathfrak{s}(D(\mathbf{X}_{1},\mathbf{X}_{2}))=\mathfrak{s}(\mathbf{X}_{2})$ by Lemma A.4(iv), to prove the invariance result, it suffices to show the following: If $r_{1},r_{2},q\in(0,1)$, and $\mathbf{X}_{1},\mathbf{X}_{2},\mathbf{W}$ are mutually independent vectors of i.i.d. geometric random variables with respective parameters $r_{1},r_{2},q$, then
$(\mathbf{U}_{1},\mathbf{U}_{2})\overset{d}{=}\bigl{(}\widetilde{D}(\mathbf{W},\mathbf{U}_{1}),\widetilde{D}(\mathbf{W},\mathbf{U}_{2})\bigr{)},\quad\text{where}$
$\mathbf{U}_{1}=\mathbf{X}_{1},\quad\text{and}\quad\mathbf{U}_{2}=\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2}).$
By independence of $\mathbf{X}_{1},\mathbf{X}_{2},\mathbf{W}$ and the Burke property in Proposition 5.2,
$(\widetilde{D}(\mathbf{W},\mathbf{X}_{1}),\widetilde{T}(\mathbf{W},\mathbf{X}_{1}),\mathbf{X}_{2})\overset{d}{=}(\mathbf{X}_{1},\mathbf{W},\mathbf{X}_{2}),$ (6.1)
and similarly,
$(\mathbf{X}_{1},\mathbf{X}_{2})\overset{d}{=}\bigl{(}\mathbf{X}_{1},\widetilde{D}(\mathbf{W},\mathbf{X}_{2})\bigr{)}$ (6.2)
By definition of $\mathbf{U}_{1}$ and $\mathbf{U}_{2}$ along with the braid relation in Theorem 1.9 (specifically the zero-temperature version of (3.3)), we have
$\bigl{(}\widetilde{D}(\mathbf{W},\mathbf{U}_{1}),\widetilde{D}(\mathbf{W},\mathbf{U}_{2})\bigr{)}$ $=\Bigl{(}\widetilde{D}(\mathbf{W},\mathbf{X}_{1}),\widetilde{D}\bigl{(}\widetilde{D}(\mathbf{W},\mathbf{X}_{1}),\widetilde{D}(\widetilde{T}(\mathbf{W},\mathbf{X}_{1}),\mathbf{X}_{2})\bigr{)}\Bigr{)}$
[…]
We turn to proving the distributional equality (1.17): By another application of the Burke property, we have
$(\mathbf{X}_{1},\mathbf{X}_{2})\overset{d}{=}\bigl{(}\widetilde{D}(\mathbf{X}_{2},\mathbf{X}_{1}),\widetilde{T}(\mathbf{X}_{2},\mathbf{X}_{1})\bigr{)}.$
Then, we have
$\bigl{(}\mathbf{X}_{1},\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})\bigr{)}\overset{d}{=}\bigl{(}\widetilde{D}(\mathbf{X}_{2},\mathbf{X}_{1}),\widetilde{D}(\widetilde{D}(\mathbf{X}_{2},\mathbf{X}_{1}),\widetilde{T}(\mathbf{X}_{2},\mathbf{X}_{1}))\bigr{)}=\bigl{(}\widetilde{D}(\mathbf{X}_{2},\mathbf{X}_{1}),\mathbf{X}_{2}\bigr{)},$
where in the last equality we used the fact that $(\mathbf{X}_{1},\mathbf{X}_{2})\mapsto(\widetilde{T}(\mathbf{X}_{2},\mathbf{X}_{1}),\widetilde{D}(\mathbf{X}_{2},\mathbf{X}_{1}))$ is an involution, which is proved in Lemma A.2. ∎
6.2. Cyclic Interval Notation
Before proving Theorem 1.13, we discuss some conventions around cyclic notation. Recall that we take $\mathbb{Z}_{N}$ to be the cyclic group of order $N$.
Definition 6.1.
For $i,j\in\mathbb{Z}_{N}$, we define the periodic intervals
$(i,j]:=\begin{cases}\varnothing&i=j\\ \{i+1,i+2,\ldots,j\}&i\neq j\end{cases}$
$[i,j):=\begin{cases}\varnothing&i=j\\ \{i,i+1,\ldots,j-1\}&i\neq j\end{cases}$
$[i,j]:=\begin{cases}i&i=j\\ \{i,i+1,\ldots,j\}&i\neq j\end{cases}$
$(i,j):=\begin{cases}\varnothing&j\in\{i,i+1\}\\ \{i+1,i+2,\ldots,j-1\}&i\neq j.\end{cases}$
Using this cyclic notation, we now prove the following lemma that will be used several times in the proof of Theorem 1.13.
Lemma 6.2.
Let $\mathbf{Y}\in\mathbb{R}^{\mathbb{Z}_{N}}$. Then, the following hold.
  1. (i)
    If $a,b\in\mathbb{Z}_{N}$ where $a\neq b$, then for any $c\in[a,b)$, $Y_{(a,c]}+Y_{[c+1,b]}=Y_{[a+1,b]}$.
  2. (ii)
    For $a,c\in\mathbb{Z}_{N}$, we have $Y_{(a,c]}+Y_{[c+1,a]}=\mathfrak{s}(\mathbf{Y})$.
Proof.
Item (i): We prove this in two cases. First, if $c=a$, then $Y_{(a,c]}=0$, so $Y_{(a,c]}+Y_{[c+1,b]}=Y_{[a+1,b]}.$
Next, if $c\in[a,b)$, but $c\neq a$, then $Y_{(a,c]}=\sum_{\ell=a+1}^{c}Y_{\ell}$ and $Y_{[c+1,b]}=\sum_{\ell=c+1}^{b}Y_{\ell},$ so
$Y_{(a,c]}+Y_{[c+1,b]}=\sum_{\ell=a+1}^{c}Y_{\ell}+\sum_{\ell=c+1}^{b}Y_{\ell}=\sum_{\ell=a+1}^{b}Y_{\ell}=Y_{[a+1,b]},$
where the sums are taken in cyclic order. The condition $c\in[a,b)$ ensures there is no double-counting.
Item (ii): We likewise prove this statement in two cases. If $a=c$, then $Y_{(a,c]}=0$ and $Y_{[c+1,a]}=Y_{[a+1,a]}=\sum_{k\in\mathbb{Z}_{N}}Y_{k}=\mathfrak{s}(\mathbf{Y})$.
On the other hand, if $a\neq c$ then
$Y_{(a,c]}+Y_{[c+1,a]}=\sum_{\ell=a+1}^{c}Y_{\ell}+\sum_{\ell=c+1}^{a}Y_{\ell}=\sum_{\ell\in\mathbb{Z}_{N}}Y_{\ell}=\mathfrak{s}({\mathbf{Y}}).\qed$
6.3. Proof of Theorem 1.13
We start with proving several algebraic lemmas that are used as inputs in the proof of the Theorem. Recall the definitions of the spaces $\mathbb{R}_{\theta}^{\mathbb{Z}_{N}},\mathbb{R}_{\theta,\geq 0}^{\mathbb{Z}_{N}}$, and $\mathcal{Z}_{\theta_{1},\theta_{2}}$ defined in (1.16) and (1.18). The first lemma below is the key to our combinatorial description of the joint law in Theorem 1.13. Specifically, it turns the problem into understanding a single quantity–namely $\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]}$.
Lemma 6.3.
Let $\mathbf{X}_{1}\in\mathbb{R}^{\mathbb{Z}_{N}}_{\theta_{1}}$ and $\mathbf{X}_{2}\in\mathbb{R}^{\mathbb{Z}_{N}}_{\theta_{2}}$ where $0\leq\theta_{1}\leq\theta_{2}$. Then, for $i\in\mathbb{Z}_{N}$,
$\widetilde{D}_{i}(\mathbf{X}_{1},\mathbf{X}_{2})=X_{1,i}+\max(0,\theta_{2}-\theta_{1}-\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]}),$
where $Y_{\ell}=X_{2,\ell+1}-X_{1,\ell}$.
Proof.
Notice that when $\theta_{2}\geq\theta_{1}$, we have that $Y_{i}+\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}=\max\bigl{(}\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]},\theta_{2}-\theta_{1}\bigr{)}$ for all $i\in\mathbb{Z}_{N}$. Thus, we have
$\widetilde{D}_{i}(\mathbf{X}_{1},\mathbf{X}_{2})$ $=X_{2,i+1}+\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}-\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]}$
$=X_{1,i}+Y_{i}+\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}-\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]}$
$=X_{1,i}+\max(\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]},\theta_{2}-\theta_{1})-\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]})$
$=X_{1,i}+\max(0,\theta_{2}-\theta_{1}-\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]})\qed$
Since $\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]}\geq Y_{(i-1,i-1]}=0$, we have the following corollary of Lemma 6.3.
Corollary 6.4.
Let $\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{R}^{\mathbb{Z}_{N}}_{\theta}$ for some $\theta\in\mathbb{R}$. Then $\widetilde{D}_{i}(\mathbf{X}_{1},\mathbf{X}_{2})=X_{1,i}$ for all $i\in\mathbb{Z}_{N}$.
Lemma 6.5.
Let $\mathbf{X}_{1}\in\mathbb{R}^{\mathbb{Z}_{N}}_{\theta_{1}}$ and $\mathbf{X}_{2}\in\mathbb{R}^{\mathbb{Z}_{N}}_{\theta_{2}}$ where $\theta_{1}\leq\theta_{2}$. Set $Y_{i}:=X_{2,i+1}-X_{1,i}$ for all $i\in\mathbb{Z}_{N}$. If $Y_{i}\leq 0$ for some $i\in\mathbb{Z}_{N}$, then $\widetilde{D}_{i+1}(\mathbf{X}_{1},\mathbf{X}_{2})=X_{1,i+1}$ and $\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}\geq\theta_{2}-\theta_{1}$.
Proof.
Observe that $Y_{(i,i-1]}=Y_{i+1}+Y_{i+2}+\ldots+Y_{i-1}\geq Y_{i}+Y_{i+1}+Y_{i+2}+\ldots+Y_{i-1}=\theta_{2}-\theta_{1}$ since $Y_{i}\leq 0$. Hence, $\theta_{2}-\theta_{1}-Y_{(i,i-1]}\leq 0$ so $\max(0,\theta_{2}-\theta_{1}-\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]})=0$. The lemma now follows from Lemma 6.3. ∎
Lemma 6.6.
Let $\mathbf{X}_{1}\in\mathbb{R}^{\mathbb{Z}_{N}}_{\theta_{1}}$ and $\mathbf{X}_{2}\in\mathbb{R}^{\mathbb{Z}_{N}}_{\theta_{2}}$ where $\theta_{1}\leq\theta_{2}$. Set $Y_{i}:=X_{2,i+1}-X_{1,i}$ and let $\widetilde{D}_{i}$ denote the $i$th component of $\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})$. If $\widetilde{D}_{i}>X_{1,i}$ and $\widetilde{D}_{i+1}=X_{1,i+1}$ for some $i\in\mathbb{Z}_{N}$, then $Y_{i}\leq 0$.
Proof.
Suppose for the sake of contradiction that $Y_{i}>0$. Then, $Y_{(i,i-1]}=\theta_{2}-\theta_{1}-Y_{i}<\theta_{2}-\theta_{1}$. Since $\widetilde{D}_{i+1}=X_{1,i+1}$, it follows by Lemma 6.3 that $\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}\geq\theta_{2}-\theta_{1}$, so there exists $m\in\mathbb{Z}_{N}$ such that $Y_{(i,m]}\geq\theta_{2}-\theta_{1}$. Since $Y_{(i,i-1]}=0<\theta_{2}-\theta_{1}$, we have $m\neq i-1$, and therefore, $Y_{(i-1,m]}=Y_{i}+Y_{(i,m]}>\theta_{2}-\theta_{1}$. By another application of Lemma 6.3, this implies that $\widetilde{D}_{i}=X_{1,i}$, a contradiction to our assumption. ∎
Lemma 6.7.
Let $\mathbf{X}_{1}\in\mathbb{R}^{\mathbb{Z}_{N}}_{\theta_{1}}$ and $\mathbf{X}_{2}\in\mathbb{R}^{\mathbb{Z}_{N}}_{\theta_{2}}$ where $\theta_{1}\leq\theta_{2}$. Set $Y_{i}:=X_{2,i+1}-X_{1,i}$ and let $\widetilde{D}_{i}$ denote the $i$th component of $\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})$. If $\widetilde{D}_{i}>X_{1,i}$, then $Y_{i}<\theta_{2}-\theta_{1}$.
Proof.
By Lemma 6.3, $\theta_{2}-\theta_{1}-\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]}>0$, so since $Y_{i}=Y_{(i-1,i]}$, we have $\theta_{2}-\theta_{1}-Y_{i}>0$. ∎
We now complete the proof of Theorem 1.13.
Proof.
of Theorem 1.13. Let $0\leq\theta_{1}<\theta_{2}$, let $\mathbf{X}_{1}$ be uniformly distributed on the set $\mathbb{Z}_{\theta_{1},\geq 0}^{\mathbb{Z}_{N}}$, and let $\mathbf{X}_{2}$ be uniformly distributed on the set $\mathbb{Z}_{\theta_{2},\geq 0}^{\mathbb{Z}_{N}}$. By Lemmas 6.3, A.3, and A.4(iv), the joint distribution of $(\mathbf{X}_{1},\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2}))$ is supported on the set
$\mathcal{Z}_{\theta_{1},\theta_{2}}=\{(\mathbf{u}_{1},\mathbf{u}_{2})\in\mathbb{Z}^{\mathbb{Z}_{N}}_{\theta_{1},\geq 0}\times\mathbb{Z}^{\mathbb{Z}_{N}}_{\theta_{2},\geq 0}\mid\mathbf{u}_{1}\leq\mathbf{u}_{2}\}.$
Observe that
$|\mathbb{Z}_{\theta_{1},\geq 0}^{\mathbb{Z}_{N}}\times\mathbb{Z}_{\theta_{2},\geq 0}^{\mathbb{Z}_{N}}|=\binom{\theta_{1}+N-1}{N-1}\binom{\theta_{2}+N-1}{N-1}$,
so for any $\mathbf{u}_{1}\in\mathbb{Z}_{\theta_{1},\geq 0}^{\mathbb{Z}_{N}}$ and $\mathbf{u}_{2}\in\mathbb{Z}_{\theta 2,+}^{\mathbb{Z}_{N}}$ with $\mathbf{u}_{1}\leq\mathbf{u}_{2}$ coordinate-wise, we have
$\mathbb{P}(\mathbf{X}_{1}=\mathbf{u}_{1},\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})=\mathbf{u}_{2})=\frac{\alpha(\mathbf{u}_{1},\mathbf{u}_{2})}{\binom{\theta_{1}+N-1}{N-1}\binom{\theta_{2}+N-1}{N-1}}$,
where $\alpha(\mathbf{u}_{1},\mathbf{u}_{2})$ is the number of vectors $\mathbf{X}_{2}\in\mathbb{Z}_{\theta_{2},\geq 0}^{\mathbb{Z}_{N}}$ satisfying $\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})=\mathbf{u}_{2}$, when $\mathbf{X}_{1}=\mathbf{u}_{1}$.
We compute the value of $\alpha(\mathbf{u}_{1},\mathbf{u}_{2})$ in several cases, depending on how many indices $i\in\mathbb{Z}_{N}$ satisfy the equality $u_{1,i}=u_{2,i}$. We make a change of coordinates $\mathbf{Y}=(Y_{i})_{i\in\mathbb{Z}_{N}}$, where $Y_{i}=X_{2,i+1}-u_{1,i}$. We immediately see that the number of vectors $\mathbf{X}_{2}$ satisfying $\mathbf{u}_{2}=\widetilde{D}(\mathbf{u}_{1},\mathbf{X}_{2})$ is in bijection with the set of corresponding vectors $\mathbf{Y}$.
Case 1: Zero Equalities: In this case, $u_{2,i}>u_{1,i}$ for all $i\in\mathbb{Z}_{N}$. Assume that $\mathbf{u}_{2}=\widetilde{D}(\mathbf{u}_{1},\mathbf{X}_{2})$ for some $\mathbf{X}_{2}\in\mathbb{Z}_{\theta_{2},\geq 0}^{\mathbb{Z}_{N}}$. By Lemma 6.3, we have $\theta_{2}-\theta_{1}>\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]}$ for all $i\in\mathbb{Z}_{N}$. By the contrapositive of Lemma 6.5, we have that $Y_{i}>0$ for all $i\in\mathbb{Z}_{N}$, so $Y_{(i-1,j]}$ is maximized when we include as many $Y_{i}$s in the sum as possible. Thus, $\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]}=Y_{(i-1,i-2]}=\theta_{2}-\theta_{1}-Y_{i-1}$. Hence, using Lemma 6.3 in the last equality below, we have
$X_{2,i}$ $=u_{1,i-1}+Y_{i-1}$
$=u_{1,i-1}+\theta_{2}-\theta_{1}-\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]}$
$=u_{1,i-1}+u_{2,i}-u_{1,i}$
On the other hand, if we define $X_{2,i}=u_{1,i-1}+u_{2,i}-u_{1,i}$, then $\mathfrak{s}(\mathbf{X}_{2})=\theta_{2}$, $X_{2,i}\geq 0$ for all $i\in\mathbb{Z}_{N}$, and $Y_{i}=u_{2,i+1}-u_{1,i+1}>0$ for all $i\in\mathbb{Z}_{N}$, and by applying the steps above in reverse, we obtain that $\mathbf{u}_{2}=\widetilde{D}(\mathbf{u}_{1},\mathbf{X}_{2})$. Thus, there is a unique $\mathbf{X}_{2}\in\mathbb{Z}_{\theta_{2},\geq 0}^{\mathbb{Z}_{N}}$ such that $\widetilde{D}(\mathbf{u}_{1},\mathbf{X}_{2})=\mathbf{u}_{2}$, and so $\alpha(\mathbf{u}_{1},\mathbf{u}_{2})=1$.
Case 2: $\mathbf{<N}$ Equalities: In this case, there exists $i\in\mathbb{Z}_{N}$ such that $u_{2,i-1}>u_{1,i-1}$ and $u_{2,i}=u_{1,i}$. By invariance of the uniform distribution on $\mathbb{Z}_{\theta,\geq 0}^{\mathbb{Z}_{N}}$ under cyclic permutations and Lemma A.6, the joint distribution of $\bigl{(}\mathbf{X}_{1},\mathbf{\widetilde{D}}\bigr{)}$ is invariant under cyclic permutations, so we may, without loss of generality, assume that $u_{2,1}=u_{1,1}$ and $u_{2,N}>u_{1,N}$. Then, starting from $i=1$, there is a collection of sequences of equalities and inequalities, when comparing $u_{1,i}$ with $u_{2,i}$. Let $k_{m}\in\mathbb{Z}_{N}$ be the ending index of the $m$th consecutive equality and $\ell_{m}\in\mathbb{Z}_{N}$ be the ending index of the $m$th consecutive inequality, as shown in Figure 5. To be precise, we write
$k_{1}=\max\{t\in\mathbb{Z}_{\geq 1}\mid u_{2,j}=u_{1,j}\quad\forall 1\leq j\leq t\}.$
Then, we can recursively define
$\ell_{m}=\max\{t\in\mathbb{Z}_{\geq k_{m}+1}\mid u_{2,j}>u_{1,j}\quad\forall k_{m}+1\leq j\leq t\}$ (6.3)
$k_{m}=\max\{t\in\mathbb{Z}_{\geq\ell_{m-1}+1}\mid u_{2,j}=u_{1,j}\quad\forall\ell_{m-1}+1\leq j\leq t\},$ (6.4)
until $m=s$, where $s$ is the integer satisfying $\ell_{s}=N$. Such an integer is guaranteed to exist since $u_{2,N}>u_{1,N}$, but $u_{1,1}=u_{2,1}$. Recalling that $\mathbf{u}_{1},\mathbf{u}_{2}$ are vectors with indices in $\mathbb{Z}_{N}$, we can embed $k_{m},\ell_{m}$ in $\mathbb{Z}_{N}$. Moreover, if we let $k_{m}=k_{m^{\prime}}$ and $\ell_{m}=\ell_{m^{\prime}}$ for all $m,m^{\prime}\in\mathbb{Z}$ such that $m\equiv m^{\prime}\pmod{s}$, then $k_{m},\ell_{m}\in\mathbb{Z}_{N}$ are defined for nonpositive $m$ as well as for $m>s$. Note that we sometimes add $N\mathbf{1}_{m=s}$ when working with the lengths or indices of vectors, since, for instance, when we are considering a vector whose length is the number of indices between $\ell_{s}$ and $k_{1},$ we must ‘wrap-around’ indices. Thus, the length of the vector would be $(N-\ell_{s})+(k_{1}-0)$ rather than $\ell_{s}-k_{1}$.
We now provide an overview of the proof strategy using lemmas that we will prove in a later subsection. Recall the change of coordinates $\mathbf{Y}$ discussed in the previous case. We claim that the following conditions are necessary and sufficient for $\mathbf{X}_{2}$ to satisfy $\widetilde{D}(\mathbf{u}_{1},\mathbf{X}_{2})=\mathbf{u}_{2}$:
$Y_{j}=u_{2,j+1}-u_{1,j+1}\text{ for all }j\in(k_{m},\ell_{m}),m\in\llbracket 1,s\rrbracket$ (6.5a)
$Y_{[\ell_{m-1},j]}\leq 0\text{ for all }j\in[\ell_{m-1},k_{m}-1],m\in\llbracket 1,s\rrbracket$ (6.5b)
$Y_{[\ell_{m-1},k_{m}]}=u_{2,k_{m}+1}-u_{1,k_{m}+1}\text{ for all }m\in\llbracket 1,s\rrbracket$ (6.5c)
$Y_{j}\geq-u_{1,j}\text{ for all }j\in\mathbb{Z}_{N}$ (6.5d)
Once this is shown, it suffices to count the number of vectors $\mathbf{Y}\in\mathbb{Z}^{\mathbb{Z}_{N}}$ satisfying the conditions above. To do this, we claim that the number of vectors $\mathbf{Y}$ satisfying (6.5) is the same as the number of $s$-tuples of vectors $(\mathbf{a}_{1},\ldots,\mathbf{a}_{s})$ satisfying (1.19). To show this, we first observe that if we choose values of $Y_{j}$ for every $j\in[\ell_{m},k_{m+1}-1]$ and $m\in\llbracket 1,s\rrbracket$, then setting $Y_{k_{m+1}}=u_{2,k_{m+1}+1}-u_{1,k_{m+1}+1}-Y_{[\ell_{m},k_{m+1}-1]}$ and $Y_{j}=u_{2,j+1}-u_{1,j+1}$ for all $j\in(k_{m},\ell_{m})$ and $m\in\llbracket 1,s\rrbracket$, fixes the other values of $Y_{j}$. Hence, if we take $a_{m,p}$ to represent $Y_{\ell_{m}+p-1}$ for $m\in\llbracket 1,s\rrbracket$ and $1\leq p\leq k_{m+1}-\ell_{m}+N\mathbf{1}_{m=s}$, we find the relevant set of equivalent constraints on the free values of $Y_{j}$ to count the total number of possible vectors $\mathbf{Y}$. Namely, we give lower bounds on the vector coordinates $a_{m,p}\geq-u_{1,\ell_{m}+p-1}$ and an upper bound of 0 on the sums of vector coordinates $a_{m,1},a_{m,1}+a_{m,2},\ldots,a_{m,1}+a_{m,2}+\ldots+a_{m,p}$ for $1\leq p\leq k_{m+1}-\ell_{m}+N1_{m=s}$.
We turn now to showing that the conditions (6.5) are necessary and sufficient for $\mathbf{X}_{2}$ to satisfy $\mathbf{u}_{2}=\widetilde{D}(\mathbf{u}_{1},\mathbf{X}_{2})$.
Proof of necessity: First, assuming $\mathbf{u}_{2}=\widetilde{D}(\mathbf{u}_{1},\mathbf{X}_{2})$, we will show that each of the conditions in (6.5) hold.
To do this, we state some intermediate lemmas, whose proofs are postponed until Section 6.3.1.
Lemma 6.8.
If $\ell_{0}=k_{1}+1$, then we have N1N-1 equalities $u_{1,i}=u_{2,i}$, and $s=1$. Moreover, the conditions (6.5) hold.
Lemma 6.9.
We have that $\max_{j\in\mathbb{Z}_{N}}Y_{(k_{m}-1,j]}=Y_{(k_{m}-1,\ell_{m-1}-1]}$ for all $m\in\llbracket 1,s\rrbracket$.
Lemma 6.10.
Whenever $\ell_{0}\neq k_{1}+1$, then we have that $\max_{j\in\mathbb{Z}_{N}}Y_{(k_{m},j]}=Y_{(k_{m},\ell_{m-1}-1]}$ for all $m\in\llbracket 1,s\rrbracket$.
Lemma 6.11.
Whenever $q\in(k_{m},\ell_{m})$, we have that $\max\limits_{j\in\mathbb{Z}_{N}}Y_{(q,j]}=Y_{(q,q-1]}$ for $m\in\llbracket 1,s\rrbracket$.
By Lemma 6.8, we may assume that $\ell_{0}\neq k_{1}+1$. Note that this implies $\ell_{m-1}\neq k_{m}+1$ for all $m\in\llbracket 1,s\rrbracket$; otherwise there would be N1N-1 equalities of the form $u_{2,i}=u_{1,i}$.
By Lemma 6.11, whenever $m\in[1,s]$ and $j\in(k_{m},\ell_{m})$ to get $\max_{i\in\mathbb{Z}_{N}}Y_{(j,i]}=Y_{(j,j-1]}=\theta_{2}-\theta_{1}-Y_{j}$. Then, by assumption on the locations of the equalities, followed by an application of Lemma 6.3,
$u_{1,j+1}<u_{2,j+1}$ $=u_{1,j+1}+\max(0,\theta_{2}-\theta_{1}-\max_{i\in\mathbb{Z}_{N}}Y_{(j,i]})$
$=u_{1,j+1}+\max(0,\theta_{2}-\theta_{1}-(\theta_{2}-\theta_{1}-Y_{j}))$
$=u_{1,j+1}+\max(0,Y_{j})$
By the inequality, we must have $\max(0,Y_{j})=Y_{j}$, and then, $u_{2,j+1}=u_{1,j+1}+Y_{j}$, giving us (6.5a).
Next, since $u_{2,k_{m}}=u_{1,k_{m}}$ by definition of $k_{m}$, Lemma 6.3 implies that $\max_{j\in\mathbb{Z}_{N}}Y_{(k_{m}-1,j]}\geq\theta_{2}-\theta_{1}$. Combined with Lemma 6.9, we have $Y_{(k_{m}-1,\ell_{m-1}-1]}\geq\theta_{2}-\theta_{1}$. Lemma 6.2(i) now tells us that
$Y_{[\ell_{m-1},k_{m}-1]}\leq 0.$ (6.6)
Then, by Lemma 6.2(ii), we have that
$Y_{[\ell_{m-1},j]}=Y_{(k_{m}-1,j]}-Y_{(k_{m}-1,\ell_{m-1}-1]}\leq 0,\quad\text{for }j\in[\ell_{m-1},k_{m}-1),$ (6.7)
and (6.6) and (6.7) together prove (6.5b).
By definition of $k_{m}$, we have $u_{2,k_{m}+1}>u_{1,k_{m}+1}$. Then, by Lemma 6.3, followed by Lemma 6.10, then Lemma 6.2(ii), we observe that, for $m\in\llbracket 1,s\rrbracket$,
$u_{2,k_{m}+1}-u_{1,k_{m}+1}=\theta_{2}-\theta_{1}-\max\limits_{j\in\mathbb{Z}_{N}}Y_{(k_{m},j]}=\theta_{2}-\theta_{1}-Y_{(k_{m},\ell_{m-1}-1]}=Y_{(\ell_{m-1}-1,k_{m}]}=Y_{[\ell_{m-1},k_{m}]},$
proving (6.5c).
Finally, observe that (6.5d) follows directly from the fact that $Y_{i}=X_{2,i+1}-u_{1,i}$ and $X_{2,i}\geq 0$ for all $i\in\mathbb{Z}_{N}$ by definition.
Proof of sufficiency: Next, we assume that (6.5) holds, and prove that $\mathbf{u}_{2}=\widetilde{D}(\mathbf{u}_{1},\mathbf{X}_{2})$. We first show that (6.5) implies that $\mathfrak{s}(\mathbf{Y})=\theta_{2}-\theta_{1}$, which is equivalent to $\mathfrak{s}(\mathbf{X}_{2})=\theta_{2}$. To see this, observe that
$\mathfrak{s}(\mathbf{Y})$ $=\sum\limits_{m\in\llbracket 1,s\rrbracket}Y_{[\ell_{m-1},\ell_{m})}$
$=\sum\limits_{m\in\llbracket 1,s\rrbracket}Y_{[\ell_{m-1},k_{m}]}+\sum\limits_{m\in\llbracket 1,s\rrbracket}Y_{(k_{m},\ell_{m})}$
$=\sum\limits_{m\in\llbracket 1,s\rrbracket}\left(u_{2,k_{m}+1}-u_{1,k_{m}+1}\right)+\sum\limits_{m\in\llbracket 1,s\rrbracket}\sum\limits_{j\in(k_{m},\ell_{m})}\left(u_{2,j+1}-u_{1,j+1}\right)$
$=\sum\limits_{m\in\llbracket 1,s\rrbracket}\sum\limits_{j\in[k_{m}+1,\ell_{m}+1)}\left(u_{2,j}-u_{1,j}\right)=\sum_{j\in\mathbb{Z}_{N}}(u_{2,j}-u_{1,j})$
$=\theta_{2}-\theta_{1},$
where the second equality follows from Lemma 6.2(i), the third equality follows from (6.5a) and (6.5c), and the final equality follows since $(\mathbf{u}_{1},\mathbf{u}_{2})\in\mathcal{Z}_{\theta_{1},\theta_{2}}$. Combined with the condition (6.5d), we have $\mathbf{X}_{2}\in\mathbb{Z}_{\theta_{2},\geq 0}^{\mathbb{Z}_{N}}$.
We now show that $\widetilde{D}_{i}(\mathbf{u}_{1},\mathbf{X}_{2})=u_{2,i}$ for $i\in(k_{m},\ell_{m}+1),m\in\llbracket 1,s\rrbracket$. In other words, we prove that the value of $\widetilde{D}(\mathbf{u}_{1},\mathbf{X}_{2})$ aligns with $\mathbf{u}_{2}$ at every index where $u_{2,i}>u_{1,i}$. This suffices to show equality at every other index since we know $\widetilde{D}_{i}(\mathbf{u}_{1},\mathbf{X}_{2})\geq u_{1,i}$ and $\mathfrak{s}(\widetilde{D}(\mathbf{u}_{1},\mathbf{X}_{2}))=\mathfrak{s}(\mathbf{X}_{2})=\mathfrak{s}(\mathbf{u}_{1})+\mathfrak{s}(\mathbf{Y})=\mathfrak{s}(\mathbf{u}_{2})$.
We note that (6.5a) gives that $Y_{j}>0$ for $j\in(k_{m},\ell_{m}),m\in[1,s]$. Hence, when trying to maximize any cyclic sum $Y_{(i,j]}$, it is strictly better to extend the bounds of the cyclic sum to include these positive values. Similarly, (6.5c) implies that it is strictly better to contain the sum $Y_{[\ell_{m},k_{m+1}]}$ in the cyclic sum if possible, while (6.5b) implies that it is never better to only add a partial component of $Y_{[\ell_{m},k_{m+1}]}$.
Therefore, for $m\in\llbracket 1,s\rrbracket$, we have that $\max\limits_{j\in\mathbb{Z}_{N}}Y_{(k_{m},j]}=Y_{(k_{m},\ell_{m-1}-1]}$, and by Lemma 6.3,
$\widetilde{D}_{k_{m}+1}(\mathbf{u}_{1},\mathbf{X}_{2})$ $=u_{1,k_{m}+1}+\max(0,\theta_{2}-\theta_{1}-\max\limits_{j\in\mathbb{Z}_{N}}Y_{(k_{m},j]})$
$=u_{1,k_{m}+1}+\max(0,\theta_{2}-\theta_{1}-Y_{(k_{m},\ell_{m-1}-1]})$
$=u_{1,k_{m}+1}+\max(0,Y_{[\ell_{m-1},k_{m}]})$
$=u_{1,k_{m}+1}+Y_{[\ell_{m-1},k_{m}]}=u_{2,k_{m}+1}.$
Moreover, for $i\in(k_{m}+1,\ell_{m}+1)$, we find that $\max\limits_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]}=Y_{(i-1,i-2]}$ and
$\widetilde{D}_{i}(\mathbf{u}_{1},\mathbf{X}_{2})$ $=u_{1,i}+\max(0,\theta_{2}-\theta_{1}-\max\limits_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]})$
$=u_{1,i}+\max(0,\theta_{2}-\theta_{1}-Y_{(i-1,i-2]})$
$=u_{1,i}+\max(0,Y_{i-1})$
$=u_{1,i}+Y_{i-1}=u_{2,i},$
as desired. ∎
6.3.1. Proofs of Lemmas 6.8-6.11
[…]
Figure 7. Illustration of the N1N-1 equalities case discussed in Lemma 6.8. Note that $k_{m}=k_{1},\ell_{m}=\ell_{1}$ for all $m$ since $s=1$.
Proof of Lemma 6.8..
It follows immediately that if $\ell_{0}=k_{1}+1$, then $u_{1,i}=u_{2,i}$ for all $i\neq N$. See Figure 7. Thus, $s=1$, and the condition (6.5a) is empty in this case. Additionally, since $i=N$ is the unique index where $u_{2,i}>u_{1,i}$, and by Lemma 6.3, we see that
$\theta_{2}-\theta_{1}$ $=\sum\limits_{i\in\mathbb{Z}_{N}}(u_{2,i}-u_{1,i})=u_{2,N}-u_{1,N}=\theta_{2}-\theta_{1}-\max\limits_{j\in\mathbb{Z}_{N}}Y_{(N-1,j]},$
Hence, $\max_{j\in\mathbb{Z}_{N}}Y_{(N-1,j]}=0$, so $Y_{(N-1,j]}\leq 0$ for all $j\in\mathbb{Z}_{N}$. Combining this result with the observation that $Y_{N}\leq 0$ by Lemma 6.6, we get (6.5b). Moreover, we get (6.5c) from the observation that $Y_{[\ell_{m},k_{m+1}]}=Y_{[N,N-1]}=\theta_{2}-\theta_{1}=u_{2,N}-u_{1,N}$. Note that (6.5d) follows since $Y_{j}=u_{2,j}-u_{1,j}\geq-u_{1,j}$. ∎
Proof of Lemma 6.9..
Let $m\in\llbracket 1,s\rrbracket$. We show that show that $Y_{(k_{m}-1,j]}\leq Y_{(k_{m}-1,\ell_{m-1}-1]}$ for all $j\in\mathbb{Z}_{N}$. We observe by definition of $\ell_{m-1}$ that
$u_{2,\ell_{m-1}}>u_{1,\ell_{m-1}},\quad\text{and}\quad u_{2,\ell_{m-1}+1}=u_{1,\ell_{m-1}+1}.$ (6.8)
We consider three cases for the value of $j\in\mathbb{Z}_{N}$.
  1. (1)
    Case 2.1: $j\in[k_{m}-1,\ell_{m-1}-1)$. This case is depicted in Figure 8. By the inequality in (6.8) and Lemma 6.3, we have that
    $Y_{(\ell_{m-1}-1,j]}\leq\max_{i\in\mathbb{Z}_{N}}Y_{(\ell_{m-1}-1,i]}<\theta_{2}-\theta_{1},$
    so $Y_{[j+1,\ell_{m-1}-1]}=\theta_{2}-\theta_{1}-Y_{(\ell_{m-1}-1,j]}>0$ by Lemma 6.2(ii). Then, an application of Lemma 6.2(i) yields
    $Y_{(k_{m}-1,\ell_{m-1}-1]}=Y_{(k_{m}-1,j]}+Y_{[j+1,\ell_{m-1}-1]}>Y_{(k_{m}-1,j]}.$
[…]
    Figure 8. Illustrations of $k_{m}-1,\ell_{m-1}-1)$ in Lemma 6.9, Case 2.1 with (a) $m\neq 1$ and (b) $m=1$ (so the interval is $k_{1}-1,\ell_{s}-1)$).
  2. (2)
    Case 2.2: $j=\ell_{m-1}$ where $k_{m}-1\neq\ell_{m-1}$. Note the case where $j=k_{m}-1$ was already handled in Case 2.1. By (6.8) and Lemma 6.6, we know that $Y_{\ell_{m-1}}\leq 0$, so $Y_{(k_{m}-1,\ell_{m-1}]}=Y_{(k_{m}-1,\ell_{m-1}-1]}+Y_{\ell_{m-1}}\leq Y_{(k_{m}-1,\ell_{m-1}-1]}$.
  3. (3)
    Case 2.3: $k_{m}-1\notin\{\ell_{m-1},\ell_{m-1}+1\}$, and $j\in(\ell_{m-1},k_{m}-1)$. This case is depicted in Figure 9. Note this case only occurs when there are at least three consecutive equalities in the string ending at index $k_{m}$. Assume for the sake of contradiction that there exists some $j\in(\ell_{m-1},k_{m}-1)$ such that $Y_{(k_{m}-1,j]}\geq Y_{(k_{m}-1,i]}$ for all $i\in\mathbb{Z}_{N}$ and $Y_{(k_{m}-1,j]}>Y_{(k_{m}-1,\ell_{m-1}-1]}$. When starting from $k_{m}$ and moving in cyclic order, take $j$ to be the first index that fulfills this condition. Then, by Cases 2.1 and 2.2, we have
    $Y_{(k_{m}-1,j]}>Y_{(k_{m}-1,i]},\quad\text{for all }i\in[k_{m},j).$ (6.9)
    On the other hand, since $Y_{(k_{m}-1,j]}\geq Y_{(k_{m}-1,i]}$ for all $i\in\mathbb{Z}_{N}$, Lemma 6.2(i) implies
    $Y_{(j,i]}=Y_{[j+1,i]}=Y_{(k_{m}-1,i]}-Y_{(k_{m}-1,j]}\leq 0\text{ for all }i\in(j,k_{m}-1).$ (6.10)
    Moreover, as $u_{2,k_{m}}=u_{1,k_{m}}$ by definition of $k_{m}$ and since $j$ was chosen to be a maximizer, Lemma 6.3 implies that $Y_{(k_{m}-1,j]}\geq\theta_{2}-\theta_{1}=\mathfrak{s}(\mathbf{Y})$, so by Lemma 6.2(ii),
    $Y_{(j,k_{m}-1]}=Y_{[j+1,k_{m}-1]}\leq 0.$ (6.11)
    Since $j\in(\ell_{m-1},k_{m}-1)$, we must have that $u_{1,j+1}=u_{2,j+1}$. Then, Lemma 6.3 implies that
    $\max_{i\in\mathbb{Z}_{N}}Y_{(j,i]}\geq\theta_{2}-\theta_{1}>0.$ (6.12)
    Hence, we have that
    $\max_{i\in\mathbb{Z}_{N}}Y_{(j,i]}=\max(0,\max_{i\in\mathbb{Z}_{N}\setminus\{j\}}Y_{(j,i]})=\max_{i\in\mathbb{Z}_{N}\setminus\{j\}}Y_{(j,i]}.$ (6.13)
    Then, using (6.13), (6.10), (6.11) and Lemma 6.2(i) in the first step below, followed by Lemma 6.2(ii), and then followed by (6.9), we obtain
    $\max_{i\in\mathbb{Z}_{N}}Y_{(j,i]}$ $=\max_{i\in\mathbb{Z}_{N}\setminus[j,k_{m}-1]}(Y_{(j,k_{m}-1]}+Y_{[k_{m},i]})$
    $=\theta_{2}-\theta_{1}-Y_{[k_{m},j]}+\max_{i\in\mathbb{Z}_{N}\setminus[j,k_{m}-1]}Y_{(k_{m}-1,i]}$
    $<\theta_{2}-\theta_{1}-Y_{[k_{m},j]}+Y_{(k_{m}-1,j]}=\theta_{2}-\theta_{1},$
    and this contradicts (6.12). ∎
[…]
Figure 9. Illustrations of the periodic interval $(\ell_{m-1},k_{m}-1)$ in Lemma 6.9, Case 2.3 with (a) $m\neq 1$ and (b) $m=1$
Proof of Lemma 6.10..
The assumption $\ell_{0}\neq k_{1}+1$ implies that $\ell_{m-1}\neq k_{m}+1$ for all $m\in\llbracket 1,s\rrbracket$ (otherwise there would be N1N-1 equalities of the form $u_{1,i}=u_{2,i}$). By definition of $k_{m}$, $u_{2,k_{m}}=u_{1,k_{m}}$ and $u_{2,k_{m}+1}>u_{1,k_{m}+1}$, so using Lemma 6.3, we have
$Y_{(k_{m},\ell_{m-1}-1]}=Y_{(k_{m}-1,\ell_{m-1}-1]}-Y_{k_{m}}\geq\theta_{2}-\theta_{1}-Y_{k_{m}}=Y_{(k_{m},k_{m}-1]}.$
Using this, followed by Lemma 6.9 in the second line below, we see that
$\max_{j\in\mathbb{Z}_{N}}Y_{(k_{m},j]}$ $=\max_{j\in\mathbb{Z}_{N}\setminus\{k_{m}-1\}}Y_{(k_{m}-1,j]}-Y_{k_{m}}$
$=Y_{(k_{m}-1,\ell_{m-1}-1]}-Y_{k_{m}}=Y_{(k_{m},\ell_{m-1}-1]}.\qed$
Proof of Lemma 6.11..
Observe that, if $\ell_{m}=k_{m}+1$, then the interval $(k_{m},\ell_{m})$ is empty, so our statement is vacuously true. From now on, we can assume that $\ell_{m}\neq k_{m}+1$.
We now show that $Y_{(q,j]}\leq Y_{(q,q-1]}$ for all $j\in\mathbb{Z}_{N}$. To do so, observe that $u_{2,q}>u_{1,q}$ for $q\in(k_{m},\ell_{m})$, which implies that $\max\limits_{j\in\mathbb{Z}_{N}}Y_{(q-1,j]}<\theta_{2}-\theta_{1}$ by Lemma 6.3. Then, by Lemma 6.2(ii), $Y_{[j+1,q-1]}>0$ for any $j\in\mathbb{Z}_{N}$. Combined with Lemma 6.2(i), for $j\in[q,q-1)$ (equivalently $j\neq q-1$),
$Y_{(q,j]}<Y_{(q,j]}+Y_{[j+1,q-1]}=Y_{(q,q-1]},$
as desired. ∎
Appendix A Inputs from previous work and technical inputs
A.1. Basic algebraic properties
We adapted the definition of the periodic Pitman transform from the work of [CGS24]
$\mathfrak{T}_{i}(\mathbf{\mathbf{X}_{1}},\mathbf{\mathbf{X}_{2}})=X_{1,i}+\log\Biggl{(}\frac{\sum_{j\in\mathbb{Z}_{N}}e^{X_{2,[i,j]}-X_{1,[i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{X_{2,[i+1,j]-X_{1,[i+1,j]}}}}\Biggr{)}$
$\mathfrak{D}_{i}(\mathbf{\mathbf{X}_{1}},\mathbf{\mathbf{X}_{2}})=X_{2,i}+\log\Biggl{(}\frac{\sum_{j\in\mathbb{Z}_{N}}e^{X_{2,(i,j]}-X_{1,(i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{X_{2,(i-1,j]}-X_{1,(i-1,j]}}}\Biggr{)}$
In [CGS24], these maps were denoted as $T^{N,2}$ and $D^{N,2}$, respectively. It is important to note that this is equivalent to our definition of the map up to a shift:
$D(\mathbf{X}_{1},\mathbf{X}_{2})=\mathfrak{D}(\mathbf{X}_{1},\tau_{1}\mathbf{X}_{2}),\quad\text{and}\quad T(\mathbf{X}_{1},\mathbf{X}_{2})=\tau_{-1}\mathfrak{T}(\mathbf{X}_{1},\tau_{1}\mathbf{X}_{2}).$ (A.1)
where $\tau_{k}:\mathbb{R}^{\mathbb{Z}_{N}}\rightarrow\mathbb{R}^{\mathbb{Z}_{N}}$ is the shift map: $(\tau_{k}X)_{i}=X_{i+k}$ for all $i\in\mathbb{Z}_{N}$.
Lemma A.1.
$\frac{1}{\beta}\log\Bigl{(}\sum\limits_{i=0}^{k}\exp(\beta E_{i})\Bigr{)}$ converges to $\max\limits_{i}E_{i}$, uniformly on the values of $E$ (but not k) as $\beta\to\infty$.
Proof.
$\frac{1}{\beta}\log\Bigl{(}\sum\limits_{i=0}^{k}\exp(\beta E_{i})\Bigr{)}-\max\limits_{i}E_{i}$ $=\frac{1}{\beta}\log\Bigl{(}\sum\limits_{i=0}^{k}\exp(\beta E_{i})\Bigr{)}-\frac{1}{\beta}\log\Bigl{(}\exp(\beta\max\limits_{i}E_{i})\Bigr{)}$
$=\frac{1}{\beta}\log\Biggl{(}\frac{\sum\limits_{i=0}^{k}\exp(\beta E_{i})}{\max\limits_{i}\exp(\beta E_{i})}\Biggr{)}$
From here, we bound below and above.
$\frac{1}{\beta}\log\Biggl{(}\frac{\sum\limits_{i=0}^{k}\exp(\beta E_{i})}{\max\limits_{i}\exp(\beta E_{i})}\Biggr{)}$ $\leq\frac{1}{\beta}\log\Biggl{(}\frac{k\max\limits_{i}\exp(\beta E_{i})}{\max\limits_{i}\exp(\beta E_{i})}\Biggr{)}=\frac{1}{\beta}\log(k)$
Similarly,
$\frac{1}{\beta}\log\Biggl{(}\frac{\sum\limits_{i=0}^{k}\exp(\beta E_{i})}{\max\limits_{i}\exp(\beta E_{i})}\Biggr{)}$ $\geq\frac{1}{\beta}\log\Biggl{(}\frac{\max\limits_{i}\exp(\beta E_{i})}{\max\limits_{i}\exp(\beta E_{i})}\Biggr{)}=\frac{1}{\beta}\log(1)=0$
This concludes our proof. ∎
Lemma A.2.
The maps $\mathcal{P}$ and $\widetilde{\mathcal{P}}$ define in (1.2) and (1.11) are involutions on $\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}$.
Proof.
By [CGS24, Corollary 5.9], the map $\mathcal{W}:\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}\to\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}$ defined by
$\mathcal{W}(\mathbf{X}_{1},\mathbf{X}_{2})=\bigl{(}\mathfrak{D}(\mathbf{X}_{1},\mathbf{X}_{2}),\mathfrak{T}(\mathbf{X}_{1},\mathbf{X}_{2})\bigr{)},$
is an involution. Therefore, combined with (A.1), for $(\mathbf{W}_{1},\mathbf{W}_{2})\in\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}$, we have
$\mathcal{P}(\mathcal{P}(\mathbf{W}_{1},\mathbf{W}_{2}))$ $=(T(D(\mathbf{W}_{2},\mathbf{W}_{1}),T(\mathbf{W}_{2},\mathbf{W}_{1})),D(D(\mathbf{W}_{2},\mathbf{W}_{1}),T(\mathbf{W}_{2},\mathbf{W}_{1})))$
$=(T(\mathfrak{D}(\mathbf{W}_{2},\tau_{1}\mathbf{W}_{1}),\tau_{-1}\mathfrak{T}(\mathbf{W}_{2},\tau_{1}\mathbf{W}_{1})),D(\mathfrak{D}(\mathbf{W}_{2},\tau_{1}\mathbf{W}_{1}),\tau_{-1}\mathfrak{T}(\mathbf{W}_{2},\tau_{1}\mathbf{W}_{1})))$
$=(\tau_{-1}\mathfrak{T}(\mathfrak{D}(\mathbf{W}_{2},\tau_{1}\mathbf{W}_{1}),\tau_{1}\tau_{-1}\mathfrak{T}(\mathbf{W}_{2},\tau_{1}\mathbf{W}_{1})),\mathfrak{D}(\mathfrak{D}(\mathbf{W}_{2},\tau_{1}\mathbf{W}_{1}),\tau_{1}\tau_{-1}\mathfrak{T}(\mathbf{W}_{2},\tau_{1}\mathbf{W}_{1})))$
$=(\tau_{-1}\mathfrak{T}(\mathfrak{D}(\mathbf{W}_{2},\tau_{1}\mathbf{W}_{1}),\mathfrak{T}(\mathbf{W}_{2},\tau_{1}\mathbf{W}_{1})),\mathfrak{D}(\mathfrak{D}(\mathbf{W}_{2},\tau_{1}\mathbf{W}_{1}),\mathfrak{T}(\mathbf{W}_{2},\tau_{1}\mathbf{W}_{1})))$
$=(\tau_{-1}\tau_{1}\mathbf{W}_{1},\mathbf{W}_{2})$
$=(\mathbf{W}_{1},\mathbf{W}_{2})$
Thus, $\mathcal{P}$ is an involution. Passing to zero-temperature, we get that
$\widetilde{\mathcal{P}}(\widetilde{\mathcal{P}}(\mathbf{W}_{1},\mathbf{W}_{2}))$ $=\lim_{\beta\to\infty}\frac{1}{\beta}\mathcal{P}(\beta\widetilde{\mathcal{P}}(\mathbf{W}_{1},\mathbf{W}_{2}))$
$=\lim_{\beta\to\infty}\frac{1}{\beta}\mathcal{P}(\mathcal{P}(\beta\mathbf{W}_{1},\beta\mathbf{W}_{2}))$
$=\lim_{\beta\to\infty}\frac{1}{\beta}(\beta\mathbf{W}_{1},\beta\mathbf{W}_{2})$
$=(\mathbf{W}_{1},\mathbf{W}_{2}),$
where the penultimate equality follows from $\mathcal{P}$ being an involution and the rest follow from the uniform convergence of $\mathcal{P}$ via Lemma A.1. ∎
Lemma A.3.
$\widetilde{\mathcal{P}}|_{\mathbb{R}^{\mathbb{Z}_{N}}_{\geq 0}\times\mathbb{R}^{\mathbb{Z}_{N}}_{\geq 0}}$ is an involution on $\mathbb{R}^{\mathbb{Z}_{N}}_{\geq 0}\times\mathbb{R}^{\mathbb{Z}_{N}}_{\geq 0}$.
Proof.
Since $\mathcal{P}$ is an involution on $\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}$, it suffices to show that $(\mathbf{X}_{1},\mathbf{X}_{2})\in\mathbb{R}^{\mathbb{Z}_{N}}_{\geq 0}\times\mathbb{R}^{\mathbb{Z}_{N}}_{\geq 0}$ implies $\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2})\in\mathbb{R}_{\geq 0}^{\mathbb{Z}_{N}}$ and $\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})\in\mathbb{R}^{\mathbb{Z}_{N}}_{\geq 0}$. Fix any $i\in\mathbb{Z}_{N}$, and take $\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{R}^{\mathbb{Z}_{N}}_{\geq 0}$. Set $Y_{i}=X_{2,i+1}-X_{1,i}$ for all $i\in\mathbb{Z}_{N}$. We observe that
$Y_{(i-1,j]}=\begin{cases}Y_{i}+Y_{(i,j]},&\text{ if }j\in\mathbb{Z}_{N}\setminus\{i-1\}\\ 0,&\text{ if }j=i-1\end{cases}$
It follows that:
$\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]}$ $=0\vee\left(Y_{i}+\max_{j\in\mathbb{Z}_{N}\setminus\{i-1\}}Y_{(i,j]}\right)$
$\leq 0\vee\left(Y_{i}+\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}\right)$
$\leq 0\vee\left(X_{1,i}+\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}\right)$
$\leq X_{1,i}+\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}$
Similarly, we have that $Y_{[i,j]}=\begin{cases}-Y_{i-1}+Y_{[i-1,j]},&\text{ if }j\in\mathbb{Z}_{N}\setminus\{i-1\}\\ Y_{[i-1,i-2]},&\text{ if }j=i-1\end{cases}$
Therefore, we know
$\max_{j\in\mathbb{Z}_{N}}Y_{[i,j]}$ $=Y_{[i-1,i-2]}\vee\left(-Y_{i-1}+\max_{j\in\mathbb{Z}_{N}\setminus\{i-1\}}Y_{[i-1,j]}\right)$
$\leq\max_{j\in\mathbb{Z}_{N}}Y_{[i-1,j]}\vee\left(X_{2,i-1}+\max_{j\in\mathbb{Z}_{N}}Y_{[i-1,j]}\right)$
$\leq X_{2,i-1}+\max_{j\in\mathbb{Z}_{N}}Y_{[i-1,j]}$
By the definitions in (1.11), we see that $\widetilde{T}_{i}(\mathbf{X}_{1},\mathbf{X}_{2})\geq 0$ and $\widetilde{D}_{i}(\mathbf{X}_{1},\mathbf{X}_{2})\geq 0$ for all $i\in\mathbb{Z}_{N}$. ∎
Lemma A.4.
Let $\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{R}^{\mathbb{Z}_{N}}$. Then,
  1. (i)
    $\mathfrak{s}(T(\mathbf{X}_{1},\mathbf{X}_{2}))=\mathfrak{s}(\mathbf{X}_{1})$
  2. (ii)
    $\mathfrak{s}(D(\mathbf{X}_{1},\mathbf{X}_{2}))=\mathfrak{s}(\mathbf{X}_{2})$
  3. (iii)
    $\mathfrak{s}(\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2}))=\mathfrak{s}(\mathbf{X}_{1})$
  4. (iv)
    $\mathfrak{s}(\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2}))=\mathfrak{s}(\mathbf{X}_{2})$
Proof.
These all follow immediately from the definitions (1.2) and (1.11) since the sums are telescoping. ∎
Lemma A.5.
[CGS24, Lemma 5.10] Let $\mathbf{X}_{1},\mathbf{X_{2}}\in\mathbb{R}^{\mathbb{Z}_{N}}$. Then, for $i\in\mathbb{Z}_{N}$,
  1. (i)
    $D_{i}(\mathbf{X}_{1},\mathbf{X_{2}})+T_{i}(\mathbf{X}_{1},\mathbf{X_{2}})=X_{1,i}+X_{2,i}$
  2. (ii)
    $e^{-D_{i}(\mathbf{X}_{1},\mathbf{X_{2}})}+e^{-T_{i+1}(\mathbf{X}_{1},\mathbf{X_{2}})}=e^{-X_{1,i}}+e^{-X_{2,i+1}}$
  3. (iii)
    $\widetilde{D}_{i}(\mathbf{X}_{1},\mathbf{X_{2}})+\widetilde{T}_{i}(\mathbf{X}_{1},\mathbf{X_{2}})=X_{1,i}+X_{2,i}$.
Proof.
(i)-(ii) were proved in shifted form in [CGS24]. (iii) is the zero-temperature limit of (i). ∎
Lemma A.6.
[CGS24, Lemma B.3] Let $\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{R}^{\mathbb{Z}_{N}}$.
  1. (i)
    If $(\mathbf{U}_{1},\mathbf{U}_{2})=\mathcal{P}(\mathbf{X}_{1},\mathbf{X}_{2})$, then $\mathcal{P}(\tau_{1}\mathbf{X}_{1},\tau_{1}\mathbf{X}_{2})=(\tau_{1}\mathbf{U}_{1},\tau_{1}\mathbf{U}_{2})$
  2. (ii)
    If $(\mathbf{U}_{1},\mathbf{U}_{2})=\widetilde{\mathcal{P}}(\mathbf{X}_{1},\mathbf{X}_{2})$, then $\widetilde{\mathcal{P}}(\tau_{1}\mathbf{X}_{1},\tau_{1}\mathbf{X}_{2})=(\tau_{1}\mathbf{U}_{1},\tau_{1}\mathbf{U}_{2})$
Lemma A.7.
The determinant of the Jacobian of the transformation
$\mathcal{P}:(\mathbf{X}_{1},\mathbf{X}_{2})\to(T(\mathbf{X}_{2},\mathbf{X}_{1}),D(\mathbf{X}_{2},\mathbf{X}_{1}))$
has absolute value $1$.
Proof.
This is shown in [CGS24, Proposition 5.12]. The version of the periodic Pitman transform used there is a shifted version of the map we use, which amounts to swapping columns and rows of the Jacobian matrix, leaving the absolute value of the determinant unchanged. ∎
Lemma A.8.
For $\mathbf{X}_{1},\mathbf{X}_{2},\mathbf{X}_{3}\in\mathbb{R}^{\mathbb{Z}_{N}}$,
$D(\mathbf{X}_{1},D(\mathbf{X}_{2},\mathbf{X}_{3}))=D\bigl{(}D(\mathbf{X}_{1},\mathbf{X}_{2}),D(T(\mathbf{X}_{1},\mathbf{X}_{2}),\mathbf{X}_{3})$
Proof.
We know from [CGS24, Proposition 5.13] that
$\mathfrak{D}(\mathbf{X}_{1},\mathfrak{D}(\mathbf{X}_{2},\mathbf{X}_{3}))=\mathfrak{D}\bigl{(}\mathfrak{D}(\mathbf{X}_{1},\mathbf{X}_{2}),\mathfrak{D}(\mathfrak{T}(\mathbf{X}_{1},\mathbf{X}_{2}),\mathbf{X}_{3}\bigr{)}.$ (A.2)
By (A.1) and Lemma A.6, we have
$\tau_{1}\mathfrak{D}(\mathbf{X}_{1},\mathbf{X}_{2})=\mathfrak{D}(\tau_{1}\mathbf{X}_{1},\tau_{1}\mathbf{X}_{2}).$ (A.3)
Then, we have the following:
$\quad D\bigl{(}D(\mathbf{X}_{1},\mathbf{X}_{2}),D(T(\mathbf{X}_{1},\mathbf{X}_{2}),\mathbf{X}_{3})\bigr{)}$
[…]
as desired. ∎

### v2

@@
 We prove that, for polymers in a periodic environment, single-path and multi-path partition functions are preserved under the action of this transform on the weights in the polymer model.
-As a corollary, we prove that the discrete periodic Pitman transform satisfies the same braid relations that are satisfied for the full-line Pitman transform, shown by Biane, Bougerol, and O’Connell.
+As a corollary, we prove that the discrete periodic Pitman transform satisfies the same braid relations that are satisfied for the full-line Pitman transform shown by Biane, Bougerol, and O’Connell.
 Combined with a new inhomogeneous Burke property for the periodic Pitman transform, we prove a multi-path invariance result for the periodic inverse-gamma polymer under permutations of the column parameters.
 In the limit to the full-line case, we obtain a multi-path extension of a recent invariance result of Bates, Emrah, Martin, Seppäläinen, and the fifth author, in both positive and zero-temperature.
-Additionally, we give a combinatorial description of the distribution of the $2$-component jointly invariant measures for a discrete-time Markov chain.
 1. Introduction
@@
 The braid relations, along with the fact that each $\mathcal{P}_{k}$ is an involution (proved in [CGS24], see Lemma A.2) shows that this collection of operators generates a copy of the infinite symmetric group.
-Both of these main results pass to zero-temperature limits, stated in Theorems 1.8 and 1.9.
-We then utilize the invariance of partition functions and a new inhomogeneous Burke property (Propositions 5.1, 5.2, and 5.3) to prove a distributional invariance of the inverse-gamma polymer partition function under the permutation of the row and column parameters in both the periodic and full-line cases (Theorems 1.10 and 1.11).
+We then utilize the invariance of partition functions and a new inhomogeneous Burke property (Proposition 4.1) to prove a distributional invariance of the inverse-gamma polymer partition function under the permutation of the row and column parameters in both the periodic and full-line cases (Theorems 1.7 and 1.8).
 In the full-line case, this gives a multi-path and positive temperature extension of a recent result of Bates, Emrah, Martin, Seppäläinen, and the fifth author [BEM+25b].
@@
 See the discussion before and after Theorem 1.8 for more on the connection to other works.
-Beginning with the work of Ferrari and Martin in particle systems [FM07, FM06, FM09], and adapted to last-passage percolation, polymers, and their scaling limits in [FS20, SS23, Bus24, BSS24, BFS25, GRASS25, CGS24], Pitman transforms have been used to describe the distribution of jointly invariant or multi-type invariant measures in these models.
-Indeed, the work [CGS24] introduced the periodic Pitman transform to describe the jointly invariant measures for the KPZ equation with periodic noise.
-A similar object was previously used in [FM07] to find the stationary measure for multi-species TASEP on a ring (see [CGS24, Appendix F] for more of a discussion on this connection).
-In Theorem 1.12, we develop a similar description of the jointly invariant measures for a discrete-time Markov chain in zero-temperature.
-Then, in Theorem 1.13, we give a combinatorial description of this invariant measure.
-We can obtain these joint stationary measures in the full-line settings by taking appropriate scaling limits of the measure we study in Theorem 1.13, and we hope that our description will lead to further understanding in future work.
+Each of our main results has an analogue in zero-temperature.
+These results are stated in Section 1.6.
 1.1. The positive-temperature discrete periodic Pitman transform
 For $N\in\mathbb{N}$, let $\mathbb{R}^{\mathbb{Z}_{N}}$ denote the set of $N$-tuples, together with periodic indexing.
-That is, for $\mathbf{X}\in\mathbb{R}^{\mathbb{Z}_{N}}$, $\mathbf{X}=(X_{1},X_{2},\ldots,X_{N})$, where the indices represent elements of $\mathbb{Z}_{N}$ (see Figure 1).
+That is, for $\mathbf{X}\in\mathbb{R}^{\mathbb{Z}_{N}}$, $\mathbf{X}=(X_{1},X_{2},\ldots,X_{N})$, where the indices represent elements of $\mathbb{Z}_{N}$ with addition modulo $N$ (see Figure 1).
 For a vector $\mathbf{X}\in\mathbb{R}^{\mathbb{Z}_{N}}$ and indices $i,j\in\mathbb{Z}_{N}$, we set
@@
 Figure 1. Representation of indices $[i,j]$, demonstrating periodicity.
-We define the positive-temperature discrete periodic Pitman transform as the map $\mathcal{P}:\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}\to\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}$ defined by
+The positive-temperature discrete periodic Pitman transform is the map $\mathcal{P}:\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}\to\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}$ defined by
 $\mathcal{P}(\mathbf{W}_{1},\mathbf{W}_{2})$ $=(T(\mathbf{W}_{2},\mathbf{W}_{1}),D(\mathbf{W}_{2},\mathbf{W}_{1})),\quad\text{where}$ (1.2)
@@
 For real-valued weights $\mathbf{W}=\bigl{(}\mathbf{W}_{i}\in\mathbb{R}^{\mathbb{Z}}\bigr{)}_{i\in\mathbb{Z}}=\bigl{(}W_{(i,j)}\in\mathbb{R}\mid i,j\in\mathbb{Z}\bigr{)}$ corresponding to the vertices of the directed lattice $\mathbb{Z}^{2}$, we define the polymer partition function with inverse temperature $\beta>0$ between the initial point $(i,j)$ and endpoint $(n,m)$ in $\mathbb{Z}^{2}$ as
-$Z^{\beta,\mathbf{W}}(n,m\mid i,j):=\sum_{\pi\in\Pi_{(i,j),(n,m)}}\prod_{(r,\ell)\in\pi}e^{\beta W_{(r,\ell)}}$ (1.7)
+$Z^{\beta,\mathbf{W}}(n,m\mid i,j):=\sum_{\pi\in\Pi_{(i,j),(n,m)}}\prod_{(r,\ell)\in\pi}e^{\beta W_{(r,\ell)}}.$ (1.7)
 The partition function has a natural multi-path extension. For $(U,V)\in\Psi$, we define the multi-path partition function between $U$ and $V$ to be
-$Z^{\beta,\mathbf{W}}(V\mid U):=\sum_{\pi\in\Pi_{U,V}}\prod_{(r,\ell)\in\pi}e^{\beta W_{(r,\ell)}}$ (1.8)
+$Z^{\beta,\mathbf{W}}(V\mid U):=\sum_{\pi\in\Pi_{U,V}}\prod_{(r,\ell)\in\pi}e^{\beta W_{(r,\ell)}}.$ (1.8)
 We typically take $\beta=1$, in which case we omit the superscript $\beta$ in the notation.
@@
 Let $N\in\mathbb{N}$, and let $\mathbf{W}=(\mathbf{W}_{k}\in\mathbb{R}^{\mathbb{Z}_{N}}\mid k\in\mathbb{Z})=(W_{k,\ell}\mid k\in\mathbb{Z},\ell\in\mathbb{Z}_{N})$ be a collection of real-valued weights.
-Extend these periodically to weights on $\mathbb{Z}\times\mathbb{Z}$ by the condition $W_{k,\ell}=W_{j,\ell}$ whenever $k\equiv j\mod N$.
+Extend these periodically to weights on $\mathbb{Z}\times\mathbb{Z}$ by the condition $W_{k,j}=W_{k,\ell}$ whenever $j\equiv\ell\mod N$.
 If $k_{1},k_{2},\ldots,k_{m}\in\mathbb{Z}$ and the sets of initial points $U=\bigl{(}\bigl{(}a_{i},b_{i})_{i\in\llbracket 1,m\rrbracket}\bigr{)}$ and end points $V=\bigl{(}(c_{i},d_{i})_{i\in\llbracket 1,m\rrbracket}\bigr{)}$ satisfy $(U,V)\in\Psi$ and, for all $i,j\in\llbracket 1,m\rrbracket$, $a_{i}\neq k_{j}+1$ and $c_{i}\neq k_{j}$, then,
@@
 In this example, we take $m=1$ and consider the action of the operator $\mathcal{P}_{3}$, which replaces the weights in columns $3$ and $4$ with $T(\mathbf{W}_{4},\mathbf{W}_{3})$ and $D(\mathbf{W}_{4},\mathbf{W}_{3})$, respectively.
-The partition function for each pair of points and the multi-path partition function for this collection of $3$ pairs are both preserved under this transformation of the weights.
+The partition function for each pair of points and the multi-path partition function for this collection of 33 pairs are both preserved under this transformation of the weights.
 Theorem 1.4 is proved in Section 2.
@@
 With that as our starting point, we iterate the matrix relation in Corollary 2.5 in two different ways to obtain and use the interpretation of the matrix product as the polymer partition function to obtain the other two equalities.
-1.5. Zero-temperature limits
-We also define the zero-temperature discrete periodic Pitman transform as the map $\widetilde{\mathcal{P}}:\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}\to\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}$ defined by
-$\widetilde{\mathcal{P}}(\mathbf{W}_{1},\mathbf{W}_{2})$ $=\bigl{(}\widetilde{T}(\mathbf{W}_{2},\mathbf{W}_{1}),\widetilde{D}(\mathbf{W}_{2},\mathbf{W}_{1})\bigr{)},\quad\text{where}$ (1.11)
-$\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2})$ $=X_{1,i-1}+\max_{j\in\mathbb{Z}_{N}}Y_{[i-1,j]}-\max_{j\in\mathbb{Z}_{N}}Y_{[i,j]},\quad\text{and}$
-$\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})$ $=X_{2,i+1}+\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}-\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]},\quad\text{where}\quad Y_{\ell}=X_{2,\ell+1}-X_{1,\ell}.$
-One immediately sees that we can obtain $\widetilde{\mathcal{P}}$ as a limiting case of the positive-temperature discrete periodic Pitman transform:
-$\widetilde{\mathcal{P}}(\mathbf{W}_{1},\mathbf{W}_{2})=\lim_{\beta\to\infty}\frac{1}{\beta}\mathcal{P}(\beta\mathbf{W}_{1},\beta\mathbf{W}_{2})$
-Furthermore, this convergence is uniform as a function of the vectors $\mathbf{W}_{1},\mathbf{W}_{2}$. See Lemma A.1 for details.
-Definition 1.7.
-For real-valued weights $\mathbf{W}=\bigl{(}W_{(i,j)}\in\mathbb{R}\mid i,j\in\mathbb{Z}\bigr{)}$ corresponding to the vertices of the directed lattice $\mathbb{Z}^{2}$, we define the last-passage percolation (LPP) time between the points $(i,j)$ and $(n,m)$ in $\mathbb{Z}^{2}$ as the maximal sum over weights along an up-right path between $(i,j)$ and $(n,m)$.
-That is,
-$G^{\mathbf{W}}(n,m\mid i,j)\coloneqq\max\limits_{\pi\in\Pi_{(i,j),(n,m)}}\sum_{(r,\ell)\in\pi}W_{(r,\ell)}$
-We similarly define multi-path last-passage times analogously as in (1.8):
-$G^{\mathbf{W}}(V\mid U)=\max_{\pi\in\Pi_{U,V}}\sum_{(r,\ell)\in\pi}W_{(r,\ell)}.$
-With this notation defined, we get the following zero-temperature analogues of Theorems 1.4 and 1.6:
-Theorem 1.8.
-The same invariance in Theorem 1.4 holds, $Z^{\mathbf{W}}$ is replaced by $G^{\mathbf{W}}$, and $T,D$ are replaced by $\widetilde{T},\widetilde{D}$.
-Theorem 1.9.
-For $k\in\mathbb{Z}$, define the map $\widetilde{\mathcal{P}}_{k}:(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}\to(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$ by (1.9), with $T,D$ replaced by $\widetilde{T},\widetilde{D}$.
-Then, the same braid relations in Theorem 1.6 hold.
-The details of the limit transitions to Theorems 1.8 and 1.9 are handled in Section 4.
+Specifically, Lemma 3.1 is the key innovation that allows us to complete the proof.
 1.5. Permutation invariance
@@
 Let $\lambda>0$ and $\mathbf{a}\in\mathbb{R}^{\mathbb{Z}}$ and $\mathbf{b}\in\mathbb{R}^{\mathbb{Z}_{N}}$ be sequences of real numbers such that $a_{i}+b_{j}>0$ for $i\in\mathbb{Z}$ and $j\in\mathbb{Z}_{N}$.
-We take independent weights $\mathbf{W}^{\mathbf{a},\mathbf{b}}=\bigl{(}W_{(i,j)}\mid i\in\mathbb{Z},\ j\in\mathbb{Z}_{N}\bigr{)}$ such that $W_{(i,j)}\sim\text{Log-Inv-Gamma}(a_{i}+b_{j},\lambda)$ for $i\in\mathbb{Z}$ and $j\in\mathbb{Z}_{N}$ (see the notation section below).
+We take independent weights $\mathbf{W}^{\mathbf{a},\mathbf{b}}=\bigl{(}W_{(i,j)}\mid i\in\mathbb{Z},\ j\in\mathbb{Z}_{N}\bigr{)}$ such that $W_{(i,j)}\sim\text{Log-Inv-Gamma}(a_{i}+b_{j},\lambda)$ for $i\in\mathbb{Z}$ and $j\in\mathbb{Z}_{N}$ (see Definition 1.15 below).
 Extend these periodically to weights on $\mathbb{Z}\times\mathbb{Z}$.
@@
 where $\sigma\mathbf{a}=(a_{\sigma(i)})_{i\in\mathbb{Z}}$.
-The same invariance holds for last-passage percolation if the weights satisfy either $W_{(i,j)}\sim\text{Exp}(a_{i}+b_{j})$ with $a_{i}+b_{j}>0$ for all $i,j$, or $W_{(i,j)}\sim\text{Geom}(q_{i}r_{j})$ for all $i,j$, where $q_{i}r_{j}\in(0,1)$ for all $i,j$.
 The condition (1.11) ensures that, between the pairs of points in $(U,V)$, the column rates are either permuted among themselves or are left unchanged. See Figure 4 for an example.
@@
 The following is an extension of the recent result [BEM+25a, Theorem 3.1] to multi-paths and to the positive-temperature case.
+A sketch of the proof for the extension to multi-paths in zero temperature was previously discussed in [BEM+25a, Remark 4.3].
 The technique follows the same outline as developed in [BEM+25a], as we use a coupling of the weights given by applying a Pitman transform to neighboring rows and columns.
 However, the coupling we use in this proof is different because we treat the weights as periodic outside a large box and apply the periodic Pitman transform.
+The algebra involved in this positive-temperature case is also different.
 Theorem 1.8.
@@
 where $\sigma\mathbf{a}=(a_{\sigma(i)})_{i\in\mathbb{Z}}$ and $\tau\mathbf{b}=(b_{\tau(j)})_{j\in\mathbb{Z}}$.
-The same invariance holds for last-passage percolation if the weights satisfy either $W_{(i,j)}\sim\text{Exp}(a_{i}+b_{j})$ for all $i,j$ or $W_{(i,j)}\sim\text{Geom}(q_{i}r_{j})$ for all $i,j$, where $q_{i}r_{j}\in(0,1)$ for all $i,j$.
-Theorems 1.10 and 1.11 are proved in Section 5.1.
+Theorems 1.7 and 1.8 are proved in Section 4.
 When the initial point is fixed, invariances of this flavor can be deduced from the symmetry of Schur functions and Whittaker functions (see, for example, [BP08, COSZ14, OSZ14, BC14]).
@@
 Connections between these results are discussed further in [BEM+25a, Section 3].
-1.7. Distribution of the jointly invariant measures for a discrete-time Markov chain
-The original motivation for introducing the map $D$ in [CGS24] was to understand the jointly invariant measures for the KPZ equation with periodic noise.
-This was done by first finding jointly invariant measures for the periodic O’Connell-Yor polymer.
-As a by-product of the analysis, it was shown that the same measures are jointly invariant for a discrete-time Markov chain (see [CGS24, Section 4]).
-In this section, we study the zero-temperature analogue of that Markov chain.
-The state space is $\mathbb{R}_{\geq 0}^{\mathbb{Z}_{N}}\times\mathbb{R}_{\geq 0}^{\mathbb{Z}_{N}}$, which is preserved by the chain by Lemma A.3.
-If the state of the Markov chain at time $m$ is $(\mathbf{U}_{1}^{(m)},\mathbf{U}_{2}^{(m)})$, and $\mathbf{W}\in\mathbb{R}^{\mathbb{Z}_{N}}_{\geq 0}$ is independent of the chain up to time $m$, then the time $m+1$ state is defined as
-$\bigl{(}\mathbf{U}^{(m+1)}_{1},\mathbf{U}^{(m+1)}_{2}\bigr{)}=\bigl{(}\widetilde{D}(\mathbf{W},\mathbf{U}_{1}^{(m)}),\widetilde{D}(\mathbf{W},\mathbf{U}_{2}^{(m)})\bigr{)}.$ (1.14)
-Marginally, $\bigl{(}\mathbf{U}_{1}^{(m)}\bigr{)}_{m\geq 0}$ and $\bigl{(}\mathbf{U}_{2}^{(m)}\bigr{)}_{m\geq 0}$ are each Markov chains with the same dynamics.
-We call the invariant measures for this coupled Markov chain jointly invariant measures because their joint distribution is preserved under the Markov chain where each component evolves using the same driving weights.
-For $\mathbf{X}\in\mathbb{R}^{\mathbb{Z}_{N}}$, define
-$\mathfrak{s}(\mathbf{X})=\sum_{i\in\mathbb{Z}_{N}}X_{i}.$ (1.15)
-For $\theta\in\mathbb{R}$, define
-$\mathbb{R}^{\mathbb{Z}_{N}}_{\theta}$ $=\{\mathbf{X}\in\mathbb{R}^{\mathbb{Z}_{N}}\mid\mathfrak{s}(\mathbf{X})=\theta\},\quad\text{and}$ (1.16)
-$\mathbb{R}^{\mathbb{Z}_{N}}_{\theta,\geq 0}$ $=\{\mathbf{X}\in\mathbb{R}^{\mathbb{Z}_{N}}_{\theta}\mid X_{i}\geq 0\text{ for }i\in\mathbb{Z}_{N}\}$
-We also define $\mathbb{Z}^{\mathbb{Z}_{N}}_{\theta}=\mathbb{R}_{\theta}^{\mathbb{Z}_{N}}\cap\mathbb{Z}^{\mathbb{Z}_{N}}$ and $\mathbb{Z}^{\mathbb{Z}_{N}}_{\theta,\geq 0}:=\mathbb{R}^{\mathbb{Z}_{N}}_{\theta,\geq 0}\cap\mathbb{Z}^{\mathbb{Z}_{N}}$.
-We now state the following theorem.
+1.6. Zero-temperature limits
+We also define the zero-temperature discrete periodic Pitman transform as the map $\widetilde{\mathcal{P}}:\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}\to\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}$ defined by
+$\widetilde{\mathcal{P}}(\mathbf{W}_{1},\mathbf{W}_{2})$ $=\bigl{(}\widetilde{T}(\mathbf{W}_{2},\mathbf{W}_{1}),\widetilde{D}(\mathbf{W}_{2},\mathbf{W}_{1})\bigr{)},\quad\text{where}$ (1.13)
+$\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2})$ $=X_{1,i-1}+\max_{j\in\mathbb{Z}_{N}}Y_{[i-1,j]}-\max_{j\in\mathbb{Z}_{N}}Y_{[i,j]},\quad\text{and}$
+$\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})$ $=X_{2,i+1}+\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}-\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]},\quad\text{where}\quad Y_{\ell}=X_{2,\ell+1}-X_{1,\ell}.$
+The mapping $\widetilde{D}$ is similar to, but not equivalent to a mapping used to describe the invariant measures for multi-type TASEP by Ferrari and Martin [FM07].
+See [CGS24, Appendix F] for more on this connection.
+One immediately sees that we can obtain $\widetilde{\mathcal{P}}$ as a limiting case of the positive-temperature discrete periodic Pitman transform:
+$\widetilde{\mathcal{P}}(\mathbf{W}_{1},\mathbf{W}_{2})=\lim_{\beta\to\infty}\frac{1}{\beta}\mathcal{P}(\beta\mathbf{W}_{1},\beta\mathbf{W}_{2})$
+Furthermore, this convergence is uniform as a function of the vectors $\mathbf{W}_{1},\mathbf{W}_{2}$. See Lemma A.1 for details.
+Definition 1.9.
+For real-valued weights $\mathbf{W}=\bigl{(}W_{(i,j)}\in\mathbb{R}\mid i,j\in\mathbb{Z}\bigr{)}$ corresponding to the vertices of the directed lattice $\mathbb{Z}^{2}$, we define the last-passage percolation (LPP) time for both a single pair of points and multi-paths analogously as in (1.7)-(1.8):
+$G^{\mathbf{W}}(V\mid U)=\max_{\pi\in\Pi_{U,V}}\sum_{(r,\ell)\in\pi}W_{(r,\ell)}.$
+For $k\in\mathbb{Z}$, define the map $\widetilde{\mathcal{P}}_{k}:(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}\to(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$ by (1.9), with $T,D$ replaced by $\widetilde{T},\widetilde{D}$.
+With this notation defined, we get the following zero-temperature analogues of our main theorems.
 Theorem 1.10.
-Let $\theta_{1},\theta_{2}\geq 0$ and $q\in(0,1)$.
-At each step $m$, let the driving weights $\mathbf{W}$ in the evolution of the Markov chain (1.14) be an i.i.d.
-sequence of $\text{Geom}(q)$ random variables.
-Let $\mathbf{X}_{1}$ be uniformly distributed on the set $\mathbb{Z}^{\mathbb{Z}_{N}}_{\theta_{1},\geq 0}$, and let $\mathbf{X}_{2}$ be uniformly distributed on the set $\mathbb{Z}^{\mathbb{Z}_{N}}_{\theta_{2},\geq 0}$, then the law of $\bigl{(}\mathbf{X}_{1},\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})\bigr{)}$ is a jointly invariant measure for the Markov chain.
-Furthermore, we have the distributional equality
-$\bigl{(}\mathbf{X}_{1},\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})\bigr{)}\overset{d}{=}\bigl{(}\widetilde{D}(\mathbf{X}_{2},\mathbf{X}_{1}),\mathbf{X}_{2}\bigr{)}$ (1.17)
-The same result holds when $\mathbf{W}$ is an i.i.d.
-sequence of $\text{Exp}(\alpha)$ random variables for any $\alpha>0$, if we instead take $\mathbf{X}_{1}$ to be uniformly distributed on $\mathbb{R}_{\theta_{1},\geq 0}^{\mathbb{Z}_{N}}$ and $\mathbf{X}_{2}$ to be uniformly distributed on $\mathbb{R}_{\theta_{2},\geq 0}^{\mathbb{Z}_{N}}$.
-Theorems 1.12, along with Theorem 1.13 below, is proved in Section 6.
-In fact, we prove that the same invariance holds when $\mathbf{X}_{1},\mathbf{X}_{2}$ are sequences of i.i.d.
-geometric random variables.
-The uniform distribution is obtained by conditioning on $\mathfrak{s}(\mathbf{X}_{1})$ and $\mathfrak{s}(\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2}))$.
-These vector sums are preserved by the map $\widetilde{D}$ and hence also the Markov chain (Lemma A.4), so the extremal invariant measures come when we consider $\mathbf{X}_{1},\mathbf{X}_{2}$ to be uniform over all nonnegative vectors with a given sum.
-As in [CGS24], we can also consider the joint Markov chain with more than two components.
-In this case, the invariant measures will be built by iterating the map $\widetilde{D}$.
-Specifically, the invariant initial conditions will be of the form $(\mathbf{X}_{1},\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2}),\widetilde{D}(\mathbf{X}_{1},\widetilde{D}(\mathbf{X}_{2},\mathbf{X}_{3})),\ldots$ For brevity, we include only the case of two components to motivate the next theorem.
-In this next result, we give an explicit combinatorial description of these jointly invariant measures from Theorem 1.12.
-We leave it as an open problem to develop a similar combinatorial description for the jointly invariant measures with more than $2$ components.
-In analogy with [CGS24], when $\mathbf{X}_{1}$ is uniform on $\mathbb{R}_{\theta_{1},\geq 0}^{\mathbb{Z}_{N}}$ and $\mathbf{X}_{2}$ is uniform on $\mathbb{R}_{\theta_{2},\geq 0}^{\mathbb{Z}_{N}}$, the law of $(\mathbf{X}_{1},\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2}))$ should be a jointly invariant measure for Brownian last-passage percolation in a periodic environment of Brownian motions.
-It can also be shown (see [CGS24, Appendix F]) that as $N\to\infty$, the invariant measure from Theorem 1.12 converges to the jointly invariant measure for i.i.d.
-exponential last-passage percolation studied in [FS20], which in turn, converges under diffusive scaling to the $2$-component marginal of the stationary horizon [Bus24]–a coupling of Brownian motions with drift that describes the jointly invariant measures for the KPZ fixed point [BSS24].
-Below, we take $\theta_{2}>\theta_{1}$, since we have the distributional equality $\bigl{(}\mathbf{X}_{1},\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})\bigr{)}\overset{d}{=}\bigl{(}\widetilde{D}(\mathbf{X}_{2},\mathbf{X}_{1}),\mathbf{X}_{2}\bigr{)}$ in Theorem 1.12.
-When $\theta_{1}=\theta_{2}$, we have $\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})=\mathbf{X}_{1}$ via Corollary 6.4 below, so the joint distribution is trivial in this case.
-Below, for $\theta_{1}\leq\theta_{2}$, we define the space
-$\mathcal{Z}_{\theta_{1},\theta_{2}}:=\{(\mathbf{u}_{1},\mathbf{u}_{2})\in\mathbb{Z}^{\mathbb{Z}_{N}}_{\theta_{1},\geq 0}\times\mathbb{Z}^{\mathbb{Z}_{N}}_{\theta_{2},\geq 0}\mid\mathbf{u}_{1}\leq\mathbf{u}_{2}\},$ (1.18)
-where $\mathbf{u}_{1}\leq\mathbf{u}_{2}$ means $u_{1,i}\leq u_{2,i}$ for all $i\in\mathbb{Z}_{N}$.
+The same invariance in Theorem 1.4 holds when $Z^{\mathbf{W}}$ is replaced by $G^{\mathbf{W}}$, and the maps $\mathcal{P}_{k_{j}}$ are replaced by $\widetilde{\mathcal{P}}_{k_{j}}$.
 Theorem 1.11.
-[…]
-$\mathbb{P}(\mathbf{X}_{1}=\mathbf{u}_{1},\mathbf{\widetilde{D}}=\mathbf{u}_{2})=\frac{\alpha(\mathbf{u}_{1},\mathbf{u}_{2})}{\binom{\theta_{1}+N-1}{N-1}\binom{\theta_{2}+N-1}{N-1}}$
-where $\alpha(\mathbf{u}_{1},\mathbf{u}_{2})$ is the following integer, split into two cases:
-  1. (i)
-    If $u_{1,i}<u_{2,i}$ for all $i\in\mathbb{Z}_{N}$, then $\alpha(\mathbf{u}_{1},\mathbf{u}_{2})=1$.
-  2. (ii)
-[…]
-    $\alpha(\mathbf{u}_{1},\mathbf{u}_{2})=\prod_{n=1}^{s}\sum_{a_{1}=-u_{1,\ell_{m}}}^{0}\sum_{a_{2}=-u_{1,\ell_{m}+1}}^{-a_{1}}\sum_{a_{3}=-u_{1,\ell_{m}+2}}^{-(a_{1}+a_{2})}\cdots\sum_{a_{k_{m+1}-\ell_{m}+N\mathbf{1}_{m=s}}=-u_{1,k_{m+1}-1+N1_{m=s}}}^{-(a_{1}+a_{2}+\cdots+a_{k_{m+1}-\ell_{m}-1+N1_{m=s}})}1.$
-    Alternatively, $\alpha(\mathbf{u}_{1},\mathbf{u}_{2})$ is the number of $s$-tuples of vectors $(\mathbf{a}_{1},\mathbf{a}_{2},\ldots,\mathbf{a}_{s})$, where, for $m\in\llbracket 1,s\rrbracket$, the vector $\mathbf{a}_{m}$ lies in the set $\mathbb{Z}^{k_{m+1}-\ell_{m}+N1_{m=s}}$ and satisfies the following conditions:
-    $a_{m,1}\leq 0$ (1.19)
-    $a_{m,1}+a_{m,2}\leq 0$
-    $\vdots$
-    $a_{m,1}+a_{m,2}+\cdots+a_{m,k_{m+1}-\ell_{m}+N1_{m=s}}\leq 0,\text{ and}$
-    $a_{m,p}\geq-u_{1,\ell_{m}+p-1}\text{ for }1\leq p\leq k_{m+1}-\ell_{m}+N1_{m=s}.$
-[…]
-Figure 5.
-Comparison of $\mathbf{u}_{1}$ to $\mathbf{u}_{2}$ with $s$ alternating equality/inequality segments.
-We place an equal sign at index $i$ if $u_{1,i}=u_{2,i}$ and an inequality sign at index $j$ if $u_{1,j}<u_{2,j}$.
-Example 1.14.
-We demonstrate a sample application of Theorem 1.13 in the case of $N=2$ and $\theta_{1},\theta_{2}\in\mathbb{Z}_{\geq 0}$ where $\theta_{1}<\theta_{2}$.
-Take some $x_{1},x_{2},y_{1},y_{2}\in\mathbb{Z}_{\geq 0}$ satisfying $x_{1}\leq y_{1},x_{2}\leq y_{2}$ and $x_{1}+x_{2}=\theta_{1},y_{1}+y_{2}=\theta_{2}$.
-If $x_{1}=y_{1}$, then $x_{2}<y_{2}$.
-In the notation of Theorem 1.13, $s=1,k_{1}=1$, and $\ell_{1}=2$, and the only relevant condition is $x_{2}=-u_{1,2}\leq a_{1,1}\leq 0,$ giving $x_{2}+1$ possible values for $a_{1,1}$.
-We can symmetrically describe the possible values when $x_{1}<y_{1},x_{2}=y_{2}$.
-Then, we observe that
-$\mathbb{P}(\mathbf{X}_{1}=(x_{1},x_{2}),\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})=(y_{1},y_{2}))=\begin{cases}\frac{x_{1}+1}{(\theta_{1}+1)(\theta_{2}+1)}&x_{1}<y_{1}\text{ and }x_{2}=y_{2}\\ \frac{x_{2}+1}{(\theta_{1}+1)(\theta_{2}+1)}&x_{1}=y_{1}\text{ and }x_{2}<y_{2}\\ \frac{1}{(\theta_{1}+1)(\theta_{2}+1)}&x_{1}<y_{1}\text{ and }x_{2}<y_{2}\end{cases}$
-[…]
-Example 1.15.
-We similarly demonstrate a computation in the case of $N=3$ and $\theta_{1}<\theta_{2}\in\mathbb{Z}_{\geq 0}$.
-Take $x_{1},x_{2},x_{3},y_{1},y_{2},y_{3}\in\mathbb{Z}_{\geq 0}$ satisfying $x_{1}\leq y_{1}$,  $x_{2}\leq y_{2},\ x_{3}\leq y_{3}$ and $x_{1}+x_{2}+x_{3}=\theta_{1},\ y_{1}+y_{2}+y_{3}=\theta_{2}$.
-Write $\mathbf{u}_{1}=(x_{1},x_{2},x_{3}),\mathbf{u}_{2}=(y_{1},y_{2},y_{3})$.
-Then, we consider the cases in which the only equality is $x_{1}=y_{1}$ and when the only equalities are $x_{1}=y_{1},\ x_{2}=y_{2}$.
-We can achieve all other sets of equalities and inequalities through cyclic permutations of $\mathbf{u}_{1},\mathbf{u}_{2}$ (Lemma A.6).
-When we have that $x_{1}=y_{1},\ x_{2}<y_{2},\ x_{3}<y_{3}$, then we have that $k_{1}=1,\ell_{1}=3$, so the conditions (1.19) are exactly $x_{3}=-u_{1,3}\leq a_{1,1}\leq 0$, giving $x_{3}+1$ distinct possible values.
-If $x_{1}=y_{1},\ x_{2}=y_{2},\ x_{3}<y_{3}$, then we have that $k_{1}=2,\ell_{1}=3$, so the conditions (1.19) are $x_{3}=-u_{1,3}\leq a_{1,1}\leq 0,\ a_{1,1}+a_{1,2}\leq 0,\ -x_{1}=-u_{1,1}\leq a_{1,2}$.
-This gives a count of
-$\alpha(\mathbf{u}_{1},\mathbf{u}_{2})$ $=\sum_{a_{1}=-x_{3}}^{0}\sum_{a_{2}=-x_{1}}^{-a_{1}}1=\sum_{a_{1}=-x_{3}}^{0}(x_{1}-a_{1}+1)=(x_{3}+1)(x_{1}+1)+\frac{x_{3}(x_{3}+1)}{2}.$
-Noting that there are $\binom{\theta_{1}+2}{2}\binom{\theta_{2}+2}{2}$ elements in $\mathbb{Z}_{\theta_{1},+}^{\mathbb{Z}_{N}}\times\mathbb{Z}_{\theta_{2},+}^{\mathbb{Z}_{N}}$, we have the computation
-($\binom{\theta_{1}+2}{2}\binom{\theta_{2}+2}{2}\cdot\mathbb{P}$( $\mathbf{X}_{1}=(x_{1},x_{2},x_{3}),\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})=(y_{1},y_{2},y_{3}))$
-$=\begin{cases}(x_{3}+1)(x_{1}+1)+\frac{x_{3}(x_{3}+1)}{2}&x_{1}=y_{1},x_{2}=y_{2},\text{ and }x_{3}<y_{3}\\ \hfil x_{3}+1&x_{1}=y_{1},x_{2}<y_{2},\text{ and }x_{3}<y_{3}\\ \hfil 1&x_{1}<y_{1},x_{2}<y_{2},\text{ and }x_{3}<y_{3}\end{cases}$
+The same braid relations in Theorem 1.6 hold for the maps $\widetilde{\mathcal{P}}_{k}$.
+Theorem 1.12.
+The analogues of Theorems 1.7 and 1.8 hold, when $Z^{\mathbf{W}}$ is replaced by $G^{\mathbf{W}}$ and, instead of log-inverse-gamma weights, the weights satisfy either $W_{(i,j)}\sim\text{Geom}(q_{i}r_{j})$ for all $i,j$, where $q_{i}r_{j}\in(0,1)$, or $W_{(i,j)}\sim\text{Exp}(a_{i}+b_{j})$ for all $i,j$, where $a_{i}+b_{j}>0$ (see Definitions 1.13-1.14 below).
+The details of the proofs of these theorems are given in Appendix B.
 1.7. Organization of the paper
@@
 The braid relations in Theorem 1.6 are proved in Section 3.
-The details of the zero-temperature limit transitions are handled in Section 4.
-Section 5 proves inhomogeneous Burke properties for the positive and zero-temperature periodic Pitman transforms and then uses those to prove the permutation invariance in Theorems 1.10 and 1.11.
-Theorems 1.12 and 1.13 are proved in Section 6.
-The Appendix contains some prior results from the work [CGS24] and some auxiliary technical results.
+Section 4 proves an inhomogeneous Burke property for the periodic Pitman transform and uses it to prove Theorems 1.7 and 1.8.
+Appendix A contains some prior results from the work [CGS24] and some auxiliary technical results.
+The necessary details for the zero-temperature analogues in Theorems 1.10-1.12 are proved in Appendix B.
 1.8. Notation and conventions
@@
 2.1. Matrix encoding
-In order to prove our polymer partition function invariance, we follow the matrix encoding method introduced at Noumi and Yamada in [NY04]. We will work with infinite matrices.
+In order to prove our polymer partition function invariance, we follow the matrix encoding method introduced at Noumi and Yamada in [NY04].
+The key difference is that we will work with infinite matrices.
 Definition 2.1.
@@
 Proof.
-First, we justify that this matrix product even makes sense, in that it is finite valued.
+First, we justify that this matrix product has finite entries.
 Because these $H$ matrices (as well as the $E$ matrices) are upper triangular, each individual index in the product consists of a finite sum of products of elements in the respective $H$ matrices.
@@
 $=(\Delta E(-T(\mathbf{W}_{2},\mathbf{W}_{1}))^{-1}\Delta)(\Delta E(-D(\mathbf{W}_{2},\mathbf{W}_{1}))^{-1}\Delta)$
-$=H(T(\mathbf{W}_{2},\mathbf{W}_{1}))H(D(\mathbf{W}_{2},\mathbf{W}_{1}))\qed$
+$=H(T(\mathbf{W}_{2},\mathbf{W}_{1}))H(D(\mathbf{W}_{2},\mathbf{W}_{1})).\qed$
 2.2. Relation to the periodic polymer
@@
 $Z^{\mathbf{W}}(\ell+1,j\mid k,i)$ $=\sum\limits_{r=i}^{j}\Biggl{(}Z^{\mathbf{W}}(\ell,r\mid k,i)\exp\bigl{(}\sum_{s=r}^{j}W_{(\ell+1,s)}\bigr{)}\Biggr{)}$
-$=\sum\limits_{r=i}^{j}\Biggl{(}Z^{\mathbf{W}}(\ell,r\mid k,i)\Bigl{(}H(\mathbf{W}_{\ell+1})\Bigr{)}_{(r,j)}\Biggr{)}$
+$=\sum\limits_{r=i}^{j}\Biggl{(}Z^{\mathbf{W}}(\ell,r\mid k,i)\Bigl{(}H(\mathbf{W}_{\ell+1})\Bigr{)}_{(r,j)}\Biggr{)}.$
 By our inductive hypothesis, we know that this equals
-$\sum\limits_{r=i}^{j}\Biggl{(}\Bigl{(}H(\mathbf{W}_{k})H(\mathbf{W}_{k+1})\ldots H(\mathbf{W}_{\ell})\Bigr{)}_{(i,r)}\Bigl{(}H(\mathbf{W}_{\ell+1})\Bigr{)}_{(r,j)}\Biggr{)}$
+$\sum\limits_{r=i}^{j}\Biggl{(}\Bigl{(}H(\mathbf{W}_{k})H(\mathbf{W}_{k+1})\ldots H(\mathbf{W}_{\ell})\Bigr{)}_{(i,r)}\Bigl{(}H(\mathbf{W}_{\ell+1})\Bigr{)}_{(r,j)}\Biggr{)}.$
 Remembering now that $\bigl{(}H(\mathbf{W}_{\ell+1})\bigr{)}_{(r,j)}=0$ for $r>j$, and $\bigl{(}H(\mathbf{W}_{k})H(\mathbf{W}_{k+1})\ldots H(\mathbf{W}_{\ell})\bigr{)}_{(i,r)}=0$ for $r<i$, we see that
-$Z^{\mathbf{W}}(\ell+1,j\mid k,i)$ $=\sum\limits_{r=-\infty}^{\infty}\Biggl{(}\Bigl{(}H(\mathbf{W}_{k})H(\mathbf{W}_{k+1})\ldots H(\mathbf{W}_{\ell})\Bigr{)}_{(i,r)}\Bigl{(}H(\mathbf{W}_{\ell+1})\Bigr{)}_{(r,j)}\Biggr{)}$
+$\quad\;Z^{\mathbf{W}}(\ell+1,j\mid k,i)$
+$=\sum\limits_{r=-\infty}^{\infty}\Biggl{(}\Bigl{(}H(\mathbf{W}_{k})H(\mathbf{W}_{k+1})\ldots H(\mathbf{W}_{\ell})\Bigr{)}_{(i,r)}\Bigl{(}H(\mathbf{W}_{\ell+1})\Bigr{)}_{(r,j)}\Biggr{)}$
 $=\Bigl{(}H(\mathbf{W}_{k})H(\mathbf{W}_{k+1})\ldots H(\mathbf{W}_{\ell})H(\mathbf{W}_{\ell+1})\Bigr{)}_{(i,j)}.\qed$
@@
 Retaining the periodic weight vectors $\mathbf{W}=(\mathbf{W}_{r}\in\mathbb{R}^{\mathbb{Z}_{N}}:r\in\mathbb{Z})$ as before, for $k\in\mathbb{Z}$, recall the operator $\mathcal{P}_{k}:(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}\to(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$ where
-$\mathcal{P}_{k}\mathbf{W}=(\ldots,\mathbf{W}_{k-1},T(\mathbf{W}_{k+1},\mathbf{W}_{k}),D(\mathbf{W}_{k+1},\mathbf{W}_{k}),\mathbf{W}_{k+2,\ldots})$
+$\mathcal{P}_{k}\mathbf{W}=(\ldots,\mathbf{W}_{k-1},T(\mathbf{W}_{k+1},\mathbf{W}_{k}),D(\mathbf{W}_{k+1},\mathbf{W}_{k}),\mathbf{W}_{k+2,\ldots}).$
 Proposition 2.7.
@@
 Then, for all $b,d\in\mathbb{Z}$,
-$Z^{\mathcal{P}_{k_{m}}\cdots\mathcal{P}_{k_{1}}\mathbf{W}}(c,d\mid a,b)=Z^{\mathbf{W}}(c,d\mid a,b)$
+$Z^{\mathcal{P}_{k_{m}}\cdots\mathcal{P}_{k_{1}}\mathbf{W}}(c,d\mid a,b)=Z^{\mathbf{W}}(c,d\mid a,b).$
 Proof.
@@
 $Z^{\mathbf{W}}(c,d\mid a,b)=\bigl{(}H(\mathbf{W}_{a})H(\mathbf{W}_{a+1})\cdots H(\mathbf{W}_{c})\bigr{)}_{(b,d)},$
-which we claim is preserved by $\mathcal{P}_{k_{m}}\ldots\mathcal{P}_{k_{1}}$ for any $b,d$.
+which we claim is preserved by the action $\mathcal{P}_{k_{m}}\ldots\mathcal{P}_{k_{1}}$ for any $b,d\in\mathbb{Z}$.
 In our base case $n=1$, we split into three sub-cases: $k_{1}<a-1$, $k_{1}>c$, and $a\leq k_{1}<c$.
@@
 […]
-$=Z^{\mathcal{P}_{k_{1}}\mathbf{W}}(c,d\mid a,b)$
+$=Z^{\mathcal{P}_{k_{1}}\mathbf{W}}(c,d\mid a,b).$
 Now, we assume by way of induction that the invariance holds for m1m-1 transforms.
@@
 $=Z^{\mathcal{P}_{k_{m}}\mathbf{W}^{\prime}}(c,d\mid a,b)$
-$=Z^{\mathcal{P}_{k_{m}}\mathcal{P}_{k_{m-1}}\ldots\mathcal{P}_{k_{1}}\mathbf{W}}(c,d\mid a,b)$
+$=Z^{\mathcal{P}_{k_{m}}\mathcal{P}_{k_{m-1}}\ldots\mathcal{P}_{k_{1}}\mathbf{W}}(c,d\mid a,b).$
 This concludes our proof, noting that the second equality follows from the case $m=1$. ∎
@@
 Write the path matrix $M=\bigl{(}Z^{\mathbf{W}}(c_{j},d_{j}\mid a_{i},b_{i})\bigr{)}_{(i,j)\in\llbracket 1,m\rrbracket}$. By LGV, we have
-$\det M=\sum\limits_{P=(P_{1}\ldots P_{m})\in\Pi_{U,V_{\sigma}}\ \forall\sigma\in S_{m}}\operatorname{sgn}(\sigma)\prod\limits_{(i,j)\in P}e^{W_{(i,j)}}$
+$\det M=\sum\limits_{P=(P_{1}\ldots P_{m})\in\Pi_{U,V_{\sigma}}\ \forall\sigma\in S_{m}}\operatorname{sgn}(\sigma)\prod\limits_{(i,j)\in P}e^{W_{(i,j)}}.$
 By definition of $\Psi$, for $\sigma\neq\text{Id}$, $\Pi_{U,V_{\sigma}}=\varnothing$. Because $\operatorname{sgn}(\text{Id})=1$, we have that
-$\det M=\sum\limits_{P=(P_{1}\ldots P_{m})\in\Pi_{U,V}}\prod\limits_{(i,j)\in P}e^{W_{(i,j)}}=Z^{\mathbf{W}}(V\mid U)$
+$\det M=\sum\limits_{P=(P_{1}\ldots P_{m})\in\Pi_{U,V}}\prod\limits_{(i,j)\in P}e^{W_{(i,j)}}=Z^{\mathbf{W}}(V\mid U).$
 By Proposition 2.7, each $M_{i,j}=Z^{\mathbf{W}}(c_{j},d_{j}\mid a_{i},b_{i})$ remains invariant under $\mathcal{P}_{k_{m}}\cdots\mathcal{P}_{k_{1}}$, so $M$ is unchanged, and so $\det M$ is invariant.
@@
 […]
-$\qquad H\Bigl{(}D\bigl{(}D(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),D\bigl{(}T(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),\mathbf{W}_{k}\bigr{)}\bigr{)}\Bigr{)}$ (3.4)
+$\qquad H\Bigl{(}D\bigl{(}D(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),D\bigl{(}T(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),\mathbf{W}_{k}\bigr{)}\bigr{)}\Bigr{)}.$ (3.4)
 On the other hand, we can also apply the partition function invariance under the operator $\mathcal{P}_{i}\mathcal{P}_{i+1}\mathcal{P}_{i}$:
@@
 […]
-$\qquad H\Bigl{(}D\bigl{(}T\bigl{(}\mathbf{W}_{k+2},D(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)},T(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)}\Bigr{)}H\Bigl{(}D(\mathbf{W}_{k+2},D(\mathbf{W}_{k+1},\mathbf{W}_{k}))\Bigr{)}$ (3.5)
+$\qquad H\Bigl{(}D\bigl{(}T\bigl{(}\mathbf{W}_{k+2},D(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)},T(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)}\Bigr{)}H\Bigl{(}D(\mathbf{W}_{k+2},D(\mathbf{W}_{k+1},\mathbf{W}_{k}))\Bigr{)}.$ (3.5)
 From Lemma A.8, we know that the last terms in (3.4) and (3.5) are equal. Because $H$ is invertible (see Lemma 2.2), we can equate both equations and cancel out the last term, obtaining that:
 $H\Bigl{(}T\bigl{(}T(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),\mathbf{W}_{k}\bigr{)}\Bigr{)}H\Bigl{(}T\bigl{(}D(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),D\bigl{(}T(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),\mathbf{W}_{k}\bigr{)}\Bigr{)}$
-$=H\Bigl{(}T\bigl{(}T\bigl{(}\mathbf{W}_{k+2},D(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)},T(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)}\Bigr{)}H\Bigl{(}D\bigl{(}T\bigl{(}\mathbf{W}_{k+2},D(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)},T(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)}\Bigr{)}$
+$=H\Bigl{(}T\bigl{(}T\bigl{(}\mathbf{W}_{k+2},D(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)},T(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)}\Bigr{)}H\Bigl{(}D\bigl{(}T\bigl{(}\mathbf{W}_{k+2},D(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)},T(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)}\Bigr{)}.$
 By repeated application of Lemma A.4, we have that
@@
 $=e^{Y_{2,N}}\left(Z^{\mathbf{Y}}(1,N\mid 1,1)+Z^{\mathbf{Y}}(2,N-1\mid 1,1)\right)$
-$=e^{Y_{2,N}}\Biggl{(}\prod_{i=1}^{N}e^{Y_{1,i}}+\left(H(\mathbf{Y}_{1})H(\mathbf{Y}_{2})\right)_{1,N-1}\Biggr{)}$ (3.6)
+$=e^{Y_{2,N}}\Biggl{(}\prod_{i=1}^{N}e^{Y_{1,i}}+\left(H(\mathbf{Y}_{1})H(\mathbf{Y}_{2})\right)_{1,N-1}\Biggr{)}.$ (3.6)
 Similarly, we conclude that:
@@
 and
-$(H(\mathbf{Y}^{\prime}_{1})H(\mathbf{Y}_{2}^{\prime}))_{(N,N)}=Z^{\mathbf{Y}^{\prime}}(2,N\mid 1,N)=e^{Y^{\prime}_{1,N}+Y^{\prime}_{2,N}}$ (3.9)
+$(H(\mathbf{Y}^{\prime}_{1})H(\mathbf{Y}_{2}^{\prime}))_{(N,N)}=Z^{\mathbf{Y}^{\prime}}(2,N\mid 1,N)=e^{Y^{\prime}_{1,N}+Y^{\prime}_{2,N}}.$ (3.9)
 To finish the proof, simply notice that the proof of the above only required $\prod_{i=1}^{N}e^{Y^{\prime}_{i}}=\prod_{i=1}^{N}e^{Y_{i}}$.
@@
 $\prod_{i=1}^{k}e^{Y_{1,i}}=\mathfrak{s}(\mathbf{Y}_{1})\Bigl{(}\prod_{i=k+1}^{N}e^{Y_{1,i}}\Bigr{)}^{-1}.\qed$
-4. Zero-temperature analogues: proofs of Theorems 1.8 and 1.9
-Because polymer models have a strong connection to last-passage percolation (LPP), our algebraic results will naturally generalize to zero-temperature equivalents, through limiting arguments.
-In these limits, the $(+,\times)$ algebra is swapped for the $(\max,+)$ algebra, and our maps reflect this change.
-As introduced in [CGS24], the ‘correct’ map to look at here is $\widetilde{D}$, the uniform limit as $\beta$ goes to infinity of $\frac{1}{\beta}D(\beta\mathbf{X}_{1},\beta\mathbf{X}_{2})$, as well as $\widetilde{T}$, which is reached through an identical limit.
-In this section, we use our algebraic results about the behavior of the polymer model under the periodic Pitman transform (as well as the aforementioned limiting arguments) to prove analogous facts about LPP under the zero-temperature periodic Pitman transform.
-The following intermediate result is needed.
-Lemma 4.1.
-The limit as $\beta\to\infty$ of $\frac{1}{\beta}\log\bigl{(}Z^{\beta,\mathbf{W}}(V\mid U)\bigr{)}$ is $G^{\mathbf{W}}(V\mid U)$, the (multi-path)-LPP time between $U$ and $V$.
-Further, this convergence is uniform as a function of $\mathbf{W}$.
-Proof.
-By Lemma A.1, this converges uniformly, since there are finitely many multi-paths between any two sets of points. ∎
-Note that in the special case where $U,V$ are points, this shows that the single-path polymer partition function converges to single-path LPP.
-Also, because the convergence to LPP is uniform on the weights, we know that the zero-temperature limit is well-behaved up to weights which depend on $\beta$.
-If the weights $\mathbf{W}_{\beta}$ converge to $\mathbf{\widetilde{\mathcal{W}}}$, then $\frac{1}{\beta}\log(Z^{\beta,\mathbf{W}_{\beta}}(V\mid U))$ will converge to $G^{\mathbf{\widetilde{\mathcal{W}}}}(V\mid U)$.
-With this in mind, we can now prove Theorem 1.8.
-Proof of Theorem 1.8..
-As in Theorem 1.4, we take $k_{1},k_{2},\ldots,k_{m}\in\mathbb{Z},U=\bigl{(}a_{i},b_{i})_{i\in\llbracket 1,m\rrbracket}\bigr{)}$, and $V=\bigl{(}(c_{i},d_{i})_{i\in\llbracket 1,m\rrbracket}\bigr{)}$ satisfying $(U,V)\in\Psi$ and, for all $i,j\in\llbracket 1,n\rrbracket$, $a_{i}\neq k_{j}+1$ and $c_{i}\neq k_{j}$.
-To prove Theorem 1.8, we must show that $G^{\mathbf{W}}(V\mid U)=G^{\widetilde{\mathcal{P}}_{k_{n}}\ldots\widetilde{\mathcal{P}}_{k_{1}}\mathbf{W}}(V\mid U)$, where the operators $\mathcal{P}_{k}$ are defined in 1.9, with $T,D$ replaced by $\widetilde{T},\widetilde{D}$.
-Now, for $\beta>0$, let $\mathcal{P}_{k}^{\beta}:(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}\to(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$ be operators defined by their action:
-$\quad\,\mathcal{P}_{k}^{\beta}(\ldots,\mathbf{W}_{k-1},\mathbf{W}_{k},\mathbf{W}_{k+1},\mathbf{W}_{k+2},\ldots)$ (4.1)
-$=(\ldots,\mathbf{W}_{k-1},\frac{1}{\beta}T(\beta\mathbf{W}_{k+1},\beta\mathbf{W}_{k}),\frac{1}{\beta}D(\beta\mathbf{W}_{k+1},\beta\mathbf{W}_{k}),\mathbf{W}_{k+2},\ldots).$
-By Theorem 1.4, we know that $Z^{\mathbf{W}}(V\mid U)=Z^{\mathcal{P}_{k_{m}}\ldots\mathcal{P}_{k_{1}}\mathbf{W}}(V\mid U)$.
-Then, by definition of the polymer model with inverse temperature $\beta$ (1.8), for all $\beta>0$,
-$Z^{\beta,\mathbf{W}}(V\mid U)=Z^{\beta\mathbf{W}}(V\mid U)=Z^{\mathcal{P}_{k_{1}}(\beta\mathbf{W})}(V\mid U)=Z^{\beta(\frac{1}{\beta}\mathcal{P}_{k_{1}}(\beta\mathbf{W}))}(V\mid U)=Z^{\beta,\mathcal{P}_{k_{1}}^{\beta}(\mathbf{W})}(V\mid U).$ (4.2)
-By induction, we have $Z^{\beta,\mathbf{W}}(V\mid U)=Z^{\beta,\mathcal{P}_{k_{n}}^{\beta}\ldots\mathcal{P}_{k_{1}}^{\beta}\mathbf{W}}(V\mid U)$.
-By Lemma 4, we know that $\frac{1}{\beta}\log\bigl{(}Z^{\beta,\mathbf{W}}(V\mid U)\bigr{)}$ converges to $G^{\mathbf{W}}(V\mid U)$ uniformly on the weights, so $G^{\mathbf{W}}(V\mid U)$ is preserved under the limiting operator $\lim\limits_{\beta\to\infty}\mathcal{P}_{k_{n}}^{\beta}\ldots\mathcal{P}_{k_{1}}^{\beta}$.
-We claim that each $\mathcal{P}_{k}^{\beta}$ converges to $\widetilde{\mathcal{P}}_{k}$ uniformly as a function of the weights, which implies by induction that
-$G^{\mathbf{W}}(V\mid U)=G^{\lim\limits_{\beta\to\infty}\mathcal{P}_{k_{m}}^{\beta}\ldots\mathcal{P}_{k_{1}}^{\beta}\mathbf{W}}(V\mid U)=G^{\widetilde{\mathcal{P}}_{k_{m}}\ldots\widetilde{\mathcal{P}}_{k_{1}}\mathbf{W}}(V\mid U),$ (4.3)
-proving the theorem. The necessary uniform convergence is shown component-wise:
-$\quad\,\frac{1}{\beta}D_{i}(\beta\mathbf{X}_{1},\beta\mathbf{X}_{2})-\widetilde{D}_{i}(\mathbf{X}_{1},\mathbf{X}_{2})$
-$=X_{2,i+1}+\frac{1}{\beta}\log\Biggl{(}\frac{\sum_{j\in\mathbb{Z}_{N}}e^{\beta Y_{(i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{\beta Y_{(i-1,j]}}}\Biggr{)}-(X_{2,i+1}+\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}-\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]})$
-$=\Bigl{(}\frac{1}{\beta}\log\bigl{(}\sum_{j\in\mathbb{Z}_{N}}e^{\beta Y_{(i,j]}}\bigr{)}-\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}\Bigr{)}-\Bigl{(}\frac{1}{\beta}\log\bigl{(}\sum_{j\in\mathbb{Z}_{N}}e^{\beta Y_{(i-1,j]}}\bigr{)}-\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]}\Bigr{)}$
-By Lemma A.1, we know that $\frac{1}{\beta}\log(N)\leq\frac{1}{\beta}D_{i}(\beta\mathbf{X}_{1},\beta\mathbf{X}_{2})-\widetilde{D}_{i}(\mathbf{X}_{1},\mathbf{X}_{2})\leq\frac{1}{\beta}\log(N)$, proving uniform convergence.
-An identical argument holds for $T_{i}-\widetilde{T}_{i}$.
-∎
-We next show the braid relations pass to zero-temperature.
-Proof of Theorem 1.9..
-Defining $\mathcal{P}_{k}^{\beta}$ as in (4.1), we see that $\mathcal{P}_{k}^{\beta}\mathcal{P}_{k+1}^{\beta}\mathcal{P}_{k}^{\beta}=\mathcal{P}_{k+1}^{\beta}\mathcal{P}_{k}^{\beta}\mathcal{P}_{k+1}^{\beta}$, as an immediate consequence of Theorem 1.6.
-As noted in the proof of Theorem 1.8, $\mathcal{P}_{k}^{\beta}$ converges to $\widetilde{\mathcal{P}}_{k}$ uniformly.
-This immediately applies the braid relation for $\widetilde{\mathcal{P}}_{k}$, since we have
-$\widetilde{\mathcal{P}}_{k}\widetilde{\mathcal{P}}_{k+1}\widetilde{\mathcal{P}}_{k}=\lim\limits_{\beta\to\infty}\mathcal{P}_{k}^{\beta}\mathcal{P}_{k+1}^{\beta}\mathcal{P}_{k}^{\beta}=\lim\limits_{\beta\to\infty}\mathcal{P}_{k+1}^{\beta}\mathcal{P}_{k}^{\beta}\mathcal{P}_{k+1}^{\beta}=\widetilde{\mathcal{P}}_{k+1}\widetilde{\mathcal{P}}_{k}\widetilde{\mathcal{P}}_{k+1}.\qed$
-5. Burke properties and permutation invariance
-In this section, we generalize the Burke property in [CGS24, Proposition 5.12] to an inhomogeneous environment and provide geometric and exponential analogues of other Burke property for the zero-temperature discrete periodic Pitman transform.
+4. Burke property and permutation invariance
+In this section, we prove Theorems 1.7 and 1.8. We first need two intermediate lemmas. In the first, we generalize the Burke property in [CGS24, Proposition 5.12] to an inhomogeneous environment.
 Proposition 4.1 (Log-inverse-gamma Burke property).
@@
 Proof.
-The joint density $\mathbb{P}(\mathbf{X}_{1}\in d\mathbf{x}_{1},\mathbf{X}_{2}\in d\mathbf{x}_{2})$ is proportional to
+Using Lemmas A.4-A.5 in the penultimate equality below, the joint density $\mathbb{P}(\mathbf{X}_{1}\in d\mathbf{x}_{1},\mathbf{X}_{2}\in d\mathbf{x}_{2})$ is proportional to
 $\exp\Biggl{(}-\sum_{i\in\mathbb{Z}_{N}}\Bigl{(}(a_{1}+b_{i})x_{1,i}+(a_{2}+b_{i})x_{2,i}\Bigr{)}-\lambda\sum_{i\in\mathbb{Z}_{N}}(e^{-x_{1,i}}+e^{-x_{2,i}})\Biggr{)}$
@@
 $=\exp\Biggl{(}-\sum_{i\in\mathbb{Z}_{N}}\Bigl{(}(a_{1}+b_{i})T_{i}(\mathbf{x}_{1},\mathbf{x}_{2})+(a_{2}+b_{i})D_{i}(\mathbf{x}_{1},\mathbf{x}_{2})\Bigr{)}-\lambda\sum_{i\in\mathbb{Z}_{N}}(e^{-T_{i}(\mathbf{x}_{1},\mathbf{x}_{2})}+e^{-D_{i}(\mathbf{x}_{1},\mathbf{x}_{2})})\Biggr{)},$
-where we used Lemmas A.4-A.5 in the penultimate equality.
-Thus, the density is preserved under our transformation, so all we must show is that the determinant of the Jacobian matrix of the transformation has absolute value 1.
-This follows from Lemma A.7.
-∎
-Proposition 5.2 (Geometric Burke property).
-Take $\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{Z}_{\geq 0}^{\mathbb{Z}_{N}}$ to be independent sequences of independent random variables such that $X_{1,i}\sim\text{Geom}(p_{1}q_{i})$ and $X_{2,i}\sim\text{Geom}(p_{2}q_{i})$ for all $i\in\mathbb{Z}_{N}$.
-Then,
-$(\mathbf{X}_{1},\mathbf{X}_{2})\overset{d}{=}(\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2}),\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2}))$
-Proof.
-Letting $\mathbf{k}_{1},\mathbf{k}_{2}\in\mathbb{Z}_{\geq 0}^{N}$ where $k_{m,i}$ is the $i$th component of $\mathbf{k}_{m}$, we have that
-$\mathbb{P}(\mathbf{X}_{1}=\mathbf{k}_{1},\mathbf{X}_{2}=\mathbf{k}_{2})$ $=\mathbb{P}\left(X_{1,1}=k_{1,1}\right)\cdots\mathbb{P}\left(X_{1,N}=k_{1,N}\right)\mathbb{P}\left(X_{2,1}=k_{2,1}\right)\cdots\mathbb{P}\left(X_{2,N}=k_{2,N}\right)$
-$=\left(\prod\limits_{i=1}^{N}(1-p_{1}q_{i})(1-p_{2}q_{i})\right)p_{1}^{\mathfrak{s}(\mathbf{k}_{1})}p_{2}^{\mathfrak{s}(\mathbf{k}_{2})}\prod_{i=1}^{N}q_{i}^{k_{1,i}+k_{2,i}}$
-$=\left(\prod\limits_{i=1}^{N}(1-p_{1}q_{i})(1-p_{2}q_{i})\right)p_{1}^{\mathfrak{s}(\widetilde{D}(\mathbf{k}_{2},\mathbf{k}_{1}))}p_{2}^{\mathfrak{s}(\widetilde{T}(\mathbf{k}_{2},\mathbf{k}_{1}))}\prod_{i=1}^{N}q_{i}^{\widetilde{D}_{i}(\mathbf{k}_{2},\mathbf{k}_{1})+\widetilde{T}_{i}(\mathbf{k}_{2},\mathbf{k}_{1})}$
-$=\prod\limits_{i=1}^{N}(1-p_{1}q_{i})(p_{1}q_{i})^{\widetilde{D}_{i}(\mathbf{k}_{2},\mathbf{k}_{1})}\prod\limits_{i=1}^{N}(1-p_{2}q_{i})(p_{2}q_{i})^{\widetilde{T}_{i}(\mathbf{k}_{2},\mathbf{k}_{1})}$
-$=\mathbb{P}(\mathbf{X}_{1}=\widetilde{D}(\mathbf{k}_{2},\mathbf{k}_{1}),\mathbf{X}_{2}=\widetilde{T}(\mathbf{k}_{2},\mathbf{k}_{1}))$
-$=\mathbb{P}(\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})=\mathbf{k}_{2},\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2})=\mathbf{k}_{1})$
-where the first equality follows from the independence of $\mathbf{X}_{1},\mathbf{X}_{2}$ and $X_{m,i},X_{m,j}$ for $m\in\{1,2\}$ and $i,j\in\mathbb{Z}_{N}$ where $i\neq j$; the second equality follows from our definition of the geometric distribution; the third equality follows from Lemmas A.4(iii)-(iv) and A.5(i); and the final equality follows from the bijection in Lemma A.3.
-∎
-Proposition 5.3 (Exponential Burke property).
-Take $\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{Z}_{\geq 0}^{\mathbb{Z}_{N}}$ to be independent sequences such that $X_{1,i}\sim\text{Exp}(a_{1}+b_{i})$ and $X_{2,i}\sim\text{Exp}(a_{2}+b_{i})$ for all $i\in\mathbb{Z}_{N}$ where $a_{m}+b_{i}>0$ for $m\in\{1,2\}$ and $i\in\mathbb{Z}_{N}$.
-Then,
-$(\mathbf{X}_{1},\mathbf{X}_{2})\overset{d}{=}(\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2}),\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2}))$
-Proof.
-For $m\in\{1,2\}$ and $i\in\mathbb{Z}_{N}$, define
-$X_{m,i}^{\prime}\sim\text{Geom}\Biggl{(}\left(1-\frac{a_{m}}{n}\right)\left(1-\frac{b_{i}}{n}\right)\Biggr{)}$
-Then, since $\left(1-\frac{a_{m}}{n}\right)\left(1-\frac{b_{i}}{n}\right)=1-\frac{a_{m}+b_{i}}{n}+O(n^{-2})$, we have that
-$\lim\limits_{n\to\infty}\frac{X_{m,i}^{\prime}}{n}\overset{d}{=}X_{m,i}\sim\text{Exp}(a_{m}+b_{i}),\quad\text{for }m\in\{1,2\},\text{ and }i\in\mathbb{Z}_{N}$
-Observe from the definitions of $\widetilde{T}$ and $\widetilde{D}$ (1.11) that
-$\frac{1}{n}\widetilde{T}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime})=\widetilde{T}\Bigl{(}\frac{\mathbf{X}_{1}^{\prime}}{n},\frac{\mathbf{X}_{2}^{\prime}}{n}\Bigr{)},\quad\text{and}\quad\frac{1}{n}\widetilde{D}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime})=\widetilde{D}\Bigl{(}\frac{\mathbf{X}_{1}^{\prime}}{n},\frac{\mathbf{X}_{2}^{\prime}}{n}\Bigr{)}$
-By Proposition 5.2, we know that $(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime})\overset{d}{=}(\widetilde{T}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime}),\widetilde{D}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime}))$.
-Since $\max$, addition, and subtraction are continuous, $\widetilde{D}$ and $\widetilde{T}$ are continuous maps.
-We then have that
-$\Bigl{(}\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2}),\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})\Bigr{)}\overset{d}{=}\lim\limits_{n\to\infty}\frac{1}{n}\Bigl{(}\widetilde{T}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime}),\widetilde{D}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime})\Bigr{)}\overset{d}{=}\lim\limits_{n\to\infty}\frac{1}{n}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime})\overset{d}{=}(\mathbf{X}_{1},\mathbf{X}_{2}).\qed$
-5.1. Proofs of Theorems 1.10 and 1.11
-We first prove a lemma that connects the condition stated in Theorem 1.10 to the condition in Theorem 1.4.
+Thus, the density is preserved under our transformation, so all we must show is that the determinant of the Jacobian matrix of the transformation has absolute value 1. This follows from Lemma A.7. ∎
+We next prove a lemma that connects the condition stated in Theorem 1.7 to the condition in Theorem 1.4.
 Lemma 4.2.
@@
 Since $\sigma_{1}$ only permutes elements of $\mathbb{Z}_{<x}$, we have that $k_{\ell_{t}}+1<x$ for $1\leq t\leq r$.
-Since $\sigma_{2}$ only permutes elements of $\mathbb{Z}_{\geq x}$, $k_{p_{t}}+1>x$ for $1\leq t\leq n$.
+Since $\sigma_{2}$ only permutes elements of $\mathbb{Z}_{\geq x}$, we have $k_{p_{t}}+1>x$ for $1\leq t\leq n$.
 Hence, $x\neq k_{\ell}+1$ for $1\leq\ell\leq m$.
@@
 Take this decomposition to have the minimal number of terms.
-Then, we consider the action of the map $\mathcal{P}_{k_{m}}\cdots\mathcal{P}_{k_{1}}$ on the weights $\mathbf{W}$, as in Theorem 1.4.
+Then, we consider the action of the map $\mathcal{P}_{k_{m}}\cdots\mathcal{P}_{k_{1}}$ on the weights $\mathbf{W}^{\mathbf{a},\mathbf{b}}$, as in Theorem 1.4.
 By Lemma 4.2, the condition (1.11) and the minimality of the decomposition ensures that the conditions of Theorem 1.4 are met, so
-$Z^{\mathcal{P}_{k_{m}}\cdots\mathcal{P}_{k_{1}}\mathbf{W}}(V\mid U)=Z^{\mathbf{W}}(V\mid U),\quad\text{for all }(U,V)\in\mathcal{U}^{\sigma}.$
-Furthermore, by successively applying the Burke property in Proposition 5.1, we see that each operator $\mathcal{P}_{k}$ swaps the column parameters in columns $k$ and $k+1$, so we have that $\mathcal{P}_{k_{m}}\cdots\mathcal{P}_{k_{1}}\mathbf{W}\overset{d}{=}\mathbf{W}^{\sigma\mathbf{a},\mathbf{b}}$.
+$Z^{\mathcal{P}_{k_{m}}\cdots\mathcal{P}_{k_{1}}\mathbf{W}^{\mathbf{a},\mathbf{b}}}(V\mid U)=Z^{\mathbf{W}^{\mathbf{a},\mathbf{b}}}(V\mid U),\quad\text{for all }(U,V)\in\mathcal{U}^{\sigma}.$
+Furthermore, by successively applying the Burke property in Proposition 4.1, we see that each operator $\mathcal{P}_{k}$ swaps the column parameters in columns $k$ and $k+1$, so we have that $\mathcal{P}_{k_{m}}\cdots\mathcal{P}_{k_{1}}\mathbf{W}^{\mathbf{a},\mathbf{b}}\overset{d}{=}\mathbf{W}^{\sigma\mathbf{a},\mathbf{b}}$.
 This completes the proof.
-In the last-passage percolation setting, the proof is the same, instead using the invariance statement in Theorem 1.8 and the Burke properties in Propositions 5.2 and 5.3. ∎
+∎
 Proof of Theorem 1.8..
@@
 ∎
-6. Jointly invariant measures: proofs of Theorems 1.12 and 1.13
-We prove two theorems in this section, starting with Theorem 1.12.
-6.1. Proof of Theorem 1.12
-Proof of Theorem 1.12..
-We prove the case when the driving weights are geometric random variables, as the exponential case is proved the same way.
-The key is the Burke property in Proposition 5.2.
-The proof of the case when the driving weights are exponential is the same, instead using the Burke property in Proposition 5.3.
-One can readily see that, for any $q\in(0,1)$, if $\mathbf{X}\in\mathbb{Z}_{\geq 0}^{\mathbb{Z}_{N}}$ is an i.i.d.
-collection of $\text{Geom}(r)$ random variables, then the conditional law of $\mathbf{X}$, given $\mathfrak{s}(\mathbf{X})=\theta$ is the uniform distribution on $\mathbb{Z}_{\theta,\geq 0}^{\mathbb{Z}_{N}}$.
-Since $\mathfrak{s}(D(\mathbf{X}_{1},\mathbf{X}_{2}))=\mathfrak{s}(\mathbf{X}_{2})$ by Lemma A.4(iv), to prove the invariance result, it suffices to show the following: If $r_{1},r_{2},q\in(0,1)$, and $\mathbf{X}_{1},\mathbf{X}_{2},\mathbf{W}$ are mutually independent vectors of i.i.d.
-geometric random variables with respective parameters $r_{1},r_{2},q$, then
-$(\mathbf{U}_{1},\mathbf{U}_{2})\overset{d}{=}\bigl{(}\widetilde{D}(\mathbf{W},\mathbf{U}_{1}),\widetilde{D}(\mathbf{W},\mathbf{U}_{2})\bigr{)},\quad\text{where}$
-$\mathbf{U}_{1}=\mathbf{X}_{1},\quad\text{and}\quad\mathbf{U}_{2}=\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2}).$
-By independence of $\mathbf{X}_{1},\mathbf{X}_{2},\mathbf{W}$ and the Burke property in Proposition 5.2,
-$(\widetilde{D}(\mathbf{W},\mathbf{X}_{1}),\widetilde{T}(\mathbf{W},\mathbf{X}_{1}),\mathbf{X}_{2})\overset{d}{=}(\mathbf{X}_{1},\mathbf{W},\mathbf{X}_{2}),$ (6.1)
-and similarly,
-$(\mathbf{X}_{1},\mathbf{X}_{2})\overset{d}{=}\bigl{(}\mathbf{X}_{1},\widetilde{D}(\mathbf{W},\mathbf{X}_{2})\bigr{)}$ (6.2)
-By definition of $\mathbf{U}_{1}$ and $\mathbf{U}_{2}$ along with the braid relation in Theorem 1.9 (specifically the zero-temperature version of (3.3)), we have
-$\bigl{(}\widetilde{D}(\mathbf{W},\mathbf{U}_{1}),\widetilde{D}(\mathbf{W},\mathbf{U}_{2})\bigr{)}$ $=\Bigl{(}\widetilde{D}(\mathbf{W},\mathbf{X}_{1}),\widetilde{D}\bigl{(}\widetilde{D}(\mathbf{W},\mathbf{X}_{1}),\widetilde{D}(\widetilde{T}(\mathbf{W},\mathbf{X}_{1}),\mathbf{X}_{2})\bigr{)}\Bigr{)}$
-[…]
-[…]
-We turn to proving the distributional equality (1.17): By another application of the Burke property, we have
-$(\mathbf{X}_{1},\mathbf{X}_{2})\overset{d}{=}\bigl{(}\widetilde{D}(\mathbf{X}_{2},\mathbf{X}_{1}),\widetilde{T}(\mathbf{X}_{2},\mathbf{X}_{1})\bigr{)}.$
-Then, we have
-$\bigl{(}\mathbf{X}_{1},\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})\bigr{)}\overset{d}{=}\bigl{(}\widetilde{D}(\mathbf{X}_{2},\mathbf{X}_{1}),\widetilde{D}(\widetilde{D}(\mathbf{X}_{2},\mathbf{X}_{1}),\widetilde{T}(\mathbf{X}_{2},\mathbf{X}_{1}))\bigr{)}=\bigl{(}\widetilde{D}(\mathbf{X}_{2},\mathbf{X}_{1}),\mathbf{X}_{2}\bigr{)},$
-where in the last equality we used the fact that $(\mathbf{X}_{1},\mathbf{X}_{2})\mapsto(\widetilde{T}(\mathbf{X}_{2},\mathbf{X}_{1}),\widetilde{D}(\mathbf{X}_{2},\mathbf{X}_{1}))$ is an involution, which is proved in Lemma A.2.
-∎
-6.2. Cyclic Interval Notation
-Before proving Theorem 1.13, we discuss some conventions around cyclic notation. Recall that we take $\mathbb{Z}_{N}$ to be the cyclic group of order $N$.
-Definition 6.1.
-For $i,j\in\mathbb{Z}_{N}$, we define the periodic intervals
-$(i,j]:=\begin{cases}\varnothing&i=j\\ \{i+1,i+2,\ldots,j\}&i\neq j\end{cases}$
-$[i,j):=\begin{cases}\varnothing&i=j\\ \{i,i+1,\ldots,j-1\}&i\neq j\end{cases}$
-$[i,j]:=\begin{cases}i&i=j\\ \{i,i+1,\ldots,j\}&i\neq j\end{cases}$
-$(i,j):=\begin{cases}\varnothing&j\in\{i,i+1\}\\ \{i+1,i+2,\ldots,j-1\}&i\neq j.\end{cases}$
-Using this cyclic notation, we now prove the following lemma that will be used several times in the proof of Theorem 1.13.
-Lemma 6.2.
-Let $\mathbf{Y}\in\mathbb{R}^{\mathbb{Z}_{N}}$. Then, the following hold.
-  1. (i)
-    If $a,b\in\mathbb{Z}_{N}$ where $a\neq b$, then for any $c\in[a,b)$, $Y_{(a,c]}+Y_{[c+1,b]}=Y_{[a+1,b]}$.
-  2. (ii)
-    For $a,c\in\mathbb{Z}_{N}$, we have $Y_{(a,c]}+Y_{[c+1,a]}=\mathfrak{s}(\mathbf{Y})$.
-Proof.
-Item (i): We prove this in two cases. First, if $c=a$, then $Y_{(a,c]}=0$, so $Y_{(a,c]}+Y_{[c+1,b]}=Y_{[a+1,b]}.$
-Next, if $c\in[a,b)$, but $c\neq a$, then $Y_{(a,c]}=\sum_{\ell=a+1}^{c}Y_{\ell}$ and $Y_{[c+1,b]}=\sum_{\ell=c+1}^{b}Y_{\ell},$ so
-$Y_{(a,c]}+Y_{[c+1,b]}=\sum_{\ell=a+1}^{c}Y_{\ell}+\sum_{\ell=c+1}^{b}Y_{\ell}=\sum_{\ell=a+1}^{b}Y_{\ell}=Y_{[a+1,b]},$
-where the sums are taken in cyclic order. The condition $c\in[a,b)$ ensures there is no double-counting.
-Item (ii): We likewise prove this statement in two cases. If $a=c$, then $Y_{(a,c]}=0$ and $Y_{[c+1,a]}=Y_{[a+1,a]}=\sum_{k\in\mathbb{Z}_{N}}Y_{k}=\mathfrak{s}(\mathbf{Y})$.
-On the other hand, if $a\neq c$ then
-$Y_{(a,c]}+Y_{[c+1,a]}=\sum_{\ell=a+1}^{c}Y_{\ell}+\sum_{\ell=c+1}^{a}Y_{\ell}=\sum_{\ell\in\mathbb{Z}_{N}}Y_{\ell}=\mathfrak{s}({\mathbf{Y}}).\qed$
-6.3. Proof of Theorem 1.13
-We start with proving several algebraic lemmas that are used as inputs in the proof of the Theorem.
-Recall the definitions of the spaces $\mathbb{R}_{\theta}^{\mathbb{Z}_{N}},\mathbb{R}_{\theta,\geq 0}^{\mathbb{Z}_{N}}$, and $\mathcal{Z}_{\theta_{1},\theta_{2}}$ defined in (1.16) and (1.18).
-The first lemma below is the key to our combinatorial description of the joint law in Theorem 1.13.
-Specifically, it turns the problem into understanding a single quantity–namely $\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]}$.
-Lemma 6.3.
-Let $\mathbf{X}_{1}\in\mathbb{R}^{\mathbb{Z}_{N}}_{\theta_{1}}$ and $\mathbf{X}_{2}\in\mathbb{R}^{\mathbb{Z}_{N}}_{\theta_{2}}$ where $0\leq\theta_{1}\leq\theta_{2}$. Then, for $i\in\mathbb{Z}_{N}$,
-$\widetilde{D}_{i}(\mathbf{X}_{1},\mathbf{X}_{2})=X_{1,i}+\max(0,\theta_{2}-\theta_{1}-\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]}),$
-where $Y_{\ell}=X_{2,\ell+1}-X_{1,\ell}$.
-Proof.
-Notice that when $\theta_{2}\geq\theta_{1}$, we have that $Y_{i}+\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}=\max\bigl{(}\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]},\theta_{2}-\theta_{1}\bigr{)}$ for all $i\in\mathbb{Z}_{N}$.
-Thus, we have
-$\widetilde{D}_{i}(\mathbf{X}_{1},\mathbf{X}_{2})$ $=X_{2,i+1}+\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}-\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]}$
-$=X_{1,i}+Y_{i}+\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}-\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]}$
-$=X_{1,i}+\max(\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]},\theta_{2}-\theta_{1})-\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]})$
-$=X_{1,i}+\max(0,\theta_{2}-\theta_{1}-\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]})\qed$
-Since $\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]}\geq Y_{(i-1,i-1]}=0$, we have the following corollary of Lemma 6.3.
-Corollary 6.4.
-Let $\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{R}^{\mathbb{Z}_{N}}_{\theta}$ for some $\theta\in\mathbb{R}$. Then $\widetilde{D}_{i}(\mathbf{X}_{1},\mathbf{X}_{2})=X_{1,i}$ for all $i\in\mathbb{Z}_{N}$.
-Lemma 6.5.
-Let $\mathbf{X}_{1}\in\mathbb{R}^{\mathbb{Z}_{N}}_{\theta_{1}}$ and $\mathbf{X}_{2}\in\mathbb{R}^{\mathbb{Z}_{N}}_{\theta_{2}}$ where $\theta_{1}\leq\theta_{2}$.
-Set $Y_{i}:=X_{2,i+1}-X_{1,i}$ for all $i\in\mathbb{Z}_{N}$.
-If $Y_{i}\leq 0$ for some $i\in\mathbb{Z}_{N}$, then $\widetilde{D}_{i+1}(\mathbf{X}_{1},\mathbf{X}_{2})=X_{1,i+1}$ and $\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}\geq\theta_{2}-\theta_{1}$.
-Proof.
-Observe that $Y_{(i,i-1]}=Y_{i+1}+Y_{i+2}+\ldots+Y_{i-1}\geq Y_{i}+Y_{i+1}+Y_{i+2}+\ldots+Y_{i-1}=\theta_{2}-\theta_{1}$ since $Y_{i}\leq 0$.
-Hence, $\theta_{2}-\theta_{1}-Y_{(i,i-1]}\leq 0$ so $\max(0,\theta_{2}-\theta_{1}-\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]})=0$.
-The lemma now follows from Lemma 6.3.
-∎
-Lemma 6.6.
-Let $\mathbf{X}_{1}\in\mathbb{R}^{\mathbb{Z}_{N}}_{\theta_{1}}$ and $\mathbf{X}_{2}\in\mathbb{R}^{\mathbb{Z}_{N}}_{\theta_{2}}$ where $\theta_{1}\leq\theta_{2}$.
-Set $Y_{i}:=X_{2,i+1}-X_{1,i}$ and let $\widetilde{D}_{i}$ denote the $i$th component of $\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})$.
-If $\widetilde{D}_{i}>X_{1,i}$ and $\widetilde{D}_{i+1}=X_{1,i+1}$ for some $i\in\mathbb{Z}_{N}$, then $Y_{i}\leq 0$.
-Proof.
-Suppose for the sake of contradiction that $Y_{i}>0$.
-Then, $Y_{(i,i-1]}=\theta_{2}-\theta_{1}-Y_{i}<\theta_{2}-\theta_{1}$.
-Since $\widetilde{D}_{i+1}=X_{1,i+1}$, it follows by Lemma 6.3 that $\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}\geq\theta_{2}-\theta_{1}$, so there exists $m\in\mathbb{Z}_{N}$ such that $Y_{(i,m]}\geq\theta_{2}-\theta_{1}$.
-Since $Y_{(i,i-1]}=0<\theta_{2}-\theta_{1}$, we have $m\neq i-1$, and therefore, $Y_{(i-1,m]}=Y_{i}+Y_{(i,m]}>\theta_{2}-\theta_{1}$.
-By another application of Lemma 6.3, this implies that $\widetilde{D}_{i}=X_{1,i}$, a contradiction to our assumption.
-∎
-Lemma 6.7.
-Let $\mathbf{X}_{1}\in\mathbb{R}^{\mathbb{Z}_{N}}_{\theta_{1}}$ and $\mathbf{X}_{2}\in\mathbb{R}^{\mathbb{Z}_{N}}_{\theta_{2}}$ where $\theta_{1}\leq\theta_{2}$.
-Set $Y_{i}:=X_{2,i+1}-X_{1,i}$ and let $\widetilde{D}_{i}$ denote the $i$th component of $\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})$.
-If $\widetilde{D}_{i}>X_{1,i}$, then $Y_{i}<\theta_{2}-\theta_{1}$.
-Proof.
-By Lemma 6.3, $\theta_{2}-\theta_{1}-\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]}>0$, so since $Y_{i}=Y_{(i-1,i]}$, we have $\theta_{2}-\theta_{1}-Y_{i}>0$. ∎
-We now complete the proof of Theorem 1.13.
-Proof.
-of Theorem 1.13.
-Let $0\leq\theta_{1}<\theta_{2}$, let $\mathbf{X}_{1}$ be uniformly distributed on the set $\mathbb{Z}_{\theta_{1},\geq 0}^{\mathbb{Z}_{N}}$, and let $\mathbf{X}_{2}$ be uniformly distributed on the set $\mathbb{Z}_{\theta_{2},\geq 0}^{\mathbb{Z}_{N}}$.
-By Lemmas 6.3, A.3, and A.4(iv), the joint distribution of $(\mathbf{X}_{1},\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2}))$ is supported on the set
-$\mathcal{Z}_{\theta_{1},\theta_{2}}=\{(\mathbf{u}_{1},\mathbf{u}_{2})\in\mathbb{Z}^{\mathbb{Z}_{N}}_{\theta_{1},\geq 0}\times\mathbb{Z}^{\mathbb{Z}_{N}}_{\theta_{2},\geq 0}\mid\mathbf{u}_{1}\leq\mathbf{u}_{2}\}.$
-Observe that
-$|\mathbb{Z}_{\theta_{1},\geq 0}^{\mathbb{Z}_{N}}\times\mathbb{Z}_{\theta_{2},\geq 0}^{\mathbb{Z}_{N}}|=\binom{\theta_{1}+N-1}{N-1}\binom{\theta_{2}+N-1}{N-1}$,
-so for any $\mathbf{u}_{1}\in\mathbb{Z}_{\theta_{1},\geq 0}^{\mathbb{Z}_{N}}$ and $\mathbf{u}_{2}\in\mathbb{Z}_{\theta 2,+}^{\mathbb{Z}_{N}}$ with $\mathbf{u}_{1}\leq\mathbf{u}_{2}$ coordinate-wise, we have
-$\mathbb{P}(\mathbf{X}_{1}=\mathbf{u}_{1},\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})=\mathbf{u}_{2})=\frac{\alpha(\mathbf{u}_{1},\mathbf{u}_{2})}{\binom{\theta_{1}+N-1}{N-1}\binom{\theta_{2}+N-1}{N-1}}$,
-where $\alpha(\mathbf{u}_{1},\mathbf{u}_{2})$ is the number of vectors $\mathbf{X}_{2}\in\mathbb{Z}_{\theta_{2},\geq 0}^{\mathbb{Z}_{N}}$ satisfying $\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})=\mathbf{u}_{2}$, when $\mathbf{X}_{1}=\mathbf{u}_{1}$.
-We compute the value of $\alpha(\mathbf{u}_{1},\mathbf{u}_{2})$ in several cases, depending on how many indices $i\in\mathbb{Z}_{N}$ satisfy the equality $u_{1,i}=u_{2,i}$.
-We make a change of coordinates $\mathbf{Y}=(Y_{i})_{i\in\mathbb{Z}_{N}}$, where $Y_{i}=X_{2,i+1}-u_{1,i}$.
-We immediately see that the number of vectors $\mathbf{X}_{2}$ satisfying $\mathbf{u}_{2}=\widetilde{D}(\mathbf{u}_{1},\mathbf{X}_{2})$ is in bijection with the set of corresponding vectors $\mathbf{Y}$.
-Case 1: Zero Equalities: In this case, $u_{2,i}>u_{1,i}$ for all $i\in\mathbb{Z}_{N}$.
-Assume that $\mathbf{u}_{2}=\widetilde{D}(\mathbf{u}_{1},\mathbf{X}_{2})$ for some $\mathbf{X}_{2}\in\mathbb{Z}_{\theta_{2},\geq 0}^{\mathbb{Z}_{N}}$.
-By Lemma 6.3, we have $\theta_{2}-\theta_{1}>\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]}$ for all $i\in\mathbb{Z}_{N}$.
-By the contrapositive of Lemma 6.5, we have that $Y_{i}>0$ for all $i\in\mathbb{Z}_{N}$, so $Y_{(i-1,j]}$ is maximized when we include as many $Y_{i}$s in the sum as possible.
-Thus, $\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]}=Y_{(i-1,i-2]}=\theta_{2}-\theta_{1}-Y_{i-1}$.
-Hence, using Lemma 6.3 in the last equality below, we have
-$X_{2,i}$ $=u_{1,i-1}+Y_{i-1}$
-$=u_{1,i-1}+\theta_{2}-\theta_{1}-\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]}$
-$=u_{1,i-1}+u_{2,i}-u_{1,i}$
-On the other hand, if we define $X_{2,i}=u_{1,i-1}+u_{2,i}-u_{1,i}$, then $\mathfrak{s}(\mathbf{X}_{2})=\theta_{2}$, $X_{2,i}\geq 0$ for all $i\in\mathbb{Z}_{N}$, and $Y_{i}=u_{2,i+1}-u_{1,i+1}>0$ for all $i\in\mathbb{Z}_{N}$, and by applying the steps above in reverse, we obtain that $\mathbf{u}_{2}=\widetilde{D}(\mathbf{u}_{1},\mathbf{X}_{2})$.
-Thus, there is a unique $\mathbf{X}_{2}\in\mathbb{Z}_{\theta_{2},\geq 0}^{\mathbb{Z}_{N}}$ such that $\widetilde{D}(\mathbf{u}_{1},\mathbf{X}_{2})=\mathbf{u}_{2}$, and so $\alpha(\mathbf{u}_{1},\mathbf{u}_{2})=1$.
-Case 2: $\mathbf{<N}$ Equalities: In this case, there exists $i\in\mathbb{Z}_{N}$ such that $u_{2,i-1}>u_{1,i-1}$ and $u_{2,i}=u_{1,i}$.
-By invariance of the uniform distribution on $\mathbb{Z}_{\theta,\geq 0}^{\mathbb{Z}_{N}}$ under cyclic permutations and Lemma A.6, the joint distribution of $\bigl{(}\mathbf{X}_{1},\mathbf{\widetilde{D}}\bigr{)}$ is invariant under cyclic permutations, so we may, without loss of generality, assume that $u_{2,1}=u_{1,1}$ and $u_{2,N}>u_{1,N}$.
-Then, starting from $i=1$, there is a collection of sequences of equalities and inequalities, when comparing $u_{1,i}$ with $u_{2,i}$.
-Let $k_{m}\in\mathbb{Z}_{N}$ be the ending index of the $m$th consecutive equality and $\ell_{m}\in\mathbb{Z}_{N}$ be the ending index of the $m$th consecutive inequality, as shown in Figure 5.
-To be precise, we write
-$k_{1}=\max\{t\in\mathbb{Z}_{\geq 1}\mid u_{2,j}=u_{1,j}\quad\forall 1\leq j\leq t\}.$
-Then, we can recursively define
-$\ell_{m}=\max\{t\in\mathbb{Z}_{\geq k_{m}+1}\mid u_{2,j}>u_{1,j}\quad\forall k_{m}+1\leq j\leq t\}$ (6.3)
-$k_{m}=\max\{t\in\mathbb{Z}_{\geq\ell_{m-1}+1}\mid u_{2,j}=u_{1,j}\quad\forall\ell_{m-1}+1\leq j\leq t\},$ (6.4)
-until $m=s$, where $s$ is the integer satisfying $\ell_{s}=N$.
-Such an integer is guaranteed to exist since $u_{2,N}>u_{1,N}$, but $u_{1,1}=u_{2,1}$.
-Recalling that $\mathbf{u}_{1},\mathbf{u}_{2}$ are vectors with indices in $\mathbb{Z}_{N}$, we can embed $k_{m},\ell_{m}$ in $\mathbb{Z}_{N}$.
-Moreover, if we let $k_{m}=k_{m^{\prime}}$ and $\ell_{m}=\ell_{m^{\prime}}$ for all $m,m^{\prime}\in\mathbb{Z}$ such that $m\equiv m^{\prime}\pmod{s}$, then $k_{m},\ell_{m}\in\mathbb{Z}_{N}$ are defined for nonpositive $m$ as well as for $m>s$.
-Note that we sometimes add $N\mathbf{1}_{m=s}$ when working with the lengths or indices of vectors, since, for instance, when we are considering a vector whose length is the number of indices between $\ell_{s}$ and $k_{1},$ we must ‘wrap-around’ indices.
-Thus, the length of the vector would be $(N-\ell_{s})+(k_{1}-0)$ rather than $\ell_{s}-k_{1}$.
-We now provide an overview of the proof strategy using lemmas that we will prove in a later subsection.
-Recall the change of coordinates $\mathbf{Y}$ discussed in the previous case.
-We claim that the following conditions are necessary and sufficient for $\mathbf{X}_{2}$ to satisfy $\widetilde{D}(\mathbf{u}_{1},\mathbf{X}_{2})=\mathbf{u}_{2}$:
-$Y_{j}=u_{2,j+1}-u_{1,j+1}\text{ for all }j\in(k_{m},\ell_{m}),m\in\llbracket 1,s\rrbracket$ (6.5a)
-$Y_{[\ell_{m-1},j]}\leq 0\text{ for all }j\in[\ell_{m-1},k_{m}-1],m\in\llbracket 1,s\rrbracket$ (6.5b)
-$Y_{[\ell_{m-1},k_{m}]}=u_{2,k_{m}+1}-u_{1,k_{m}+1}\text{ for all }m\in\llbracket 1,s\rrbracket$ (6.5c)
-$Y_{j}\geq-u_{1,j}\text{ for all }j\in\mathbb{Z}_{N}$ (6.5d)
-Once this is shown, it suffices to count the number of vectors $\mathbf{Y}\in\mathbb{Z}^{\mathbb{Z}_{N}}$ satisfying the conditions above.
-To do this, we claim that the number of vectors $\mathbf{Y}$ satisfying (6.5) is the same as the number of $s$-tuples of vectors $(\mathbf{a}_{1},\ldots,\mathbf{a}_{s})$ satisfying (1.19).
-To show this, we first observe that if we choose values of $Y_{j}$ for every $j\in[\ell_{m},k_{m+1}-1]$ and $m\in\llbracket 1,s\rrbracket$, then setting $Y_{k_{m+1}}=u_{2,k_{m+1}+1}-u_{1,k_{m+1}+1}-Y_{[\ell_{m},k_{m+1}-1]}$ and $Y_{j}=u_{2,j+1}-u_{1,j+1}$ for all $j\in(k_{m},\ell_{m})$ and $m\in\llbracket 1,s\rrbracket$, fixes the other values of $Y_{j}$.
-Hence, if we take $a_{m,p}$ to represent $Y_{\ell_{m}+p-1}$ for $m\in\llbracket 1,s\rrbracket$ and $1\leq p\leq k_{m+1}-\ell_{m}+N\mathbf{1}_{m=s}$, we find the relevant set of equivalent constraints on the free values of $Y_{j}$ to count the total number of possible vectors $\mathbf{Y}$.
-Namely, we give lower bounds on the vector coordinates $a_{m,p}\geq-u_{1,\ell_{m}+p-1}$ and an upper bound of 0 on the sums of vector coordinates $a_{m,1},a_{m,1}+a_{m,2},\ldots,a_{m,1}+a_{m,2}+\ldots+a_{m,p}$ for $1\leq p\leq k_{m+1}-\ell_{m}+N1_{m=s}$.
-We turn now to showing that the conditions (6.5) are necessary and sufficient for $\mathbf{X}_{2}$ to satisfy $\mathbf{u}_{2}=\widetilde{D}(\mathbf{u}_{1},\mathbf{X}_{2})$.
-Proof of necessity: First, assuming $\mathbf{u}_{2}=\widetilde{D}(\mathbf{u}_{1},\mathbf{X}_{2})$, we will show that each of the conditions in (6.5) hold.
-To do this, we state some intermediate lemmas, whose proofs are postponed until Section 6.3.1.
-Lemma 6.8.
-If $\ell_{0}=k_{1}+1$, then we have N1N-1 equalities $u_{1,i}=u_{2,i}$, and $s=1$. Moreover, the conditions (6.5) hold.
-Lemma 6.9.
-We have that $\max_{j\in\mathbb{Z}_{N}}Y_{(k_{m}-1,j]}=Y_{(k_{m}-1,\ell_{m-1}-1]}$ for all $m\in\llbracket 1,s\rrbracket$.
-Lemma 6.10.
-Whenever $\ell_{0}\neq k_{1}+1$, then we have that $\max_{j\in\mathbb{Z}_{N}}Y_{(k_{m},j]}=Y_{(k_{m},\ell_{m-1}-1]}$ for all $m\in\llbracket 1,s\rrbracket$.
-Lemma 6.11.
-Whenever $q\in(k_{m},\ell_{m})$, we have that $\max\limits_{j\in\mathbb{Z}_{N}}Y_{(q,j]}=Y_{(q,q-1]}$ for $m\in\llbracket 1,s\rrbracket$.
-By Lemma 6.8, we may assume that $\ell_{0}\neq k_{1}+1$.
-Note that this implies $\ell_{m-1}\neq k_{m}+1$ for all $m\in\llbracket 1,s\rrbracket$; otherwise there would be N1N-1 equalities of the form $u_{2,i}=u_{1,i}$.
-By Lemma 6.11, whenever $m\in[1,s]$ and $j\in(k_{m},\ell_{m})$ to get $\max_{i\in\mathbb{Z}_{N}}Y_{(j,i]}=Y_{(j,j-1]}=\theta_{2}-\theta_{1}-Y_{j}$.
-Then, by assumption on the locations of the equalities, followed by an application of Lemma 6.3,
-$u_{1,j+1}<u_{2,j+1}$ $=u_{1,j+1}+\max(0,\theta_{2}-\theta_{1}-\max_{i\in\mathbb{Z}_{N}}Y_{(j,i]})$
-$=u_{1,j+1}+\max(0,\theta_{2}-\theta_{1}-(\theta_{2}-\theta_{1}-Y_{j}))$
-$=u_{1,j+1}+\max(0,Y_{j})$
-By the inequality, we must have $\max(0,Y_{j})=Y_{j}$, and then, $u_{2,j+1}=u_{1,j+1}+Y_{j}$, giving us (6.5a).
-Next, since $u_{2,k_{m}}=u_{1,k_{m}}$ by definition of $k_{m}$, Lemma 6.3 implies that $\max_{j\in\mathbb{Z}_{N}}Y_{(k_{m}-1,j]}\geq\theta_{2}-\theta_{1}$.
-Combined with Lemma 6.9, we have $Y_{(k_{m}-1,\ell_{m-1}-1]}\geq\theta_{2}-\theta_{1}$.
-Lemma 6.2(i) now tells us that
-$Y_{[\ell_{m-1},k_{m}-1]}\leq 0.$ (6.6)
-Then, by Lemma 6.2(ii), we have that
-$Y_{[\ell_{m-1},j]}=Y_{(k_{m}-1,j]}-Y_{(k_{m}-1,\ell_{m-1}-1]}\leq 0,\quad\text{for }j\in[\ell_{m-1},k_{m}-1),$ (6.7)
-and (6.6) and (6.7) together prove (6.5b).
-By definition of $k_{m}$, we have $u_{2,k_{m}+1}>u_{1,k_{m}+1}$. Then, by Lemma 6.3, followed by Lemma 6.10, then Lemma 6.2(ii), we observe that, for $m\in\llbracket 1,s\rrbracket$,
-$u_{2,k_{m}+1}-u_{1,k_{m}+1}=\theta_{2}-\theta_{1}-\max\limits_{j\in\mathbb{Z}_{N}}Y_{(k_{m},j]}=\theta_{2}-\theta_{1}-Y_{(k_{m},\ell_{m-1}-1]}=Y_{(\ell_{m-1}-1,k_{m}]}=Y_{[\ell_{m-1},k_{m}]},$
-proving (6.5c).
-Finally, observe that (6.5d) follows directly from the fact that $Y_{i}=X_{2,i+1}-u_{1,i}$ and $X_{2,i}\geq 0$ for all $i\in\mathbb{Z}_{N}$ by definition.
-Proof of sufficiency: Next, we assume that (6.5) holds, and prove that $\mathbf{u}_{2}=\widetilde{D}(\mathbf{u}_{1},\mathbf{X}_{2})$.
-We first show that (6.5) implies that $\mathfrak{s}(\mathbf{Y})=\theta_{2}-\theta_{1}$, which is equivalent to $\mathfrak{s}(\mathbf{X}_{2})=\theta_{2}$.
-To see this, observe that
-$\mathfrak{s}(\mathbf{Y})$ $=\sum\limits_{m\in\llbracket 1,s\rrbracket}Y_{[\ell_{m-1},\ell_{m})}$
-$=\sum\limits_{m\in\llbracket 1,s\rrbracket}Y_{[\ell_{m-1},k_{m}]}+\sum\limits_{m\in\llbracket 1,s\rrbracket}Y_{(k_{m},\ell_{m})}$
-$=\sum\limits_{m\in\llbracket 1,s\rrbracket}\left(u_{2,k_{m}+1}-u_{1,k_{m}+1}\right)+\sum\limits_{m\in\llbracket 1,s\rrbracket}\sum\limits_{j\in(k_{m},\ell_{m})}\left(u_{2,j+1}-u_{1,j+1}\right)$
-$=\sum\limits_{m\in\llbracket 1,s\rrbracket}\sum\limits_{j\in[k_{m}+1,\ell_{m}+1)}\left(u_{2,j}-u_{1,j}\right)=\sum_{j\in\mathbb{Z}_{N}}(u_{2,j}-u_{1,j})$
-$=\theta_{2}-\theta_{1},$
-where the second equality follows from Lemma 6.2(i), the third equality follows from (6.5a) and (6.5c), and the final equality follows since $(\mathbf{u}_{1},\mathbf{u}_{2})\in\mathcal{Z}_{\theta_{1},\theta_{2}}$.
-Combined with the condition (6.5d), we have $\mathbf{X}_{2}\in\mathbb{Z}_{\theta_{2},\geq 0}^{\mathbb{Z}_{N}}$.
-We now show that $\widetilde{D}_{i}(\mathbf{u}_{1},\mathbf{X}_{2})=u_{2,i}$ for $i\in(k_{m},\ell_{m}+1),m\in\llbracket 1,s\rrbracket$.
-In other words, we prove that the value of $\widetilde{D}(\mathbf{u}_{1},\mathbf{X}_{2})$ aligns with $\mathbf{u}_{2}$ at every index where $u_{2,i}>u_{1,i}$.
-This suffices to show equality at every other index since we know $\widetilde{D}_{i}(\mathbf{u}_{1},\mathbf{X}_{2})\geq u_{1,i}$ and $\mathfrak{s}(\widetilde{D}(\mathbf{u}_{1},\mathbf{X}_{2}))=\mathfrak{s}(\mathbf{X}_{2})=\mathfrak{s}(\mathbf{u}_{1})+\mathfrak{s}(\mathbf{Y})=\mathfrak{s}(\mathbf{u}_{2})$.
-We note that (6.5a) gives that $Y_{j}>0$ for $j\in(k_{m},\ell_{m}),m\in[1,s]$.
-Hence, when trying to maximize any cyclic sum $Y_{(i,j]}$, it is strictly better to extend the bounds of the cyclic sum to include these positive values.
-Similarly, (6.5c) implies that it is strictly better to contain the sum $Y_{[\ell_{m},k_{m+1}]}$ in the cyclic sum if possible, while (6.5b) implies that it is never better to only add a partial component of $Y_{[\ell_{m},k_{m+1}]}$.
-Therefore, for $m\in\llbracket 1,s\rrbracket$, we have that $\max\limits_{j\in\mathbb{Z}_{N}}Y_{(k_{m},j]}=Y_{(k_{m},\ell_{m-1}-1]}$, and by Lemma 6.3,
-$\widetilde{D}_{k_{m}+1}(\mathbf{u}_{1},\mathbf{X}_{2})$ $=u_{1,k_{m}+1}+\max(0,\theta_{2}-\theta_{1}-\max\limits_{j\in\mathbb{Z}_{N}}Y_{(k_{m},j]})$
-$=u_{1,k_{m}+1}+\max(0,\theta_{2}-\theta_{1}-Y_{(k_{m},\ell_{m-1}-1]})$
-$=u_{1,k_{m}+1}+\max(0,Y_{[\ell_{m-1},k_{m}]})$
-$=u_{1,k_{m}+1}+Y_{[\ell_{m-1},k_{m}]}=u_{2,k_{m}+1}.$
-Moreover, for $i\in(k_{m}+1,\ell_{m}+1)$, we find that $\max\limits_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]}=Y_{(i-1,i-2]}$ and
-$\widetilde{D}_{i}(\mathbf{u}_{1},\mathbf{X}_{2})$ $=u_{1,i}+\max(0,\theta_{2}-\theta_{1}-\max\limits_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]})$
-$=u_{1,i}+\max(0,\theta_{2}-\theta_{1}-Y_{(i-1,i-2]})$
-$=u_{1,i}+\max(0,Y_{i-1})$
-$=u_{1,i}+Y_{i-1}=u_{2,i},$
-as desired. ∎
-6.3.1. Proofs of Lemmas 6.8-6.11
-[…]
-Figure 7. Illustration of the N1N-1 equalities case discussed in Lemma 6.8. Note that $k_{m}=k_{1},\ell_{m}=\ell_{1}$ for all $m$ since $s=1$.
-Proof of Lemma 6.8..
-It follows immediately that if $\ell_{0}=k_{1}+1$, then $u_{1,i}=u_{2,i}$ for all $i\neq N$.
-See Figure 7.
-Thus, $s=1$, and the condition (6.5a) is empty in this case.
-Additionally, since $i=N$ is the unique index where $u_{2,i}>u_{1,i}$, and by Lemma 6.3, we see that
-$\theta_{2}-\theta_{1}$ $=\sum\limits_{i\in\mathbb{Z}_{N}}(u_{2,i}-u_{1,i})=u_{2,N}-u_{1,N}=\theta_{2}-\theta_{1}-\max\limits_{j\in\mathbb{Z}_{N}}Y_{(N-1,j]},$
-Hence, $\max_{j\in\mathbb{Z}_{N}}Y_{(N-1,j]}=0$, so $Y_{(N-1,j]}\leq 0$ for all $j\in\mathbb{Z}_{N}$.
-Combining this result with the observation that $Y_{N}\leq 0$ by Lemma 6.6, we get (6.5b).
-Moreover, we get (6.5c) from the observation that $Y_{[\ell_{m},k_{m+1}]}=Y_{[N,N-1]}=\theta_{2}-\theta_{1}=u_{2,N}-u_{1,N}$.
-Note that (6.5d) follows since $Y_{j}=u_{2,j}-u_{1,j}\geq-u_{1,j}$.
-∎
-Proof of Lemma 6.9..
-Let $m\in\llbracket 1,s\rrbracket$. We show that show that $Y_{(k_{m}-1,j]}\leq Y_{(k_{m}-1,\ell_{m-1}-1]}$ for all $j\in\mathbb{Z}_{N}$. We observe by definition of $\ell_{m-1}$ that
-$u_{2,\ell_{m-1}}>u_{1,\ell_{m-1}},\quad\text{and}\quad u_{2,\ell_{m-1}+1}=u_{1,\ell_{m-1}+1}.$ (6.8)
-We consider three cases for the value of $j\in\mathbb{Z}_{N}$.
-  1. (1)
-    Case 2.1: $j\in[k_{m}-1,\ell_{m-1}-1)$. This case is depicted in Figure 8. By the inequality in (6.8) and Lemma 6.3, we have that
-    $Y_{(\ell_{m-1}-1,j]}\leq\max_{i\in\mathbb{Z}_{N}}Y_{(\ell_{m-1}-1,i]}<\theta_{2}-\theta_{1},$
-    so $Y_{[j+1,\ell_{m-1}-1]}=\theta_{2}-\theta_{1}-Y_{(\ell_{m-1}-1,j]}>0$ by Lemma 6.2(ii). Then, an application of Lemma 6.2(i) yields
-    $Y_{(k_{m}-1,\ell_{m-1}-1]}=Y_{(k_{m}-1,j]}+Y_{[j+1,\ell_{m-1}-1]}>Y_{(k_{m}-1,j]}.$
-[…]
-    Figure 8. Illustrations of $k_{m}-1,\ell_{m-1}-1)$ in Lemma 6.9, Case 2.1 with (a) $m\neq 1$ and (b) $m=1$ (so the interval is $k_{1}-1,\ell_{s}-1)$).
-  2. (2)
-    Case 2.2: $j=\ell_{m-1}$ where $k_{m}-1\neq\ell_{m-1}$.
-Note the case where $j=k_{m}-1$ was already handled in Case 2.1.
-By (6.8) and Lemma 6.6, we know that $Y_{\ell_{m-1}}\leq 0$, so $Y_{(k_{m}-1,\ell_{m-1}]}=Y_{(k_{m}-1,\ell_{m-1}-1]}+Y_{\ell_{m-1}}\leq Y_{(k_{m}-1,\ell_{m-1}-1]}$.
-  3. (3)
-    Case 2.3: $k_{m}-1\notin\{\ell_{m-1},\ell_{m-1}+1\}$, and $j\in(\ell_{m-1},k_{m}-1)$.
-This case is depicted in Figure 9.
-Note this case only occurs when there are at least three consecutive equalities in the string ending at index $k_{m}$.
-Assume for the sake of contradiction that there exists some $j\in(\ell_{m-1},k_{m}-1)$ such that $Y_{(k_{m}-1,j]}\geq Y_{(k_{m}-1,i]}$ for all $i\in\mathbb{Z}_{N}$ and $Y_{(k_{m}-1,j]}>Y_{(k_{m}-1,\ell_{m-1}-1]}$.
-When starting from $k_{m}$ and moving in cyclic order, take $j$ to be the first index that fulfills this condition.
-Then, by Cases 2.1 and 2.2, we have
-    $Y_{(k_{m}-1,j]}>Y_{(k_{m}-1,i]},\quad\text{for all }i\in[k_{m},j).$ (6.9)
-    On the other hand, since $Y_{(k_{m}-1,j]}\geq Y_{(k_{m}-1,i]}$ for all $i\in\mathbb{Z}_{N}$, Lemma 6.2(i) implies
-    $Y_{(j,i]}=Y_{[j+1,i]}=Y_{(k_{m}-1,i]}-Y_{(k_{m}-1,j]}\leq 0\text{ for all }i\in(j,k_{m}-1).$ (6.10)
-    Moreover, as $u_{2,k_{m}}=u_{1,k_{m}}$ by definition of $k_{m}$ and since $j$ was chosen to be a maximizer, Lemma 6.3 implies that $Y_{(k_{m}-1,j]}\geq\theta_{2}-\theta_{1}=\mathfrak{s}(\mathbf{Y})$, so by Lemma 6.2(ii),
-    $Y_{(j,k_{m}-1]}=Y_{[j+1,k_{m}-1]}\leq 0.$ (6.11)
-    Since $j\in(\ell_{m-1},k_{m}-1)$, we must have that $u_{1,j+1}=u_{2,j+1}$. Then, Lemma 6.3 implies that
-    $\max_{i\in\mathbb{Z}_{N}}Y_{(j,i]}\geq\theta_{2}-\theta_{1}>0.$ (6.12)
-    Hence, we have that
-    $\max_{i\in\mathbb{Z}_{N}}Y_{(j,i]}=\max(0,\max_{i\in\mathbb{Z}_{N}\setminus\{j\}}Y_{(j,i]})=\max_{i\in\mathbb{Z}_{N}\setminus\{j\}}Y_{(j,i]}.$ (6.13)
-    Then, using (6.13), (6.10), (6.11) and Lemma 6.2(i) in the first step below, followed by Lemma 6.2(ii), and then followed by (6.9), we obtain
-    $\max_{i\in\mathbb{Z}_{N}}Y_{(j,i]}$ $=\max_{i\in\mathbb{Z}_{N}\setminus[j,k_{m}-1]}(Y_{(j,k_{m}-1]}+Y_{[k_{m},i]})$
-    $=\theta_{2}-\theta_{1}-Y_{[k_{m},j]}+\max_{i\in\mathbb{Z}_{N}\setminus[j,k_{m}-1]}Y_{(k_{m}-1,i]}$
-    $<\theta_{2}-\theta_{1}-Y_{[k_{m},j]}+Y_{(k_{m}-1,j]}=\theta_{2}-\theta_{1},$
-    and this contradicts (6.12). ∎
-[…]
-Figure 9. Illustrations of the periodic interval $(\ell_{m-1},k_{m}-1)$ in Lemma 6.9, Case 2.3 with (a) $m\neq 1$ and (b) $m=1$
-Proof of Lemma 6.10..
-The assumption $\ell_{0}\neq k_{1}+1$ implies that $\ell_{m-1}\neq k_{m}+1$ for all $m\in\llbracket 1,s\rrbracket$ (otherwise there would be N1N-1 equalities of the form $u_{1,i}=u_{2,i}$).
-By definition of $k_{m}$, $u_{2,k_{m}}=u_{1,k_{m}}$ and $u_{2,k_{m}+1}>u_{1,k_{m}+1}$, so using Lemma 6.3, we have
-$Y_{(k_{m},\ell_{m-1}-1]}=Y_{(k_{m}-1,\ell_{m-1}-1]}-Y_{k_{m}}\geq\theta_{2}-\theta_{1}-Y_{k_{m}}=Y_{(k_{m},k_{m}-1]}.$
-Using this, followed by Lemma 6.9 in the second line below, we see that
-$\max_{j\in\mathbb{Z}_{N}}Y_{(k_{m},j]}$ $=\max_{j\in\mathbb{Z}_{N}\setminus\{k_{m}-1\}}Y_{(k_{m}-1,j]}-Y_{k_{m}}$
-$=Y_{(k_{m}-1,\ell_{m-1}-1]}-Y_{k_{m}}=Y_{(k_{m},\ell_{m-1}-1]}.\qed$
-Proof of Lemma 6.11..
-Observe that, if $\ell_{m}=k_{m}+1$, then the interval $(k_{m},\ell_{m})$ is empty, so our statement is vacuously true. From now on, we can assume that $\ell_{m}\neq k_{m}+1$.
-We now show that $Y_{(q,j]}\leq Y_{(q,q-1]}$ for all $j\in\mathbb{Z}_{N}$.
-To do so, observe that $u_{2,q}>u_{1,q}$ for $q\in(k_{m},\ell_{m})$, which implies that $\max\limits_{j\in\mathbb{Z}_{N}}Y_{(q-1,j]}<\theta_{2}-\theta_{1}$ by Lemma 6.3.
-Then, by Lemma 6.2(ii), $Y_{[j+1,q-1]}>0$ for any $j\in\mathbb{Z}_{N}$.
-Combined with Lemma 6.2(i), for $j\in[q,q-1)$ (equivalently $j\neq q-1$),
-$Y_{(q,j]}<Y_{(q,j]}+Y_{[j+1,q-1]}=Y_{(q,q-1]},$
-as desired. ∎
 Appendix A Inputs from previous work and technical inputs
 A.1. Basic algebraic properties
-We adapted the definition of the periodic Pitman transform from the work of [CGS24]
-$\mathfrak{T}_{i}(\mathbf{\mathbf{X}_{1}},\mathbf{\mathbf{X}_{2}})=X_{1,i}+\log\Biggl{(}\frac{\sum_{j\in\mathbb{Z}_{N}}e^{X_{2,[i,j]}-X_{1,[i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{X_{2,[i+1,j]-X_{1,[i+1,j]}}}}\Biggr{)}$
-$\mathfrak{D}_{i}(\mathbf{\mathbf{X}_{1}},\mathbf{\mathbf{X}_{2}})=X_{2,i}+\log\Biggl{(}\frac{\sum_{j\in\mathbb{Z}_{N}}e^{X_{2,(i,j]}-X_{1,(i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{X_{2,(i-1,j]}-X_{1,(i-1,j]}}}\Biggr{)}$
-In [CGS24], these maps were denoted as $T^{N,2}$ and $D^{N,2}$, respectively. It is important to note that this is equivalent to our definition of the map up to a shift:
+In the present work, we use a different formulation of the maps $D$ and $T$ than those introduced in [CGS24]; there, the relevant mappings were defined as
+$\mathfrak{T}_{i}(\mathbf{\mathbf{X}_{1}},\mathbf{\mathbf{X}_{2}})$ $=X_{1,i}+\log\Biggl{(}\frac{\sum_{j\in\mathbb{Z}_{N}}e^{X_{2,[i,j]}-X_{1,[i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{X_{2,[i+1,j]-X_{1,[i+1,j]}}}}\Biggr{)},\quad\text{and}$
+$\mathfrak{D}_{i}(\mathbf{\mathbf{X}_{1}},\mathbf{\mathbf{X}_{2}})$ $=X_{2,i}+\log\Biggl{(}\frac{\sum_{j\in\mathbb{Z}_{N}}e^{X_{2,(i,j]}-X_{1,(i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{X_{2,(i-1,j]}-X_{1,(i-1,j]}}}\Biggr{)}.$
+In [CGS24], these maps were denoted as $T^{N,2}$ and $D^{N,2}$, respectively. It is important to note that this is equivalent to our definition of the maps defined in (1.2) up to a shift:
 $D(\mathbf{X}_{1},\mathbf{X}_{2})=\mathfrak{D}(\mathbf{X}_{1},\tau_{1}\mathbf{X}_{2}),\quad\text{and}\quad T(\mathbf{X}_{1},\mathbf{X}_{2})=\tau_{-1}\mathfrak{T}(\mathbf{X}_{1},\tau_{1}\mathbf{X}_{2}).$ (A.1)
-where $\tau_{k}:\mathbb{R}^{\mathbb{Z}_{N}}\rightarrow\mathbb{R}^{\mathbb{Z}_{N}}$ is the shift map: $(\tau_{k}X)_{i}=X_{i+k}$ for all $i\in\mathbb{Z}_{N}$.
+where, for $k\in\mathbb{Z}_{N}$, $\tau_{k}:\mathbb{R}^{\mathbb{Z}_{N}}\rightarrow\mathbb{R}^{\mathbb{Z}_{N}}$ is the shift map: $(\tau_{k}X)_{i}=X_{i+k}$ for all $i\in\mathbb{Z}_{N}$.
 Lemma A.1.
-$\frac{1}{\beta}\log\Bigl{(}\sum\limits_{i=0}^{k}\exp(\beta E_{i})\Bigr{)}$ converges to $\max\limits_{i}E_{i}$, uniformly on the values of $E$ (but not k) as $\beta\to\infty$.
+As $\beta\to\infty$, $\frac{1}{\beta}\log\Bigl{(}\sum\limits_{i=0}^{k}\exp(\beta E_{i})\Bigr{)}$ converges to $\max\limits_{i}E_{i}$, uniformly as a function of the $E_{i}$ (but not as a function of $k$)
 Proof.
+First, observe that
 $\frac{1}{\beta}\log\Bigl{(}\sum\limits_{i=0}^{k}\exp(\beta E_{i})\Bigr{)}-\max\limits_{i}E_{i}$ $=\frac{1}{\beta}\log\Bigl{(}\sum\limits_{i=0}^{k}\exp(\beta E_{i})\Bigr{)}-\frac{1}{\beta}\log\Bigl{(}\exp(\beta\max\limits_{i}E_{i})\Bigr{)}$
-$=\frac{1}{\beta}\log\Biggl{(}\frac{\sum\limits_{i=0}^{k}\exp(\beta E_{i})}{\max\limits_{i}\exp(\beta E_{i})}\Biggr{)}$
+$=\frac{1}{\beta}\log\Biggl{(}\frac{\sum\limits_{i=0}^{k}\exp(\beta E_{i})}{\max\limits_{i}\exp(\beta E_{i})}\Biggr{)}.$
 From here, we bound below and above.
-$\frac{1}{\beta}\log\Biggl{(}\frac{\sum\limits_{i=0}^{k}\exp(\beta E_{i})}{\max\limits_{i}\exp(\beta E_{i})}\Biggr{)}$ $\leq\frac{1}{\beta}\log\Biggl{(}\frac{k\max\limits_{i}\exp(\beta E_{i})}{\max\limits_{i}\exp(\beta E_{i})}\Biggr{)}=\frac{1}{\beta}\log(k)$
+$\frac{1}{\beta}\log\Biggl{(}\frac{\sum\limits_{i=0}^{k}\exp(\beta E_{i})}{\max\limits_{i}\exp(\beta E_{i})}\Biggr{)}$ $\leq\frac{1}{\beta}\log\Biggl{(}\frac{k\max\limits_{i}\exp(\beta E_{i})}{\max\limits_{i}\exp(\beta E_{i})}\Biggr{)}=\frac{1}{\beta}\log(k).$
 Similarly,
-$\frac{1}{\beta}\log\Biggl{(}\frac{\sum\limits_{i=0}^{k}\exp(\beta E_{i})}{\max\limits_{i}\exp(\beta E_{i})}\Biggr{)}$ $\geq\frac{1}{\beta}\log\Biggl{(}\frac{\max\limits_{i}\exp(\beta E_{i})}{\max\limits_{i}\exp(\beta E_{i})}\Biggr{)}=\frac{1}{\beta}\log(1)=0$
+$\frac{1}{\beta}\log\Biggl{(}\frac{\sum\limits_{i=0}^{k}\exp(\beta E_{i})}{\max\limits_{i}\exp(\beta E_{i})}\Biggr{)}$ $\geq\frac{1}{\beta}\log\Biggl{(}\frac{\max\limits_{i}\exp(\beta E_{i})}{\max\limits_{i}\exp(\beta E_{i})}\Biggr{)}=\frac{1}{\beta}\log(1)=0.$
 This concludes our proof. ∎
 Lemma A.2.
-The maps $\mathcal{P}$ and $\widetilde{\mathcal{P}}$ define in (1.2) and (1.11) are involutions on $\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}$.
+The maps $\mathcal{P}$ and $\widetilde{\mathcal{P}}$ defined in (1.2) and (1.13) are involutions on $\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}$.
 Proof.
@@
 $=(\tau_{-1}\tau_{1}\mathbf{W}_{1},\mathbf{W}_{2})$
-$=(\mathbf{W}_{1},\mathbf{W}_{2})$
+$=(\mathbf{W}_{1},\mathbf{W}_{2}).$
 Thus, $\mathcal{P}$ is an involution. Passing to zero-temperature, we get that
@@
 $\leq 0\vee\left(X_{1,i}+\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}\right)$
-$\leq X_{1,i}+\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}$
+$\leq X_{1,i}+\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}.$
 Similarly, we have that $Y_{[i,j]}=\begin{cases}-Y_{i-1}+Y_{[i-1,j]},&\text{ if }j\in\mathbb{Z}_{N}\setminus\{i-1\}\\ Y_{[i-1,i-2]},&\text{ if }j=i-1\end{cases}$
@@
 $\leq\max_{j\in\mathbb{Z}_{N}}Y_{[i-1,j]}\vee\left(X_{2,i-1}+\max_{j\in\mathbb{Z}_{N}}Y_{[i-1,j]}\right)$
-$\leq X_{2,i-1}+\max_{j\in\mathbb{Z}_{N}}Y_{[i-1,j]}$
+$\leq X_{2,i-1}+\max_{j\in\mathbb{Z}_{N}}Y_{[i-1,j]}.$
 By the definitions in (1.13), we see that $\widetilde{T}_{i}(\mathbf{X}_{1},\mathbf{X}_{2})\geq 0$ and $\widetilde{D}_{i}(\mathbf{X}_{1},\mathbf{X}_{2})\geq 0$ for all $i\in\mathbb{Z}_{N}$. ∎
@@
   1. (i)
-    If $(\mathbf{U}_{1},\mathbf{U}_{2})=\mathcal{P}(\mathbf{X}_{1},\mathbf{X}_{2})$, then $\mathcal{P}(\tau_{1}\mathbf{X}_{1},\tau_{1}\mathbf{X}_{2})=(\tau_{1}\mathbf{U}_{1},\tau_{1}\mathbf{U}_{2})$
+    If $(\mathbf{U}_{1},\mathbf{U}_{2})=\mathcal{P}(\mathbf{X}_{1},\mathbf{X}_{2})$, then $\mathcal{P}(\tau_{1}\mathbf{X}_{1},\tau_{1}\mathbf{X}_{2})=(\tau_{1}\mathbf{U}_{1},\tau_{1}\mathbf{U}_{2})$.
   2. (ii)
-    If $(\mathbf{U}_{1},\mathbf{U}_{2})=\widetilde{\mathcal{P}}(\mathbf{X}_{1},\mathbf{X}_{2})$, then $\widetilde{\mathcal{P}}(\tau_{1}\mathbf{X}_{1},\tau_{1}\mathbf{X}_{2})=(\tau_{1}\mathbf{U}_{1},\tau_{1}\mathbf{U}_{2})$
+    If $(\mathbf{U}_{1},\mathbf{U}_{2})=\widetilde{\mathcal{P}}(\mathbf{X}_{1},\mathbf{X}_{2})$, then $\widetilde{\mathcal{P}}(\tau_{1}\mathbf{X}_{1},\tau_{1}\mathbf{X}_{2})=(\tau_{1}\mathbf{U}_{1},\tau_{1}\mathbf{U}_{2})$.
 Lemma A.7.
@@
 For $\mathbf{X}_{1},\mathbf{X}_{2},\mathbf{X}_{3}\in\mathbb{R}^{\mathbb{Z}_{N}}$,
-$D(\mathbf{X}_{1},D(\mathbf{X}_{2},\mathbf{X}_{3}))=D\bigl{(}D(\mathbf{X}_{1},\mathbf{X}_{2}),D(T(\mathbf{X}_{1},\mathbf{X}_{2}),\mathbf{X}_{3})$
+$D(\mathbf{X}_{1},D(\mathbf{X}_{2},\mathbf{X}_{3}))=D\bigl{(}D(\mathbf{X}_{1},\mathbf{X}_{2}),D(T(\mathbf{X}_{1},\mathbf{X}_{2}),\mathbf{X}_{3}).$
 Proof.
@@
 as desired. ∎
+Appendix B Proofs of the zero-temperature analogues in Theorems 1.10-1.12
+Because polymer models have a strong connection to last-passage percolation (LPP), our algebraic results will naturally generalize to zero-temperature equivalents, through limiting arguments.
+In these limits, the $(+,\times)$ algebra is swapped for the $(\max,+)$ algebra, and our maps reflect this change.
+As introduced in [CGS24], the ‘correct’ map to look at here is $\widetilde{D}$, the uniform limit as $\beta$ goes to infinity of $\frac{1}{\beta}D(\beta\mathbf{X}_{1},\beta\mathbf{X}_{2})$, as well as $\widetilde{T}$, which is reached through an identical limit.
+In this section, we use our algebraic results about the behavior of the polymer model under the periodic Pitman transform (as well as the aforementioned limiting arguments) to prove analogous facts about LPP under the zero-temperature periodic Pitman transform.
+The following intermediate result is needed.
+Lemma B.1.
+For any $k\in\mathbb{N}$ and any pair of $k$-tuples of ordered pairs $U=\bigl{(}(a_{i},b_{i})\bigr{)}_{i\in\llbracket 1,k\rrbracket}$, $V=\bigl{(}(c_{i},d_{i})\bigr{)}_{i\in\llbracket 1,k\rrbracket}$, as $\beta\to\infty$, $\frac{1}{\beta}\log\bigl{(}Z^{\beta,\mathbf{W}}(V\mid U)\bigr{)}$ converges uniformly (as a function of $\mathbf{W}$) to $G^{\mathbf{W}}(V\mid U)$.
+Proof.
+This follows as a direct consequence of Lemma A.1 since there are finitely many multi-paths between any two sets of points. ∎
+Note that in the special case where $U,V$ are points, this shows that the single-path polymer partition function converges to single-path LPP.
+Also, because the convergence to LPP is uniform on the weights, we know that the zero-temperature limit is well-behaved up to weights which depend on $\beta$.
+If the weights $\mathbf{W}_{\beta}$ converge to $\mathbf{\widetilde{\mathcal{W}}}$, then $\frac{1}{\beta}\log(Z^{\beta,\mathbf{W}_{\beta}}(V\mid U))$ will converge to $G^{\mathbf{\widetilde{\mathcal{W}}}}(V\mid U)$.
+With this in mind, we can now prove Theorem 1.10.
+Proof of Theorem 1.10..
+As in Theorem 1.4, we take $k_{1},k_{2},\ldots,k_{m}\in\mathbb{Z},U=\bigl{(}a_{i},b_{i})_{i\in\llbracket 1,m\rrbracket}\bigr{)}$, and $V=\bigl{(}(c_{i},d_{i})_{i\in\llbracket 1,m\rrbracket}\bigr{)}$ satisfying $(U,V)\in\Psi$ and, for all $i,j\in\llbracket 1,n\rrbracket$, $a_{i}\neq k_{j}+1$ and $c_{i}\neq k_{j}$.
+To prove Theorem 1.10, we must show that $G^{\mathbf{W}}(V\mid U)=G^{\widetilde{\mathcal{P}}_{k_{n}}\ldots\widetilde{\mathcal{P}}_{k_{1}}\mathbf{W}}(V\mid U)$, where the operators $\mathcal{P}_{k}$ are defined in 1.9, with $T,D$ replaced by $\widetilde{T},\widetilde{D}$.
+Now, for $\beta>0$, let $\mathcal{P}_{k}^{\beta}:(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}\to(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$ be operators defined by their action:
+$\quad\,\mathcal{P}_{k}^{\beta}(\ldots,\mathbf{W}_{k-1},\mathbf{W}_{k},\mathbf{W}_{k+1},\mathbf{W}_{k+2},\ldots)$ (B.1)
+$=(\ldots,\mathbf{W}_{k-1},\frac{1}{\beta}T(\beta\mathbf{W}_{k+1},\beta\mathbf{W}_{k}),\frac{1}{\beta}D(\beta\mathbf{W}_{k+1},\beta\mathbf{W}_{k}),\mathbf{W}_{k+2},\ldots).$
+By Theorem 1.4, we know that $Z^{\mathbf{W}}(V\mid U)=Z^{\mathcal{P}_{k_{m}}\ldots\mathcal{P}_{k_{1}}\mathbf{W}}(V\mid U)$.
+Then, by definition of the polymer model with inverse temperature $\beta$ (1.8), for all $\beta>0$,
+$Z^{\beta,\mathbf{W}}(V\mid U)=Z^{\beta\mathbf{W}}(V\mid U)=Z^{\mathcal{P}_{k_{1}}(\beta\mathbf{W})}(V\mid U)=Z^{\beta(\frac{1}{\beta}\mathcal{P}_{k_{1}}(\beta\mathbf{W}))}(V\mid U)=Z^{\beta,\mathcal{P}_{k_{1}}^{\beta}(\mathbf{W})}(V\mid U).$ (B.2)
+By induction, we have $Z^{\beta,\mathbf{W}}(V\mid U)=Z^{\beta,\mathcal{P}_{k_{n}}^{\beta}\ldots\mathcal{P}_{k_{1}}^{\beta}\mathbf{W}}(V\mid U)$.
+By Lemma B, we know that $\frac{1}{\beta}\log\bigl{(}Z^{\beta,\mathbf{W}}(V\mid U)\bigr{)}$ converges to $G^{\mathbf{W}}(V\mid U)$ uniformly on the weights, so $G^{\mathbf{W}}(V\mid U)$ is preserved under the limiting operator $\lim\limits_{\beta\to\infty}\mathcal{P}_{k_{n}}^{\beta}\ldots\mathcal{P}_{k_{1}}^{\beta}$.
+We claim that each $\mathcal{P}_{k}^{\beta}$ converges to $\widetilde{\mathcal{P}}_{k}$ uniformly as a function of the weights, which implies by induction that
+$G^{\mathbf{W}}(V\mid U)=G^{\lim\limits_{\beta\to\infty}\mathcal{P}_{k_{m}}^{\beta}\ldots\mathcal{P}_{k_{1}}^{\beta}\mathbf{W}}(V\mid U)=G^{\widetilde{\mathcal{P}}_{k_{m}}\ldots\widetilde{\mathcal{P}}_{k_{1}}\mathbf{W}}(V\mid U),$ (B.3)
+proving the theorem. The necessary uniform convergence is shown component-wise:
+$\quad\,\frac{1}{\beta}D_{i}(\beta\mathbf{X}_{1},\beta\mathbf{X}_{2})-\widetilde{D}_{i}(\mathbf{X}_{1},\mathbf{X}_{2})$
+$=X_{2,i+1}+\frac{1}{\beta}\log\Biggl{(}\frac{\sum_{j\in\mathbb{Z}_{N}}e^{\beta Y_{(i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{\beta Y_{(i-1,j]}}}\Biggr{)}-(X_{2,i+1}+\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}-\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]})$
+$=\Bigl{(}\frac{1}{\beta}\log\bigl{(}\sum_{j\in\mathbb{Z}_{N}}e^{\beta Y_{(i,j]}}\bigr{)}-\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}\Bigr{)}-\Bigl{(}\frac{1}{\beta}\log\bigl{(}\sum_{j\in\mathbb{Z}_{N}}e^{\beta Y_{(i-1,j]}}\bigr{)}-\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]}\Bigr{)}.$
+By Lemma A.1, we know that $\frac{1}{\beta}\log(N)\leq\frac{1}{\beta}D_{i}(\beta\mathbf{X}_{1},\beta\mathbf{X}_{2})-\widetilde{D}_{i}(\mathbf{X}_{1},\mathbf{X}_{2})\leq\frac{1}{\beta}\log(N)$, proving uniform convergence.
+An identical argument holds for $T_{i}-\widetilde{T}_{i}$.
+∎
+We next show the braid relations pass to zero-temperature.
+Proof of Theorem 1.11..
+Defining $\mathcal{P}_{k}^{\beta}$ as in (B.1), we see that $\mathcal{P}_{k}^{\beta}\mathcal{P}_{k+1}^{\beta}\mathcal{P}_{k}^{\beta}=\mathcal{P}_{k+1}^{\beta}\mathcal{P}_{k}^{\beta}\mathcal{P}_{k+1}^{\beta}$, as an immediate consequence of Theorem 1.6.
+As noted in the proof of Theorem 1.10, $\mathcal{P}_{k}^{\beta}$ converges to $\widetilde{\mathcal{P}}_{k}$ uniformly.
+This implies the braid relation for $\widetilde{\mathcal{P}}_{k}$, since we have
+$\widetilde{\mathcal{P}}_{k}\widetilde{\mathcal{P}}_{k+1}\widetilde{\mathcal{P}}_{k}=\lim\limits_{\beta\to\infty}\mathcal{P}_{k}^{\beta}\mathcal{P}_{k+1}^{\beta}\mathcal{P}_{k}^{\beta}=\lim\limits_{\beta\to\infty}\mathcal{P}_{k+1}^{\beta}\mathcal{P}_{k}^{\beta}\mathcal{P}_{k+1}^{\beta}=\widetilde{\mathcal{P}}_{k+1}\widetilde{\mathcal{P}}_{k}\widetilde{\mathcal{P}}_{k+1}.\qed$
+To prove Theorem 1.12, we first need analogous Burke properties in the geometric and exponential cases.
+Proposition B.2 (Geometric Burke property).
+Take $\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{Z}_{\geq 0}^{\mathbb{Z}_{N}}$ to be independent sequences of independent random variables such that $X_{1,i}\sim\text{Geom}(p_{1}q_{i})$ and $X_{2,i}\sim\text{Geom}(p_{2}q_{i})$ for all $i\in\mathbb{Z}_{N}$.
+Then,
+$(\mathbf{X}_{1},\mathbf{X}_{2})\overset{d}{=}(\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2}),\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})).$
+Proof.
+Letting $\mathbf{k}_{1},\mathbf{k}_{2}\in\mathbb{Z}_{\geq 0}^{N}$ where $k_{m,i}$ is the $i$th component of $\mathbf{k}_{m}$, we have that
+$\mathbb{P}(\mathbf{X}_{1}=\mathbf{k}_{1},\mathbf{X}_{2}=\mathbf{k}_{2})$ $=\mathbb{P}\left(X_{1,1}=k_{1,1}\right)\cdots\mathbb{P}\left(X_{1,N}=k_{1,N}\right)\mathbb{P}\left(X_{2,1}=k_{2,1}\right)\cdots\mathbb{P}\left(X_{2,N}=k_{2,N}\right)$
+$=\left(\prod\limits_{i=1}^{N}(1-p_{1}q_{i})(1-p_{2}q_{i})\right)p_{1}^{\mathfrak{s}(\mathbf{k}_{1})}p_{2}^{\mathfrak{s}(\mathbf{k}_{2})}\prod_{i=1}^{N}q_{i}^{k_{1,i}+k_{2,i}}$
+$=\left(\prod\limits_{i=1}^{N}(1-p_{1}q_{i})(1-p_{2}q_{i})\right)p_{1}^{\mathfrak{s}(\widetilde{D}(\mathbf{k}_{2},\mathbf{k}_{1}))}p_{2}^{\mathfrak{s}(\widetilde{T}(\mathbf{k}_{2},\mathbf{k}_{1}))}\prod_{i=1}^{N}q_{i}^{\widetilde{D}_{i}(\mathbf{k}_{2},\mathbf{k}_{1})+\widetilde{T}_{i}(\mathbf{k}_{2},\mathbf{k}_{1})}$
+$=\prod\limits_{i=1}^{N}(1-p_{1}q_{i})(p_{1}q_{i})^{\widetilde{D}_{i}(\mathbf{k}_{2},\mathbf{k}_{1})}\prod\limits_{i=1}^{N}(1-p_{2}q_{i})(p_{2}q_{i})^{\widetilde{T}_{i}(\mathbf{k}_{2},\mathbf{k}_{1})}$
+$=\mathbb{P}(\mathbf{X}_{1}=\widetilde{D}(\mathbf{k}_{2},\mathbf{k}_{1}),\mathbf{X}_{2}=\widetilde{T}(\mathbf{k}_{2},\mathbf{k}_{1}))$
+$=\mathbb{P}(\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})=\mathbf{k}_{2},\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2})=\mathbf{k}_{1})$
+where the first equality follows from the independence of $\mathbf{X}_{1},\mathbf{X}_{2}$ and $X_{m,i},X_{m,j}$ for $m\in\{1,2\}$ and $i,j\in\mathbb{Z}_{N}$ where $i\neq j$; the second equality follows from our definition of the geometric distribution; the third equality follows from Lemmas A.4(iii)-(iv) and A.5(i); and the final equality follows from the bijection in Lemma A.3.
+∎
+Proposition B.3 (Exponential Burke property).
+Let $\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{Z}_{\geq 0}^{\mathbb{Z}_{N}}$ be independent sequences of independent random variables such that $X_{1,i}\sim\text{Exp}(a_{1}+b_{i})$ and $X_{2,i}\sim\text{Exp}(a_{2}+b_{i})$ for all $i\in\mathbb{Z}_{N}$ where $a_{m}+b_{i}>0$ for $m\in\{1,2\}$ and $i\in\mathbb{Z}_{N}$.
+Then,
+$(\mathbf{X}_{1},\mathbf{X}_{2})\overset{d}{=}(\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2}),\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})).$
+Proof.
+For $m\in\{1,2\}$ and $i\in\mathbb{Z}_{N}$, define
+$X_{m,i}^{\prime}\sim\text{Geom}\Biggl{(}\left(1-\frac{a_{m}}{n}\right)\left(1-\frac{b_{i}}{n}\right)\Biggr{)}.$
+Then, since $\left(1-\frac{a_{m}}{n}\right)\left(1-\frac{b_{i}}{n}\right)=1-\frac{a_{m}+b_{i}}{n}+O(n^{-2})$, we have that
+$\lim\limits_{n\to\infty}\frac{X_{m,i}^{\prime}}{n}\overset{d}{=}X_{m,i}\sim\text{Exp}(a_{m}+b_{i}),\quad\text{for }m\in\{1,2\},\text{ and }i\in\mathbb{Z}_{N}$
+Observe from the definitions of $\widetilde{T}$ and $\widetilde{D}$ (1.13) that
+$\frac{1}{n}\widetilde{T}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime})=\widetilde{T}\Bigl{(}\frac{\mathbf{X}_{1}^{\prime}}{n},\frac{\mathbf{X}_{2}^{\prime}}{n}\Bigr{)},\quad\text{and}\quad\frac{1}{n}\widetilde{D}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime})=\widetilde{D}\Bigl{(}\frac{\mathbf{X}_{1}^{\prime}}{n},\frac{\mathbf{X}_{2}^{\prime}}{n}\Bigr{)}.$
+By Proposition B.2, we know that $(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime})\overset{d}{=}(\widetilde{T}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime}),\widetilde{D}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime}))$.
+Since $\max$, addition, and subtraction are continuous, $\widetilde{D}$ and $\widetilde{T}$ are continuous maps.
+We then have that
+$\Bigl{(}\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2}),\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})\Bigr{)}\overset{d}{=}\lim\limits_{n\to\infty}\frac{1}{n}\Bigl{(}\widetilde{T}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime}),\widetilde{D}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime})\Bigr{)}\overset{d}{=}\lim\limits_{n\to\infty}\frac{1}{n}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime})\overset{d}{=}(\mathbf{X}_{1},\mathbf{X}_{2}).\qed$
+Proof of Theorem 1.12.
+The proof is the same as the proof of Theorems 1.7 and 1.8: replace the use of Theorem 1.4 with Theorem 1.10 and replace the Burke property in Propositions 4.1 with the Burke property in Propositions B.2 and B.3.
+∎

### v3

@@
 We develop the theory of the discrete periodic Pitman transform, first introduced by Corwin, Gu, and the fifth author.
+We prove that the discrete periodic Pitman transform satisfies the same braid relations that are satisfied for the full-line Pitman transform shown by Biane, Bougerol, and O’Connell.
+This defines a group action of the infinite symmetric group on sequences of vectors in $\mathbb{R}^{\mathbb{Z}_{N}}$.
 We prove that, for polymers in a periodic environment, single-path and multi-path partition functions are preserved under the action of this transform on the weights in the polymer model.
-As a corollary, we prove that the discrete periodic Pitman transform satisfies the same braid relations that are satisfied for the full-line Pitman transform shown by Biane, Bougerol, and O’Connell.
 Combined with a new inhomogeneous Burke property for the periodic Pitman transform, we prove a multi-path invariance result for the periodic inverse-gamma polymer under permutations of the column parameters.
@@
 In the semi-discrete case, the positive-temperature lifting of the full-line Pitman transform was introduced by [MY01, O’C03] in the context of Brownian queuing process and later by [O’C12] in the setting of semi-discrete directed polymers.
-In the fully discrete case, the discrete geometric Pitman transform was connected to the geometric lifting of the classical Robinson-Schensted-Knuth (RSK) correspondence [Kir01] and studied by [NY04, COSZ14, OSZ14, Cor21] in the context of discrete directed polymer models.
-Despite arising in different contexts, these Pitman transforms share a rich set of both algebraic and probabilistic properties that have enabled significant progress in the study of stochastic integrable systems, including directed percolation, random polymers, the KPZ equation, and so on.
-For example, the Burke property–the analogous product measure preservation property of Burke’s theorem for $M/M/1$ queues [Bur56]–has been proved in several contexts (see [HB76, OY01, OY02, KOR02, Mar09, DMO05, AD95] as well as [EJS24, BEM+25a] in an inhomogeneous setting) and has been used to determine the fluctuation exponents in several last-passage and polymer settings [CG06, BCS06, SV10, Sep12, EJS23, LNS23].
-Moreover, the joint invariance of last-passage values/polymer partition functions under the zero-temperature/positive-temperature Pitman transforms provides key tools to the construction of central limiting objects in the KPZ universality class [DOV22] and for proving convergence of many different models to these limiting objects [DV21, Wu23, ACH24, Zha25].
-The purpose of this paper is to demonstrate the rich algebraic and probabilistic structure of a periodic version of the Pitman transform, recently introduced in [CGS24].
-Our first main result, Theorem 1.4, gives an invariance result for directed polymers in a periodic environment under a periodic Pitman transformation of the column weights.
+In the fully discrete case, the discrete geometric Pitman transform was connected to the geometric lifting of the classical Robinson-Schensted-Knuth (RSK) correspondence [KIR01] and was studied in [NY04, COS+14, OSZ14, COR02] in the context of discrete directed polymer models.
+Despite arising in different contexts, these Pitman transforms share a rich set of algebraic and probabilistic properties that have enabled significant progress in the study of stochastic integrable systems, including directed percolation, random polymers, and the KPZ equation.
+For example, the Burke property–the analogous product measure preservation property of Burke’s theorem for $M/M/1$ queues [BUR56]–has been proved in several contexts (see [HB76, OY01, OY02, KOR02, MAR09, DMO05, AD95] as well as [EJS25, BEM+25] in an inhomogeneous setting) and has been used to compute the fluctuation exponents for several last-passage percolation and polymer models [CG06, BCS06, SV10, SEP12, EJS23, LNS23].
+Moreover, the joint invariance of last-passage values and polymer partition functions under the zero-temperature and positive-temperature Pitman transforms provides key tools for constructing central limiting objects in the KPZ universality class [DOV22b] and for proving convergence of many different models to these limiting objects [DV21, WU23, ACH24, ZHA25].
+The purpose of this paper is to demonstrate the rich algebraic and probabilistic structure of a periodic version of the Pitman transform, recently introduced in [CGS26].
+Our first main result, Theorem 1.1 proves that the periodic Pitman transform is an involution and proves braid relations for these transforms, analogous to the relations enjoyed by the full-line Pitman transform [BBO05] (see also [DNV22a, DV24]).
+More specifically, we consider a collection of operators on the space $(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$ of bi-infinite sequences of $N$-periodic vectors and operators $\mathcal{P}_{k}$, which replace the $k$th and k+1k+1st vector with their Pitman transform.
+These relations show that this collection of operators generates a copy of the infinite symmetric group, which defines a group action on the space of infinite sequences of vectors in $\mathbb{R}^{\mathbb{Z}_{N}}$.
+Our next main result, Theorem 1.4, shows that single-path and multi-path partition functions for directed polymers in a periodic environment are preserved under this group action on the columns of weights for the polymer.
 Previous results of this flavor in a non-periodic setting have appeared in several contexts and are discussed after the statement of the theorem.
 The key machinery is the encoding of geometric RSK in terms of matrix relations from [NY04].
-Corwin [Cor21] previously used these tools to prove the analogous invariance for full-space polymers.
-The next main result, Theorem 1.6, uses this polymer invariance result to prove braid relations for the periodic Pitman transform, analogous to the relations enjoyed by the full-line Pitman transform [BBO05].
-More specifically, we consider a collection of operators on the space $(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$ of bi-infinite sequences of $N$-periodic vectors and operators $\mathcal{P}_{k}$, which replace the $k$th and k+1k+1st vector with their Pitman transform.
-The braid relations, along with the fact that each $\mathcal{P}_{k}$ is an involution (proved in [CGS24], see Lemma A.2) shows that this collection of operators generates a copy of the infinite symmetric group.
-We then utilize the invariance of partition functions and a new inhomogeneous Burke property (Proposition 4.1) to prove a distributional invariance of the inverse-gamma polymer partition function under the permutation of the row and column parameters in both the periodic and full-line cases (Theorems 1.7 and 1.8).
-In the full-line case, this gives a multi-path and positive temperature extension of a recent result of Bates, Emrah, Martin, Seppäläinen, and the fifth author [BEM+25b].
+Corwin [COR02] previously used these tools to prove the analogous invariance for full-space polymers.
+We then utilize the invariance of partition functions and a new inhomogeneous Burke property (Proposition 4.1) to prove a distributional invariance of the inverse-gamma polymer partition function under the permutation of the row and column parameters in both the periodic and full-space cases (Theorems 1.5 and 1.6).
+In the full-line case, this gives a multi-path and positive-temperature extension of a recent result of Bates, Emrah, Martin, Seppäläinen, and the fifth author [BEM+25].
 This result bears similarity to an invariance found previously by Dauvergne [DAU22].
@@
 Each of our main results has an analogue in zero-temperature.
-These results are stated in Section 1.6.
+These are stated in Section 1.5.
 1.1. The positive-temperature discrete periodic Pitman transform
-For $N\in\mathbb{N}$, let $\mathbb{R}^{\mathbb{Z}_{N}}$ denote the set of $N$-tuples, together with periodic indexing.
-That is, for $\mathbf{X}\in\mathbb{R}^{\mathbb{Z}_{N}}$, $\mathbf{X}=(X_{1},X_{2},\ldots,X_{N})$, where the indices represent elements of $\mathbb{Z}_{N}$ with addition modulo $N$ (see Figure 1).
+For $N\in\mathbb{N}$, let $\mathbb{R}^{\mathbb{Z}_{N}}$ denote the set of $N$-tuples of real numbers, together with periodic indexing.
+That is, for $\mathbf{X}\in\mathbb{R}^{\mathbb{Z}_{N}}$, we may write $\mathbf{X}=(X_{1},X_{2},\ldots,X_{N})$, where the indices represent elements of $\mathbb{Z}_{N}$ with addition modulo $N$.
 For a vector $\mathbf{X}\in\mathbb{R}^{\mathbb{Z}_{N}}$ and indices $i,j\in\mathbb{Z}_{N}$, we set
 $\mathbf{X}_{(i,j]}:=\sum_{\ell=i+1}^{j}X_{\ell},\quad\text{and}\quad\mathbf{X}_{[i,j]}:=\sum_{\ell=i}^{j}X_{\ell},$ (1.1)
-where the sums are taken in cyclic order. We follow the convention that $\mathbf{X}_{(i,i]}=0$.
+where the sums are taken in cyclic order (see Figure 1). We follow the convention that $\mathbf{X}_{(i,i]}=0$.
 j][i
@@
 The positive-temperature discrete periodic Pitman transform is the map $\mathcal{P}:\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}\to\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}$ defined by
-$\mathcal{P}(\mathbf{W}_{1},\mathbf{W}_{2})$ $=(T(\mathbf{W}_{2},\mathbf{W}_{1}),D(\mathbf{W}_{2},\mathbf{W}_{1})),\quad\text{where}$ (1.2)
-$T(\mathbf{X}_{1},\mathbf{X}_{2})$ $=X_{1,i-1}+\log\Biggl{(}\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{[i-1,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{[i,j]}}}\Biggr{)},\quad\text{and}$
-$D(\mathbf{X}_{1},\mathbf{X}_{2})$ $=X_{2,i+1}+\log\Biggl{(}\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{(i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{(i-1,j]}}}\Biggr{)},\quad\text{where}\quad Y_{\ell}=X_{2,\ell+1}-X_{1,\ell}.$
-This map is a shifted version of the map studied in [CGS24]; see Appendix A for the precise connection.
-Note that we reverse the order of the inputs $\mathbf{W}_{1},\mathbf{W}_{2}$ in the application of $T$ and $D$.
-Also, note that the maps $\mathcal{P},T,D$ depend on the length $N$ of the periodicity; we will subsume dependence on this value in the notation and fix $N$, unless otherwise noted.
+$\mathcal{P}(\mathbf{X}_{1},\mathbf{X}_{2})$ $=(T(\mathbf{X}_{1},\mathbf{X}_{2}),D(\mathbf{X}_{1},\mathbf{X}_{2})),\quad\text{where}$ (1.2)
+$T_{i}(\mathbf{X}_{1},\mathbf{X}_{2})$ $=X_{2,i-1}+\log\Biggl(\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{[i-1,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{[i,j]}}}\Biggr),\quad\text{and}$
+$D_{i}(\mathbf{X}_{1},\mathbf{X}_{2})$ $=X_{1,i+1}+\log\Biggl(\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{(i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{(i-1,j]}}}\Biggr),\quad\text{where}\quad Y_{\ell}=X_{1,\ell+1}-X_{2,\ell}.$
+This map is a shifted version of the map studied in [CGS26] (see Appendix A.1).
+Note that the maps $\mathcal{P},T,D$ depend on the length $N$ of the periodicity; we will subsume dependence on this value in the notation and fix the positive integer $N$, unless otherwise noted.
+Let $(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$ be the set of bi-infinite sequences of elements of $\mathbb{R}^{\mathbb{Z}_{N}}$.
+For $k\in\mathbb{Z}$, we define the operator $\mathcal{P}_{k}:(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}\to(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$ by its action:
+$\mathcal{P}_{k}(\ldots,\mathbf{X}_{k-1},\mathbf{X}_{k},\mathbf{X}_{k+1},\mathbf{X}_{k+2},\ldots)=(\ldots,\mathbf{X}_{k-1},T(\mathbf{X}_{k},\mathbf{X}_{k+1}),D(\mathbf{X}_{k},\mathbf{X}_{k+1}),\mathbf{X}_{k+2},\ldots).$ (1.3)
+Our first main result is the following:
+Theorem 1.1.
+The operators $\mathcal{P}_{k}$ satisfy the following:
+  1. (i)
+    For $k\in\mathbb{Z}$, the map $\mathcal{P}_{k}$ is an involution.
+  2. (ii)
+    For $k\in\mathbb{Z}$,
+    $\mathcal{P}_{k}\mathcal{P}_{k+1}\mathcal{P}_{k}=\mathcal{P}_{k+1}\mathcal{P}_{k}\mathcal{P}_{k+1}.$
+  3. (iii)
+    For |kj|>1|k-j|>1,
+    $\mathcal{P}_{j}\mathcal{P}_{k}=\mathcal{P}_{k}\mathcal{P}_{j}.$
+The commutativity in Item (iii) of Theorem 1.1 is immediate from the definition.
+Item (i) was previously proved in [CGS26, Corollary 5.9], although it is stated slightly differently there for a shifted version of the operator $\mathcal{P}$ (see Appendix A.1).
+In the present paper, we give an alternative, self-contained proof of Item (i) in Section 2.2.
+Item (ii) is called the braid relation, and it is also proved in Section 2.2.
+Crucial to the proof is a matrix relation in Proposition 2.7, which is also used to prove Theorem 1.4.
+This matrix relation comes from adapting the matrix representation of geometric RSK from the work of Noumi and Yamada [NY04] to the periodic case.
+The braid relation consists of three equalities ((2.11)-(2.13)), the third of which was previously proved in [CGS26, Proposition 5.13].
+From there, we iterate the matrix relation in Proposition 2.7 in two different ways and use an interpretation of the matrix product as a polymer partition function (lemma 3.1) to obtain the other two equalities.
+Specifically, Lemma 2.8 is the key innovation that allows us to complete the proof.
+We call a permutation $\sigma:\mathbb{Z}\to\mathbb{Z}$ a finite permutation if $\sigma$ fixes all but finitely many coordinates.
+Let $\mathcal{S}_{\mathbb{Z}}$ denote the group of these finite permutations under composition of functions.
+The relations satisfied by the maps $\mathcal{P}_{k}$ in Theorem 1.1 are the same as those satisfied by the nearest-neighbor transpositions $(k,k+1)$, which generate the group $\mathcal{S}_{\mathbb{Z}}$.
+Thus, if $\sigma$ is any such finite permutation, if we decompose it in two different ways as
+$\sigma=(k_{1},k_{1}+1)(k_{2},k_{2}+1)\cdots(k_{n},k_{n}+1)=(\ell_{1},\ell_{1}+1)(\ell_{2},\ell_{2}+1)\cdots(\ell_{r},\ell_{r}+1),$
+then Theorem 1.1 implies that we have the following equality of operators:
+$\mathcal{P}_{k_{1}}\mathcal{P}_{k_{2}}\cdots\mathcal{P}_{k_{n}}=\mathcal{P}_{\ell_{1}}\mathcal{P}_{\ell_{2}}\cdots\mathcal{P}_{\ell_{r}}.$
+For any finite permutation $\sigma$, we then define the operator $\mathcal{P}_{\sigma}$ by the product above.
+The operators $\mathcal{P}_{\sigma}$ therefore define a group action of $\mathcal{S}_{\mathbb{Z}}$ on the space $(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$.
 1.2. Polymers on $\mathbb{Z}^{2}$ and periodic environments
-Definition 1.1.
 Consider the 2-dimensional integer lattice $\mathbb{Z}^{2}$. We say that an up-right path of length $L\in\mathbb{N}$ is an ordered collection of points
-$\pi:=\left\{(a_{i},b_{i})_{i\in\llbracket 1,L\rrbracket}\mid(a_{i+1},b_{i+1})-(a_{i},b_{i})\in\{(1,0),(0,1)\}\ \forall 1\leq i<L\right\}$ (1.3)
-where $(a_{i},b_{i})\in\mathbb{Z}^{2}\ \forall i$ and the notation $\llbracket a,b\rrbracket$ denotes the set of $k\in\mathbb{Z}$ satisfying $a\leq k\leq b$.
+$\pi:=\left\{(a_{i},b_{i})_{i\in\llbracket 1,L\rrbracket}\mid(a_{i+1},b_{i+1})-(a_{i},b_{i})\in\{(1,0),(0,1)\}\ \forall 1\leq i<L\right\},$ (1.4)
+where $(a_{i},b_{i})\in\mathbb{Z}^{2}\ \forall i$.
+Here, the notation $\llbracket a,b\rrbracket$ denotes the set of $k\in\mathbb{Z}$ satisfying $a\leq k\leq b$.
 We also define the set of all up-right paths to be $\Pi_{\mathbb{Z}^{2}}$.
@@
 $\Pi_{(i,j),(n,m)}:=\{\pi\in\Pi_{\mathbb{Z}^{2}}\mid(a_{1},b_{1})=(i,j)\ ,(a_{L},b_{L})=(n,m)\}$ (1.5)
-to be the set of paths starting at $(i,j)$ and ending at $(n,m)$, noting $\Pi_{(i,j),(n,m)}\neq\varnothing\iff m\geq j$ and $n\geq i$.
+to be the set of paths starting at $(i,j)$ and ending at $(n,m)$, noting that $\Pi_{(i,j),(n,m)}\neq\varnothing$ if and only if $m\geq j$ and $n\geq i$.
 In this case, L=m+njiL=m+n-j-i for any $\pi$ in $\Pi_{(i,j),(n,m)}$.
 We say that two paths $\pi_{1},\pi_{2}\in\Pi_{\mathbb{Z}^{2}}$ are non-intersecting if $\pi_{1}\cap\pi_{2}=\varnothing$.
-Also of interest to us are sets of non-intersecting paths on our graph.
-We define a multi-path $(\pi_{1},\ldots,\pi_{k})$ to be a non-intersecting $k$-tuple of paths.
+We say that a non-intersecting $k$-tuple of paths $(\pi_{1},\ldots,\pi_{k})$ is called a multi-path.
 See Figure 2 for a visual example.
@@
 For $k\in\mathbb{N}$ and $k$-tuples of ordered pairs $U=\bigl((a_{i},b_{i})\bigr)_{i\in\llbracket 1,k\rrbracket}$, $V=\bigl((c_{i},d_{i})\bigr)_{i\in\llbracket 1,k\rrbracket}$, we define the set of multi-paths between them to be
-$\Pi_{U,V}:=\{\pi=(\pi_{1},\dots,\pi_{k})\mid\pi_{i}\in\Pi_{(a_{i},b_{i}),(c_{i},d_{i})}\ \forall i\in\llbracket 1,k\rrbracket,\ \pi_{i}\cap\pi_{j}=\varnothing\ \forall i\neq j\in\llbracket 1,k\rrbracket\}$ (1.5)
+$\Pi_{U,V}:=\{\pi=(\pi_{1},\dots,\pi_{k})\mid\pi_{i}\in\Pi_{(a_{i},b_{i}),(c_{i},d_{i})}\ \forall i\in\llbracket 1,k\rrbracket,\ \pi_{i}\cap\pi_{j}=\varnothing\ \forall i\neq j\in\llbracket 1,k\rrbracket\}.$ (1.6)
 We will often require that there is a unique pairing of initial points in $U$ and end points in $V$ such that a non-intersecting multi-path exists.
-In other words, we require that if $U$ and $V$ each have $k$ elements, then there does not exist a nontrivial permutation $\sigma$ in the $k$th-order symmetric group $S_{k}$ such that $\Pi_{U,V_{\sigma}}\neq\varnothing$ where $V_{\sigma}:=\bigl{(}(c_{\sigma(i)},d_{\sigma(i)})\bigr{)}_{i\in\llbracket 1,k\rrbracket}$.
+In other words, we require that if $U$ and $V$ each have $k$ elements, then there does not exist a nontrivial permutation $\varphi$ in the symmetric group on $k$ elements, $\mathcal{S}_{k}$, such that $\Pi_{U,V_{\varphi}}\neq\varnothing$ where $V_{\varphi}:=\bigl((c_{\varphi(i)},d_{\varphi(i)})\bigr)_{i\in\llbracket 1,k\rrbracket}$.
 We denote the pairs of $k$-tuples of ordered pairs satisfying this property as
-$\Psi:=\bigl{\{}(U,V)\mid U,V\subseteq\mathbb{Z}^{2},\ |U|=|V|,\ \Pi_{U,V}\neq\varnothing,\ \text{ and }\Pi_{U,V_{\sigma}}=\varnothing\ \forall\sigma\in S_{|U|}\setminus\{\text{Id}\}\bigr{\}}$ (1.6)
+$\Psi:=\bigl\{(U,V)\mid U,V\subseteq\mathbb{Z}^{2},\ |U|=|V|,\ \Pi_{U,V}\neq\varnothing,\ \text{ and }\Pi_{U,V_{\varphi}}=\varnothing\ \forall\varphi\in\mathcal{S}_{|U|}\setminus\{\text{Id}\}\bigr\}.$ (1.7)
+As an example, if $U=\{(0,0),(2,0)\}$ and $V=\{(2,4),(4,4)\}$ (as depicted in Figure 2), then $(U,V)\in\Psi$.
+More generally, a pair $(U,V)$ lies in $\Psi$ if and only if $\Pi_{U,V}\neq\varnothing$, the points of $U$ lie along some down-right path, and the points of $V$ lie along a down-right path.
+See Figure 3 for an example of a pair $(U,V)\notin\Psi$.
 […]
 Figure 2. Example multi-path in $\mathbb{Z}^{2}$ from $\{(0,0),(2,0)\}$ to $\{(2,4),(4,4)\}$
+[…]
+Figure 3.
+Two multipaths in $\mathbb{Z}^{2}$ from $U:=\{u_{1},u_{2}\}$ (green) to $V:=\{v_{1},v_{2}\}$ (orange), where $(U,V)\notin\Psi$.
+To see this, note that there exists a multipath from $(u_{1},u_{2})$ to $(v_{1},v_{2})$ (denoted here in blue/thick) as well as a multipath from $(u_{1},u_{2})$ to $(v_{2},v_{1})$ (denoted here in red/thin).
 Definition 1.2.
-For real-valued weights $\mathbf{W}=\bigl{(}\mathbf{W}_{i}\in\mathbb{R}^{\mathbb{Z}}\bigr{)}_{i\in\mathbb{Z}}=\bigl{(}W_{(i,j)}\in\mathbb{R}\mid i,j\in\mathbb{Z}\bigr{)}$ corresponding to the vertices of the directed lattice $\mathbb{Z}^{2}$, we define the polymer partition function with inverse temperature $\beta>0$ between the initial point $(i,j)$ and endpoint $(n,m)$ in $\mathbb{Z}^{2}$ as
-$Z^{\beta,\mathbf{W}}(n,m\mid i,j):=\sum_{\pi\in\Pi_{(i,j),(n,m)}}\prod_{(r,\ell)\in\pi}e^{\beta W_{(r,\ell)}}.$ (1.7)
+For real-valued weights $\mathbf{X}=\bigl(\mathbf{X}_{i}\in\mathbb{R}^{\mathbb{Z}}\bigr)_{i\in\mathbb{Z}}=\bigl(X_{(i,j)}\in\mathbb{R}\mid i,j\in\mathbb{Z}\bigr)$ corresponding to the vertices of the directed lattice $\mathbb{Z}^{2}$, we define the polymer partition function with inverse temperature $\beta>0$ between the initial point $(i,j)$ and end point $(n,m)$ in $\mathbb{Z}^{2}$ as
+$Z^{\beta,\mathbf{X}}(n,m\mid i,j):=\sum_{\pi\in\Pi_{(i,j),(n,m)}}\prod_{(r,\ell)\in\pi}e^{\beta X_{(r,\ell)}}.$ (1.8)
 The partition function has a natural multi-path extension. For $(U,V)\in\Psi$, we define the multi-path partition function between $U$ and $V$ to be
-$Z^{\beta,\mathbf{W}}(V\mid U):=\sum_{\pi\in\Pi_{U,V}}\prod_{(r,\ell)\in\pi}e^{\beta W_{(r,\ell)}}.$ (1.8)
+$Z^{\beta,\mathbf{X}}(V\mid U):=\sum_{\pi\in\Pi_{U,V}}\prod_{(r,\ell)\in\pi}e^{\beta X_{(r,\ell)}}.$ (1.9)
 We typically take $\beta=1$, in which case we omit the superscript $\beta$ in the notation.
@@
 1.3. Invariance of periodic polymer partition functions
-Let $(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$ be the set of bi-infinite sequences of elements of $\mathbb{R}^{\mathbb{Z}_{N}}$.
-For $k\in\mathbb{Z}$, we define the operator $\mathcal{P}_{k}:(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}\to(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$ by its action:
-$\mathcal{P}_{k}(\ldots,\mathbf{W}_{k-1},\mathbf{W}_{k},\mathbf{W}_{k+1},\mathbf{W}_{k+2},\ldots)=(\ldots,\mathbf{W}_{k-1},T(\mathbf{W}_{k+1},\mathbf{W}_{k}),D(\mathbf{W}_{k+1},\mathbf{W}_{k}),\mathbf{W}_{k+2},\ldots)$ (1.9)
-Our first main result gives the invariance of the polymer partition functions under the iterated Pitman transform defined in (1.2).
+Our next main result gives the invariance of the polymer partition functions under the iterated Pitman transform defined in (1.2).
+For $\sigma\in\mathcal{S}_{\mathbb{Z}}$, we let $\mathcal{W}^{\sigma}$ be the set of tuples $(U,V)\in\Psi$, such that
+[…]
+See Figure 4 for an example.
+Figure 4.
+An example of an pair $(U,V)\in\mathcal{W}^{\sigma}$, where $U$ and $V$ each contain 33 elements.
+Each column corresponds to a periodic vector of weights $\mathbf{X}_{k}$.
+The gray horizontal lines are used to indicate a periodic environment with $N=4$.
+In this figure, the column parameters of columns $3,5$, and 66 (blue) are permuted, and the column parameters in columns 88 and 99 (red) are permuted.
+The condition is that, for each of the three end points $(x,y)$ in $U$ (lower left), $\sigma(\mathbb{Z}_{<x})=\mathbb{Z}_{<x}$, while for each of the three end points $(x,y)$ in $V$ (upper right), $\sigma(\mathbb{Z}_{>x})=\mathbb{Z}_{>x}$.
+The multi-path partition function for this pair is preserved after changing the weights from $\mathbf{X}$ to $\sigma\mathbf{X}$, and the distributional equality in Theorem 1.5 holds jointly over all end point pairs $(U,V)\in\Psi$ satisfying this property.
 Theorem 1.4.
-Let $N\in\mathbb{N}$, and let $\mathbf{W}=(\mathbf{W}_{k}\in\mathbb{R}^{\mathbb{Z}_{N}}\mid k\in\mathbb{Z})=(W_{k,\ell}\mid k\in\mathbb{Z},\ell\in\mathbb{Z}_{N})$ be a collection of real-valued weights.
-Extend these periodically to weights on $\mathbb{Z}\times\mathbb{Z}$ by the condition $W_{k,j}=W_{k,\ell}$ whenever $j\equiv\ell\mod N$.
-If $k_{1},k_{2},\ldots,k_{m}\in\mathbb{Z}$ and the sets of initial points $U=\bigl{(}\bigl{(}a_{i},b_{i})_{i\in\llbracket 1,m\rrbracket}\bigr{)}$ and end points $V=\bigl{(}(c_{i},d_{i})_{i\in\llbracket 1,m\rrbracket}\bigr{)}$ satisfy $(U,V)\in\Psi$ and, for all $i,j\in\llbracket 1,m\rrbracket$, $a_{i}\neq k_{j}+1$ and $c_{i}\neq k_{j}$, then,
-$Z^{\mathcal{P}_{k_{m}}\ldots\mathcal{P}_{k_{1}}\mathbf{W}}(V\mid U)=Z^{\mathbf{W}}(V\mid U).$ (1.10)
-Remark 1.5.
-See Figure 3 for a visual depiction of the invariance in Theorem 1.4.
-In Theorem 1.6 below, we show braid relations for the operators $\mathcal{P}_{k}$.
-Along with the fact that each $\mathcal{P}_{k}$ is an involution (see Lemma A.2), these relations imply that, if we have the equality of products of transpositions $(k_{m},k_{m}+1)(k_{m-1},k_{m-1}+1)\cdots(k_{1},k_{1}+1)=(k_{m}^{\prime},k_{m}^{\prime}+1)(k_{m-1}^{\prime},k_{m-1}^{\prime}+1)\cdots(k_{1}^{\prime},k_{1}^{\prime}+1)$, then $\mathcal{P}_{k_{m}}\ldots\mathcal{P}_{k_{1}}\mathbf{W}=\mathcal{P}_{k_{m}}^{\prime}\ldots\mathcal{P}_{k_{1}}^{\prime}\mathbf{W}$.
-Hence, we consider the action of $\mathcal{P}_{k_{m}}\ldots\mathcal{P}_{k_{1}}$ on $\mathbf{W}$ as the action of a general permutation, independent of its choice of decomposition.
-In this case, (1.11) gives an alternate condition for the set of endpoints whose partition function is preserved.
-An example is given in Figure 4 below.
-This connection is made precise in Lemma 4.2.
-Figure 3.
-A depiction of the polymer invariance in Theorem 1.4.
-Each column is associated to a periodic vector of weights, $\mathbf{W}_{k}$.
-The horizontal lines are used to indicate a periodic environment for the polymer with $N=4$.
-In this example, we take $m=1$ and consider the action of the operator $\mathcal{P}_{3}$, which replaces the weights in columns $3$ and $4$ with $T(\mathbf{W}_{4},\mathbf{W}_{3})$ and $D(\mathbf{W}_{4},\mathbf{W}_{3})$, respectively.
-The partition function for each pair of points and the multi-path partition function for this collection of 33 pairs are both preserved under this transformation of the weights.
+Let $N\in\mathbb{N}$, and let $\mathbf{X}=(\mathbf{X}_{k}\in\mathbb{R}^{\mathbb{Z}_{N}}\mid k\in\mathbb{Z})=(X_{k,\ell}\mid k\in\mathbb{Z},\ell\in\mathbb{Z}_{N})$ be a collection of real-valued weights corresponding to the columns of the lattice.
+Extend these periodically to weights on $\mathbb{Z}\times\mathbb{Z}$ by the condition that $X_{k,j}=X_{k,\ell}$ whenever $j\equiv\ell\mod N$.
+Then, if $\sigma\in\mathcal{S}_{\mathbb{Z}}$ and $(U,V)\in\mathcal{W}^{\sigma}$,
+$Z^{\mathcal{P}_{\sigma}\mathbf{X}}(V\mid U)=Z^{\mathbf{X}}(V\mid U).$ (1.11)
 Theorem 1.4 is proved in Section 3.
@@
 The concept can be traced back to queuing theory from the works [WEB79, TW87], where it was shown that the law of the departures process from two $\cdot/M/1$ queues in series is invariant under swapping of the service rates of the two queues.
-In the polymer context, this type of result can be traced back to the work of Noumi and Yamada [NY04], who gave a description of the geometric Robinson-Schendsted-Knuth correspondence in terms of certain matrix product relations.
-Corwin [Cor21] used this to give an analogous result to our Theorem 1.4 for a full-line version of the discrete Pitman transform.
+In the polymer context, Corwin [COR02] used the work of Noumi and Yamada [NY04] to prove an analogous result to our Theorem 1.4 for a full-line version of the discrete Pitman transform.
 The proof of Theorem 1.4 likewise uses the machinery from [NY04], but we instead work with $\mathbb{Z}\times\mathbb{Z}$ periodic matrices instead of finite-dimensional matrices.
+The matrix relations in Proposition 2.7 that are key to proving Theorem 1.1(ii) are also the key to the proof of Theorem 1.4.
 Theorem 1.4 limits to a zero-temperature analogue in Theorem 1.9.
-This zero-temperature limit is the periodic analogue of invariances shown in [DOV22, DNV22, DV24, BEM+25a].
-The result in [DOV22] is particularly significant, as it was key in constructing the directed landscape, the central limiting object of the KPZ universality class.
-It would be interesting to see if Theorem 1.4 could be the starting point for constructing and proving convergence to the periodic directed landscape, but we note that the action of this transform is qualitatively different than that in [DOV22].
+This zero-temperature limit is the periodic analogue of invariances shown in [DOV22b, DNV22a, DV24, BEM+25].
+The result in [DOV22b] is particularly significant, as it was key in constructing the directed landscape–the central limiting object of the KPZ universality class.
+It would be interesting to see if Theorem 1.4 could be the starting point for constructing and proving convergence to the periodic directed landscape.
+However, we note that the action of this transform is qualitatively different than that in [DOV22b].
 Specifically, in the full-line case, the action of the iterated Pitman transforms and independent collection of Brownian motions and transforms them to the Brownian melon–a set of nonintersecting paths whose geodesic structure is more amenable to asymptotic analysis.
@@
 with the inverse-gamma distribution, the transformed weights are equal in distribution to the original weights.
-This is the Burke property shown in [CGS24, Proposition 5.12], and extended to an inhomogeneous setting in Proposition 4.1.
+This is the Burke property shown in [CGS26, Proposition 5.12], and extended to an inhomogeneous setting in Proposition 4.1.
 As a corollary, we obtain a permutation invariance result for the multi-path inhomogeneous inverse-gamma polymer below in Theorem 1.5.
-This extends the previous result of [BEM+25a, Theorem 3.1] to a positive-temperature and multi-path setting.
-In this sense, the periodic Pitman transform is more analogous to a limiting case of the full-line Pitman transform, where the initial point of a random walk is sent to -$\infty$ and the output of the transform is recentered at 0 (see [DNV22, DV24]).
+This extends the previous result of [BEM+25, Theorem 3.1] to a positive-temperature and multi-path setting.
+In this sense, the periodic Pitman transform is more analogous to a limiting case of the full-line Pitman transform, where the initial point of a random walk is sent to -$\infty$ and the output of the transform is recentered at 0 (see [DNV22a, DV24]).
 This limiting full-line setting is where we see analogous Burke properties, as in [OY01, OY02].
-1.4. Braid relations
-We now describe our next result, which gives a braid relation for the Pitman operators defined in (1.9).
-The next theorem gives a braid relation for these operators, which is the same as that satisfied for the full-line Pitman transform [BBO05].
-See also [DV24].
-Combined with the fact that the map $\mathcal{P}_{k}$ is an involution (see Lemma A.2), the operators $\mathcal{P}_{k}$ generate a representation of the infinite symmetric group (note that the same braid relations are satisfied by the nearest-neighbor transpositions $(k,k+1)$).
-Theorem 1.6.
-For $k\in\mathbb{Z}$,
-$\mathcal{P}_{k}\mathcal{P}_{k+1}\mathcal{P}_{k}=\mathcal{P}_{k+1}\mathcal{P}_{k}\mathcal{P}_{k+1}.$
-Theorem 1.6 is proved in Section 3.
-The key is the matrix relation in Corollary 2.5 below, which is also used to prove Theorem 1.4.
-When written out, the braid relation consists of three equalities ((3.1)-(3.3) below).
-The third of these equalities was previously proved in [CGS24, Proposition 5.13].
-With that as our starting point, we iterate the matrix relation in Corollary 2.5 in two different ways to obtain and use the interpretation of the matrix product as the polymer partition function to obtain the other two equalities.
-Specifically, Lemma 3.1 is the key innovation that allows us to complete the proof.
 1.4. Permutation invariance
 In this section, we consider an inhomogeneous, periodic inverse-gamma polymer.
-Let $\lambda>0$ and $\mathbf{a}\in\mathbb{R}^{\mathbb{Z}}$ and $\mathbf{b}\in\mathbb{R}^{\mathbb{Z}_{N}}$ be sequences of real numbers such that $a_{i}+b_{j}>0$ for $i\in\mathbb{Z}$ and $j\in\mathbb{Z}_{N}$.
-We take independent weights $\mathbf{W}^{\mathbf{a},\mathbf{b}}=\bigl{(}W_{(i,j)}\mid i\in\mathbb{Z},\ j\in\mathbb{Z}_{N}\bigr{)}$ such that $W_{(i,j)}\sim\text{Log-Inv-Gamma}(a_{i}+b_{j},\lambda)$ for $i\in\mathbb{Z}$ and $j\in\mathbb{Z}_{N}$ (see Definition 1.15 below).
+Let $\lambda>0$, and let and $\mathbf{a}\in\mathbb{R}^{\mathbb{Z}}$ and $\mathbf{b}\in\mathbb{R}^{\mathbb{Z}_{N}}$ be sequences of real numbers such that $a_{i}+b_{j}>0$ for $i\in\mathbb{Z}$ and $j\in\mathbb{Z}_{N}$.
+We take mutually independent weights $\mathbf{X}^{\mathbf{a},\mathbf{b}}=\bigl(X_{(i,j)}\mid i\in\mathbb{Z},\ j\in\mathbb{Z}_{N}\bigr)$ such that $X_{(i,j)}\sim\text{Log-Inv-Gamma}(a_{i}+b_{j},\lambda)$ for $i\in\mathbb{Z}$ and $j\in\mathbb{Z}_{N}$ (see Definition 1.13 below).
 Extend these periodically to weights on $\mathbb{Z}\times\mathbb{Z}$.
-We call a permutation $\sigma:\mathbb{Z}\to\mathbb{Z}$ a finite permutation if $\sigma$ fixes all but finitely many coordinates.
-For a finite permutation $\sigma$, we let $\mathcal{U}^{\sigma}$ be the set of tuples $(U,V)\in\Psi$, such that
+Theorem 1.5.
+For vectors $\mathbf{a}\in\mathbb{R}^{\mathbb{Z}}$ and $\mathbf{b}\in\mathbb{R}^{\mathbb{Z}_{N}}$ satisfying $a_{i}+b_{j}>0$ for $i\in\mathbb{Z}$ and $j\in\mathbb{Z}_{N}$, define periodic weights $\mathbf{X}^{\mathbf{a},\mathbf{b}}$ as above and extend them to $\mathbb{Z}\times\mathbb{Z}$.
+Then, for any finite permutation $\sigma\in\mathcal{S}_{\mathbb{Z}}$, we have
+$\Bigl(Z^{\mathbf{X}^{\mathbf{a},\mathbf{b}}}(V\mid U):(U,V)\in\mathcal{W}^{\sigma}\Bigr)\overset{d}{=}\Bigl(Z^{\mathbf{X}^{\sigma\mathbf{a},\mathbf{b}}}(V\mid U):(U,V)\in\mathcal{W}^{\sigma}\Bigr),$
+where $\sigma\mathbf{a}=(a_{\sigma(i)})_{i\in\mathbb{Z}}$.
+The condition (1.10) ensures that, between the pairs of points in $(U,V)$, the column rates are either permuted among themselves or are left unchanged. See Figure 4.
+We can also permute the $\mathbf{b}$ parameters as well, with doubly periodic weights $(X_{(i,j)}\mid i\in\mathbb{Z}_{M},j\in\mathbb{Z}_{N})$ for $M,N\in\mathbb{N}$, when we restrict to points that lie inside the box $\llbracket 0,M\rrbracket\times\llbracket 0,N\rrbracket$.
+In this case, the model isn’t actually periodic at all; it is simply a restriction of the full-space polymer to a finite grid.
+For finite permutations $\sigma,\tau\in\mathcal{S}_{\mathbb{Z}}$, define the set $\mathcal{W}^{\sigma,\tau}$ to be the set of tuples $(U,V)$, where $U,V$ are each sets of end point pairs in $\mathbb{Z}\times\mathbb{Z}$ satisfying the condition (1.10) for $\sigma$ as well as the condition
 […]
-Theorem 1.7.
-For vectors $\mathbf{a}\in\mathbb{R}^{\mathbb{Z}}$ and $\mathbf{b}\in\mathbb{R}^{\mathbb{Z}_{N}}$ satisfying $a_{i}+b_{j}>0$ for $i\in\mathbb{Z}$ and $j\in\mathbb{Z}_{N}$, define periodic weights $\mathbf{W}^{\mathbf{a},\mathbf{b}}$ as above and extend them to $\mathbb{Z}\times\mathbb{Z}$.
-Then, for any finite permutation $\sigma:\mathbb{Z}\to\mathbb{Z}$, we have
-$\Bigl{(}Z^{\mathbf{W}^{\mathbf{a},\mathbf{b}}}(V\mid U):(U,V)\in\mathcal{U}^{\sigma}\Bigr{)}\overset{d}{=}\Bigl{(}Z^{\mathbf{W}^{\sigma\mathbf{a},\mathbf{b}}}(V\mid U):(U,V)\in\mathcal{U}^{\sigma}\Bigr{)},$
-where $\sigma\mathbf{a}=(a_{\sigma(i)})_{i\in\mathbb{Z}}$.
-The condition (1.11) ensures that, between the pairs of points in $(U,V)$, the column rates are either permuted among themselves or are left unchanged. See Figure 4 for an example.
-Figure 4.
-An example of an element $(U,V)\in\mathcal{U}^{\sigma}$, where $U$ and $V$ each contain 33 elements.
-The gray horizontal lines are used to indicate a periodic environment with $N=4$.
-In this figure, the column parameters of columns $3,5$, and 66 (blue) are permuted, and the column parameters in columns $8$ and 99 (red) are permuted.
-The condition is that, for each of the three endpoints $(x,y)$ in $U$ (lower left), $\sigma(\mathbb{Z}_{<x})=\mathbb{Z}_{<x}$, while for each of the three endpoints $(x,y)$ in $V$ (upper right), $\sigma(\mathbb{Z}_{>x})=\mathbb{Z}_{>x}$.
-The distributional equality in Theorem 1.7 holds jointly over all endpoint pairs $(U,V)\in\Psi$ satisfying this property.
-We can allow for permuting the $\mathbf{b}$ parameters as well, with doubly periodic weights $(W_{(i,j)}\mid i\in\mathbb{Z}_{M},j\in\mathbb{Z}_{N})$ for $M,N\in\mathbb{N}$ when we restrict to points that lie inside the box $\llbracket 0,M\rrbracket\times\llbracket 0,N\rrbracket$.
-In this case, the model isn’t actually periodic at all; it is just a restriction of the full-space polymer to a finite grid.
-For finite permutations $\sigma,\tau:\mathbb{Z}\to\mathbb{Z}$, define the set $\mathcal{U}^{\sigma,\tau}$ to be the set of tuples $(U,V)$, where $U,V$ are each sets of endpoint pairs in $\mathbb{Z}\times\mathbb{Z}$ satisfying (1.11) as well as
-[…]
-The following is an extension of the recent result [BEM+25a, Theorem 3.1] to multi-paths and to the positive-temperature case.
-A sketch of the proof for the extension to multi-paths in zero temperature was previously discussed in [BEM+25a, Remark 4.3].
-The technique follows the same outline as developed in [BEM+25a], as we use a coupling of the weights given by applying a Pitman transform to neighboring rows and columns.
+The following is an extension of the recent result [BEM+25, Theorem 3.1] to multi-paths and to the positive-temperature case.
+A sketch of the proof for the extension to multi-paths in zero-+temperature was previously discussed in [BEM+25, Remark 4.3].
+The technique follows the same outline as developed in [BEM+25], as we use a coupling of the weights given by applying a Pitman transform to neighboring rows and columns.
 However, the coupling we use in this proof is different because we treat the weights as periodic outside a large box and apply the periodic Pitman transform.
@@
 Theorem 1.6.
-For vectors $\mathbf{a}\in\mathbb{R}^{\mathbb{Z}}$ and $\mathbf{b}\in\mathbb{R}^{\mathbb{Z}}$ satisfying $a_{i}+b_{j}>0$ for $i,j\in\mathbb{Z}$ define independent weights $\mathbf{W}^{\mathbf{a},\mathbf{b}}=\bigl{(}W_{(i,j)}:i,j\in\mathbb{Z}\bigr{)}$ such that $W_{(i,j)}\sim\text{Log-Inv-Gamma}(a_{i}+b_{j},\lambda)$ for all $i,j$.
-Then, for any finite permutations $\sigma,\tau:\mathbb{Z}\to\mathbb{Z}$, we have
-$\Bigl{(}Z^{\mathbf{W}^{\mathbf{a},\mathbf{b}}}(V\mid U):(U,V)\in\mathcal{U}^{\sigma,\tau}\Bigr{)}\overset{d}{=}\Bigl{(}Z^{\mathbf{W}^{\sigma\mathbf{a},\tau\mathbf{b}}}(V\mid U):(U,V)\in\mathcal{U}^{\sigma,\tau}\Bigr{)},$
+For vectors $\mathbf{a}\in\mathbb{R}^{\mathbb{Z}}$ and $\mathbf{b}\in\mathbb{R}^{\mathbb{Z}}$ satisfying $a_{i}+b_{j}>0$ for $i,j\in\mathbb{Z}$ define independent weights $\mathbf{X}^{\mathbf{a},\mathbf{b}}=\bigl(X_{(i,j)}:i,j\in\mathbb{Z}\bigr)$ such that $X_{(i,j)}\sim\text{Log-Inv-Gamma}(a_{i}+b_{j},\lambda)$ for all $i,j$.
+Then, for any finite permutations $\sigma,\tau\in\mathcal{S}_{\mathbb{Z}}$, we have
+$\Bigl(Z^{\mathbf{X}^{\mathbf{a},\mathbf{b}}}(V\mid U):(U,V)\in\mathcal{W}^{\sigma,\tau}\Bigr)\overset{d}{=}\Bigl(Z^{\mathbf{X}^{\sigma\mathbf{a},\tau\mathbf{b}}}(V\mid U):(U,V)\in\mathcal{W}^{\sigma,\tau}\Bigr),$
 where $\sigma\mathbf{a}=(a_{\sigma(i)})_{i\in\mathbb{Z}}$ and $\tau\mathbf{b}=(b_{\tau(j)})_{j\in\mathbb{Z}}$.
-Theorems 1.7 and 1.8 are proved in Section 4.
-When the initial point is fixed, invariances of this flavor can be deduced from the symmetry of Schur functions and Whittaker functions (see, for example, [BP08, COSZ14, OSZ14, BC14]).
-A similar type of invariance for last-passage percolation was discovered in the work of Dauvergne [Dau22, Theorem 1.5], motivated by a shift-invariance result previously proved in Borodin, Gorin, and Wheeler [BGW22].
-Connections between these results are discussed further in [BEM+25a, Section 3].
+Theorems 1.5 and 1.6 are proved in Section 4.
+When the initial point is fixed, invariances of this flavor can be deduced from the symmetry of Schur functions and Whittaker functions (see, for example, [BP08, COS+14, OSZ14, BC14]).
+A similar type of invariance for last-passage percolation was discovered in the work of Dauvergne [DAU22, Theorem 1.5], motivated by a shift-invariance result previously proved by Borodin, Gorin, and Wheeler [BGW22].
+Connections between these results are discussed further in [BEM+25, Section 3].
 1.5. Zero-temperature limits
-We also define the zero-temperature discrete periodic Pitman transform as the map $\widetilde{\mathcal{P}}:\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}\to\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}$ defined by
-$\widetilde{\mathcal{P}}(\mathbf{W}_{1},\mathbf{W}_{2})$ $=\bigl{(}\widetilde{T}(\mathbf{W}_{2},\mathbf{W}_{1}),\widetilde{D}(\mathbf{W}_{2},\mathbf{W}_{1})\bigr{)},\quad\text{where}$ (1.13)
-$\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2})$ $=X_{1,i-1}+\max_{j\in\mathbb{Z}_{N}}Y_{[i-1,j]}-\max_{j\in\mathbb{Z}_{N}}Y_{[i,j]},\quad\text{and}$
-$\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})$ $=X_{2,i+1}+\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}-\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]},\quad\text{where}\quad Y_{\ell}=X_{2,\ell+1}-X_{1,\ell}.$
-The mapping $\widetilde{D}$ is similar to, but not equivalent to a mapping used to describe the invariant measures for multi-type TASEP by Ferrari and Martin [FM07].
-See [CGS24, Appendix F] for more on this connection.
+We define the zero-temperature discrete periodic Pitman transform as the map $\widetilde{\mathcal{P}}:\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}\to\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}$ defined by
+$\widetilde{\mathcal{P}}(\mathbf{X}_{1},\mathbf{X}_{2})$ $=\bigl(\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2}),\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})\bigr),\quad\text{where}$ (1.13)
+$\widetilde{T}_{i}(\mathbf{X}_{1},\mathbf{X}_{2})$ $=X_{2,i-1}+\max_{j\in\mathbb{Z}_{N}}Y_{[i-1,j]}-\max_{j\in\mathbb{Z}_{N}}Y_{[i,j]},\quad\text{and}$
+$\widetilde{D}_{i}(\mathbf{X}_{1},\mathbf{X}_{2})$ $=X_{1,i+1}+\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}-\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]},\quad\text{where}\quad Y_{\ell}=X_{1,\ell+1}-X_{2,\ell}.$
+The mapping $\widetilde{D}$ is similar but not equivalent to a mapping used to describe the invariant measures for multi-type TASEP by Ferrari and Martin [FM07].
+See [CGS26, Appendix F] for more on this connection.
 One immediately sees that we can obtain $\widetilde{\mathcal{P}}$ as a limiting case of the positive-temperature discrete periodic Pitman transform:
-$\widetilde{\mathcal{P}}(\mathbf{W}_{1},\mathbf{W}_{2})=\lim_{\beta\to\infty}\frac{1}{\beta}\mathcal{P}(\beta\mathbf{W}_{1},\beta\mathbf{W}_{2})$
-Furthermore, this convergence is uniform as a function of the vectors $\mathbf{W}_{1},\mathbf{W}_{2}$. See Lemma A.1 for details.
+$\widetilde{\mathcal{P}}(\mathbf{X}_{1},\mathbf{X}_{2})=\lim_{\beta\to\infty}\frac{1}{\beta}\mathcal{P}(\beta\mathbf{X}_{1},\beta\mathbf{X}_{2}).$
 Definition 1.7.
-For real-valued weights $\mathbf{W}=\bigl{(}W_{(i,j)}\in\mathbb{R}\mid i,j\in\mathbb{Z}\bigr{)}$ corresponding to the vertices of the directed lattice $\mathbb{Z}^{2}$, we define the last-passage percolation (LPP) time for both a single pair of points and multi-paths analogously as in (1.7)-(1.8):
-$G^{\mathbf{W}}(V\mid U)=\max_{\pi\in\Pi_{U,V}}\sum_{(r,\ell)\in\pi}W_{(r,\ell)}.$
-For $k\in\mathbb{Z}$, define the map $\widetilde{\mathcal{P}}_{k}:(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}\to(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$ by (1.9), with $T,D$ replaced by $\widetilde{T},\widetilde{D}$.
+For real-valued weights $\mathbf{X}=\bigl(X_{(i,j)}\in\mathbb{R}\mid i,j\in\mathbb{Z}\bigr)$ corresponding to the vertices of the directed lattice $\mathbb{Z}^{2}$, we define the last-passage percolation (LPP) time for both a single pair of points and multi-paths analogously as in (1.8)-(1.9):
+$G^{\mathbf{X}}(V\mid U)=\max_{\pi\in\Pi_{U,V}}\sum_{(r,\ell)\in\pi}X_{(r,\ell)}.$
+For $k\in\mathbb{Z}$, we can also define the map $\widetilde{\mathcal{P}}_{k}:(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}\to(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$ by (1.3), with $T,D$ replaced by $\widetilde{T},\widetilde{D}$.
 With this notation defined, we get the following zero-temperature analogues of our main theorems.
 Theorem 1.8.
-The same invariance in Theorem 1.4 holds when $Z^{\mathbf{W}}$ is replaced by $G^{\mathbf{W}}$, and the maps $\mathcal{P}_{k_{j}}$ are replaced by $\widetilde{\mathcal{P}}_{k_{j}}$.
+The same relations in Theorem 1.1 hold for the maps $\widetilde{\mathcal{P}}_{k}$.
+Theorem 1.8 similarly allows us to define maps $\widetilde{\mathcal{P}}_{\sigma}$ for a general finite permutation $\sigma\in\mathcal{S}_{\mathbb{Z}}$.
 Theorem 1.9.
-The same braid relations in Theorem 1.6 hold for the maps $\widetilde{\mathcal{P}}_{k}$.
+The same invariance in Theorem 1.4 holds when $Z^{\mathbf{X}}$ is replaced by $G^{\mathbf{X}}$, and the maps $\mathcal{P}_{\sigma}$ are replaced by $\widetilde{\mathcal{P}}_{\sigma}$.
 Theorem 1.10.
-The analogues of Theorems 1.7 and 1.8 hold, when $Z^{\mathbf{W}}$ is replaced by $G^{\mathbf{W}}$ and, instead of log-inverse-gamma weights, the weights satisfy either $W_{(i,j)}\sim\text{Geom}(q_{i}r_{j})$ for all $i,j$, where $q_{i}r_{j}\in(0,1)$, or $W_{(i,j)}\sim\text{Exp}(a_{i}+b_{j})$ for all $i,j$, where $a_{i}+b_{j}>0$ (see Definitions 1.13-1.14 below).
-The details of the proofs of these theorems are given in Appendix B.
+The analogues of Theorems 1.5 and 1.6 hold, when $Z^{\mathbf{X}}$ is replaced by $G^{\mathbf{X}}$ and, instead of log-inverse-gamma weights, the weights satisfy either $X_{(i,j)}\sim\text{Geom}(q_{i}r_{j})$ for all $i,j$, where $q_{i}r_{j}\in(0,1)$, or $X_{(i,j)}\sim\text{Exp}(a_{i}+b_{j})$ for all $i,j$, where $a_{i}+b_{j}>0$ (see Definitions 1.11-1.12 below).
+The details of the limit transitions for the proofs of these theorems are given in Appendix A.2.
 1.6. Organization of the paper
+Section 2 develops the algebraic structure of the periodic Pitman transform, culminating in the proof of Theorem 1.1 in Section 2.2.
 Section 3 is focused on proving the partition function invariance in Theorem 1.4.
-The braid relations in Theorem 1.6 are proved in Section 3.
-Section 4 proves an inhomogeneous Burke property for the periodic Pitman transform and uses it to prove Theorems 1.7 and 1.8.
-Appendix A contains some prior results from the work [CGS24] and some auxiliary technical results.
-The necessary details for the zero-temperature analogues in Theorems 1.10-1.12 are proved in Appendix B.
+Section 4 proves an inhomogeneous Burke property for the periodic Pitman transform and uses it to prove Theorems 1.5 and 1.6.
+Appendix A.1 contains some prior results from the work [CGS26] and some auxiliary technical results.
+The necessary details for the zero-temperature analogues in Theorems 1.9-1.10 are proved in Appendix A.2.
 1.7. Notation and conventions
-We use $\overset{d}{=}$ to indicate that two random variables are equal in distribution. Below, we define our notation for geometric, exponential, and log-inverse-gamma random variables.
+We use $\overset{d}{=}$ to indicate that two random variables or stochastic processes are equal in distribution.
+Below, we define our notation for geometric, exponential, and log-inverse-gamma random variables.
 Definition 1.11 (Geometric Random Variable).
@@
 E.R.E.
-was supported by Princeton University’s Office of Undergraduate Research OURSIP Program through the Frances Lane Summer Research Fund.
-We also thank Ivan Corwin and Amol Aggarwal for helpful discussions related to this project.
-2. Invariance of polymer partition function
-In this section, we prove Theorem 1.4.
-The key to our proof is a matrix encoding of polymer partition functions, analogous to that developed in [NY04].
-Section 2.1 describes the necessary algebraic properties of these matrices, and then Section 2.2 makes the connection to the polymer partition function and completes the proof.
+was supported by Princeton University’s Office of Undergraduate Research OURSIP Program through the Frances Lane Summer Research Fund and by Princeton University’s Office of Undergraduate Research Undergraduate Fund for Academic Conferences through the Hewlett Foundation Fund.
+This material is based upon work supported by the National Science Foundation under Grant No.
+2015553.
+We thank Ivan Corwin and Amol Aggarwal for helpful discussions related to this project.
+We also thank the anonymous referee for helpful comments on an earlier version of this paper that have greatly improved the presentation of our results.
+2. Algebraic Properties of the Pitman Transform
+The end goal of this section is to prove Theorem 1.1, which is done in in Section 2.2.
+We start by developing the necessary algebraic background.
+Versions of the following two results are proved in [CGS26] for shifted versions of the maps $D$ and $T$.
+We provide self-contained proofs in this section.
+For this lemma and below, we define the notation
+$\mathfrak{s}(\mathbf{X})=\sum_{i\in\mathbb{Z}_{N}}X_{i},\quad\text{for }\mathbf{X}\in\mathbb{R}^{\mathbb{Z}_{N}}.$
+Lemma 2.1.
+Let $\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{R}^{\mathbb{Z}_{N}}$. Then,
+  1. (i)
+    $\mathfrak{s}(T(\mathbf{X}_{1},\mathbf{X}_{2}))=\mathfrak{s}(\mathbf{X}_{2})$, and $\mathfrak{s}(D(\mathbf{X}_{1},\mathbf{X}_{2}))=\mathfrak{s}(\mathbf{X}_{1})$.
+  2. (ii)
+    $\mathfrak{s}(\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2}))=\mathfrak{s}(\mathbf{X}_{2})$, and $\mathfrak{s}(\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2}))=\mathfrak{s}(\mathbf{X}_{1})$.
+Proof.
+These follow immediately from the definitions (1.2) and (1.13), since the sums are telescoping. ∎
+Lemma 2.2.
+Let $\mathbf{X}_{1},\mathbf{X_{2}}\in\mathbb{R}^{\mathbb{Z}_{N}}$. Then, for $i\in\mathbb{Z}_{N}$,
+  1. (i)
+    $D_{i}(\mathbf{X}_{1},\mathbf{X_{2}})+T_{i}(\mathbf{X}_{1},\mathbf{X_{2}})=X_{1,i}+X_{2,i}$,
+  2. (ii)
+    $e^{-D_{i}(\mathbf{X}_{1},\mathbf{X_{2}})}+e^{-T_{i+1}(\mathbf{X}_{1},\mathbf{X_{2}})}=e^{-X_{2,i}}+e^{-X_{1,i+1}}$, and
+  3. (iii)
+    $\widetilde{D}_{i}(\mathbf{X}_{1},\mathbf{X_{2}})+\widetilde{T}_{i}(\mathbf{X}_{1},\mathbf{X_{2}})=X_{1,i}+X_{2,i}$.
+Proof.
+Item (i): From the definition (1.2), we observe that
+$T_{i}(\mathbf{X}_{1},\mathbf{X}_{2})$ $=X_{2,i-1}+\log\Biggl(\frac{e^{Y_{i-1}}\sum_{j\in\mathbb{Z}_{N}}e^{Y_{(i-1,j]}}}{e^{Y_{i}}\sum_{j\in\mathbb{Z}_{N}}e^{Y_{(i,j]}}}\Biggr)$
+$=X_{2,i-1}+Y_{i-1}-Y_{i}-\log\Biggl(\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{(i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{(i-1,j]}}}\Biggr)$
+$=X_{2,i}+X_{1,i}-D_{i}(\mathbf{X}_{1},\mathbf{X}_{2}).$
+Item (ii): We compute this as follows:
+$e^{-D_{i}(\mathbf{\mathbf{X}_{1}},\mathbf{\mathbf{X}_{2}})}+e^{-T_{i+1}(\mathbf{\mathbf{X}_{1}},\mathbf{\mathbf{X}_{2}})}$ $=e^{-X_{1,i+1}}\left(\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{(i-1,j]}}}{\sum_{j\in\mathbb{Z}_{N}}{e^{Y_{(i,j]}}}}\right)+e^{-X_{2,i}}\left(\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{[i+1,j]}}}{\sum_{j\in\mathbb{Z}_{N}}{e^{Y_{[i,j]}}}}\right)$
+$=e^{-X_{2,i}}\left(\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{(i-1,j]}}}{e^{Y_{i}}\sum_{j\in\mathbb{Z}_{N}}{e^{Y_{(i,j]}}}}\right)+e^{-X_{1,i+1}}\left(\frac{e^{Y_{i}}\sum_{j\in\mathbb{Z}_{N}}e^{Y_{[i+1,j]}}}{\sum_{j\in\mathbb{Z}_{N}}{e^{Y_{[i,j]}}}}\right)$
+$=e^{-X_{2,i}}\left(1+\frac{1-e^{\sum_{j\in\mathbb{Z}_{N}}Y_{j}}}{\sum_{j\in\mathbb{Z}_{N}}{e^{Y_{[i,j]}}}}\right)+e^{-X_{1,i+1}}\left(1+\frac{e^{Y_{i}}(e^{\sum_{j\in\mathbb{Z}_{N}}Y_{j}}-1)}{\sum_{j\in\mathbb{Z}_{N}}{e^{Y_{[i,j]}}}}\right)$
+$=e^{-X_{2,i}}+e^{-X_{1,i+1}}+e^{-X_{2,i}}\left(\frac{1-e^{\sum_{j\in\mathbb{Z}_{N}}Y_{j}}}{\sum_{j\in\mathbb{Z}_{N}}{e^{Y_{[i,j]}}}}+\frac{e^{{\sum_{j\in\mathbb{Z}_{N}}}Y_{j}}-1}{\sum_{j\in\mathbb{Z}_{N}}{e^{Y_{[i,j]}}}}\right)$
+$=e^{-X_{2,i}}+e^{-X_{1,i+1}}.$
+Item (iii): This follows as the zero-temperature limit of Item (i). ∎
 2.1. Matrix encoding
-In order to prove our polymer partition function invariance, we follow the matrix encoding method introduced at Noumi and Yamada in [NY04].
-The key difference is that we will work with infinite matrices.
+The proofs of Theorems 1.1 and 1.4 follow by using the algebraic properties above to prove certain matrix relations.
+For this, we adapt the framework of Noumi and Yamada [NY04] to infinite matrices which have a periodic structure.
 Definition 2.3.
@@
 All of the same properties that hold for $\mathbb{Z}$ trivially also hold for $\mathbb{Z}_{N}$, via the inclusion map $i:\mathbb{R}^{\mathbb{Z}_{N}}\to\mathbb{R}^{\mathbb{Z}}$.
-The intuition behind this approach is that matrix multiplication $H(\mathbf{x}_{1})\cdots H(\mathbf{x}_{n})$ naturally encodes the partition function across a certain diagram with vertex weights given by the arrays $\mathbf{x}_{1},\ldots,\mathbf{x}_{n}$ as discussed in Lemma 2.6.
-Meanwhile, the outputs of map $E$ are more algebraically straightforward and closely related to the matrix inverse of $H$, which we prove in the next Lemma 2.2 as an infinite-dimensional analogue of [NY04, Equation 1.5].
+The intuition behind this approach is that matrix multiplication $H(\mathbf{x}_{1})\cdots H(\mathbf{x}_{n})$ naturally encodes the partition function across a certain diagram with vertex weights given by the arrays $\mathbf{x}_{1},\ldots,\mathbf{x}_{n}$ as discussed later in Lemma 3.1.
+Meanwhile, the outputs of map $E$ are more algebraically straightforward and closely related to the matrix inverse of map $H$, which we prove in the next Lemma 2.4 as an infinite-dimensional analogue of [NY04, Equation 1.5].
 Lemma 2.4.
 For all $\mathbf{x}\in\mathbb{R}^{\mathbb{Z}}$, we have
-$\Delta H(\mathbf{x})\Delta E(-\mathbf{x})=E(-\mathbf{x})\Delta H(\mathbf{x})\Delta=\text{Id}_{\mathbb{Z}},$ (2.2)
-where $\Delta=\operatorname{diag}((-1)^{i-1})_{i\in\mathbb{Z}}$ and $\text{Id}_{\mathbb{Z}}=\operatorname{diag}(1)_{i\in\mathbb{Z}}$.
+$\Delta H(\mathbf{x})\Delta E(-\mathbf{x})=E(-\mathbf{x})\Delta H(\mathbf{x})\Delta=\mathrm{Id},$ (2.2)
+where $\Delta=\operatorname{diag}((-1)^{i-1})_{i\in\mathbb{Z}}$ and $\mathrm{Id}=\operatorname{diag}(1)_{i\in\mathbb{Z}}$.
 Proof.
@@
 $=H(\mathbf{x})\sum_{j\in\mathbb{Z}}(-1)^{j-1}e^{-x_{j}}E_{j,j}+H(\mathbf{x})\sum_{j\in\mathbb{Z}}(-1)^{j-1}E_{j,j+1}$
-$=\sum_{i\in\mathbb{Z}}(-1)^{i-1}E_{i,i}+\sum_{i\leq j}\bigl{(}(-1)^{j-1}e^{x_{i}+\cdots+x_{j-1}}+(-1)^{j-2}e^{x_{i}+\cdots+x_{j-1}}\bigr{)}E_{i,k}$
+$=\sum_{i\in\mathbb{Z}}(-1)^{i-1}E_{i,i}+\sum_{j\geq i}\bigl((-1)^{j-1}e^{x_{i}+\cdots+x_{j-1}}+(-1)^{j-2}e^{x_{i}+\cdots+x_{j-1}}\bigr)E_{i,j}$
 $=\Delta\ ,$
@@
 A similar calculation shows that $E(-\mathbf{x})\Delta H(\mathbf{x})=\Delta$.
-This concludes the proof, since one can easily check that $\Delta^{2}=Id$.
+This concludes the proof, since one can easily check that $\Delta^{2}=\mathrm{Id}$.
 ∎
@@
 Lemma 2.6.
-For $\mathbf{W}_{1},\mathbf{W}_{2}\in\mathbb{R}^{\mathbb{Z}_{N}}$, we have that
-$E(-\mathbf{W}_{2})E(-\mathbf{W}_{1})=E(-D(\mathbf{W}_{2},\mathbf{W}_{1}))E(-T(\mathbf{W}_{2},\mathbf{W}_{1})).$ (2.4)
+For $\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{R}^{\mathbb{Z}_{N}}$, we have that
+$E(-\mathbf{X}_{2})E(-\mathbf{X}_{1})=E(-D(\mathbf{X}_{1},\mathbf{X}_{2}))E(-T(\mathbf{X}_{1},\mathbf{X}_{2})).$ (2.4)
 Proof.
-By direct computation and applying Lemma A.5 in the third equality below, we see that
-$\quad\,E(-D(\mathbf{W}_{2},\mathbf{W}_{1}))E(-T(\mathbf{W}_{2},\mathbf{W}_{1}))$
-$=\left(\sum_{i\in\mathbb{Z}}e^{-D_{i}(\mathbf{W}_{2},\mathbf{W}_{1})}E_{i,i}+\sum_{i\in\mathbb{Z}}E_{i,i+1}\right)\left(\sum_{i\in\mathbb{Z}}e^{-T_{i}(\mathbf{W}_{2},\mathbf{W}_{1})}E_{i,i}+\sum_{i\in\mathbb{Z}}E_{i,i+1}\right)$
-$=\sum_{i\in\mathbb{Z}}e^{-(D_{i}(\mathbf{W}_{2},\mathbf{W}_{1})+T_{i}(\mathbf{W}_{2},\mathbf{W}_{1}))}E_{i,i}+\sum_{i\in\mathbb{Z}}\left(e^{-D_{i}(\mathbf{W}_{2},\mathbf{W}_{1})}+e^{-T_{i+1}(\mathbf{W}_{2},\mathbf{W}_{1})}\right)E_{i,i+1}+\sum_{i\in\mathbb{Z}}E_{i,i+2}$
-$=\sum_{i\in\mathbb{Z}}e^{-(W_{1,i}+W_{2,i})}E_{i,i}+\sum_{i\in\mathbb{Z}}\left(e^{-W_{1,i+1}}+e^{-W_{2,i}}\right)E_{i,i+1}+\sum_{i\in\mathbb{Z}}E_{i,i+2}$
-$=\left(\sum_{i\in\mathbb{Z}}e^{-W_{2,i}}E_{i,i}+\sum_{i\in\mathbb{Z}}E_{i,i+1}\right)\left(\sum_{i\in\mathbb{Z}}e^{-W_{1,i}}E_{i,i}+\sum_{i\in\mathbb{Z}}E_{i,i+1}\right)$
-$=E(-\mathbf{W}_{2})E(-\mathbf{W}_{1}).\qed$
-Corollary 2.5.
-For $\mathbf{W}_{1},\mathbf{W}_{2}\in\mathbb{R}^{\mathbb{Z}_{N}}$, we have
-$H(\mathbf{W}_{1})H(\mathbf{W}_{2})=H(T(\mathbf{W}_{2},\mathbf{W}_{1}))H(D(\mathbf{W}_{2},\mathbf{W}_{1})).$ (2.5)
+By direct computation and applying Lemma 2.2 in the third equality below, we see that
+$\quad\,E(-D(\mathbf{X}_{1},\mathbf{X}_{2}))E(-T(\mathbf{X}_{1},\mathbf{X}_{2}))$
+$=\left(\sum_{i\in\mathbb{Z}}e^{-D_{i}(\mathbf{X}_{1},\mathbf{X}_{2})}E_{i,i}+\sum_{i\in\mathbb{Z}}E_{i,i+1}\right)\left(\sum_{i\in\mathbb{Z}}e^{-T_{i}(\mathbf{X}_{1},\mathbf{X}_{2})}E_{i,i}+\sum_{i\in\mathbb{Z}}E_{i,i+1}\right)$
+$=\sum_{i\in\mathbb{Z}}e^{-(D_{i}(\mathbf{X}_{1},\mathbf{X}_{2})+T_{i}(\mathbf{X}_{1},\mathbf{X}_{2}))}E_{i,i}+\sum_{i\in\mathbb{Z}}\left(e^{-D_{i}(\mathbf{X}_{1},\mathbf{X}_{2})}+e^{-T_{i+1}(\mathbf{X}_{1},\mathbf{X}_{2})}\right)E_{i,i+1}+\sum_{i\in\mathbb{Z}}E_{i,i+2}$
+$=\sum_{i\in\mathbb{Z}}e^{-(X_{1,i}+X_{2,i})}E_{i,i}+\sum_{i\in\mathbb{Z}}\left(e^{-X_{1,i+1}}+e^{-X_{2,i}}\right)E_{i,i+1}+\sum_{i\in\mathbb{Z}}E_{i,i+2}$
+$=\left(\sum_{i\in\mathbb{Z}}e^{-X_{2,i}}E_{i,i}+\sum_{i\in\mathbb{Z}}E_{i,i+1}\right)\left(\sum_{i\in\mathbb{Z}}e^{-X_{1,i}}E_{i,i}+\sum_{i\in\mathbb{Z}}E_{i,i+1}\right)=E(-\mathbf{X}_{2})E(-\mathbf{X}_{1}).\qed$
+The following is the key step that will allow us to prove both Theorem 1.1 and Theorem 1.4.
+Proposition 2.7.
+For $\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{R}^{\mathbb{Z}_{N}}$, we have
+$H(\mathbf{X}_{1})H(\mathbf{X}_{2})=H(T(\mathbf{X}_{1},\mathbf{X}_{2}))H(D(\mathbf{X}_{1},\mathbf{X}_{2})).$ (2.5)
 Proof.
-First, we justify that this matrix product has finite entries.
-Because these $H$ matrices (as well as the $E$ matrices) are upper triangular, each individual index in the product consists of a finite sum of products of elements in the respective $H$ matrices.
-While each row and column has infinitely many non-zero terms, there are only a finite number of non-zero terms in the dot product.
-Further, the product is upper triangular, so by repeating this argument, we can see that an arbitrary product of $H$ (or $E$) matrices is finite valued.
+First, we justify that this matrix product has finite non-negative real entries.
+Because the outputs of the matrix map $H$ (as well as matrix map $E$) are upper triangular, each individual index in the product consists of a finite sum of products of elements in the respective $H$ matrices.
+While each row and column has infinitely many non-zero terms, there are only a finite number of non-zero terms in each dot product.
+Further, the resulting matrix is upper triangular, so by repeating this argument, we can see that an arbitrary product of $H$ (or $E$) matrices has finite non-negative real entries.
 With this in mind, the proof follows from Lemma 2.4, since:
 […]
-$=\Delta E(-\mathbf{W}_{1})^{-1}E(-\mathbf{W}_{2})^{-1}\Delta$
-$=\Delta\bigl{(}E(-\mathbf{W}_{2})E(-\mathbf{W}_{1})\bigr{)}^{-1}\Delta$
+$=\Delta E(-\mathbf{X}_{1})^{-1}E(-\mathbf{X}_{2})^{-1}\Delta$
+$=\Delta\bigl(E(-\mathbf{X}_{2})E(-\mathbf{X}_{1})\bigr)^{-1}\Delta$
 […]
-$=(\Delta E(-T(\mathbf{W}_{2},\mathbf{W}_{1}))^{-1}\Delta)(\Delta E(-D(\mathbf{W}_{2},\mathbf{W}_{1}))^{-1}\Delta)$
-$=H(T(\mathbf{W}_{2},\mathbf{W}_{1}))H(D(\mathbf{W}_{2},\mathbf{W}_{1})).\qed$
-2.2. Relation to the periodic polymer
-We now show the following relation between the $H$ matrices and the polymer partition function:
+$=(\Delta E(-T(\mathbf{X}_{1},\mathbf{X}_{2}))^{-1}\Delta)(\Delta E(-D(\mathbf{X}_{1},\mathbf{X}_{2}))^{-1}\Delta)$
+$=H(T(\mathbf{X}_{1},\mathbf{X}_{2}))H(D(\mathbf{X}_{1},\mathbf{X}_{2})).\qed$
+2.2. Proof of Theorem 1.1
+As observed previously, Item (iii) is immediate from the definition. We prove Items (i) and (ii) separately.
+2.2.1. Proof of Item (i):
+Our goal is to show that
+$\mathcal{P}(\mathcal{P}(\mathbf{X}_{1},\mathbf{X}_{2}))=(T(T(\mathbf{X}_{1},\mathbf{X}_{2}),D(\mathbf{X}_{1},\mathbf{X}_{2})),D(T(\mathbf{X}_{1},\mathbf{X}_{2}),D(\mathbf{X}_{1},\mathbf{X}_{2})))=(\mathbf{X}_{1},\mathbf{X}_{2}).$ (2.6)
+Firstly, with the convention $Y_{\ell}=X_{1,\ell+1}-X_{2,\ell}$ and $Y^{\prime}_{\ell}=T_{\ell+1}(\mathbf{X}_{1},\mathbf{X}_{2})-D_{\ell}(\mathbf{X}_{1},\mathbf{X}_{2})$, notice that
+$T_{i}(T(\mathbf{X}_{1},\mathbf{X}_{2}),D(\mathbf{X}_{1},\mathbf{X}_{2}))$ $=D_{i-1}(\mathbf{X}_{1},\mathbf{X}_{2})+\log\Biggl(\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y^{\prime}_{[i-1,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{Y^{\prime}_{[i,j]}}}\Biggr)$
+$=X_{1,i}+\log\Biggl(\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{(i-1,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{(i-2,j]}}}\Biggr)+\log\Biggl(\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y^{\prime}_{[i-1,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{Y^{\prime}_{[i,j]}}}\Biggr),$
+and this equals $X_{1,i}$ if and only if
+$\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y^{\prime}_{[i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{(i-1,j]}}}=\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y^{\prime}_{[i-1,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{(i-2,j]}}}.$
+Hence, it is sufficient to prove that the quantity $\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y^{\prime}_{[i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{(i-1,j]}}}$ is the same for all $i\in\mathbb{Z}_{N}$.
+To show this, first observe that
+$e^{Y^{\prime}_{i}}$ $=e^{X_{2,i}-X_{1,i+1}}\Biggl(\frac{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{[i,\ell]}}}{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{[i+1,\ell]}}}\Biggr)\Biggl(\frac{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{(i-1,\ell]}}}{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{(i,\ell]}}}\Biggr)$ (2.7)
+$\Longrightarrow e^{Y^{\prime}_{[i,j]}}$ $=e^{-Y_{[i,j]}}\Biggl(\frac{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{[i,\ell]}}}{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{[j+1,\ell]}}}\Biggr)\Biggl(\frac{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{(i-1,\ell]}}}{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{(j,\ell]}}}\Biggr)$
+$\Longrightarrow\sum_{j\in\mathbb{Z}_{N}}e^{Y^{\prime}_{[i,j]}}$ $=\sum_{j\in\mathbb{Z}_{N}}e^{-Y_{(i,j]}}\Biggl(\frac{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{(i,\ell]}}}{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{[j+1,\ell]}}}\Biggr)\Biggl(\frac{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{(i-1,\ell]}}}{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{(j,\ell]}}}\Biggr)$
+$\Longrightarrow\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y^{\prime}_{[i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{(i-1,j]}}}$ $=\Biggl(\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{(i,\ell]}}\Biggr)\sum_{j\in\mathbb{Z}_{N}}\frac{e^{-Y_{(i,j]}}}{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{[j+1,\ell]}}\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{(j,\ell]}}}.$
+Now, we note that $\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{[j+1,\ell]}}-\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{(j,\ell]}}=e^{\mathfrak{s}(Y)}-1$ for each $j\in\mathbb{Z}_{N}$.
+Then, multiplying and dividing by this term, we obtain
+$\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y^{\prime}_{[i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{(i-1,j]}}}$ $=\Biggl(\frac{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{(i,\ell]}}}{e^{\mathfrak{s}(Y)}-1}\Biggr)\sum_{j\in\mathbb{Z}_{N}}\frac{e^{-Y_{(i,j]}}\bigl(\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{[j+1,\ell]}}-\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{(j,\ell]}}\bigr)}{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{[j+1,\ell]}}\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{(j,\ell]}}}$ (2.8)
+$=\Biggl(\frac{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{(i,\ell]}}}{e^{\mathfrak{s}(Y)}-1}\Biggr)\sum_{j\in\mathbb{Z}_{N}}\Biggl(\frac{e^{-Y_{(i,j]}}}{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{(j,\ell]}}}-\frac{e^{-Y_{(i,j]}-Y_{j+1}}}{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{(j+1,\ell]}}}\Biggr)$
+$=\Biggl(\frac{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{(i,\ell]}}}{e^{\mathfrak{s}(Y)}-1}\Biggr)\Biggl(\frac{1}{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{(i,\ell]}}}-\frac{e^{-\mathfrak{s}(Y)}}{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{(i,\ell]}}}\Biggr)=e^{-\mathfrak{s}(Y)},$
+where in the penultimate step, we start the sum at $j=i$, note that the sum is telescoping, and use the fact that
+$Y_{(i,j]}+Y_{j+1}=\begin{cases}Y_{(i,j+1]},&\text{if }j\not\equiv i-1\\ \mathfrak{s}(Y),&\text{if }j\equiv i-1\end{cases}.$
+Our final expression does not depend on $i$, as we wanted to prove. Similarly, we notice that
+$D_{i}(T(\mathbf{X}_{1},\mathbf{X}_{2}),D(\mathbf{X}_{1},\mathbf{X}_{2}))$ $=T_{i+1}(\mathbf{X}_{1},\mathbf{X}_{2})+\log\Biggl(\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y^{\prime}_{(i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{Y^{\prime}_{(i-1,j]}}}\Biggr)$ (2.9)
+$=X_{2,i}+\log\Biggl(\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{[i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{[i+1,j]}}}\Biggr)+\log\Biggl(\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y^{\prime}_{(i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{Y^{\prime}_{(i-1,j]}}}\Biggr),$
+and this equals $X_{2,i}$ if and only if
+$\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y^{\prime}_{(i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{[i+1,j]}}}=\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y^{\prime}_{[i-1,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{(i-2,j]}}}.$
+Hence, it is sufficient to prove that the quantity $\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y^{\prime}_{(i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{[i+1,j]}}}$ is the same for all $i\in\mathbb{Z}_{N}$.
+From our previous observations, we know that
+$e^{Y^{\prime}_{(i,j]}}$ $=e^{-Y_{(i,j]}}\Biggl(\frac{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{[i+1,\ell]}}}{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{[j+1,\ell]}}}\Biggr)\Biggl(\frac{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{(i,\ell]}}}{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{(j,\ell]}}}\Biggr)$ (2.10)
+$\Longrightarrow\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y^{\prime}_{(i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{[i+1,j]}}}$ $=\Biggl(\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{(i,\ell]}}\Biggr)\sum_{j\in\mathbb{Z}_{N}}\frac{e^{-Y_{(i,j]}}}{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{[j+1,\ell]}}\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{(j,\ell]}}},$
+which we have already shown to be equal to $e^{-\mathfrak{s}(Y)}$, which does not depend on $i$. This completes the proof.
+2.2.2. Proof of Item (ii)
+Notice that the relation $\mathcal{P}_{k}\mathcal{P}_{k+1}\mathcal{P}_{k}=\mathcal{P}_{k+1}\mathcal{P}_{k}\mathcal{P}_{k+1}$ translates into the following three equalities:
+$T\Bigl(T(\mathbf{X}_{k},\mathbf{X}_{k+1}),T\bigl(D(\mathbf{X}_{k},\mathbf{X}_{k+1}),\mathbf{X}_{k+2}\bigr)\biggr)=T\Bigl(\mathbf{X}_{k},T(\mathbf{X}_{k+1},\mathbf{X}_{k+2})\Bigr),$ (2.11)
+$D\Bigl(T(\mathbf{X}_{k},\mathbf{X}_{k+1}),T\bigl(D(\mathbf{X}_{k},\mathbf{X}_{k+1}),\mathbf{X}_{k+2}\bigr)\Bigr)=T\Bigl(D\bigl(\mathbf{X}_{k},T(\mathbf{X}_{k+1},\mathbf{X}_{k+2})\bigr),D(\mathbf{X}_{k+1},\mathbf{X}_{k+2})\Bigr),$ (2.12)
+$D\Bigl(D(\mathbf{X}_{k},\mathbf{X}_{k+1}),\mathbf{X}_{k+2}\Bigr)=D\Bigl(D\bigl(\mathbf{X}_{k},T(\mathbf{X}_{k+1},\mathbf{X}_{k+2})\bigr),D(\mathbf{X}_{k+1},\mathbf{X}_{k+2})\Bigr).$ (2.13)
+The last of these, equation (2.13) was proved previously as [CGS26, Proposition 5.13], recorded in Lemma A.2 of the appendix.
+We now use the matrix relation of Proposition 2.7 to demonstrate that (2.13) implies both (2.11) and (2.12).
+The main tool will be the following lemma:
 Lemma 2.8.
-Let $\mathbf{W}=(\mathbf{W}_{i})_{i\in\mathbb{Z}}\in(\mathbb{Z}_{N})^{\mathbb{Z}}$ and $k,\ell\in\mathbb{Z}$ such that $\ell\geq k$, then:
-$\Bigl{(}H(\mathbf{W}_{k})H(\mathbf{W}_{k+1})\ldots H(\mathbf{W}_{\ell})\Bigr{)}_{(i,j)}=Z^{\mathbf{W}}(\ell,j\mid k,i)\quad\forall i,j\in\mathbb{Z}.$
-In other words, the partition function across any two points $(k,i)$ and $(\ell,j)$ on the directed lattice $\mathbb{Z}^{2}$ with periodic weights $\mathbf{W}$ is given by the $(i,j)$th entry in the matrix product of the $H$ matrices.
+For $\mathbf{W}=(\mathbf{W}_{i})_{i\in\mathbb{Z}},\mathbf{W}^{\prime}=(\mathbf{W}^{\prime}_{i})_{i\in\mathbb{Z}}\in(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$ satisfying $\mathfrak{s}(\mathbf{W}_{1})=\mathfrak{s}(\mathbf{W}^{\prime}_{1})$ we have that $H(\mathbf{W}_{1})H(\mathbf{W}_{2})=H(\mathbf{W}^{\prime}_{1})H(\mathbf{W}_{2}^{\prime})$ if and only if $\mathbf{W}_{1}=\mathbf{W}^{\prime}_{1}$ and $\mathbf{W}_{2}=\mathbf{W}_{2}^{\prime}$.
+Before proving Lemma 2.8, we first use Lemmas 2.8 and A.2 to complete the proof of Theorem 1.1 (ii):
+Proof of Theorem 1.1(ii)..
+We apply Proposition 2.7 three times: first to the last two terms, then the first two, then the last two again:
+$H\bigl(\mathbf{X}_{k}\bigr)H\bigl(\mathbf{X}_{k+1}\bigr)H\bigl(\mathbf{X}_{k+2}\bigr)=$
+[…]
+[…]
+[…]
+$\qquad H\Bigl(D\bigl(D\bigl(\mathbf{X}_{k},T(\mathbf{X}_{k+1},\mathbf{X}_{k+2})\bigr),D(\mathbf{X}_{k+1},\mathbf{X}_{k+2})\bigr)\Bigr).$ (2.14)
+Alternatively, we apply Proposition 2.7 to the first two terms, then the last two terms, then again the first two terms to get
+$H\bigl(\mathbf{X}_{k}\bigr)H\bigl(\mathbf{X}_{k+1}\bigr)H\bigl(\mathbf{X}_{k+2}\bigr)=$
+[…]
+[…]
+[…]
+$\qquad H\Bigl(D\bigl(T(\mathbf{X}_{k},\mathbf{X}_{k+1}),T\bigl(D(\mathbf{X}_{k},\mathbf{X}_{k+1}),\mathbf{X}_{k+2}\bigr)\big)\Bigr)H\Bigl(D\bigl(D(\mathbf{X}_{k},\mathbf{X}_{k+1}),\mathbf{X}_{k+2}\bigr)\Bigr).$ (2.15)
+From Lemma A.2, we know that the last terms in (2.14) and (2.15) are equal. Because $H$ is invertible (see Lemma 2.4), we can equate both equations and cancel out the last term, obtaining that:
+$H\Bigl(T\bigl(\mathbf{X}_{k},T(\mathbf{X}_{k+1},\mathbf{X}_{k+2})\bigr))\Bigr)H\Bigl(T\bigl(D\bigl(\mathbf{X}_{k},T(\mathbf{X}_{k+1},\mathbf{X}_{k+2})\bigr),D(\mathbf{X}_{k+1},\mathbf{X}_{k+2})\bigr)\Bigr)$
+$=H\Bigl(T\bigl(T(\mathbf{X}_{k},\mathbf{X}_{k+1}),T\bigl(D(\mathbf{X}_{k},\mathbf{X}_{k+1}),\mathbf{X}_{k+2}\bigr)\big)\Bigr)H\Bigl(D\bigl(T(\mathbf{X}_{k},\mathbf{X}_{k+1}),T\bigl(D(\mathbf{X}_{k},\mathbf{X}_{k+1}),\mathbf{X}_{k+2}\bigr)\big)\Bigr).$
+By repeated application of Lemma 2.1, we have that
+$\mathfrak{s}(T(\mathbf{X}_{k},T(\mathbf{X}_{k+1},\mathbf{X}_{k+2})))=\mathfrak{s}(\mathbf{X}_{k+2})=\mathfrak{s}(T(T(\mathbf{X}_{k},\mathbf{X}_{k+1}),T(D(\mathbf{X}_{k},\mathbf{X}_{k+1}),\mathbf{X}_{k+2}))).$
+Applying Lemma 2.8, we conclude (2.11) and (2.12), as we wanted to prove. ∎
+Proof of Lemma 2.8..
+Immediately, the backwards direction is clear.
+Then, it remains to show that, if $H(\mathbf{W}_{1})H(\mathbf{W}_{2})=H(\mathbf{W}_{1}^{\prime})H(\mathbf{W}_{2}^{\prime})$, then $\mathbf{W}_{1}=\mathbf{W}_{1}^{\prime}$ and $\mathbf{W}_{2}=\mathbf{W}_{2}^{\prime}$.
+Let us prove by backwards induction on $k=N,\ldots,1$ that $W_{1,k}=W_{1,k}^{\prime}$ and $W_{2,k}=W_{2,k}^{\prime}$.
+We start with $k=N$ as our base case. From Lemma 3.1 below, we have that
+$\left(H(\mathbf{W}_{1})H(\mathbf{W}_{2})\right)_{1,N}$ $=Z^{\mathbf{W}}(2,N\mid 1,1)$
+$=e^{W_{2,N}}\left(Z^{\mathbf{W}}(1,N\mid 1,1)+Z^{\mathbf{W}}(2,N-1\mid 1,1)\right)$
+$=e^{W_{2,N}}\Biggl(\prod_{i=1}^{N}e^{W_{1,i}}+\left(H(\mathbf{W}_{1})H(\mathbf{W}_{2})\right)_{1,N-1}\Biggr).$ (2.16)
+Similarly, we conclude that
+$\left(H(\mathbf{W}^{\prime}_{1})H(\mathbf{W}^{\prime}_{2})\right)_{1,N}$ $=Z^{\mathbf{W}^{\prime}}(2,N\mid 1,1)$
+$=e^{W^{\prime}_{2,N}}\left(Z^{\mathbf{W}^{\prime}}(1,N\mid 1,1)+Z^{\mathbf{W}^{\prime}}(2,N-1\mid 1,1)\right)$
+$=e^{W^{\prime}_{2,N}}\Biggl(\prod_{i=1}^{N}e^{W^{\prime}_{1,i}}+\left(H(\mathbf{W}^{\prime}_{1})H(\mathbf{W}^{\prime}_{2})\right)_{1,N-1}\Biggr).$ (2.17)
+By assumption, we have $H(\mathbf{W}_{1})H(\mathbf{W}_{2})=H(\mathbf{W}^{\prime}_{1})H(\mathbf{W}_{2}^{\prime})$ and $\prod_{i=1}^{N}e^{W_{1,i}}=e^{\mathfrak{s}(W_{1})}=e^{\mathfrak{s}(W^{\prime}_{1})}=\prod_{i=1}^{N}e^{W^{\prime}_{1,i}}$.
+Equating (2.16) and (2.17) and canceling the nonzero expressions in parentheses yields $W_{2,N}=W^{\prime}_{2,N}$.
+Next, we see that $W^{\prime}_{1,N}=W_{1,N}$, since:
+$(H(\mathbf{W}_{1})H(\mathbf{W}_{2}))_{(N,N)}=Z^{\mathbf{W}}(2,N\mid 1,N)=e^{W_{1,N}+W_{2,N}},$ (2.18)
+and
+$(H(\mathbf{W}^{\prime}_{1})H(\mathbf{W}_{2}^{\prime}))_{(N,N)}=Z^{\mathbf{W}^{\prime}}(2,N\mid 1,N)=e^{W^{\prime}_{1,N}+W^{\prime}_{2,N}}.$ (2.19)
+To finish the proof, simply notice that the proof of the above only required $\prod_{i=1}^{N}e^{W^{\prime}_{1,i}}=\prod_{i=1}^{N}e^{W_{1,i}}$.
+Indeed, if we have $\prod_{i=1}^{k}e^{W_{1,i}}=\prod_{i=1}^{k}e^{W_{2,i}}$ for some $k\in\llbracket 1,N\rrbracket$, then the same proof shows that $W_{1,k}=W_{1,k}^{\prime}$ and $W_{2,k}=W_{2,k}^{\prime}$.
+This follows by backwards induction on $k$, since
+$\prod_{i=1}^{k}e^{W_{1,i}}=e^{\mathfrak{s}(\mathbf{W}_{1})}\Biggl(\prod_{i=k+1}^{N}e^{W_{1,i}}\Biggr)^{-1}.\qed$
+3. Invariance of polymer partition functions
+In this section, we prove Theorem 1.4.
+The key to our proof is a matrix encoding of polymer partition functions, analogous to that developed in [NY04].
+In our setting, this is given by the following lemma.
+Lemma 3.1.
+Let $\mathbf{X}=(\mathbf{X}_{i})_{i\in\mathbb{Z}}\in(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$ and $k,\ell\in\mathbb{Z}$ such that $\ell\geq k$, then:
+$\Bigl(H(\mathbf{X}_{k})H(\mathbf{X}_{k+1})\cdots H(\mathbf{X}_{\ell})\Bigr)_{(i,j)}=Z^{\mathbf{X}}(\ell,j\mid k,i)\quad\forall i,j\in\mathbb{Z}.$
+In other words, the partition function across any two points $(k,i)$ and $(\ell,j)$ on the directed lattice $\mathbb{Z}^{2}$ with periodic weights $\mathbf{X}$ is given by the $(i,j)$th entry in the matrix product of the $H$ matrices.
 Proof.
 We prove the result by induction on $\ell\geq k$. For the base case, $\ell=k$, by definition we have that
-$H(\mathbf{W}_{k})_{(i,j)}=\begin{cases}\exp\Biggl{(}{\sum\limits_{r=i}^{j}}W_{(k,r)}\Biggr{)},\quad&j\geq i\\ 0,\ &\text{otherwise}\end{cases}$
+$H(\mathbf{X}_{k})_{(i,j)}=\begin{cases}\exp\Biggl({\sum\limits_{r=i}^{j}}X_{(k,r)}\Biggr),\quad&j\geq i\\ 0,\ &\text{otherwise}\end{cases}$
 Similarly, on a single column, we either have a unique up/right path between $(k,i)$ and $(k,j)$ picking up all the weights between $i$ and $j$ or have no path between them, depending on whether $j\geq i$.
 Thus,
-$Z^{W}(k,j\mid k,i)=\begin{cases}\exp\Biggl{(}{\sum\limits_{r=i}^{j}}W_{(k,r)}\Biggr{)},\quad&j\geq i\\ 0,\ &\text{otherwise}\end{cases}=H(\mathbf{W}_{k})_{(i,j)}$
+$Z^{X}(k,j\mid k,i)=\begin{cases}\exp\Biggl({\sum\limits_{r=i}^{j}}X_{(k,r)}\Biggr),\quad&j\geq i\\ 0,\ &\text{otherwise}\end{cases}=H(\mathbf{X}_{k})_{(i,j)}$
 This proves our base case. Now, let us assume that it holds for some $\ell$, and prove it holds for $\ell+1$. We split into two cases: $j\geq i$ and $j<i$.
@@
 Thus,
-$Z^{\mathbf{W}}(\ell+1,j\mid k,i)$ $=\sum\limits_{r=i}^{j}\Biggl{(}Z^{\mathbf{W}}(\ell,r\mid k,i)\exp\bigl{(}\sum_{s=r}^{j}W_{(\ell+1,s)}\bigr{)}\Biggr{)}$
-$=\sum\limits_{r=i}^{j}\Biggl{(}Z^{\mathbf{W}}(\ell,r\mid k,i)\Bigl{(}H(\mathbf{W}_{\ell+1})\Bigr{)}_{(r,j)}\Biggr{)}.$
+$Z^{\mathbf{X}}(\ell+1,j\mid k,i)$ $=\sum\limits_{r=i}^{j}\Biggl(Z^{\mathbf{X}}(\ell,r\mid k,i)\exp\Bigl(\sum_{s=r}^{j}X_{(\ell+1,s)}\Bigr)\Biggr)$
+$=\sum\limits_{r=i}^{j}\Biggl(Z^{\mathbf{X}}(\ell,r\mid k,i)\Bigl(H(\mathbf{X}_{\ell+1})\Bigr)_{(r,j)}\Biggr).$
 By our inductive hypothesis, we know that this equals
-$\sum\limits_{r=i}^{j}\Biggl{(}\Bigl{(}H(\mathbf{W}_{k})H(\mathbf{W}_{k+1})\ldots H(\mathbf{W}_{\ell})\Bigr{)}_{(i,r)}\Bigl{(}H(\mathbf{W}_{\ell+1})\Bigr{)}_{(r,j)}\Biggr{)}.$
-Remembering now that $\bigl{(}H(\mathbf{W}_{\ell+1})\bigr{)}_{(r,j)}=0$ for $r>j$, and $\bigl{(}H(\mathbf{W}_{k})H(\mathbf{W}_{k+1})\ldots H(\mathbf{W}_{\ell})\bigr{)}_{(i,r)}=0$ for $r<i$, we see that
-$\quad\;Z^{\mathbf{W}}(\ell+1,j\mid k,i)$
-$=\sum\limits_{r=-\infty}^{\infty}\Biggl{(}\Bigl{(}H(\mathbf{W}_{k})H(\mathbf{W}_{k+1})\ldots H(\mathbf{W}_{\ell})\Bigr{)}_{(i,r)}\Bigl{(}H(\mathbf{W}_{\ell+1})\Bigr{)}_{(r,j)}\Biggr{)}$
-$=\Bigl{(}H(\mathbf{W}_{k})H(\mathbf{W}_{k+1})\ldots H(\mathbf{W}_{\ell})H(\mathbf{W}_{\ell+1})\Bigr{)}_{(i,j)}.\qed$
+$\sum\limits_{r=i}^{j}\Biggl(\Bigl(H(\mathbf{X}_{k})H(\mathbf{X}_{k+1})\ldots H(\mathbf{X}_{\ell})\Bigr)_{(i,r)}\Bigl(H(\mathbf{X}_{\ell+1})\Bigr)_{(r,j)}\Biggr).$
+Remembering now that $\bigl(H(\mathbf{X}_{\ell+1})\bigr)_{(r,j)}=0$ for $r>j$, and $\bigl(H(\mathbf{X}_{k})H(\mathbf{X}_{k+1})\ldots H(\mathbf{X}_{\ell})\bigr)_{(i,r)}=0$ for $r<i$, we see that
+$Z^{\mathbf{X}}(\ell+1,j\mid k,i)$ $=\sum\limits_{r=-\infty}^{\infty}\Biggl(\Bigl(H(\mathbf{X}_{k})H(\mathbf{X}_{k+1})\ldots H(\mathbf{X}_{\ell})\Bigr)_{(i,r)}\Bigl(H(\mathbf{X}_{\ell+1})\Bigr)_{(r,j)}\Biggr)$
+$=\Bigl(H(\mathbf{X}_{k})H(\mathbf{X}_{k+1})\ldots H(\mathbf{X}_{\ell})H(\mathbf{X}_{\ell+1})\Bigr)_{(i,j)}.\qed$
 We now show the invariance of polymer partition function under the periodic Pitman transform.
-Retaining the periodic weight vectors $\mathbf{W}=(\mathbf{W}_{r}\in\mathbb{R}^{\mathbb{Z}_{N}}:r\in\mathbb{Z})$ as before, for $k\in\mathbb{Z}$, recall the operator $\mathcal{P}_{k}:(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}\to(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$ where
-$\mathcal{P}_{k}\mathbf{W}=(\ldots,\mathbf{W}_{k-1},T(\mathbf{W}_{k+1},\mathbf{W}_{k}),D(\mathbf{W}_{k+1},\mathbf{W}_{k}),\mathbf{W}_{k+2,\ldots}).$
+Retaining the periodic weight vectors $\mathbf{X}=(\mathbf{X}_{r}\in\mathbb{R}^{\mathbb{Z}_{N}}:r\in\mathbb{Z})$ as before, for $k\in\mathbb{Z}$, recall the operator $\mathcal{P}_{k}:(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}\to(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$ where
+$\mathcal{P}_{k}\mathbf{X}=(\ldots,\mathbf{X}_{k-1},T(\mathbf{X}_{k},\mathbf{X}_{k+1}),D(\mathbf{X}_{k},\mathbf{X}_{k+1}),\mathbf{X}_{k+2,\ldots}).$
 Proposition 3.2.
@@
 Then, for all $b,d\in\mathbb{Z}$,
-$Z^{\mathcal{P}_{k_{m}}\cdots\mathcal{P}_{k_{1}}\mathbf{W}}(c,d\mid a,b)=Z^{\mathbf{W}}(c,d\mid a,b).$
+$Z^{\mathcal{P}_{k_{m}}\cdots\mathcal{P}_{k_{1}}\mathbf{X}}(c,d\mid a,b)=Z^{\mathbf{X}}(c,d\mid a,b).$
 Proof.
 We will prove this by induction on $m$. By Lemma 3.1,
-$Z^{\mathbf{W}}(c,d\mid a,b)=\bigl{(}H(\mathbf{W}_{a})H(\mathbf{W}_{a+1})\cdots H(\mathbf{W}_{c})\bigr{)}_{(b,d)},$
+$Z^{\mathbf{X}}(c,d\mid a,b)=\bigl(H(\mathbf{X}_{a})H(\mathbf{X}_{a+1})\cdots H(\mathbf{X}_{c})\bigr)_{(b,d)},$
 which we claim is preserved by the action $\mathcal{P}_{k_{m}}\ldots\mathcal{P}_{k_{1}}$ for any $b,d\in\mathbb{Z}$.
-In our base case $n=1$, we split into three sub-cases: $k_{1}<a-1$, $k_{1}>c$, and $a\leq k_{1}<c$.
-In the first two sub-cases, $\mathbf{W}_{a}\ldots\mathbf{W}_{c}$ are unaffected by $\mathcal{P}_{k_{1}}$, so our invariance holds trivially.
-If $a\leq k_{1}<c$, then Corollary 2.5 gives that
-$Z^{\mathbf{W}}(c,d\mid a,b)$ $=\bigl{(}H(\mathbf{W}_{a})H(\mathbf{W}_{a+1})\ldots H(\mathbf{W}_{k_{1}})H(\mathbf{W}_{k_{1}+1})\ldots H(\mathbf{W}_{c})\bigr{)}_{(b,d)}$
+In our base case $m=1$, we split into three sub-cases: $k_{1}<a-1$, $k_{1}>c$, and $a\leq k_{1}<c$.
+In the first two sub-cases, $\mathbf{X}_{a}\ldots\mathbf{X}_{c}$ are unaffected by $\mathcal{P}_{k_{1}}$, so our invariance holds trivially.
+If $a\leq k_{1}<c$, then Proposition 2.7 gives that
+$Z^{\mathbf{X}}(c,d\mid a,b)$ $=\bigl(H(\mathbf{X}_{a})H(\mathbf{X}_{a+1})\ldots H(\mathbf{X}_{k_{1}})H(\mathbf{X}_{k_{1}+1})\ldots H(\mathbf{X}_{c})\bigr)_{(b,d)}$
 […]
-$=Z^{\mathcal{P}_{k_{1}}\mathbf{W}}(c,d\mid a,b).$
+$=Z^{\mathcal{P}_{k_{1}}\mathbf{X}}(c,d\mid a,b).$
 Now, we assume by way of induction that the invariance holds for m1m-1 transforms.
 Again, if $k_{m}<a-1$ or $k_{m}>c$ our equivalence holds trivially.
-Otherwise, let $\mathbf{W}^{\prime}=\mathcal{P}_{k_{m-1}}\ldots\mathcal{P}_{k_{1}}\mathbf{W}$.
+Otherwise, let $\mathbf{X}^{\prime}=\mathcal{P}_{k_{m-1}}\ldots\mathcal{P}_{k_{1}}\mathbf{X}$.
 Then
-$Z^{\mathbf{W}}(c,d\mid a,b)$ $=Z^{\mathbf{W}^{\prime}}(c,d\mid a,b)$
-$=Z^{\mathcal{P}_{k_{m}}\mathbf{W}^{\prime}}(c,d\mid a,b)$
-$=Z^{\mathcal{P}_{k_{m}}\mathcal{P}_{k_{m-1}}\ldots\mathcal{P}_{k_{1}}\mathbf{W}}(c,d\mid a,b).$
+$Z^{\mathbf{X}}(c,d\mid a,b)$ $=Z^{\mathbf{X}^{\prime}}(c,d\mid a,b)$
+$=Z^{\mathcal{P}_{k_{m}}\mathbf{X}^{\prime}}(c,d\mid a,b)$
+$=Z^{\mathcal{P}_{k_{m}}\mathcal{P}_{k_{m-1}}\ldots\mathcal{P}_{k_{1}}\mathbf{X}}(c,d\mid a,b).$
 This concludes our proof, noting that the second equality follows from the case $m=1$. ∎
-With Proposition 2.7 proved, we now prove Theorem 1.4 by extending to the multi-path result.
-Proof of Theorem 1.4..
-We use the Lindström-Gessel-Viennot (LGV) lemma (cf.
-[AZ18, Chapter 32], originally from [Lin73, GV85]) to prove the theorem.
-Because LGV applies only to edge-weighted graphs, we need to transfer our vertex weights to edge weights in the following way: for each vertex $(r,s)$, split it into $(r,s)_{\text{in}}$ and $(r,s)_{\text{out}}$, with edge $\bigl{(}(r,s)_{\text{in}},(r,s)_{\text{out}}\bigr{)}$ given the weight $\omega_{(r,s)}:=e^{W_{(r,s)}}$, and edges $\bigl{(}(r,s)_{\text{out}},(r,s+1)_{\text{in}}\bigr{)}$ and $\bigl{(}(r,s)_{\text{out}},(r+1,s)_{\text{in}}\bigr{)}$ given the weight 1.
-See Figure 5.
-Any path which started at $(r,s)$ starts at $(r,s)_{\text{in}}$, while any path ending at $(r,s)$ ends at $(r,s)_{\text{out}}$.
-Under this graph transformation, our partition function on the pre-transformed graph equals the sum over weight products along possible paths on the transformed diagram, so we can apply LGV.
-[…]
-Figure 5.
-Local transformation of a vertex-weighted node (left) with weight $W_{(r,s)}$ into an edge-weighted directed graph (right), where the internal edge carries weight $\omega_{(r,s)}=\exp(W_{(r,s)})$ and external edges carry weight 1.
-Paths in the original graph correspond to paths from “in” nodes to “out” nodes in the transformed graph, preserving path weights.
-Write the path matrix $M=\bigl{(}Z^{\mathbf{W}}(c_{j},d_{j}\mid a_{i},b_{i})\bigr{)}_{(i,j)\in\llbracket 1,m\rrbracket}$. By LGV, we have
-$\det M=\sum\limits_{P=(P_{1}\ldots P_{m})\in\Pi_{U,V_{\sigma}}\ \forall\sigma\in S_{m}}\operatorname{sgn}(\sigma)\prod\limits_{(i,j)\in P}e^{W_{(i,j)}}.$
-By definition of $\Psi$, for $\sigma\neq\text{Id}$, $\Pi_{U,V_{\sigma}}=\varnothing$. Because $\operatorname{sgn}(\text{Id})=1$, we have that
-$\det M=\sum\limits_{P=(P_{1}\ldots P_{m})\in\Pi_{U,V}}\prod\limits_{(i,j)\in P}e^{W_{(i,j)}}=Z^{\mathbf{W}}(V\mid U).$
-By Proposition 2.7, each $M_{i,j}=Z^{\mathbf{W}}(c_{j},d_{j}\mid a_{i},b_{i})$ remains invariant under $\mathcal{P}_{k_{m}}\cdots\mathcal{P}_{k_{1}}$, so $M$ is unchanged, and so $\det M$ is invariant.
-Thus, $Z^{\mathbf{W}}(V\mid U)=\det M$ is invariant.
-∎
-3. Braid relations: proof of Theorem 1.6
-In this section, we prove the braid relation stated in Theorem 1.6. Notice that this relation translates into the following three equalities:
-$T\Bigl{(}T\bigl{(}\mathbf{W}_{k+2},D(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)},T(\mathbf{W}_{k+1},\mathbf{W}_{k})\biggr{)}=T\Bigl{(}T(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),\mathbf{W}_{k}\Bigr{)},$ (3.1)
-$D\Bigl{(}T\bigl{(}\mathbf{W}_{k+2},D(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)},T(\mathbf{W}_{k+1},\mathbf{W}_{k})\Bigr{)}=T\Bigl{(}D(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),D\bigl{(}T(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),\mathbf{W}_{k}\bigr{)}\Bigr{)},$ (3.2)
-$D\Bigl{(}\mathbf{W}_{k+2},D(\mathbf{W}_{k+1},\mathbf{W}_{k})\Bigr{)}=D\Bigl{(}D(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),D\bigl{(}T(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),\mathbf{W}_{k}\bigr{)}\Bigr{)}.$ (3.3)
-Equation (3.3) follows from [CGS24, Proposition 5.13], as explained in Lemma A.8 of the appendix.
-We now use the partition function invariance, Theorem 1.4, to demonstrate that (3.1) implies both (3.2) and (3.3).
-The main tool will be the following lemma:
+With Proposition 3.2 proved, we now give one more intermediate lemma before proving Theorem 1.4.
 Lemma 3.3.
-For $\mathbf{Y}=(\mathbf{Y}_{i})_{i\in\mathbb{Z}},\mathbf{Y}^{\prime}=(\mathbf{Y}^{\prime}_{i})_{i\in\mathbb{Z}}\in(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$ satisfying $\mathfrak{s}(\mathbf{Y}_{1})=\mathfrak{s}(\mathbf{Y}^{\prime}_{1})$, $\mathfrak{s}(\mathbf{Y}_{2})=\mathfrak{s}(\mathbf{Y}_{2}^{\prime})$, we have that $H(\mathbf{Y}_{1})H(\mathbf{Y}_{2})=H(\mathbf{Y}^{\prime}_{1})H(\mathbf{Y}_{2}^{\prime})$ if and only if $\mathbf{Y}_{1}=\mathbf{Y}^{\prime}_{1},\mathbf{Y}_{2}=\mathbf{Y}_{2}^{\prime}$.
-Before proving Lemma 3.1, we first use Lemmas A.8 and 3.1 to complete the proof of Theorem 1.6:
-Proof of Theorem 1.6..
-Using the partition function invariance under the operator $\mathcal{P}_{i+1}\mathcal{P}_{i}\mathcal{P}_{i+1}$ yields:
-$H\bigl{(}\mathbf{W}_{k}\bigr{)}H\bigl{(}\mathbf{W}_{k+1}\bigr{)}H\bigl{(}\mathbf{W}_{k+2}\bigr{)}=$
-[…]
-[…]
-[…]
-$\qquad H\Bigl{(}D\bigl{(}D(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),D\bigl{(}T(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),\mathbf{W}_{k}\bigr{)}\bigr{)}\Bigr{)}.$ (3.4)
-On the other hand, we can also apply the partition function invariance under the operator $\mathcal{P}_{i}\mathcal{P}_{i+1}\mathcal{P}_{i}$:
-$H\bigl{(}\mathbf{W}_{k}\bigr{)}H\bigl{(}\mathbf{W}_{k+1}\bigr{)}H\bigl{(}\mathbf{W}_{k+2}\bigr{)}=$
-[…]
-[…]
-[…]
-$\qquad H\Bigl{(}D\bigl{(}T\bigl{(}\mathbf{W}_{k+2},D(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)},T(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)}\Bigr{)}H\Bigl{(}D(\mathbf{W}_{k+2},D(\mathbf{W}_{k+1},\mathbf{W}_{k}))\Bigr{)}.$ (3.5)
-From Lemma A.8, we know that the last terms in (3.4) and (3.5) are equal. Because $H$ is invertible (see Lemma 2.2), we can equate both equations and cancel out the last term, obtaining that:
-$H\Bigl{(}T\bigl{(}T(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),\mathbf{W}_{k}\bigr{)}\Bigr{)}H\Bigl{(}T\bigl{(}D(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),D\bigl{(}T(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),\mathbf{W}_{k}\bigr{)}\Bigr{)}$
-$=H\Bigl{(}T\bigl{(}T\bigl{(}\mathbf{W}_{k+2},D(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)},T(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)}\Bigr{)}H\Bigl{(}D\bigl{(}T\bigl{(}\mathbf{W}_{k+2},D(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)},T(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)}\Bigr{)}.$
-By repeated application of Lemma A.4, we have that
-$\mathfrak{s}(T(T(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),\mathbf{W}_{k}))=\mathfrak{s}(\mathbf{W}_{k+2})$
-$=\mathfrak{s}(T(T(\mathbf{W}_{k+2},D(\mathbf{W}_{k+1},\mathbf{W}_{k})),T(\mathbf{W}_{k+1},\mathbf{W}_{k}))),\quad\text{and}$
-$\mathfrak{s}(T(D(T(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),\mathbf{W}_{k})),D(\mathbf{W}_{k+2},\mathbf{W}_{k+1})))$
-$=\mathfrak{s}(\mathbf{W}_{k+1})=\mathfrak{s}(T\bigl{(}D(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),D\bigl{(}T(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),\mathbf{W}_{k}\bigr{)})).$
-Applying Lemma 3.1, we conclude (3.2) and (3.3), as we wanted to prove. ∎
-Proof of Lemma 3.1..
-Immediately, the backwards direction is clear.
-Then, it remains to show that, under the assumption that $H(\mathbf{Y}_{1})H(\mathbf{Y}_{2})=H(\mathbf{Y}^{\prime}_{1})H(\mathbf{Y}_{2}^{\prime})\implies\mathbf{Y}_{1}=\mathbf{Y}^{\prime}_{1},\mathbf{Y}_{2}=\mathbf{Y}_{2}^{\prime}$.
-Let us prove by backwards induction on $k=N,\ldots,1$ that $Y_{1,k}=Y^{\prime}_{1,k}$ and $Y_{2,k}=Y^{\prime}_{2,k}$.
-As a base case, let’s show this for $k=N$. From Lemma 2.6, note that
-$\left(H(\mathbf{Y}_{1})H(\mathbf{Y}_{2})\right)_{1,N}$ $=Z^{\mathbf{Y}}(2,N\mid 1,1)$
-$=e^{Y_{2,N}}\left(Z^{\mathbf{Y}}(1,N\mid 1,1)+Z^{\mathbf{Y}}(2,N-1\mid 1,1)\right)$
-$=e^{Y_{2,N}}\Biggl{(}\prod_{i=1}^{N}e^{Y_{1,i}}+\left(H(\mathbf{Y}_{1})H(\mathbf{Y}_{2})\right)_{1,N-1}\Biggr{)}.$ (3.6)
-Similarly, we conclude that:
-$\left(H(\mathbf{Y}^{\prime}_{1})H(\mathbf{Y}^{\prime}_{2})\right)_{1,N}$ $=Z^{\mathbf{Y}^{\prime}}(2,N\mid 1,1)$
-$=e^{Y^{\prime}_{2,N}}\left(Z^{\mathbf{Y}^{\prime}}(1,N\mid 1,1)+Z^{\mathbf{Y}^{\prime}}(2,N-1\mid 1,1)\right)$
-$=e^{Y^{\prime}_{2,N}}\Biggl{(}\prod_{i=1}^{N}e^{Y^{\prime}_{1,i}}+\left(H(\mathbf{Y}^{\prime}_{1})H(\mathbf{Y}^{\prime}_{2})\right)_{1,N-1}\Biggr{)}.$ (3.7)
-By assumption, we have $H(\mathbf{Y}_{1})H(\mathbf{Y}_{2})=H(\mathbf{Y}^{\prime}_{1})H(\mathbf{Y}_{2}^{\prime})$ and $\prod_{i=1}^{N}e^{Y_{1,i}}=e^{\mathfrak{s}(Y_{1})}=e^{\mathfrak{s}(Y^{\prime}_{1})}=\prod_{i=1}^{N}e^{Y^{\prime}_{1,i}}$.
-Equating (3.6) and (3.7) and canceling the nonzero expressions in parenthesis yields $Y_{2,N}=Y^{\prime}_{2,N}$.
-Next, we see that $Y^{\prime}_{1,N}=Y_{1,N}$, since:
-$(H(\mathbf{Y}_{1})H(\mathbf{Y}_{2}))_{(N,N)}=Z^{\mathbf{Y}}(2,N\mid 1,N)=e^{Y_{1,N}+Y_{2,N}},$ (3.8)
-and
-$(H(\mathbf{Y}^{\prime}_{1})H(\mathbf{Y}_{2}^{\prime}))_{(N,N)}=Z^{\mathbf{Y}^{\prime}}(2,N\mid 1,N)=e^{Y^{\prime}_{1,N}+Y^{\prime}_{2,N}}.$ (3.9)
-To finish the proof, simply notice that the proof of the above only required $\prod_{i=1}^{N}e^{Y^{\prime}_{i}}=\prod_{i=1}^{N}e^{Y_{i}}$.
-Alternatively, if we have $\prod_{i=1}^{k}e^{Y_{1,i}}=\prod_{i=1}^{k}e^{Y_{2,i}}$ for some $k\in\llbracket 1,N\rrbracket$, then the same proof shows that $Y_{1,k}=Y_{1,k}^{\prime}$ and $Y_{2,k}=Y_{2,k}^{\prime}$.
-This follows by backwards induction, since
-$\prod_{i=1}^{k}e^{Y_{1,i}}=\mathfrak{s}(\mathbf{Y}_{1})\Bigl{(}\prod_{i=k+1}^{N}e^{Y_{1,i}}\Bigr{)}^{-1}.\qed$
-4. Burke property and permutation invariance
-In this section, we prove Theorems 1.7 and 1.8. We first need two intermediate lemmas. In the first, we generalize the Burke property in [CGS24, Proposition 5.12] to an inhomogeneous environment.
-Proposition 4.1 (Log-inverse-gamma Burke property).
-Let $\lambda>0$, and let $\mathbf{a},\mathbf{b}\in\mathbb{R}^{\mathbb{Z}_{N}}$ be sequences such that $a_{i}+b_{j}>0$ for $i,j\in\mathbb{Z}_{N}$.
-Let $\mathbf{X}_{1},\mathbf{X}_{2}$ be independent sequences of independent random variables such that $X_{i,j}\sim(a_{i}+b_{j},\lambda)$ (see Definition 1.15).
-Then,
-$(\mathbf{X}_{1},\mathbf{X}_{2})\overset{d}{=}(T(\mathbf{X}_{1},\mathbf{X}_{2}),D(\mathbf{X}_{1},\mathbf{X}_{2})).$
-Proof.
-Using Lemmas A.4-A.5 in the penultimate equality below, the joint density $\mathbb{P}(\mathbf{X}_{1}\in d\mathbf{x}_{1},\mathbf{X}_{2}\in d\mathbf{x}_{2})$ is proportional to
-$\exp\Biggl{(}-\sum_{i\in\mathbb{Z}_{N}}\Bigl{(}(a_{1}+b_{i})x_{1,i}+(a_{2}+b_{i})x_{2,i}\Bigr{)}-\lambda\sum_{i\in\mathbb{Z}_{N}}(e^{-x_{1,i}}+e^{-x_{2,i}})\Biggr{)}$
-$=\exp\Biggl{(}-a_{1}\sum_{i\in\mathbb{Z}_{N}}x_{1,i}-a_{2}\sum_{i\in\mathbb{Z}_{N}}x_{2,i}-\sum_{i\in\mathbb{Z}_{N}}b_{i}(x_{1,i}+x_{2,i})-\lambda\sum_{i\in\mathbb{Z}_{N}}(e^{-x_{1,i}}+e^{-x_{2,i+1}})\Biggr{)}$
-$=\exp\Biggl{(}-a_{1}\sum_{i\in\mathbb{Z}_{N}}T_{i}(\mathbf{x}_{1},\mathbf{x}_{2})-a_{2}\sum_{i\in\mathbb{Z}_{N}}D_{i}(\mathbf{x}_{1},\mathbf{x}_{2})-\sum_{i\in\mathbb{Z}_{N}}b_{i}(T_{i}(\mathbf{x}_{1},\mathbf{x}_{2})+D_{i}(\mathbf{x}_{1},\mathbf{x}_{2}))$
-$\qquad\qquad\qquad\qquad-\lambda\sum_{i\in\mathbb{Z}_{N}}(e^{-T_{i+1}(\mathbf{x}_{1},\mathbf{x}_{2})}+e^{-D_{i}(\mathbf{x}_{1},\mathbf{x}_{2})})\Biggr{)}$
-$=\exp\Biggl{(}-\sum_{i\in\mathbb{Z}_{N}}\Bigl{(}(a_{1}+b_{i})T_{i}(\mathbf{x}_{1},\mathbf{x}_{2})+(a_{2}+b_{i})D_{i}(\mathbf{x}_{1},\mathbf{x}_{2})\Bigr{)}-\lambda\sum_{i\in\mathbb{Z}_{N}}(e^{-T_{i}(\mathbf{x}_{1},\mathbf{x}_{2})}+e^{-D_{i}(\mathbf{x}_{1},\mathbf{x}_{2})})\Biggr{)},$
-Thus, the density is preserved under our transformation, so all we must show is that the determinant of the Jacobian matrix of the transformation has absolute value 1. This follows from Lemma A.7. ∎
-We next prove a lemma that connects the condition stated in Theorem 1.7 to the condition in Theorem 1.4.
-Lemma 4.2.
-Let $\sigma:\mathbb{Z}\to\mathbb{Z}$ be a finite permutation, decomposed as a product of nearest-neighbor transpositions $(k_{m},k_{m}+1)(k_{m-1},k_{m-1}+1)\cdots(k_{1},k_{1}+1)$, where $k_{1},\ldots,k_{n}\in\mathbb{Z}$.
+Let $\sigma:\mathbb{Z}\to\mathbb{Z}$ be a finite permutation, decomposed as a product of nearest-neighbor transpositions $(k_{m},k_{m}+1)(k_{m-1},k_{m-1}+1)\cdots(k_{1},k_{1}+1)$, where $k_{1},\ldots,k_{m}\in\mathbb{Z}$.
 Assume that this decomposition has the minimal number of terms.
-Then, if $x\in\mathbb{Z}$ satisfies $\sigma(\mathbb{Z}_{<x})=\mathbb{Z}_{<x}$, then $x\neq k_{\ell}+1$ for all $1\leq\ell\leq n$.
+Then, if $x\in\mathbb{Z}$ satisfies $\sigma(\mathbb{Z}_{<x})=\mathbb{Z}_{<x}$, then $x\neq k_{\ell}+1$ for all $1\leq\ell\leq m$.
 Similarly, if $\sigma(\mathbb{Z}_{>x})=\mathbb{Z}_{>x}$, then $x\neq k_{\ell}$ for all $1\leq\ell\leq m$.
@@
 Proof of Theorem 1.4..
-For an arbitrary finite permutation $\sigma$, decompose it as a product of nearest-neighbor transpositions $\sigma=(k_{m},k_{m}+1)(k_{m-1},k_{m-1}+1)\cdots(k_{1},k_{1}+1)$.
-Take this decomposition to have the minimal number of terms.
-Then, we consider the action of the map $\mathcal{P}_{k_{m}}\cdots\mathcal{P}_{k_{1}}$ on the weights $\mathbf{W}^{\mathbf{a},\mathbf{b}}$, as in Theorem 1.4.
-By Lemma 4.2, the condition (1.11) and the minimality of the decomposition ensures that the conditions of Theorem 1.4 are met, so
-$Z^{\mathcal{P}_{k_{m}}\cdots\mathcal{P}_{k_{1}}\mathbf{W}^{\mathbf{a},\mathbf{b}}}(V\mid U)=Z^{\mathbf{W}^{\mathbf{a},\mathbf{b}}}(V\mid U),\quad\text{for all }(U,V)\in\mathcal{U}^{\sigma}.$
-Furthermore, by successively applying the Burke property in Proposition 4.1, we see that each operator $\mathcal{P}_{k}$ swaps the column parameters in columns $k$ and $k+1$, so we have that $\mathcal{P}_{k_{m}}\cdots\mathcal{P}_{k_{1}}\mathbf{W}^{\mathbf{a},\mathbf{b}}\overset{d}{=}\mathbf{W}^{\sigma\mathbf{a},\mathbf{b}}$.
+Let $\sigma\in\mathcal{S}_{\mathbb{Z}}$, and let $(U,V)\in\mathcal{W}^{\sigma}$ (recalling the definition of this set in (1.10)).
+Denote the points of $U$ and $V$ as $U=\bigl((a_{i},b_{i})\bigr)_{i\in\llbracket 1,k\rrbracket}$, $V=\bigl((c_{i},d_{i})\bigr)_{i\in\llbracket 1,k\rrbracket}$.
+Let $\sigma\in\mathcal{S}_{\mathbb{Z}}$.
+By assumption that $(U,V)\in\mathcal{W}^{\sigma}$, Lemma 3.3 implies that the conditions of Proposition 3.2 are satisfied so that, for each $i\in\llbracket 1,k\rrbracket$,
+$Z^{\mathbf{X}}(c_{j},d_{j}\mid a_{i},b_{i})=Z^{\mathcal{P}_{\sigma}\mathbf{X}}(c_{j},d_{j}\mid a_{i},b_{i}).$ (3.1)
+To extend this to the multi-path result, we use the Lindström-Gessel-Viennot (LGV) lemma (cf.
+[AZ18, Chapter 32], originally from [LIN73, GV85]) to prove the theorem.
+Because LGV applies only to edge-weighted graphs, we need to transfer our vertex weights to edge weights in the following way: for each vertex $(r,s)$, split it into $(r,s)_{\text{in}}$ and $(r,s)_{\text{out}}$, with edge $\bigl((r,s)_{\text{in}},(r,s)_{\text{out}}\bigr)$ given the weight $\omega_{(r,s)}:=e^{X_{(r,s)}}$, and edges $\bigl((r,s)_{\text{out}},(r,s+1)_{\text{in}}\bigr)$ and $\bigl((r,s)_{\text{out}},(r+1,s)_{\text{in}}\bigr)$ given the weight 1.
+See Figure 5.
+Any path which started at $(r,s)$ starts at $(r,s)_{\text{in}}$, while any path ending at $(r,s)$ ends at $(r,s)_{\text{out}}$.
+Under this graph transformation, our partition function on the pre-transformed graph equals the sum over weight products along possible paths on the transformed diagram, so we can apply LGV.
+[…]
+Figure 5.
+Local transformation of a vertex-weighted node (left) with weight $X_{(r,s)}$ into an edge-weighted directed graph (right), where the internal edge carries weight $\omega_{(r,s)}=\exp(X_{(r,s)})$ and external edges carry weight 1.
+Paths in the original graph correspond to paths from “in” nodes to “out” nodes in the transformed graph, preserving path weights.
+Define the matrix $M^{\mathbf{X}}:=\bigl(Z^{\mathbf{X}}(c_{j},d_{j}\mid a_{i},b_{i})\bigr)_{(i,j)\in\llbracket 1,m\rrbracket}$. By LGV, we have
+$\det M^{\mathbf{X}}=\sum\limits_{P=(P_{1}\ldots P_{m})\in\Pi_{U,V_{\tau}}\ \forall\varphi\in\mathcal{S}_{m}}\operatorname{sgn}(\varphi)\prod\limits_{(i,j)\in P}e^{X_{(i,j)}}.$
+Since $\mathcal{W}^{\sigma}\subseteq\Psi$, we have that, by definition of $\Psi$ (1.7), for $\varphi\neq\text{Id}$, $\Pi_{U,V_{\varphi}}=\varnothing$. Then, because $\operatorname{sgn}(\text{Id})=1$,
+$\det M^{\mathbf{X}}=\sum\limits_{P=(P_{1}\ldots P_{m})\in\Pi_{U,V}}\prod\limits_{(i,j)\in P}e^{X_{(i,j)}}=Z^{\mathbf{X}}(V\mid U).$ (3.2)
+By (3.1), for each $i,j$, we have $M_{i,j}^{\mathbf{X}}=M_{i,j}^{\mathcal{P}_{\sigma}\mathbf{X}}$. Therefore, by (3.2), we also have
+$Z^{\mathcal{P}_{\sigma}\mathbf{X}}(V\mid U)=\det M^{\mathcal{P}_{\sigma}\mathbf{X}}=\det M^{\mathbf{X}}=Z^{\mathbf{X}}(V\mid U).\qed$
+4. Burke property and permutation invariance
+In this section, we prove Theorems 1.5 and 1.6. We first need two intermediate lemmas. In the first, we generalize the Burke property in [CGS26, Proposition 5.12] to an inhomogeneous environment.
+Proposition 4.1 (Log-inverse-gamma Burke property).
+Let $\lambda>0$, let $a_{1},a_{2}\in\mathbb{R}$, and let $\mathbf{b}\in\mathbb{R}^{\mathbb{Z}_{N}}$ be a sequence such that $a_{m}+b_{i}>0$ for $m\in\{1,2\}$ and $i\in\mathbb{Z}_{N}$.
+Let $\mathbf{X}_{1},\mathbf{X}_{2}$ be independent sequences of independent random variables such that $X_{m,i}\sim\text{Log-Inv-Gamma}(a_{m}+b_{i},\lambda)$ (see Definition 1.13).
+Then,
+$(\mathbf{X}_{1},\mathbf{X}_{2})\overset{d}{=}(D(\mathbf{X}_{1},\mathbf{X}_{2}),T(\mathbf{X}_{1},\mathbf{X}_{2})).$
+Proof.
+Using Lemmas 2.1-2.2 in the penultimate equality below, the joint density $\mathbb{P}(\mathbf{X}_{1}\in d\mathbf{x}_{1},\mathbf{X}_{2}\in d\mathbf{x}_{2})$ is proportional to
+$\exp\Biggl(-\sum_{i\in\mathbb{Z}_{N}}\Bigl((a_{1}+b_{i})x_{1,i}+(a_{2}+b_{i})x_{2,i}\Bigr)-\lambda\sum_{i\in\mathbb{Z}_{N}}(e^{-x_{1,i}}+e^{-x_{2,i}})\Biggr)$
+$=\exp\Biggl(-a_{1}\sum_{i\in\mathbb{Z}_{N}}x_{1,i}-a_{2}\sum_{i\in\mathbb{Z}_{N}}x_{2,i}-\sum_{i\in\mathbb{Z}_{N}}b_{i}(x_{1,i}+x_{2,i})-\lambda\sum_{i\in\mathbb{Z}_{N}}(e^{-x_{2,i}}+e^{-x_{1,i+1}})\Biggr)$
+$=\exp\Biggl(-a_{1}\sum_{i\in\mathbb{Z}_{N}}D_{i}(\mathbf{x}_{1},\mathbf{x}_{2})-a_{2}\sum_{i\in\mathbb{Z}_{N}}T_{i}(\mathbf{x}_{1},\mathbf{x}_{2})-\sum_{i\in\mathbb{Z}_{N}}b_{i}(T_{i}(\mathbf{x}_{1},\mathbf{x}_{2})+D_{i}(\mathbf{x}_{1},\mathbf{x}_{2}))$
+$\qquad\qquad\qquad\qquad-\lambda\sum_{i\in\mathbb{Z}_{N}}(e^{-T_{i+1}(\mathbf{x}_{1},\mathbf{x}_{2})}+e^{-D_{i}(\mathbf{x}_{1},\mathbf{x}_{2})})\Biggr)$
+$=\exp\Biggl(-\sum_{i\in\mathbb{Z}_{N}}\Bigl((a_{1}+b_{i})D_{i}(\mathbf{x}_{1},\mathbf{x}_{2})+(a_{2}+b_{i})T_{i}(\mathbf{x}_{1},\mathbf{x}_{2})\Bigr)-\lambda\sum_{i\in\mathbb{Z}_{N}}(e^{-D_{i}(\mathbf{x}_{1},\mathbf{x}_{2})}+e^{-T_{i}(\mathbf{x}_{1},\mathbf{x}_{2})})\Biggr).$
+Thus, the density is preserved under our transformation, so all we must show is that the determinant of the Jacobian matrix of the transformation has absolute value 1.
+This was shown previously in [CGS26], recorded in the present paper as Lemma A.1.
+∎
+Proof of Theorem 1.5..
+For an arbitrary finite permutation $\sigma$, Theorem 1.4 implies that
+$Z^{\mathcal{P}_{\sigma}\mathbf{X}^{\mathbf{a},\mathbf{b}}}(V\mid U)=Z^{\mathbf{X}^{\mathbf{a},\mathbf{b}}}(V\mid U),\quad\text{for all }(U,V)\in\mathcal{W}^{\sigma}.$
+Furthermore, by successively applying the Burke property in Proposition 4.1, we see that each operator $\mathcal{P}_{k}$ swaps the column parameters in columns $k$ and $k+1$, so we have that $\mathcal{P}_{\sigma}\mathbf{X}^{\mathbf{a},\mathbf{b}}\overset{d}{=}\mathbf{X}^{\sigma\mathbf{a},\mathbf{b}}$.
 This completes the proof.
@@
 Consider the independent weights in the finite grid $\llbracket-M,M\rrbracket\times\llbracket-M,M\rrbracket$, and extend these periodically.
-Consider first endpoint pairs $(U,V)$ that consist of points lying in the finite grid $\llbracket-M,M\rrbracket\times\llbracket-M,M\rrbracket$, and apply the partition function invariance in Theorem 1.4 and the Burke property in Proposition 4.1 (with Lemma 4.2 to connect the two conditions) to both rows and columns, with $N=2M+1$.
+Consider first end point pairs $(U,V)$ that consist of points lying in the finite grid $\llbracket-M,M\rrbracket\times\llbracket-M,M\rrbracket$, and apply the partition function invariance in Theorem 1.4 and the Burke property in Proposition 4.1 to both rows and columns, with $N=2M+1$.
 The result follows by sending $M\to\infty$.
 ∎
-Appendix A Inputs from previous work and technical inputs
-A.1. Basic algebraic properties
-In the present work, we use a different formulation of the maps $D$ and $T$ than those introduced in [CGS24]; there, the relevant mappings were defined as
-$\mathfrak{T}_{i}(\mathbf{\mathbf{X}_{1}},\mathbf{\mathbf{X}_{2}})$ $=X_{1,i}+\log\Biggl{(}\frac{\sum_{j\in\mathbb{Z}_{N}}e^{X_{2,[i,j]}-X_{1,[i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{X_{2,[i+1,j]-X_{1,[i+1,j]}}}}\Biggr{)},\quad\text{and}$
-$\mathfrak{D}_{i}(\mathbf{\mathbf{X}_{1}},\mathbf{\mathbf{X}_{2}})$ $=X_{2,i}+\log\Biggl{(}\frac{\sum_{j\in\mathbb{Z}_{N}}e^{X_{2,(i,j]}-X_{1,(i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{X_{2,(i-1,j]}-X_{1,(i-1,j]}}}\Biggr{)}.$
-In [CGS24], these maps were denoted as $T^{N,2}$ and $D^{N,2}$, respectively. It is important to note that this is equivalent to our definition of the maps defined in (1.2) up to a shift:
-$D(\mathbf{X}_{1},\mathbf{X}_{2})=\mathfrak{D}(\mathbf{X}_{1},\tau_{1}\mathbf{X}_{2}),\quad\text{and}\quad T(\mathbf{X}_{1},\mathbf{X}_{2})=\tau_{-1}\mathfrak{T}(\mathbf{X}_{1},\tau_{1}\mathbf{X}_{2}).$ (A.1)
-where, for $k\in\mathbb{Z}_{N}$, $\tau_{k}:\mathbb{R}^{\mathbb{Z}_{N}}\rightarrow\mathbb{R}^{\mathbb{Z}_{N}}$ is the shift map: $(\tau_{k}X)_{i}=X_{i+k}$ for all $i\in\mathbb{Z}_{N}$.
+Appendix A Inputs from previous results and zero-temperature analogues
+A.1. Additional algebraic properties
+In [CGS26], index-wise shifted versions of our $T$ and $D$ maps were denoted by $T^{N,2}$ and $D^{N,2}$ respectively. To be precise, the definition of the mappings in that paper were
+$T_{i}^{N,2}(\mathbf{\mathbf{X}_{1}},\mathbf{\mathbf{X}_{2}})$ $=X_{1,i}+\log\Biggl(\frac{\sum_{j\in\mathbb{Z}_{N}}e^{X_{2,[i,j]}-X_{1,[i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{X_{2,[i+1,j]}-X_{1,[i+1,j]}}}\Biggr),\quad\text{and}$
+$D_{i}^{N,2}(\mathbf{\mathbf{X}_{1}},\mathbf{\mathbf{X}_{2}})$ $=X_{2,i}+\log\Biggl(\frac{\sum_{j\in\mathbb{Z}_{N}}e^{X_{2,(i,j]}-X_{1,(i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{X_{2,(i-1,j]}-X_{1,(i-1,j]}}}\Biggr).$
+In the present paper, we subsumed the dependence on $N$. From the definitions given in the present paper (1.2), we can see that
+$T(\mathbf{X}_{1},\mathbf{X}_{2})=\tau^{-1}T^{N,2}(\mathbf{X}_{2},\tau\mathbf{X}_{1}),\quad\text{and}\quad D(\mathbf{X}_{1},\mathbf{X}_{2})=D^{N,2}(\mathbf{X}_{2},\tau\mathbf{X}_{1}),$ (A.1)
+where $\tau:\mathbb{R}^{\mathbb{Z}_{N}}\to\mathbb{R}^{\mathbb{Z}_{N}}$ is the shift map defined by $(\tau\mathbf{X})_{i}=X_{i+1}$.
+With this in mind, we now cite two results from [CGS26], which follow from translating them into appropriate statements for the present paper via the connection in (A.1).
 Lemma A.1.
-As $\beta\to\infty$, $\frac{1}{\beta}\log\Bigl{(}\sum\limits_{i=0}^{k}\exp(\beta E_{i})\Bigr{)}$ converges to $\max\limits_{i}E_{i}$, uniformly as a function of the $E_{i}$ (but not as a function of $k$)
+[CGS26, Proposition 5.12] The determinant of the Jacobian of the transformation
+$\mathcal{P}:\mathbb{R}^{\mathbb{Z}_{N}}\to\mathbb{R}^{\mathbb{Z}_{N}}\text{ defined by }(\mathbf{X}_{1},\mathbf{X}_{2})\to\bigl(T(\mathbf{X}_{1},\mathbf{X}_{2}),D(\mathbf{X}_{1},\mathbf{X}_{2})\bigr)$
+has absolute value $1$.
+Lemma A.2.
+[CGS26, Proposition 5.13] For $\mathbf{X}_{1},\mathbf{X}_{2},\mathbf{X}_{3}\in\mathbb{R}^{\mathbb{Z}_{N}}$,
+$D\bigl(D(\mathbf{X}_{1},\mathbf{X}_{2}),\mathbf{X}_{3}\bigr)=D\bigl(D(\mathbf{X}_{1},T(\mathbf{X}_{2},\mathbf{X}_{3})),D(\mathbf{X}_{2},\mathbf{X}_{3})).$
+A.2. Proofs of the zero-temperature analogues in Theorems 1.8-1.10
+Because polymer models have a strong connection to last-passage percolation (LPP), our algebraic results will naturally generalize to zero-temperature equivalents, through limiting arguments.
+In these limits, the $(+,\times)$ algebra is swapped for the $(\max,+)$ algebra, and our maps reflect this change.
+As introduced in [CGS26], the ‘correct’ map to look at here is $\widetilde{D}$, the uniform limit as $\beta$ goes to infinity of $\frac{1}{\beta}D(\beta\mathbf{X}_{1},\beta\mathbf{X}_{2})$, as well as $\widetilde{T}$, which is reached through an identical limit.
+In this section, we use our algebraic results about the behavior of the polymer model under the periodic Pitman transform (as well as the aforementioned limiting arguments) to prove analogous facts about LPP under the zero-temperature periodic Pitman transform.
+The following intermediate results are needed.
+Lemma A.3.
+For $y_{1},\ldots,y_{k}\in\mathbb{R}$, as $\beta\to\infty$, $\frac{1}{\beta}\log\Bigl(\sum\limits_{i=0}^{k}\exp(\beta y_{i})\Bigr)$ converges to $\max\limits_{i}y_{i}$, uniformly as a function of the $y_{i}$.
+Lemma A.4.
+For any $k\in\mathbb{N}$ and any pair of $k$-tuples of ordered pairs $U=\bigl((a_{i},b_{i})\bigr)_{i\in\llbracket 1,k\rrbracket}$, $V=\bigl((c_{i},d_{i})\bigr)_{i\in\llbracket 1,k\rrbracket}$, as $\beta\to\infty$, $\frac{1}{\beta}\log\bigl(Z^{\beta,\mathbf{X}}(V\mid U)\bigr)$ converges uniformly (as a function of $\mathbf{X}$) to $G^{\mathbf{X}}(V\mid U)$.
 Proof.
-First, observe that
-$\frac{1}{\beta}\log\Bigl{(}\sum\limits_{i=0}^{k}\exp(\beta E_{i})\Bigr{)}-\max\limits_{i}E_{i}$ $=\frac{1}{\beta}\log\Bigl{(}\sum\limits_{i=0}^{k}\exp(\beta E_{i})\Bigr{)}-\frac{1}{\beta}\log\Bigl{(}\exp(\beta\max\limits_{i}E_{i})\Bigr{)}$
-$=\frac{1}{\beta}\log\Biggl{(}\frac{\sum\limits_{i=0}^{k}\exp(\beta E_{i})}{\max\limits_{i}\exp(\beta E_{i})}\Biggr{)}.$
-From here, we bound below and above.
-$\frac{1}{\beta}\log\Biggl{(}\frac{\sum\limits_{i=0}^{k}\exp(\beta E_{i})}{\max\limits_{i}\exp(\beta E_{i})}\Biggr{)}$ $\leq\frac{1}{\beta}\log\Biggl{(}\frac{k\max\limits_{i}\exp(\beta E_{i})}{\max\limits_{i}\exp(\beta E_{i})}\Biggr{)}=\frac{1}{\beta}\log(k).$
-Similarly,
-$\frac{1}{\beta}\log\Biggl{(}\frac{\sum\limits_{i=0}^{k}\exp(\beta E_{i})}{\max\limits_{i}\exp(\beta E_{i})}\Biggr{)}$ $\geq\frac{1}{\beta}\log\Biggl{(}\frac{\max\limits_{i}\exp(\beta E_{i})}{\max\limits_{i}\exp(\beta E_{i})}\Biggr{)}=\frac{1}{\beta}\log(1)=0.$
-This concludes our proof. ∎
-Lemma A.2.
-The maps $\mathcal{P}$ and $\widetilde{\mathcal{P}}$ defined in (1.2) and (1.13) are involutions on $\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}$.
-Proof.
-By [CGS24, Corollary 5.9], the map $\mathcal{W}:\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}\to\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}$ defined by
-$\mathcal{W}(\mathbf{X}_{1},\mathbf{X}_{2})=\bigl{(}\mathfrak{D}(\mathbf{X}_{1},\mathbf{X}_{2}),\mathfrak{T}(\mathbf{X}_{1},\mathbf{X}_{2})\bigr{)},$
-is an involution. Therefore, combined with (A.1), for $(\mathbf{W}_{1},\mathbf{W}_{2})\in\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}$, we have
-$\mathcal{P}(\mathcal{P}(\mathbf{W}_{1},\mathbf{W}_{2}))$ $=(T(D(\mathbf{W}_{2},\mathbf{W}_{1}),T(\mathbf{W}_{2},\mathbf{W}_{1})),D(D(\mathbf{W}_{2},\mathbf{W}_{1}),T(\mathbf{W}_{2},\mathbf{W}_{1})))$
-$=(T(\mathfrak{D}(\mathbf{W}_{2},\tau_{1}\mathbf{W}_{1}),\tau_{-1}\mathfrak{T}(\mathbf{W}_{2},\tau_{1}\mathbf{W}_{1})),D(\mathfrak{D}(\mathbf{W}_{2},\tau_{1}\mathbf{W}_{1}),\tau_{-1}\mathfrak{T}(\mathbf{W}_{2},\tau_{1}\mathbf{W}_{1})))$
-$=(\tau_{-1}\mathfrak{T}(\mathfrak{D}(\mathbf{W}_{2},\tau_{1}\mathbf{W}_{1}),\tau_{1}\tau_{-1}\mathfrak{T}(\mathbf{W}_{2},\tau_{1}\mathbf{W}_{1})),\mathfrak{D}(\mathfrak{D}(\mathbf{W}_{2},\tau_{1}\mathbf{W}_{1}),\tau_{1}\tau_{-1}\mathfrak{T}(\mathbf{W}_{2},\tau_{1}\mathbf{W}_{1})))$
-$=(\tau_{-1}\mathfrak{T}(\mathfrak{D}(\mathbf{W}_{2},\tau_{1}\mathbf{W}_{1}),\mathfrak{T}(\mathbf{W}_{2},\tau_{1}\mathbf{W}_{1})),\mathfrak{D}(\mathfrak{D}(\mathbf{W}_{2},\tau_{1}\mathbf{W}_{1}),\mathfrak{T}(\mathbf{W}_{2},\tau_{1}\mathbf{W}_{1})))$
-$=(\tau_{-1}\tau_{1}\mathbf{W}_{1},\mathbf{W}_{2})$
-$=(\mathbf{W}_{1},\mathbf{W}_{2}).$
-Thus, $\mathcal{P}$ is an involution. Passing to zero-temperature, we get that
-$\widetilde{\mathcal{P}}(\widetilde{\mathcal{P}}(\mathbf{W}_{1},\mathbf{W}_{2}))$ $=\lim_{\beta\to\infty}\frac{1}{\beta}\mathcal{P}(\beta\widetilde{\mathcal{P}}(\mathbf{W}_{1},\mathbf{W}_{2}))$
-$=\lim_{\beta\to\infty}\frac{1}{\beta}\mathcal{P}(\mathcal{P}(\beta\mathbf{W}_{1},\beta\mathbf{W}_{2}))$
-$=\lim_{\beta\to\infty}\frac{1}{\beta}(\beta\mathbf{W}_{1},\beta\mathbf{W}_{2})$
-$=(\mathbf{W}_{1},\mathbf{W}_{2}),$
-where the penultimate equality follows from $\mathcal{P}$ being an involution and the rest follow from the uniform convergence of $\mathcal{P}$ via Lemma A.1. ∎
+This follows as a direct consequence of Lemma A.3, which is itself a standard fact, since there are only finitely many multi-paths between any two sets of points. ∎
+We now prove Theorems 1.8-1.10.
+Proof of Theorem 1.8..
+To show that $\widetilde{\mathcal{P}}$ is an involution, we take a limit from the positive-temperature case:
+$\widetilde{\mathcal{P}}(\widetilde{\mathcal{P}}(\mathbf{X}_{1},\mathbf{X}_{2}))$ $=\lim_{\beta\to\infty}\frac{1}{\beta}\mathcal{P}(\beta\widetilde{\mathcal{P}}(\mathbf{X}_{1},\mathbf{X}_{2}))$
+$=\lim_{\beta\to\infty}\frac{1}{\beta}\mathcal{P}(\mathcal{P}(\beta\mathbf{X}_{1},\beta\mathbf{X}_{2}))$
+$=\lim_{\beta\to\infty}\frac{1}{\beta}(\beta\mathbf{X}_{1},\beta\mathbf{X}_{2})$
+$=(\mathbf{X}_{1},\mathbf{X}_{2}),$
+where the penultimate equality follows from $\mathcal{P}$ being an involution and the rest follow from the uniform convergence of $\mathcal{P}$ via Lemma A.3.
+Next, for $\beta>0$, define the operators $\mathcal{P}_{k}^{\beta}:(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}\to(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$ by their action:
+$\quad\,\mathcal{P}_{k}^{\beta}(\ldots,\mathbf{X}_{k-1},\mathbf{X}_{k},\mathbf{X}_{k+1},\mathbf{X}_{k+2},\ldots)$ (A.2)
+$=(\ldots,\mathbf{X}_{k-1},\frac{1}{\beta}T(\beta\mathbf{X}_{k},\beta\mathbf{X}_{k+1}),\frac{1}{\beta}D(\beta\mathbf{X}_{k},\beta\mathbf{X}_{k+1}),\mathbf{X}_{k+2},\ldots).$
+As an immediate consequence of Theorem 1.1(ii), we see that $\mathcal{P}_{k}^{\beta}\mathcal{P}_{k+1}^{\beta}\mathcal{P}_{k}^{\beta}=\mathcal{P}_{k+1}^{\beta}\mathcal{P}_{k}^{\beta}\mathcal{P}_{k+1}^{\beta}$.
+By Lemma A.3, we see that $\mathcal{P}_{k}^{\beta}$ converges, as $\beta\to\infty$, to $\widetilde{\mathcal{P}}_{k}$ (uniformly as a function of the input).
+Then, we have
+$\widetilde{\mathcal{P}}_{k}\widetilde{\mathcal{P}}_{k+1}\widetilde{\mathcal{P}}_{k}=\lim\limits_{\beta\to\infty}\mathcal{P}_{k}^{\beta}\mathcal{P}_{k+1}^{\beta}\mathcal{P}_{k}^{\beta}=\lim\limits_{\beta\to\infty}\mathcal{P}_{k+1}^{\beta}\mathcal{P}_{k}^{\beta}\mathcal{P}_{k+1}^{\beta}=\widetilde{\mathcal{P}}_{k+1}\widetilde{\mathcal{P}}_{k}\widetilde{\mathcal{P}}_{k+1}.\qed$
+Proof of Theorem 1.9..
+As in Theorem 1.4, we take $\sigma\in\mathcal{S}_{\mathbb{Z}}$ and $U=\bigl(a_{i},b_{i})_{i\in\llbracket 1,m\rrbracket}\bigr)$, and $V=\bigl((c_{i},d_{i})_{i\in\llbracket 1,m\rrbracket}\bigr)$ satisfying $(U,V)\in\mathcal{W}^{\sigma}$.
+To prove Theorem 1.9, we must show that $G^{\mathbf{X}}(V\mid U)=G^{\widetilde{\mathcal{P}}_{\sigma}\mathbf{X}}(V\mid U)$.
+By Theorem 1.4 and definition of the polymer model with parameter $\beta>0$ (1.9), we have
+$Z^{\beta,\mathbf{X}}(V\mid U)=Z^{\beta\mathbf{X}}(V\mid U)=Z^{\mathcal{P}_{\sigma}(\beta\mathbf{X})}(V\mid U)=Z^{\beta(\frac{1}{\beta}\mathcal{P}_{\sigma}(\beta\mathbf{X}))}(V\mid U)=Z^{\beta,\mathcal{P}_{\sigma}^{\beta}(\mathbf{X})}(V\mid U),$ (A.3)
+where in the last step, we simply define $\mathcal{P}_{\sigma}^{\beta}(\mathbf{X}):=\frac{1}{\beta}\mathcal{P}_{\sigma}(\beta\mathbf{X})$.
+As mentioned in the proof of Theorem 1.8, each $\mathcal{P}_{k}^{\beta}$ converges uniformly to $\widetilde{\mathcal{P}}_{k}$ as a function of the input.
+By writing $\mathcal{P}_{\sigma}^{\beta}$ as a composition of these operators and using induction, we have that $\mathcal{P}_{\sigma}^{\beta}$ converges to $\widetilde{\mathcal{P}}_{\sigma}$.
+By Lemma A.4, we know that $\frac{1}{\beta}\log\bigl(Z^{\beta,\mathbf{X}}(V\mid U)\bigr)$ converges to $G^{\mathbf{X}}(V\mid U)$ uniformly as a function of $\mathbf{X}$.
+Hence, using (A.3) in the middle equality below, we have
+$G^{\mathbf{X}}(V\mid U)=\lim_{\beta\to\infty}\frac{1}{\beta}\log Z^{\beta,\mathbf{X}}(V\mid U)=\lim_{\beta\to\infty}\frac{1}{\beta}\log Z^{\beta,\mathcal{P}_{\sigma}^{\beta}(\mathbf{X})}(V\mid U)=G^{\widetilde{\mathcal{P}}_{\sigma}\mathbf{X}}(V\mid U).\qed$
+Before proving Theorem 1.10, we first need prove that $\widetilde{\mathcal{P}}$, restricted in its domain and range to nonnegative entries, is an involution.
+Then we can find analogous Burke properties in the geometric and exponential cases.
 Lemma A.5.
@@
 Proof.
-Since $\mathcal{P}$ is an involution on $\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}$, it suffices to show that $(\mathbf{X}_{1},\mathbf{X}_{2})\in\mathbb{R}^{\mathbb{Z}_{N}}_{\geq 0}\times\mathbb{R}^{\mathbb{Z}_{N}}_{\geq 0}$ implies $\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2})\in\mathbb{R}_{\geq 0}^{\mathbb{Z}_{N}}$ and $\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})\in\mathbb{R}^{\mathbb{Z}_{N}}_{\geq 0}$.
+Since $\widetilde{\mathcal{P}}$ is an involution on $\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}$ by Theorem 1.8, it suffices to show that $(\mathbf{X}_{1},\mathbf{X}_{2})\in\mathbb{R}^{\mathbb{Z}_{N}}_{\geq 0}\times\mathbb{R}^{\mathbb{Z}_{N}}_{\geq 0}$ implies $\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2})\in\mathbb{R}_{\geq 0}^{\mathbb{Z}_{N}}$ and $\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})\in\mathbb{R}^{\mathbb{Z}_{N}}_{\geq 0}$.
 Fix any $i\in\mathbb{Z}_{N}$, and take $\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{R}^{\mathbb{Z}_{N}}_{\geq 0}$.
-Set $Y_{i}=X_{2,i+1}-X_{1,i}$ for all $i\in\mathbb{Z}_{N}$.
+Set $Y_{i}=X_{1,i+1}-X_{2,i}$ for all $i\in\mathbb{Z}_{N}$.
 We observe that
@@
 $\leq 0\vee\left(Y_{i}+\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}\right)$
-$\leq 0\vee\left(X_{1,i}+\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}\right)$
-$\leq X_{1,i}+\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}.$
+$\leq 0\vee\left(X_{1,i+1}+\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}\right)$
+$\leq X_{1,i+1}+\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}.$
 Similarly, we have that $Y_{[i,j]}=\begin{cases}-Y_{i-1}+Y_{[i-1,j]},&\text{ if }j\in\mathbb{Z}_{N}\setminus\{i-1\}\\ Y_{[i-1,i-2]},&\text{ if }j=i-1\end{cases}$
@@
 By the definitions in (1.13), we see that $\widetilde{T}_{i}(\mathbf{X}_{1},\mathbf{X}_{2})\geq 0$ and $\widetilde{D}_{i}(\mathbf{X}_{1},\mathbf{X}_{2})\geq 0$ for all $i\in\mathbb{Z}_{N}$. ∎
-Lemma A.4.
-Let $\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{R}^{\mathbb{Z}_{N}}$. Then,
-  1. (i)
-    $\mathfrak{s}(T(\mathbf{X}_{1},\mathbf{X}_{2}))=\mathfrak{s}(\mathbf{X}_{1})$
-  2. (ii)
-    $\mathfrak{s}(D(\mathbf{X}_{1},\mathbf{X}_{2}))=\mathfrak{s}(\mathbf{X}_{2})$
-  3. (iii)
-    $\mathfrak{s}(\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2}))=\mathfrak{s}(\mathbf{X}_{1})$
-  4. (iv)
-    $\mathfrak{s}(\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2}))=\mathfrak{s}(\mathbf{X}_{2})$
-Proof.
-These all follow immediately from the definitions (1.2) and (1.13) since the sums are telescoping. ∎
-Lemma A.5.
-[CGS24, Lemma 5.10] Let $\mathbf{X}_{1},\mathbf{X_{2}}\in\mathbb{R}^{\mathbb{Z}_{N}}$. Then, for $i\in\mathbb{Z}_{N}$,
-  1. (i)
-    $D_{i}(\mathbf{X}_{1},\mathbf{X_{2}})+T_{i}(\mathbf{X}_{1},\mathbf{X_{2}})=X_{1,i}+X_{2,i}$
-  2. (ii)
-    $e^{-D_{i}(\mathbf{X}_{1},\mathbf{X_{2}})}+e^{-T_{i+1}(\mathbf{X}_{1},\mathbf{X_{2}})}=e^{-X_{1,i}}+e^{-X_{2,i+1}}$
-  3. (iii)
-    $\widetilde{D}_{i}(\mathbf{X}_{1},\mathbf{X_{2}})+\widetilde{T}_{i}(\mathbf{X}_{1},\mathbf{X_{2}})=X_{1,i}+X_{2,i}$.
-Proof.
-(i)-(ii) were proved in shifted form in [CGS24]. (iii) is the zero-temperature limit of (i). ∎
-Lemma A.6.
-[CGS24, Lemma B.3] Let $\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{R}^{\mathbb{Z}_{N}}$.
-  1. (i)
-    If $(\mathbf{U}_{1},\mathbf{U}_{2})=\mathcal{P}(\mathbf{X}_{1},\mathbf{X}_{2})$, then $\mathcal{P}(\tau_{1}\mathbf{X}_{1},\tau_{1}\mathbf{X}_{2})=(\tau_{1}\mathbf{U}_{1},\tau_{1}\mathbf{U}_{2})$.
-  2. (ii)
-    If $(\mathbf{U}_{1},\mathbf{U}_{2})=\widetilde{\mathcal{P}}(\mathbf{X}_{1},\mathbf{X}_{2})$, then $\widetilde{\mathcal{P}}(\tau_{1}\mathbf{X}_{1},\tau_{1}\mathbf{X}_{2})=(\tau_{1}\mathbf{U}_{1},\tau_{1}\mathbf{U}_{2})$.
-Lemma A.7.
-The determinant of the Jacobian of the transformation
-$\mathcal{P}:(\mathbf{X}_{1},\mathbf{X}_{2})\to(T(\mathbf{X}_{2},\mathbf{X}_{1}),D(\mathbf{X}_{2},\mathbf{X}_{1}))$
-has absolute value $1$.
-Proof.
-This is shown in [CGS24, Proposition 5.12].
-The version of the periodic Pitman transform used there is a shifted version of the map we use, which amounts to swapping columns and rows of the Jacobian matrix, leaving the absolute value of the determinant unchanged.
-∎
-Lemma A.8.
-For $\mathbf{X}_{1},\mathbf{X}_{2},\mathbf{X}_{3}\in\mathbb{R}^{\mathbb{Z}_{N}}$,
-$D(\mathbf{X}_{1},D(\mathbf{X}_{2},\mathbf{X}_{3}))=D\bigl{(}D(\mathbf{X}_{1},\mathbf{X}_{2}),D(T(\mathbf{X}_{1},\mathbf{X}_{2}),\mathbf{X}_{3}).$
-Proof.
-We know from [CGS24, Proposition 5.13] that
-$\mathfrak{D}(\mathbf{X}_{1},\mathfrak{D}(\mathbf{X}_{2},\mathbf{X}_{3}))=\mathfrak{D}\bigl{(}\mathfrak{D}(\mathbf{X}_{1},\mathbf{X}_{2}),\mathfrak{D}(\mathfrak{T}(\mathbf{X}_{1},\mathbf{X}_{2}),\mathbf{X}_{3}\bigr{)}.$ (A.2)
-By (A.1) and Lemma A.6, we have
-$\tau_{1}\mathfrak{D}(\mathbf{X}_{1},\mathbf{X}_{2})=\mathfrak{D}(\tau_{1}\mathbf{X}_{1},\tau_{1}\mathbf{X}_{2}).$ (A.3)
-Then, we have the following:
-$\quad D\bigl{(}D(\mathbf{X}_{1},\mathbf{X}_{2}),D(T(\mathbf{X}_{1},\mathbf{X}_{2}),\mathbf{X}_{3})\bigr{)}$
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-as desired. ∎
-Appendix B Proofs of the zero-temperature analogues in Theorems 1.10-1.12
-Because polymer models have a strong connection to last-passage percolation (LPP), our algebraic results will naturally generalize to zero-temperature equivalents, through limiting arguments.
-In these limits, the $(+,\times)$ algebra is swapped for the $(\max,+)$ algebra, and our maps reflect this change.
-As introduced in [CGS24], the ‘correct’ map to look at here is $\widetilde{D}$, the uniform limit as $\beta$ goes to infinity of $\frac{1}{\beta}D(\beta\mathbf{X}_{1},\beta\mathbf{X}_{2})$, as well as $\widetilde{T}$, which is reached through an identical limit.
-In this section, we use our algebraic results about the behavior of the polymer model under the periodic Pitman transform (as well as the aforementioned limiting arguments) to prove analogous facts about LPP under the zero-temperature periodic Pitman transform.
-The following intermediate result is needed.
-Lemma B.1.
-For any $k\in\mathbb{N}$ and any pair of $k$-tuples of ordered pairs $U=\bigl{(}(a_{i},b_{i})\bigr{)}_{i\in\llbracket 1,k\rrbracket}$, $V=\bigl{(}(c_{i},d_{i})\bigr{)}_{i\in\llbracket 1,k\rrbracket}$, as $\beta\to\infty$, $\frac{1}{\beta}\log\bigl{(}Z^{\beta,\mathbf{W}}(V\mid U)\bigr{)}$ converges uniformly (as a function of $\mathbf{W}$) to $G^{\mathbf{W}}(V\mid U)$.
-Proof.
-This follows as a direct consequence of Lemma A.1 since there are finitely many multi-paths between any two sets of points. ∎
-Note that in the special case where $U,V$ are points, this shows that the single-path polymer partition function converges to single-path LPP.
-Also, because the convergence to LPP is uniform on the weights, we know that the zero-temperature limit is well-behaved up to weights which depend on $\beta$.
-If the weights $\mathbf{W}_{\beta}$ converge to $\mathbf{\widetilde{\mathcal{W}}}$, then $\frac{1}{\beta}\log(Z^{\beta,\mathbf{W}_{\beta}}(V\mid U))$ will converge to $G^{\mathbf{\widetilde{\mathcal{W}}}}(V\mid U)$.
-With this in mind, we can now prove Theorem 1.10.
-Proof of Theorem 1.10..
-As in Theorem 1.4, we take $k_{1},k_{2},\ldots,k_{m}\in\mathbb{Z},U=\bigl{(}a_{i},b_{i})_{i\in\llbracket 1,m\rrbracket}\bigr{)}$, and $V=\bigl{(}(c_{i},d_{i})_{i\in\llbracket 1,m\rrbracket}\bigr{)}$ satisfying $(U,V)\in\Psi$ and, for all $i,j\in\llbracket 1,n\rrbracket$, $a_{i}\neq k_{j}+1$ and $c_{i}\neq k_{j}$.
-To prove Theorem 1.10, we must show that $G^{\mathbf{W}}(V\mid U)=G^{\widetilde{\mathcal{P}}_{k_{n}}\ldots\widetilde{\mathcal{P}}_{k_{1}}\mathbf{W}}(V\mid U)$, where the operators $\mathcal{P}_{k}$ are defined in 1.9, with $T,D$ replaced by $\widetilde{T},\widetilde{D}$.
-Now, for $\beta>0$, let $\mathcal{P}_{k}^{\beta}:(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}\to(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$ be operators defined by their action:
-$\quad\,\mathcal{P}_{k}^{\beta}(\ldots,\mathbf{W}_{k-1},\mathbf{W}_{k},\mathbf{W}_{k+1},\mathbf{W}_{k+2},\ldots)$ (B.1)
-$=(\ldots,\mathbf{W}_{k-1},\frac{1}{\beta}T(\beta\mathbf{W}_{k+1},\beta\mathbf{W}_{k}),\frac{1}{\beta}D(\beta\mathbf{W}_{k+1},\beta\mathbf{W}_{k}),\mathbf{W}_{k+2},\ldots).$
-By Theorem 1.4, we know that $Z^{\mathbf{W}}(V\mid U)=Z^{\mathcal{P}_{k_{m}}\ldots\mathcal{P}_{k_{1}}\mathbf{W}}(V\mid U)$.
-Then, by definition of the polymer model with inverse temperature $\beta$ (1.8), for all $\beta>0$,
-$Z^{\beta,\mathbf{W}}(V\mid U)=Z^{\beta\mathbf{W}}(V\mid U)=Z^{\mathcal{P}_{k_{1}}(\beta\mathbf{W})}(V\mid U)=Z^{\beta(\frac{1}{\beta}\mathcal{P}_{k_{1}}(\beta\mathbf{W}))}(V\mid U)=Z^{\beta,\mathcal{P}_{k_{1}}^{\beta}(\mathbf{W})}(V\mid U).$ (B.2)
-By induction, we have $Z^{\beta,\mathbf{W}}(V\mid U)=Z^{\beta,\mathcal{P}_{k_{n}}^{\beta}\ldots\mathcal{P}_{k_{1}}^{\beta}\mathbf{W}}(V\mid U)$.
-By Lemma B, we know that $\frac{1}{\beta}\log\bigl{(}Z^{\beta,\mathbf{W}}(V\mid U)\bigr{)}$ converges to $G^{\mathbf{W}}(V\mid U)$ uniformly on the weights, so $G^{\mathbf{W}}(V\mid U)$ is preserved under the limiting operator $\lim\limits_{\beta\to\infty}\mathcal{P}_{k_{n}}^{\beta}\ldots\mathcal{P}_{k_{1}}^{\beta}$.
-We claim that each $\mathcal{P}_{k}^{\beta}$ converges to $\widetilde{\mathcal{P}}_{k}$ uniformly as a function of the weights, which implies by induction that
-$G^{\mathbf{W}}(V\mid U)=G^{\lim\limits_{\beta\to\infty}\mathcal{P}_{k_{m}}^{\beta}\ldots\mathcal{P}_{k_{1}}^{\beta}\mathbf{W}}(V\mid U)=G^{\widetilde{\mathcal{P}}_{k_{m}}\ldots\widetilde{\mathcal{P}}_{k_{1}}\mathbf{W}}(V\mid U),$ (B.3)
-proving the theorem. The necessary uniform convergence is shown component-wise:
-$\quad\,\frac{1}{\beta}D_{i}(\beta\mathbf{X}_{1},\beta\mathbf{X}_{2})-\widetilde{D}_{i}(\mathbf{X}_{1},\mathbf{X}_{2})$
-$=X_{2,i+1}+\frac{1}{\beta}\log\Biggl{(}\frac{\sum_{j\in\mathbb{Z}_{N}}e^{\beta Y_{(i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{\beta Y_{(i-1,j]}}}\Biggr{)}-(X_{2,i+1}+\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}-\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]})$
-$=\Bigl{(}\frac{1}{\beta}\log\bigl{(}\sum_{j\in\mathbb{Z}_{N}}e^{\beta Y_{(i,j]}}\bigr{)}-\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}\Bigr{)}-\Bigl{(}\frac{1}{\beta}\log\bigl{(}\sum_{j\in\mathbb{Z}_{N}}e^{\beta Y_{(i-1,j]}}\bigr{)}-\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]}\Bigr{)}.$
-By Lemma A.1, we know that $\frac{1}{\beta}\log(N)\leq\frac{1}{\beta}D_{i}(\beta\mathbf{X}_{1},\beta\mathbf{X}_{2})-\widetilde{D}_{i}(\mathbf{X}_{1},\mathbf{X}_{2})\leq\frac{1}{\beta}\log(N)$, proving uniform convergence.
-An identical argument holds for $T_{i}-\widetilde{T}_{i}$.
-∎
-We next show the braid relations pass to zero-temperature.
-Proof of Theorem 1.11..
-Defining $\mathcal{P}_{k}^{\beta}$ as in (B.1), we see that $\mathcal{P}_{k}^{\beta}\mathcal{P}_{k+1}^{\beta}\mathcal{P}_{k}^{\beta}=\mathcal{P}_{k+1}^{\beta}\mathcal{P}_{k}^{\beta}\mathcal{P}_{k+1}^{\beta}$, as an immediate consequence of Theorem 1.6.
-As noted in the proof of Theorem 1.10, $\mathcal{P}_{k}^{\beta}$ converges to $\widetilde{\mathcal{P}}_{k}$ uniformly.
-This implies the braid relation for $\widetilde{\mathcal{P}}_{k}$, since we have
-$\widetilde{\mathcal{P}}_{k}\widetilde{\mathcal{P}}_{k+1}\widetilde{\mathcal{P}}_{k}=\lim\limits_{\beta\to\infty}\mathcal{P}_{k}^{\beta}\mathcal{P}_{k+1}^{\beta}\mathcal{P}_{k}^{\beta}=\lim\limits_{\beta\to\infty}\mathcal{P}_{k+1}^{\beta}\mathcal{P}_{k}^{\beta}\mathcal{P}_{k+1}^{\beta}=\widetilde{\mathcal{P}}_{k+1}\widetilde{\mathcal{P}}_{k}\widetilde{\mathcal{P}}_{k+1}.\qed$
-To prove Theorem 1.12, we first need analogous Burke properties in the geometric and exponential cases.
-Proposition B.2 (Geometric Burke property).
-Take $\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{Z}_{\geq 0}^{\mathbb{Z}_{N}}$ to be independent sequences of independent random variables such that $X_{1,i}\sim\text{Geom}(p_{1}q_{i})$ and $X_{2,i}\sim\text{Geom}(p_{2}q_{i})$ for all $i\in\mathbb{Z}_{N}$.
+Proposition A.6 (Geometric Burke property).
+Take $\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{Z}_{\geq 0}^{\mathbb{Z}_{N}}$ to be independent sequences of independent random variables such that $X_{1,i}\sim\text{Geom}(p_{1}q_{i})$ and $X_{2,i}\sim\text{Geom}(p_{2}q_{i})$ for all $i\in\mathbb{Z}_{N}$, where $p_{m}q_{i}\in(0,1)$ for $m\in\{1,2\}$ and $i\in\mathbb{Z}_{N}$.
 Then,
-$(\mathbf{X}_{1},\mathbf{X}_{2})\overset{d}{=}(\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2}),\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})).$
+$(\mathbf{X}_{1},\mathbf{X}_{2})\overset{d}{=}(\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2}),\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2})).$
 Proof.
@@
 $=\left(\prod\limits_{i=1}^{N}(1-p_{1}q_{i})(1-p_{2}q_{i})\right)p_{1}^{\mathfrak{s}(\mathbf{k}_{1})}p_{2}^{\mathfrak{s}(\mathbf{k}_{2})}\prod_{i=1}^{N}q_{i}^{k_{1,i}+k_{2,i}}$
-$=\left(\prod\limits_{i=1}^{N}(1-p_{1}q_{i})(1-p_{2}q_{i})\right)p_{1}^{\mathfrak{s}(\widetilde{D}(\mathbf{k}_{2},\mathbf{k}_{1}))}p_{2}^{\mathfrak{s}(\widetilde{T}(\mathbf{k}_{2},\mathbf{k}_{1}))}\prod_{i=1}^{N}q_{i}^{\widetilde{D}_{i}(\mathbf{k}_{2},\mathbf{k}_{1})+\widetilde{T}_{i}(\mathbf{k}_{2},\mathbf{k}_{1})}$
-$=\prod\limits_{i=1}^{N}(1-p_{1}q_{i})(p_{1}q_{i})^{\widetilde{D}_{i}(\mathbf{k}_{2},\mathbf{k}_{1})}\prod\limits_{i=1}^{N}(1-p_{2}q_{i})(p_{2}q_{i})^{\widetilde{T}_{i}(\mathbf{k}_{2},\mathbf{k}_{1})}$
-$=\mathbb{P}(\mathbf{X}_{1}=\widetilde{D}(\mathbf{k}_{2},\mathbf{k}_{1}),\mathbf{X}_{2}=\widetilde{T}(\mathbf{k}_{2},\mathbf{k}_{1}))$
-$=\mathbb{P}(\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})=\mathbf{k}_{2},\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2})=\mathbf{k}_{1})$
-where the first equality follows from the independence of $\mathbf{X}_{1},\mathbf{X}_{2}$ and $X_{m,i},X_{m,j}$ for $m\in\{1,2\}$ and $i,j\in\mathbb{Z}_{N}$ where $i\neq j$; the second equality follows from our definition of the geometric distribution; the third equality follows from Lemmas A.4(iii)-(iv) and A.5(i); and the final equality follows from the bijection in Lemma A.3.
+$=\left(\prod\limits_{i=1}^{N}(1-p_{1}q_{i})(1-p_{2}q_{i})\right)p_{1}^{\mathfrak{s}(\widetilde{D}(\mathbf{k}_{1},\mathbf{k}_{2}))}p_{2}^{\mathfrak{s}(\widetilde{T}(\mathbf{k}_{1},\mathbf{k}_{2}))}\prod_{i=1}^{N}q_{i}^{\widetilde{D}_{i}(\mathbf{k}_{1},\mathbf{k}_{2})+\widetilde{T}_{i}(\mathbf{k}_{1},\mathbf{k}_{2})}$
+$=\prod\limits_{i=1}^{N}(1-p_{1}q_{i})(p_{1}q_{i})^{\widetilde{D}_{i}(\mathbf{k}_{1},\mathbf{k}_{2})}\prod\limits_{i=1}^{N}(1-p_{2}q_{i})(p_{2}q_{i})^{\widetilde{T}_{i}(\mathbf{k}_{1},\mathbf{k}_{2})}$
+$=\mathbb{P}(\mathbf{X}_{1}=\widetilde{D}(\mathbf{k}_{1},\mathbf{k}_{2}),\mathbf{X}_{2}=\widetilde{T}(\mathbf{k}_{1},\mathbf{k}_{2}))$
+$=\mathbb{P}(\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})=\mathbf{k}_{1},\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2})=\mathbf{k}_{2})$
+where the first equality follows from the independence of $\mathbf{X}_{1},\mathbf{X}_{2}$ and $X_{m,i},X_{m,j}$ for $m\in\{1,2\}$ and $i,j\in\mathbb{Z}_{N}$ where $i\neq j$; the second equality follows from our definition of the geometric distribution; the third equality follows from Lemmas 2.1 and 2.2(i); and the final equality follows from the bijection in Lemma A.5.
 ∎
-Proposition B.3 (Exponential Burke property).
+Proposition A.7 (Exponential Burke property).
 Let $\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{Z}_{\geq 0}^{\mathbb{Z}_{N}}$ be independent sequences of independent random variables such that $X_{1,i}\sim\text{Exp}(a_{1}+b_{i})$ and $X_{2,i}\sim\text{Exp}(a_{2}+b_{i})$ for all $i\in\mathbb{Z}_{N}$ where $a_{m}+b_{i}>0$ for $m\in\{1,2\}$ and $i\in\mathbb{Z}_{N}$.
 Then,
-$(\mathbf{X}_{1},\mathbf{X}_{2})\overset{d}{=}(\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2}),\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})).$
+$(\mathbf{X}_{1},\mathbf{X}_{2})\overset{d}{=}(\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2}),\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2})).$
 Proof.
@@
 $\frac{1}{n}\widetilde{T}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime})=\widetilde{T}\Bigl(\frac{\mathbf{X}_{1}^{\prime}}{n},\frac{\mathbf{X}_{2}^{\prime}}{n}\Bigr),\quad\text{and}\quad\frac{1}{n}\widetilde{D}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime})=\widetilde{D}\Bigl(\frac{\mathbf{X}_{1}^{\prime}}{n},\frac{\mathbf{X}_{2}^{\prime}}{n}\Bigr).$
-By Proposition B.2, we know that $(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime})\overset{d}{=}(\widetilde{T}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime}),\widetilde{D}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime}))$.
+By Proposition A.6, we know that $(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime})\overset{d}{=}(\widetilde{D}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime}),\widetilde{T}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime}))$.
 Since $\max$, addition, and subtraction are continuous, $\widetilde{D}$ and $\widetilde{T}$ are continuous maps.
 We then have that
-$\Bigl{(}\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2}),\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})\Bigr{)}\overset{d}{=}\lim\limits_{n\to\infty}\frac{1}{n}\Bigl{(}\widetilde{T}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime}),\widetilde{D}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime})\Bigr{)}\overset{d}{=}\lim\limits_{n\to\infty}\frac{1}{n}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime})\overset{d}{=}(\mathbf{X}_{1},\mathbf{X}_{2}).\qed$
+$\Bigl(\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2}),\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2})\Bigr)\overset{d}{=}\lim\limits_{n\to\infty}\frac{1}{n}\Bigl(\widetilde{D}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime}),\widetilde{T}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime})\Bigr)\overset{d}{=}\lim\limits_{n\to\infty}\frac{1}{n}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime})\overset{d}{=}(\mathbf{X}_{1},\mathbf{X}_{2}).\qed$
 Proof of Theorem 1.10.
-The proof is the same as the proof of Theorems 1.7 and 1.8: replace the use of Theorem 1.4 with Theorem 1.10 and replace the Burke property in Propositions 4.1 with the Burke property in Propositions B.2 and B.3.
+The proof is the same as the proof of Theorems 1.5 and 1.6: replace the use of Theorem 1.4 with Theorem 1.9 and replace the Burke property in Propositions 4.1 with the Burke property in Propositions A.6 and A.7.
 ∎